A Complete Bibliography of Publications in the Journal of Cell Biology: 1970–1974

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

15 May 2024 Version 1.01

Title word cross-reference

0 [ZTB74]. 1 [DC73]. 13 [ZTB74]. 2 [UBKB74]. 40 [KSN73, NLB73a, WWvdN74]. 5 [BH70, GN73, LA74, MW71, RS74b, SD71c, SJ72, TdPP72]. 80 [DU72, RAU71]. ^{+} [DDP73, DPD73]. ^{13} [WABG74]. ^{14} [SIS74]. ^{2′} [TML73]. 2+ [BSGS72, CGL74, Eds74]. ^{3} [BL70, BL71b, Dal71, GAGSP73, GDH73, HSH^{+}71, KH71a, Len72b, NLB73b, PP73b, RVHWU72, Raw73, SS72c, SJ72, WL71a, WL74]. ^{55} [Fed74]. ^{6} [GHP73, TML73]. _{0} [Pot71]. _{1} [FW73, FR74c, KK70, LS74, SS72c]. _{12} [CJC70, GJDVC74, JC73]. _{2} [APW^{+}73, FW73, KFP70, PR71a]. _{3} [MHBM73]. _{5} [RFBB74]. α [AJ73, BG74b, EDE72, KS74a, OS74]. b_{5} [IKY70]. c [DKC71]. g [CS74a]. γ [OK71, RL74].

-Acetate [ZTB74]. -amanitin [EDE72, KS74a]. -aminobutyric [OK71]. -Aminouracil [SD71c]. -amylase [AJ73, BG74b]. -Angstrom [DU72]. -Anti [SC72b]. -ATPase [DDP73, DPD73]. -Aziridinyl [DC73].

-Bromodeoxyuridine [BH70, LA74, MW71, RS74b]. ’-Cyclic [RF73]. -Deoxyglucose [UBKB74]. -Dibutyryl [TML73]. -Fetoprotein [LS74]. -glutamyl [RL74]. -glycerophosphate [OS74]. -Hydroxytryptamine [GN73, TdPP72]. -Hydroxytryptamine- [SJ72]. -Induced [NLB73a, RAU71]. -labeled [DKC71]. -Specific [Pot71, BSGS72]. -Tetradecanoyl-Phorbol- [ZTB74]. -Transformed [KSN73, WWvdN74].

1 [DSP70]. 100 [GG72]. 12- [ZTB74].

2- [GDH73, RVHWU72, Raw73]. 20 [GL70, GL71, LP70, TL70b]. 28S [Coo73, WLH74].

3 [FR72]. 3’-5’-Cyclic [FR72]. 37 [DBP70]. 37Rc [BD72]. 3T3 [PC71, PTF73, EMJT74].

450 [IKY70]. 48/ [RAU71]. 48/80 [KD74].

5 [AM73a, FR72, FR74c, MHSW71, RM72, Sto73a, Sto74, TML73, Wid72]. 5’-Adenosine [RM72, TML73]. 5-Bromodeoxyuridine [WOSE73]. 5’-Cyclic [AM73a, FR74c, MHSW71]. 5’-Monophosphate [Sto73a, Sto74]. 5’-Nucleotidase [Wid72]. 537A [KD74]. 5S [Mil73, WLH74].

60S [KNM74].

80S [KNM74].

A-23187 [KD74]. A-Binding [Wel74]. A-Cells [EH70]. A-Derivatized [PW73]. A-Rich [Wak74]. Aberrant [CP70]. Abnormal [FH70]. Abrupt [GG74a]. Absence [CP70, FR74a, LP71, PR70a, Phi74, SP71b]. Absorption [Dav73, FC72, OL73, WCDI72]. Absorptive [BW70a, NN72, SJ71]. ac [GL70, GL71, LP70, TL70b]. ac-20 [GL70, GL71, LP70, TL70b]. Acanthamoeba [BO72, BK73, BK74, CB71a, CB71b, Chl71, MB71a, RS74b, Rud71, SP73, SW72b, UWWK71]. Accessibility [Saa74]. Accumulation [HB71a, IS74, SSD^{+}74, UB74]. Accurate [GW71]. Acellular [Ett72]. Acer [SIS74]. Acetabularia [SL72a, WB70]. Acetate [LK71, ZTB74]. Acetylcholine [PP73a]. Acetylcholinesterase [LDP74, TB70, TBS71]. Acheta [Cav72]. Achromatic [DHN73]. Acid [AM70, BB72c, BBC72, BOW70, CH73, Det71, DGW74, FBBdD72, FF70b, HKF71, HBK70, IS74, JSF70, KFJ70, LG70, Lam72, LK71, Meb72, MS70c, Mül72, MP70d, OU70, PE72a, Rae70, RW70a, RR71, RW73, Rud71, Tid70, WH70, OK71, PH71]. Acid-Secreting [FF70b]. Acid-Thymine [OU70]. Acidic [BS72, SB72]. Acids [BY74a, LG70, RE72]. Acinar [BY71, CJP72, Han70b, Mel74]. Acquisition [KLJ^{+}74]. Acriflavin [Str72]. Acriflavine [LG70, Stu71]. Acriflavine-Induced [Stu71].

Acriflavine-Phosphotungstic [LG70]. Acrolein [CAC^{+}71]. Acrosiphonia [HW74]. Acrosomal [FR74b, THIM73]. Actin [CG73a, THIM73]. Actin-Like [CG73a]. Acting [Aro71]. Actinobolina [HC73]. Actinomycin [BL71a, BD72, BIS72, POB74, SG71, SG72, WM73a]. Actinomycin-D [BL71a]. Action [AI70, MGDT74, MGT74, RB71, SM71, WM73b, KS74a]. Actions [Mag73]. Active [CBS73, DB74, O’B72, PW73, SW72a, Ste73]. Actively [GG74a]. Activities [AESK71, CGP70, DMF74]. Activity [BL71a, BMS71, BK73, BDN72, CB72, CG70, Fah70a, FBBdD72, GG72, GC74, HPHB72, HRS74, HLP73, KLJ^{+}74, KK73b, KTC72, LB72, Meb72, Nag72, NB72, PBK74, SFK74, SG74, SKMH70, VSF70, WE70, WR74a, WL70, RL74]. Acyltransferases [HB71b]. Adaptation [BB73a, RO74]. Adapted [HC74]. Adenine [CSH71, GHP73, SS73a, SS73b]. Adenohypophysis [MJ74, TKTVJ73]. Adenosine [AM73a, FR72, FR74c, GG72, GG74a, GC74, GS73, LN74, MHSW71, Nag72, RM72, RF73, SSGJH73, Sto73a, Sto74, TML73, WP72a, WT73a]. Adenovirus [RO71]. Adenylate [MJ74]. Adhesion [Cul74, DBM74, GMS73, RdS73]. Adhesive [Arm70, RMR71a]. Adipose [BMS71, Cus70a, Cus70b, CR70b]. Adjacent [She73]. Administration [GHN74, KL72, WL74]. Adrenal [AH73, FG70a, FB70, FG72a, Mag72, NSM70, YSS^{+}72]. Adrenalectomized [LdH71]. Adrenergic [IF71]. adrenocortex [IKY70]. Adrenocortical [BB73b, KPH^{+}74]. Adult [BHM73, BB73b, BDW71, CJMB73, Gri70a, OO71, SK71, SWS74, UCB71]. Aedes [AS71b, SFK74]. Aegypti [AS71b, SFK74]. Aerobic [Key73, Wat72]. Aerogenes [Har70b]. Affected [Nan71a]. Affecting [RBK74]. After [AYB^{+}72, BD72, BLH74, BH70, BW73b, Coo73, DHN73, DFDF73, FCH73, Fed74, GAGSP73, GHN74, GP73a, Gri70c, HL74a, ISWS71, KLJ^{+}74, McC74, Nea73, SG72, SB73b, WL71a, WL74, WABG74, WL70]. Against [TKTVJ73]. Agar [HD70b]. Age [SSK74]. Agent [ZTB74]. Agents [ERB74, PR71b]. Agglutinability [WP74b]. Agglutination [NB73]. Agglutinins [NS74]. Aggregates [Arm70, LC71, Ros71a]. Aggregating [Uge71, ZTB74]. Aggregation [BKG73, EB74a, MG71, PH70]. Aging [Rud71, SS72a, SD71a]. Agranular [TH73b]. Albert [Pal71]. Albicans [BdN74, DC72]. Albumin [BT72, BCG72, GE70, IP73, RC72, SLWZ70]. Alcoholic [LK71]. Aldehyde [Bod70, Gho71]. Aldehyde-Fixed [Gho71]. Aldehydes [VHB^{+}72]. aleurone [AJ73]. Alga [BB73a, CPH74, EG71, HW74, MP70b, PA73, Pul74, RG72, SFMD72, WABG74]. Algae [BFK^{+}70]. Alkaline [BB73d]. Alleles [TGW73]. Alloantigens [NHS71, STM^{+}74]. Along [SY74, WPB72]. Alpha [BWC74, LS74]. Alteration [AYB^{+}72, AB71b, BT70a, WBSW74]. Alterations [BG74a, HD73b, PWT71]. Altered [RCBS70, Stu71]. Alveolar [FW73, KAI^{+}74, SSGJH73]. amanitin [EDE72, KS74a]. Amebae [Fli71]. Amebas [CS70]. Ameloblasts [WL71a]. Americana [ED70, dBTL73]. Amine [GHN74]. Amines [NB74]. Amino [CH73, DGW74, WL70]. Amino-Triazole [WL70]. Aminoacyl [SKMH70]. Aminobutyric [HBK70, Lam72, OK71]. Aminohippuric [BOW70]. Aminotransferase [PV70a, PV70b, SG74, TT71, WMG74]. Aminouracil [SD71c]. Ammocoetes [You71b]. Ammonia [Lee70]. Ammoniacal [MM70a]. Amniotic [OL73]. Amoeba [AAWD74, PI70, PK71, PD70b, Szu71, WF70]. Amoeboid [TCMA73]. Among [WB70]. Amp [MR74, Mag73, NFR^{+}74, WP74b]. Amphibian [DF74, Hop72, HF70, Hor72, HM74b, MI71a, MI71b, MI71c, Ros74, Sch70, SWD74, TH72]. Amplification [KG74, MABL71]. Amplitude [AT70]. Ampullary [SB73b]. Amputated [SB73a]. amylase [AJ73, BG74b]. Anaerobic [RS74a, Wat72]. Analoga [VT72]. Analyses [HFG73]. Analysis [BPHBK74, Bol74, BC71, CJP72, DJ74, DW70, EKP74, FW72, GMdlTFGGF74, GR70, HF70, JFdBJ73, LK72, OS72, Phi72a, SRA74, SSD^{+}74, Ste73, Str71a, ZFG74]. Analytical [ACBW^{+}74, ACWTS^{+}74, BACF^{+}74, BACTS^{+}74, TBT74]. Analyzing [BP73a]. Anchoring [BDW71]. And/Or [PKT74]. Androgen [NSM70]. Androgen-Treated [NSM70]. Anesthetic [HS72a, TBW74]. Anesthetic-Induced [HS72a]. Angelae [GM72a]. Angiogenesis [CSFC72]. Angiotensin [RB71]. Angstrom [DU72]. Animal [Dal72, HS71b, Lak73, MLD70, Rub71]. Animals [PS71]. Anionic [Nic73b, NP73, QP73]. Annulate [FSF72, Mau70]. Anolis [RP72]. Antagonism [Sto73a]. Antagonist [CA74]. Anterior [HF73, PP73b, Pel74, TVG70]. Anther [Spi70]. Anthoceros [MK70]. Anti [SC72b]. Antibodies [NMS71, Rod73, TKTVJ73, RFBB74]. Antibody [AAG^{+}74, KDC71, KS72, Rod70, SY71, TKTVJ73, DKC71]. Antibody-Forming [KS72]. Antiferritin [Fed74]. Antigen [Dal72, KDC71, WR74b]. Antigenic [BKG73, KP74]. Antigens [GM72b, PO73, PP74, SY71]. Antimicrotubular [ERB74]. Antimotility [SBRC70]. Anucleolate [Mil73]. Anuran [Dec74]. Aphid [OC72]. Aplysia [AGS74, AGTS74, LLP72]. Apparatus [Ben70, BL71b, BLH74, BB71, BR73, CMY70, CMPF71, Cos74, CMN71, CSR^{+}74, DHN73, GE70, HMC70, JA74, MHM^{+}70, Nic71, NNQH71, RW71a, SJ71, SdlI72, Tre70, WF70]. Apparatus-Rich [CMY70, CMPF71, MHM^{+}70]. Appearance [AA74, FCH73, PP73a, SWSK71, TB70, TBS71]. Application [GFK72, KR70]. Applied [KMS73]. Appraisal [GS70]. Arabinoside [JM74]. Araldite [TR73]. Arbacia [And70, LA70a, LA70b, SA70, TG70]. Architecture [QP73, SC72a]. Areas [Iwa71, SP74]. Arginine [GG71]. Arising [Com71]. Arm [Fri71]. Arm-Bearing [Fri71]. Arms [Tuc74]. Arranged [Chl71]. Arrangement [Har70a, Kay70, KHS74, She73]. Array [Woo74]. Arrays [GWS70, HC73, LR74, Roh74]. Arrested [Lak73]. Artemia [SGL73]. Arterial [HBM73, Iwa71]. Arthropoda [BDR70]. Artificial [SA70]. aryl [WMG74]. Ascaris [RS74a]. Ascites [HRWL71, KW74b, PKT74, RdS73]. Ascospore [LM70c]. Aspartate [PV70a, PV70b]. Aspects [BK70, CGL74, VDC71]. Asphyxia [PR71b]. Assay [KR70, RMR71a]. Assembly [CB71a, CB71b, DF74, KWR71, TG70]. Associated [All74b, Dan73, Fri71, GWT74, HC73, JEG73, Key73, KL73, MWR72, MBKM70, OD74, SHH73, SS72b, SMS72]. Association [BA74, BB73d, DFDF73, FL70, OHPW74]. Astrocytic [LR74]. Astrocytoma [NB74]. Asymmetry [NS74]. ATP [CWB71, Sum74]. ATP-Induced [Sum74]. ATPase [DDP73, DPD73]. Attached [Cul74]. Attachment [AB74, Gri73b, LKB71, YY74]. Attractive [Aro71]. Aurelia [Cum72]. Australian [Tam72b]. Autolysis [DFDF73]. Autonomy [BMD^{+}72]. Autophagic [Det71, SH72a]. Autophagy [Det71]. Autoradiographic [GAGSP73, GHN74, GP73b, JHK73, MC73, PP73b, Pel74, Raw73, WABG74]. Autoradiographs [BP73a]. Autoradiography [PB73, WHCFB74]. Avian [PP74]. Axenic [KWR71, RW70b]. Axo [Gob71]. Axo-Axonic [Gob71]. Axon [HKF71, LW74b, OHPW74, Ten70]. Axonal [AGS74, Eds74, HS72a, PR71b, PP72, TH73b]. Axonemal [Har70a, Zob73]. Axonemes [SL70]. Axonic [Gob71]. Axons [HL74b, HFK70, SB73a, YSW71]. Axoplasm [SJB70]. Axoplasmic [FBS71, Len72b]. Axostyle [MOL73, McI73, MT73]. Axostyles [BM74]. Aziridinyl [DC73]. Azotobacter [SBKH74]. Azurophil [BB74a, NHB72].

B [Kri72, AP72, AR74, BdL73, BG74a, BP73b, BG74b, CJC70, DGS^{+}74, GJDVC74, Gol72, JC73, KK70, MMM72, OLK^{+}71, PE72b, RFBB74, SBRC70, SCDS73, SDCS74, UBKB74, WP72a, WFO70, dLB74]. B-Treated [BdL73, dLB74]. Baby [Gol72, Gri73b, KKCP73]. Bacillus [MD72, Nan71a, Nan71b, vIG71, FW73]. Back [NLB73a]. Backfilling [AP74]. Backgrounds [BHLT74]. Bacterium [BT70c]. Bacterized [RW70b]. Balb [EMJT74, PTF73]. Balb/ [PTF73]. Balb/c [EMJT74]. Band [EP72, KP71, WWC72]. Banding [MCIJ74]. Bands [Hag70]. Barbipes [WLH74]. barley [AJ73]. Barnacle [HMS73]. Barriers [MPEB71]. Bary [HHC71]. Basal [AB71a, And72, And74, BDW71, Cos74, FD71, Har73, MR70, RC73, VB70, VB72]. Based [CB70b]. Basement [MS70c]. Basic [VT72]. Basis [LM70a, TCMA73, TGW73, WS74a]. Basket [Gob71]. Basophils [HS71a]. Bat [GHN74, Nak71]. Beads [SW72b]. Bean [GBN73, LKMB73, Vig70]. Bearing [Det71, Fri71]. Beef [HBL70, HB71c, Web72]. Beginnings [Pal71]. Behavior [EB72, HOB71, KBC73, RS70]. Bend [WS74a]. Bending [Rob72, Tuc74]. Benign [MHW71]. benzo [WMG74]. Benzpyrene [CB72]. Between [Aro71, BT70b, CF73, CF72, Iwa71, JSF70, Jef72, KFP70, LC72, MHSW71, McI74, MM72, Rub72, Rud71, SHP70, She70, She71b, She73, SD70b, SL72b, Spi70, TSE70, UB70, WWvdN74, WG73, OYC^{+}73]. Bhk [Gol71, Gol72, KKCP73]. Bhk-21 [Gol71, Gol72, KKCP73]. Bidirectional [BTPT71]. Biliary [SSS74]. Bilirubin [RRYT70]. Binding [BEG73, CG73a, DW70, Gil71, HPHB72, How74, KH71a, NB73, OW72, PRP74, SF74b, Sto74, TKTVJ73, Wel74, WW72, WMB74]. Binucleate [Ivá71]. Biochemical [BHdD70, BS70b, BB74a, CGP70, CGL74, CSGB70, DG71, FW72, GKB74, GC74, GK73a, GAGGS71, HFG73, IK71, KMO^{+}70, MMT^{+}74, Mül73, OLK^{+}71, POB74, TGW73, WT71]. Biochemistry [MGHA72]. Biogenesis [BB73c, BB73e, BB74b, GO70a, GO70b, HBG74, HB72b, Hig74, HZM71, SGL73, WHL70]. Biogenic [NB74]. Biological [Pal71]. Biology [Gla74]. Biosynthesis [Hop72, Lam72, LS74, RU70, RCDMM71, SL72a, WWBG74]. Bipolar [WWC72]. Birefringence [BB72a, BB72b, Col71]. Black [CHM72b]. Bladder [CBS72, DGS^{+}74, DGM^{+}74, DC70, GCF71, HK70b, MHW72, SCB72b, VZRE74, WD71]. Blastoderm [LT71]. Blastoderms [CI73, HI71]. Blastulae [TG70]. Block [HKF71]. Blocked [HN73]. Blood [Bak72, HAT71, MBW71, NMS71, NHS71, PG70b, Ros71b, SSP74, Ste74, Str71a, TM70, VBD70, ZF70, ZFG74, ZTB74]. Blowflies [SS72a]. Blue [BCGL74, BB73a, CGL74, EG71, WABG74]. Blue-Green [BB73a, EG71]. Bodies [AB71a, And74, BL71b, Con71, Cos74, Fli71, FS74, FD71, GR73, Har73, KJ70, ME71, MBKM70, RW70b, RC73, SP71b, Wak74]. Body [And72, Det71, FGO71, KWR71, Lar70, LM71, MR70, Nak71, Ros72, She71b, SMS72, YCD70]. Body-Like [FGO71]. Boletus [McL71]. Bombix [TO70]. Bone [BDH^{+}74, NU74]. Border [BW70a, TC70]. Borders [BS70b, CL72]. Both [BK74, CB70b, HK70b, TBL70]. Bound [ABS73, GF72, IP73, MM74a, RBK74, UBKB74]. Bovine [AS71a, KMO^{+}70, O’B72, Saa74, SS70b, UI72]. Brain [BBHH74, CT72, CBCT74, GJ72, HBL70, IBK74, KK73a, KHTH70, KLKK73, RKT71, TAKS74, TA70]. Branching [BBC72, Bra73]. Breakdown [GNAB70]. Breast [PP73a]. Bridges [BF71, BM74, FKZ^{+}71, HMC70, McI74, WT73b]. Bromide [HH71, KGK72, KSN73, Lei71, MHB73, SN71, UB74]. Bromodeoxyuridine [BH70, LA74, MW71, RS74b, WOSE73]. Brown [She71b]. Brucei [Stu71]. Brush [BS70b, BW70a, CL72]. Buds [Far72]. Buffer [Fah74]. Bulk [Goo74]. Bull [Sum74]. Bullata [RM71]. Bundles [LS73]. Bunodosoma [DC74].

C [SIS74, NTKM73, EMJT74]. Ca [BSGS72, CGL74, Eds74]. Calcified [BDM73]. Calcitonin [HRS74, DKC71]. Calcium [BW72, Bas71, BCGL74, BBW74, Ett72, HL74b, KD74, MCR74, OW72, OHPW74, PRP74, SB73a, SDSN71, dMHM74]. Calcium-Containing [BBW74, HL74b, MCR74]. Calcium-Free [SDSN71]. Calcium-Induced [KD74]. Californica [AGS74, AGTS74, LLP72]. Callinectes [BCGL74, CGL74]. Calmette [FW73]. Campaniform [MCE71]. Cancer [ML74, She70]. Cancerous [FC73b]. Candida [BdN74, DC72, VSHA72]. Capillaries [CF74, Lea71, SSP72, SSP73, SSP74]. Capillary [SP71a, VB70]. Capsular [RG72]. Carbamate [CPH74]. Carbohydrate [NWL71]. Carbohydrates [GA73, SJ71]. Carboxykinase [vRBvWW74]. Carcinoma [KPH^{+}74]. Cardiac [JSJ71, JLS73, LS71, MPAB70, MPBA73, Mat73, NSG71, Prz71, PLRK72, WH70]. Cardiacum [CPH74]. Cardiolipin [JGR^{+}71]. Caridina [Nak72]. Carolinensis [Com73, RP72]. Carotid [YCD70]. Carteri [Pal73]. Cartilage [AS71a, AESK71, EAK70, HD70b, KO70, KSR^{+}70, KESA71]. Castellanii [BO72, BK73, MB71a, RS74b, Rud71, SP73, UWWK71]. Castor [LKMB73, Vig70]. Cat [EH70, Gob71, PW74]. Catalase [Fah74, HF74, RASP74]. Catecholamine [MR71a]. Catecholamine-Containing [MR71a]. Catenated [WMKJ73]. Catesbeiana [VBD70]. Cation [SSD^{+}74]. Cationic [MGS^{+}74]. Cations [CSH71, GR74, KLT71, KTT72, MBP72, TLS70, TK71, TKT72]. Caudal [Nag72]. Caudatum [All71]. Caused [CHM72b]. Cavernata [DC74]. Cecropia [HAT71, WT73b]. Cell [AM73a, ABS73, AAWD74, APW^{+}73, Arm70, AP72, ABN70, Bak72, BW72, Bal70, BL70, Ben70, BLH74, BKG73, BY74a, BS70c, BS72, BB73c, BB73e, BB74b, CH73, CJP72, CB70b, Cla70, CSH71, CS73, CGB71, CMN71, Cus70a, Cus70b, CR70b, DHN73, DH74, DB74, DC72, DMF74, EB72, EB74c, ER74a, ER74b, FH70, FSH71, FG72a, FR72, FR74c, GP73a, GE70, Gol71, GM72b, GS73, Gri73b, GMS73, GWT74, HPHB72, HRWL71, HLI72, Han70b, HW73, HKKSH72, HRW72, HN73, HC72, IP73, Iwa71, JSF70, Jef72, JSP74, KG74, KL74a, KCR70, KFP70, KK73b, KPF74, KT72, Kra71, KS71, LSO72, Lag73, LW74a, LF72, LCRM71, LC72, LW74b, LA74, MCBBTG71, Mag72, MSM74, MPBA73, MMS^{+}72, MHBM73, MP70a, MR71a, MCB71, MCB73, MGHA72, MMLM74]. Cell [MBW71, NO71, Nak72, Nea73, Nic73a, NMS71, NHS71, NS74, NLB73a, NDAT71, NLB73b, NB73, OW74, OKY72, PD70a, PWT71, PR70a, PR71a, PR72, PKT74, PP71a, PP71b, PP74, PdWW72, PHDN73, PPF73, RWB70, RI71, RHW71, RK71, Ros71c, RMR71b, RE73, SB70a, SMD73, SG70a, SHDH74, SG74, SWJ^{+}71, SSHS74, STM^{+}74, SH74, SN71, SH71, SYW71, Ste74, SB72, SW72a, Sto73a, Sto74, TM70, TCK72, Tok73, TKT72, VDC71, VT74, VB72, WWC72, WHG73, WWvdN74, WE70, WR74a, WR74b, WMB74, ZS72, vRBvWW74, Rub71]. Cell-Cycle [KT72]. Cell-Rich [APW^{+}73]. Cell-Surface [GM72b, RMR71b]. Cell-To-Cell [RE73]. Cells [AH73, AJ74a, AJ74b, AM73b, AAG^{+}74, BT70a, BS70a, BD72, BT70b, BL70, Ben70, BL71b, BN71, BY71, BF71, BY74b, BDH^{+}74, BT71, BHS72, BHM73, BB73b, Blo70, BW70a, BS72, BW70b, BC71, BR73, CUMB71, CB72, CAC^{+}71, CF73, CI73, CG73a, CWB71, CJMB73, CGM73, CDMR71, Com71, CKS74, CS73, CGB71, Dal72, Dav73, DK73, DBP70, DW70, Eli72, EH70, EWBP71, Fah70a, FH70, FS74, FF70a, FF70b, FSF72, FL70, FC72, Gal74, GSH71, GN73, Gho71, GCF74, Gol72, GM72b, GS73, GAP70, HSH^{+}71, HT73, HS71b, HMT71, HDM74, HT74, HFK70, HF73, HD70b, HLP73, HWP73, ISWS71, IS74, Ivá71, JP71a, JP71b, JB72, KD74, KAI^{+}74, Kes70, Key73, KO70, KGK72, KKCP73, Kis74, KLJ^{+}74, KSN73, Kni73, KW74b, KS72, KP72]. Cells [KH71a, KTC72, Lak73, Lan73, LDP74, LM70b, LS73, LKB71, LP72, Lei71, LP71, LT71, LC71, LCRM71, LT70, LM73, LK72, MM70a, MC73, MGS^{+}74, ML74, MPAB70, MYWS72, MMS^{+}72, ML71, MJ70, MCB71, MCB73, MG71, MG73, Mel74, MM72, MF71, MR74, MGDT74, MGT74, MLD70, NWL71, NB74, NN72, OP74, OLK^{+}71, OPL72, OM73, OTE73, Pad70, Pad74, PD70a, PC71, PB73, PV72, PG70b, PR70b, PP73b, PKT74, POB71, Per72, POB74, Pla72, PE72a, PFSB72, PW73, PPF73, PTF73, Pot71, PS71, PHPO74, RdS73, RM71, RO71, RCBS70, RS70, RR71, RPK70, RAU71, RE72, Ros71a, RW73, RGR71, Rub72, RE73, RRYT70, SMD73, SJ71, SRA74, SD70a, SH70, SS70a, SM71, She71a, She70, She71b, SB70b, SHH73, Sim70, SS72b, SRS73]. Cells [SD70b, Spi70, SHM74, SS72c, SB72, SA71, SA72, SIS74, SO72b, SA73a, SA73b, SK70b, SJ72, Szu71, TA70, TBL70, TML73, TT71, TG70, TL70a, TSE70, UB70, VBD70, VR74, WRE71, Wak74, WWvdN74, War74, WR72, Wat72, WH70, WP74a, WCF72, WM73a, WS74c, WP72b, WOSE73, Wri73, YSW71, YCD70, YM73, YY74, YA74, Yoo70, ZLK74, AJ73, DKC71, EDE72, KS74a, OYC^{+}73, PH71, WMG74]. Cellular [BR72, CS70, EB74b, GK73a, GK73b, GWE74, MJP71a, MJP71b, MJP71c, MC72, NTKM73, OKY72, PGHK72, RB71, WT70]. Center [Aro71, OTE73, Tan73]. Central [DD70, PM72b, SJB70, SM73, Stu71]. Centrifugation [Bow73, BBHH74, CS74b, WP72b]. Centrifuged [And70]. Centriole [Dir71, Rat72, RP73]. Centrioles [AB71a, FD71, KBH73, MS70a, Prz71]. Centrosome [MK70]. Centrosomes [McL71]. Cerebellar [Gob71, HD73b, Lan73, SH70, SL72b]. Cerebellum [WMB74]. Cerebral [GJKL73, YM73, YY74]. Cerebrum [MBCP71]. Cerevisiae [Gho71, HPHB72, KPF74, LM70c, MR71b, PGR72, Wat72]. Certain [THIM73]. Cervical [MHW71]. Cessation [BW73b]. Chaetopterus [IBK74]. Chain [HT73]. Chang [WRE71]. Changes [AY70, BB72d, Blo70, CH73, CHM72b, Coh71, Dan73, DGM^{+}74, DSP70, ER74a, ER74b, GK73a, GAGGS71, GJKL73, GWT74, HI71, HRW72, IS74, JB73, Kis74, Lan73, LB72, LT74, LA71, MB71a, NLB73b, NB72, OPT74, PPF73, RHW71, SS72a, SB74, SSK74, SS72c, Tre70, Vig70, WR72, Wei72, Wri73, Yoo70, ZS72]. Channels [HFK70]. Chaos [Com73]. Character [SB74]. Characteristics [KBC73, NLB73a, PP71a, Uge71, WT71]. Characterization [AAWD74, BS70b, BS74, Bod70, BB74a, CGMD74, CSGB70, CT74, Cum72, DOP74, Det71, FF70b, GR73, HPHB72, HSWD72, Hol71, HH74, IK71, KSN73, KMS73, Lak73, MJ70, MR71a, PFGG73, SHDH74, Sim70, SF74a, SM74b, SF74b, UI72, VZRE74, dMHM74]. Charge [HSH70]. Charophyceae [CG71c]. Chemical [AAWD74, KPF74, LG70, MM70b, MMLM74, MSGE73, Nan71a, SM74b, DSP70]. Chemistry [Per73c]. Chemography [Mul71]. Chemostat [Har70b]. Chemotaxis [GR74]. Chick [AESK71, BPA74, BL71a, BN71, BDN72, CWMH71, CI73, Cor73, GG71, Hag70, HS70, HI71, HLI72, IK71, KDR74, KO70, KLJ^{+}74, KMS73, MM70a, MW71, NFR^{+}74, OW74, PW71, PP73a, Per73c, PP74, PdWW72, RK73, Shi71, SH71, TH73b, Tre70, TC71, Tre71, WH70, WCH71, RL74]. Chicken [JLS73, LG72, MBKM70, MBS72a, MBS72b]. Chinese [CWB71, KS74a, LT70, LM73, NDAT71, OCWS74, POB74, PPF73, RE73, SS72c, TL70a, Wri73]. Chironomus [EDE^{+}71, LWD^{+}72, RR71, EDE72]. Chlamydomonas [BL74, CBS73, Cos74, CR70a, FGO71, Goo71, HBG74, Hoo72, HS73, HN73, MM74a, MMLM74, RF73, SH73, SH72b, WWBG74, WCBR72, WCR72, DSP70, GL70, GL71, LP70, TL70b]. Chloramphenicol [ABBA72, KGK72, KSE73, SJG72]. Chlorella [Pul74]. Chlorophyll [WWBG74]. Chloroplast [BL74, CBS73, CG71c, GEP71, GS74, GP73b, GO70a, GO70b, GL70, Goo71, GS71b, HBG74, Hoo72, HS73, LP70, MM74a, MP70c, NS70, PA73, SJG72, SH72b, TL70b, GL71]. Chloroplasts [CG70, HK70a, MWR72, MP70b, Nic71, OM74, SL72a, SS74b]. Cho [LM73]. Cholesterol [DHWHW71, RU70, RVHWU72, Raw73, SSS74]. Cholesterol-1 [RVHWU72]. Cholinergic [BBW74]. Cholinesterase [SV73]. Chondrocytes [DKHL74, Hua74]. Chondrodystrophic [SFS71]. Chondrogenesis [Min73, NU74]. Chondroitin [Hua74]. Choroid [QWT73]. Chromatic [BB73a]. Chromatin [BA74, CS74b, DU72, JEG73, Lun72, NB72, PR72, PS70, SRA74, UP72, UP73]. Chromosomal [BG72, BS72, LM70a, LWD^{+}72, ST70, SB72]. Chromosome [BR73, GWT74, IWS72, Sch73]. Chromosomes [BC72, GL74, ISWS71, KT74b, OCWS74, RB74a, RR71, SCPH73, WLH74]. Chrysophycean [BFK^{+}70]. Chylomicrons [BMS71, BMS73]. Cilia [AS72, Boq70, FSH71, Ran74, SL70, TBW74]. Ciliary [GS72a, Gri70c, Har73, SS74a, Tam72a, WS74a]. Ciliate [All73, HP73, OBJ72, Tuc74]. Ciliatea [HC73]. Ciliated [And74, Dir71, SG70a]. Ciliogenesis [Dir74, MR70, SD74]. Cinematographic [Kri72]. Circadian [BS73]. Circular [KW74b, MWR72, WMKJ73, WKCF73]. Cisternae [Mol71]. Cistrons [BC72, GL74]. Clam [Wei72]. Class [FL70]. Classes [Bod70, SWJ^{+}71]. Claude [Pal71]. Clavata [Hea74]. Cleavage [BW72, SY74]. Cleaving [Szo70]. Cleft [CT74]. Clonal [KTC72, RRYT70, SSK74, TBL70]. Cloned [NFR^{+}74]. Clones [SH74, SHM74]. Close [UB72]. Closely [Woo74]. Clot [SBRC70]. Cnidarian [DC74]. CO [KFP70]. Coat [BL70, Ben70, She71a]. Coated [KK73a, MB71b, YY74]. Coats [McK74]. Cobalt [WE73]. Coccid [Rob72]. Cockroach [MCE71]. Coeruleus [HP73]. Coexistence [HR71]. Colcemid [Aro71, HPHB72, SL73b]. Colcemid-Binding [HPHB72]. Colcemid-Sensitive [Aro71]. Colchicine [AI70, Dan72, Dan73, Gil71, HRS74, KH71a, MGD71, SF74b, SSS74, Tur70, UBKB74, WW72, WM73b, BG74b]. Colchicine- [KH71a]. Colchicine-Binding [WW72]. Colchicine-Induced [SSS74]. coli [WFO70, HBS70, HN70b, Hen71, HB72a, SHNB72]. Collagen [ERB74, EB72, HD73a, PBK74, RET70, Tre71, WL74]. Collagenous [TC71]. Collagens [THT74]. Colloidal [WP72b]. Colonies [ZFG74]. Color [RP72]. Column [Dec74]. Columnar [BL70, Ben70, BL71b]. Combination [SPAR73]. Commitment [GWE74]. Commune [WE73]. Communication [TSE70]. Comparative [Bel71, HBL70, NSG71, Phi72a, SJ71, Str71a]. Comparing [TCK72]. Comparison [KMO^{+}70, KMS73, NF71, PD70a, PV72, RPK70, SD70b]. Compartment [IS74]. Compartmentalization [Klo70]. Compensation [LM70a]. Complementary [BB73a]. Complex [AB74, LG70, LT71, Mas73, Mau71, SMS72, You73]. Complexes [CT72, Fed74, HHC71, KH71b, MEW72, Mau70, MPL71, SM73, SC72b, vIG71]. Component [KPF74]. Components [Bry70b, DKHL74, DC72, EB74c, ME71, Zob73, EDE72]. Composition [Bou71, CJC70, GJDVC74, JC73, MJP71a, MJP71b, MJP71c, MC72, MMLM74, SB74, TAKS74, Zob73, DSP70]. Compound [RAU71]. Concanavalin [CT74, Gri73b, Lut73, NLB73b, NB73, PW73, RBK74, Sto74, UBKB74, WWvdN74, Wel74, WP74b, WMB74, Wri73]. Concentration [MS70c, SB73a, Str71a]. Concentrations [MGD71]. Concept [MMLM74]. Concerning [Dal72]. Concomitant [ASO^{+}71, EMJT74]. Condensation [GWT74]. Condensed [JEG73]. Condensing [JP71a]. Conditional [HN73]. Conditions [HC74]. Conductive [MPAB70]. Cone [Ten70, You71a]. Cones [Bun73, PB74, VHG74, YSW71]. Configuration [HB71c]. Conformation [EMJT74, Zob73]. Congenital [FG70a]. Conjugated [NMS71, NS74]. Conjugates [AAG^{+}74, GA73]. Conjugation [Klo70]. Connections [AA74, AP74, Spi70]. Connectives [Sch73]. Constant [HC74]. Constriction [CWT^{+}73]. Contact [CGB71, FG72a, HF71, MCB71, MCB73, Nam72, Pas72, RE73, WR74a]. Contact-Inhibited [CGB71, MCB71, MCB73]. Contact-Mediated [Nam72]. Contacts [PHDN73]. Containing [BBW74, DC72, HL74b, Mag72, MGS^{+}74, MCR74, MR71a, SS74b, WKCF73]. Content [HLP73, JGR^{+}71, KDS73, KS74b, Lam72, PKT74, SB74, SN71, WP72a]. Continuously [MBCP71]. Contractile [All73, BK73, HP73, KAI^{+}74, LDdV72, PD70b, RMM^{+}70, Sch72, Szo70, TCMA73]. Contraction [DSS72, Hag70, PD70b, SDSN71, SS73a, SS73b]. Contrast [LKM71, OYC^{+}73, TR73]. Control [BDN72, GS73, Gro71, HS73, IWM^{+}74, KL74a, KL74b, Lef74a, Lef74b, Mil73, PC71, PR70a, WWBG74]. Controlled [BS70c, Bod70, KCR70, SB70a]. Controlling [BN71, TC70]. Conventional [BdN74]. Conversion [JP71a]. Cooled [BF71]. Coprophila [BR72]. Cord [FBS71, OW74, SPAR73, VHG74]. Cord-Ganglia-Muscle [SPAR73]. Core [HP70]. Cornea [KFHF74, LN74, TC71, THT74]. Corneal [BS73, HD73a, PBK74, Tre70, Tre71]. Correlates [PAB71, WPB72]. Correlation [CF72, MHSW71]. Corresponding [Szo70]. Cortex [All71, FB70, FG72a, Gob71, GJKL73, KLKK73, Mag72, PBGCP74, SL72b, Str71a, YSS^{+}72, YM73]. Cortical [DC74, Kan70, LA70a, Mab73, Szo70]. Corticosteroid [LdH71]. Cortisol [AM73b, BL71a, GAP70, SM71]. Cortisone [KL72]. Coulter [Sim70]. Counter [Sim70]. Counting [GB72]. Coupling [AB71b, BT70b, DSS72, GS74, GP72, MLGML71, OM74, PAB71, She71b]. Couplings [SWS74]. Course [AW72, Cla70]. Crab [BCGL74, CGL74, Hin70, LS71, VT72]. Crassa [CW71, CLBW72, HZM71, KL73, LL72, WL71b]. Crayfish [FBS71, JA74, Per70, PR71b, Per73a, Per73b]. Creatine [IWM^{+}74]. Cricket [Cav72, Kay70, SM73]. Cristae [HC74, You71b, ZKH73]. Crithidia [RW72]. Critical [Gri70c]. Cross [BF71, FKZ^{+}71, GWS70]. Cross-Bridges [BF71, FKZ^{+}71]. Cross-Striated [GWS70]. Crude [KK73a]. Cruentum [GL72]. Crustacean [AP74, Bas71, POB71, SV73]. Cryoprotectant [MGK74]. Cryoprotectant-Induced [MGK74]. Cryoprotectants [PFSB72]. Cryptophyceae [GEP71]. Crystal [ABN70, Mag72, NO71]. Crystal-Containing [Mag72]. Crystalline [Bry70b, MP70b]. Crystallization [MBS72a, MBS72b]. Crystals [KH71a, MS70b, Mas71, OW73b]. Culbertsoni [RM72]. Cultivated [AR74, RA73]. Cultural [SD71a]. Culture [APW^{+}73, BS70a, BN71, BDH^{+}74, BT71, BT72, BHM73, Bra73, Bun73, CUMB71, CI73, CJMB73, Cor73, DB74, Far72, GCF74, HRS74, KL74b, KH71a, LF72, LS74, Lef74a, Lef74b, LC71, MGS^{+}74, MCS71, NFR^{+}74, NU74, OLK^{+}71, Pla72, RWB70, SB73a, SH71, TT71, TVG70, WR72, YM73, YY74, WMG74]. Cultured [AM73b, BT70a, BL71a, CG73a, EF73, FL70, HI71, HT74, Hua74, KH70, LCRM71, ML71, MBCP71, MLD70, NB74, OM73, Per73c, PdWW72, PE72a, RE72, RW73, SMD73, SO72a, SW71, SM71, SSHS74, SIS74, TH73b, TBT74, VR74, WP74a, YSW71, RL74]. Cultures [HLI72, Har70b, KNY74, LP72, MP73a, MP73b, MP73c, MP70a, PP73a, PBK74, RK73, Rud71, SM74a, SS74b, SPAR73, vRBvWW74]. Cupric [WHG73]. Cuprizone [TH73a]. Currents [JN74]. Cuticle [DC73]. Cyanoblast [Fah70b]. Cyclase [MJ74]. Cycle [BS72, CH73, FH70, GSH71, Gri73a, GWT74, HKF71, HN73, Jef72, KR70, KK73b, KT72, LS74, MMS^{+}72, MHBM73, MP70a, NDAT71, NLB73b, PD70a, PWT71, PR70a, PR71a, PR72, PPF73, SD71a, SB72, TCK72, VT74, WE70, WR74a, WR74b, OS74]. Cycle-Dependent [NLB73b]. Cycle-Specific [GWT74]. Cycles [SW71]. Cyclic [AM73a, CK73, ER74a, ER74b, FR72, FR74c, GS73, Mag73, MHSW71, MR74, NFR^{+}74, RM72, RF73, SL73a, SSGJH73, Sto73a, Sto74, WP74b]. Cyclic-Nucleotide [SL73a]. Cycloheximide [LG72]. Cyprinodon [HC74]. Cyst [GdB73]. Cysts [RM72, SGL73]. Cytochalasin [AP72, AR74, BdL73, BG74a, BP73b, DGS^{+}74, Gol72, Kri72, MMM72, MGDT74, MGT74, PE72b, RI71, SBRC70, SCDS73, SDCS74, UBKB74, WP72a, dLB74, BG74b]. Cytochalasin-B [Kri72]. Cytochemical [BK73, FBBdD72, GC74, HW73, HFK70, PV70a, RWB70, SFMD72, SK70b, TBB74, Vig70, Wid72, WF70]. Cytochemistry [Fah74, FBP74, NNQH71]. Cytochrome [SN71, DKC71, IKY70, RFBB74]. Cytokinesis [BdL73, HW74, Kri72, RI71, dLB74]. Cytolocalization [BC72]. Cytological [And70, GDH73, LA70a, RM71, SA70, SCPH73]. Cytology [Mül73]. Cytomorphometry [GFK72]. Cytophagocytosis [Cha73]. Cytoplasm [HG71, RR71, WG73, ZLK74]. Cytoplasmic [CBS73, Cla70, CS70, CLSA74, EI71, FSF72, Gol71, Gri70a, Hin70, KNM74, KWR71, Mel74, PI70, Roh74, RYR71, Sea72, SKMH70, Str71a, Szo70, TG70, Tuc74, UCB71, WE73, WW71]. Cytosine [JM74]. Cytosol [KLJ^{+}74].

Cytotoxicity [SG71].

D [BL71a, BD72, BIS72, MGDT74, MGT74, POB74, SG71, SG72, WM73a]. Dab [NS70]. Damage [BM73, FG70a, HW73]. Danica [GP73b, HP74, SJG72]. Dans [PA70]. Darkness [HC74]. Data [SSP74]. Death [OW74]. Decarboxylase [SFK74]. Deciliated [Ran74]. Deciliation [TBW74]. Decreased [Fli71]. Dedifferentiation [ZS72]. Deferens [Nea73]. Deficiency [DG71]. Deficient [CGP70, KLJ^{+}74, SS74b]. defined [EDE72]. Definition [FF70a]. Degenerating [Dec74, MHB70, RVHWU72]. Degeneration [LSO72, SB73a, SB73b]. Degenerative [SS72a]. Degradation [Coo73, SH72a, WP74a]. Degranulation [Bai73, RAU71]. Dehydratase [IP73]. Dehydrogenase [BS71]. Deiters [MM74b]. Demonstration [Aro71, CA71, DMF74, KPF74, LWD^{+}72, Roh74, SFMD72, SJ72, TT71, TM70, UB70]. Denaturation [WKG72]. Dendrites [WMB74]. Dendritic [VHG74]. Denervation [SB73b]. Dense [BL71b, Det71, EI71, HL74b, MHB73]. Densities [CBCT74]. Density [BCC74, Kni73, SP73, WE73, WP72b, Rub71]. Density-Inhibited [SP73]. Dentin [WL73a]. D’Enzymes [PA70]. Deoxycytidine [OU70, OTE73]. Deoxyglucose [UBKB74]. Deoxyribonuclease [DMF74, WE70]. Deoxyribonucleic [JSF70, KFJ70, OU70, PE72a, RW70a, WH70]. Dependence [BBC72, HAT71, LW74a, WR74a]. dependencies [OYC^{+}73]. Dependent [KT72, NLB73b, SDSN71, VT74]. Depends [SMD73]. Depleted [HRW72, NH74, RHW71]. Depletion [CHM72a, IS74, PW74]. Deposition [DKHL74, MPBA73, WL73a, WS74c]. Deposits [dMHM74]. Depression [CK73]. Deprivation [RK73]. Deprived [LM73, PV72]. Derivation [HS70]. Derivatized [PW73]. Derived [DJ74, NSG71, RR71, SM74a, SG74, SH74, ZFG74]. Dermo [KK70]. Desacetyl [AI70]. Description [Chr71]. Desmosome [Ash70, DRE70, Fri71, Ove73]. Desmosome-Like [Fri71]. Desmosomes [SM74b, SM74c]. Desquamation [KT72]. Detected [NLB73b]. Detection [BIS72, GA73, Lat74, MC73, PP73b]. Determination [BKK73, SSHS74, Sim70]. Determined [BOW70, Len72b, MF71, SHM74]. Developing [BG74a, Dir71, GFK72, HFG73, KO70, Klo70, LaV73, LT71, RS70, Spi70, Str71b, SdlI72, TSE70, UCB71, VHG74]. Development [BN71, BW70a, BDN72, CG71c, GK73b, GAGGS71, GJKL73, GBN73, HM72, JGR^{+}71, KG74, KDR74, Len70, Len72a, LA71, LM70c, MBKM70, MHT72, Mil73, MBCP71, OPT74, PW71, PC73, SFK74, SWR70, Tre70, Wak74, WT70]. Developmental [GB70, HI71, Vig70]. Dextrans [CF74, SP71a]. Diabetes [BHLT74]. Dialyzed [HS71a]. Diaminobenzidine [NS70]. Diaphragm [LG73, RK74b]. Dibutyryl [GS73, MR74, Sto73a, Sto74, TML73]. dichroitics [ZBSW70]. Dictyosome [Mol71]. Dictyostelium [BKG73, CCA70, OKY72]. Differences [KFP70, KTC72]. Different [BHLT74, MP70a, PD70a, SWJ^{+}71, SW70, TCK72]. Differential [EB74c, JM74, Nic73a, SG71, WM73a]. Differentiated [KL74b, LP72, LS74, Lef74a, Lef74b, SO72b, WH70]. Differentiating [Dec74, Kel71, LDP74, PP73a, RS70]. Differentiation [Cla70, Far72, FSH71, GS71b, GFK72, KS72, KNY74, LT71, LSP71a, LSP71b, NTKM73, NBF71, PGHK72, RI71, SGL73, SB74, SH71, VT74, WPB72, YSS^{+}72]. Differentiations [FF74, SS74a]. Diffraction [Bur70]. Diffusion [HM74b, PE72a]. Digestion [DRE70, PN70, SG73]. Dilated [FR74a]. Dilution [BKK73]. Dilution-Vinblastine [BKK73]. Dimensional [And72, JA74, NHS71, ST70]. Dimethyl [KKCP73]. Dinoflagellate [OD74, Rae70]. Diphosphate [BDN72]. Diploid [PP71b, SH74]. Direct [CG70, GB72, GCF71, Kan70, WHCFB74]. Directing [GP71]. Direction [HT73]. Disaccharidases [BSG71]. Disaccharides [GG74b]. Disaggregation [FC73a, Lun72]. Disassembly [KDS73, KS74b]. Discharge [JP71a, JP71b]. Discoideum [BKG73, OKY72]. Discontinuity [Gri73a]. Discontinuous [KW74b]. Discs [HOB71, ZR73]. Disk [ME71]. Disks [Row73]. Dispersed [AJ74a, AJ74b, PBGCP74]. Disrupted [WK74a]. Disruption [DHN73, FR74b]. Dissection [MSGE73]. Dissociated [HF73, MP73a, MP73b, MP73c, YM73, YY74]. Dissociation [CBS73]. Distinct [BBHH74, Ros72]. Distinctive [FG72a, PP71a]. Distribution [BOW70, DD70, EDE^{+}71, FCH73, HM74b, JB72, KLT71, KTT72, KP74, Len72b, MPL71, ML71, NHS71, NS74, Per73c, PP74, PS70, PM72b, SJ71, SLWZ70, VR74, WD71]. Distributions [BP73a, DJ74]. Dithiothreitol [WR74c]. Divalent [GR74, MBP72]. Dividing [MM72, ML70]. Division [AM73a, FSH71, GSH71, GS73, JSF70, Jef72, LA74, MCBBTG71, PR72, SD70a, TH73a]. DNA [BG72, BIS72, BH70, Cav72, CG72, CS73, Dal71, DJ74, DU72, GLG70, GG70, GP73b, GHP73, GP71, GMP74, HT73, HMT71, HK70a, HH71, Jef72, JFdBJ73, KG74, KK73b, KW74b, LWD^{+}72, Lat74, Lei71, LT70, LM73, LA74, MWR72, MSM74, MKPV74, MABL71, NB72, PFGG73, RP73, RW72, RK73, SO72a, SSHS74, SSK74, SD70b, SO72b, Stu71, SH72b, TL70a, UB74, VT72, WHCFB74, WL73b, WKG72, WMKJ73, WKCF73, WB70, ZBSW70, vIG71]. DNA-Membrane [vIG71]. DNA-Regions [HK70a]. DNAs [RW71b]. DNase [TPG72]. DNase-Sensitive [TPG72]. DNS [ZBSW70]. DNS-Organisation [ZBSW70]. Do [SW72a, WG73]. Dog [MPAB70]. Dogfish [PG70b]. Domestica [CD73]. Domesticus [Cav72]. Dopa [SFK74]. Dorsal [NNQH71, SWR70, TB70, Ten70]. Dosage [GMP74, LM70a]. Doses [BD72]. Double [Gho71, LG70, SSC73]. Double-Stranded [LG70, SSC73]. Dragonfly [Kes70]. Drone [Per72]. Droplets [SJ71, Wak74]. Drosophila [Bos70, DBW73, HB71a, LM70a, MS70a, OS74, PFGG73, Tok74, ZC71]. Drug [IF71]. Drug-Free [IF71]. Dry [BOW70]. Drying [Gri70c]. Duck [TVG70]. Duct [Bow73]. Ductuli [Nea73]. Ductus [Nea73]. Duodenal [Ben70, BL71b, LG72, MHT72]. Duration [KFJ70]. Durations [FH70]. During [AA74, Bai73, BL74, BR72, BHB73, BN71, Blo70, BdL73, BW70a, BS72, CH73, CBWS74, CG72, DF74, Det71, Esp71, FC72, GJ72, GK73a, GK73b, GAGGS71, Gri73a, GJKL73, GWE74, HM72, HR73, HF71, Jef72, KDR74, KTT72, KT74a, KR70, Kra71, Lag73, LKB71, LS74, Len72b, LA71, LM70c, MI71a, MI71b, MI71c, MHT72, MB71a, MMS^{+}72, MHBM73, ML71, McL71, Mil73, NDAT71, OPT74, PWT71, PR70a, PC73, PPF73, Prz71, SD71a, SY74, SGL73, SB74, SCB72a, SS72c, SD74, SB72, SP73, SW72b, SAE73, TH73a, Tid70, Tre70, VT72, WE70, WT70, WR72, WE73, Wei72, WM72, Yoo70, ZS72, dW71, dLB74, DSP70]. Dye [MCIJ74]. Dye-Nucleoprotein [MCIJ74]. Dynamics [BKG73, HP74, Mel74]. Dynein [GF72]. Dyskinetoplastic [Stu71, RW71b]. Dystrophic [LRL71]. Dystrophy [BH71b].

Early [BW72, CWMH71, CG72, Det71, KG74, Mil73, PC73, Pog72a, Pot71, Tan73, WCH71]. Earthworm [Ros72]. Ebner [Han70a]. Eccrine [SMS72]. Ecdysone [SFK74]. Ecdysone-Mediated [SFK74]. Ecdysterone [RM71]. Echinoderm [THIM73]. Echinosphaerium [Har70a]. Echinostelium [HHC71]. Ectodermal [TG70]. Edulis [AS72]. Eelworm [ZKH73]. Effect [BG74b, CG70, DC70, Fah74, FR72, HT70, HF74, HG71, Hua74, KSE73, Kni73, KNY74, LB72, Lei71, Lut73, LA74, MHBM73, PH71, RM71, RCDMM71, RL74, RS74b, RE73, SSC73, SBRC70, SCDS73, Sto74, WD71, WM73a, Zob73]. Effects [ABBA72, AM73b, BMS73, BD71, BBW74, BR73, CDMR71, CPH74, CSR^{+}74, DG71, DGS^{+}74, DW72, Eds74, EDE72, ERB74, EMJT74, FTCM74, FC72, GDH73, Gib70, Gol72, Goo71, GMP74, HT74, HRS74, HA70, JM74, KZ70, KL72, Kri72, LA70b, MT70, MW71, MGD71, NFR^{+}74, OM73, PV70a, RPK70, SB73a, SJG72, SN71, SPAR73, Str74, SG73, TCK72, TGW73, Tur70, UBKB74, VHB^{+}72, YCD70, GL71]. Efferentes [Nea73]. Egg [BH71c, Kan70, Mab73, SY74, TH72]. Eggs [And70, Aro73, BdL73, CWT^{+}73, Con71, DC74, LA70a, RYR71, Szo70, Wei72, dLB74]. Ehrlich [PKT74]. Either [RK73]. Elaboration [NWL71, WL71a]. Electric [WPB72]. Electrical [BT70b, PR71b, RKT71, She71b, TSE70]. Electrically [PW74]. Electrokinetic [GKB74]. Electromechanical [GP72]. Electron [AP74, BP73a, Bod70, BD71, BH71c, BJ70, Cha73, Chr71, Coh71, Com71, CW74, EAK70, GHN74, Gla74, HHP70, HKF71, HN70a, HL74b, HFK70, HWP73, KD74, Kis74, KNM74, Len72b, LC72, MM70a, MC70, MC73, MHB73, MM74b, MBCP71, Nac72, NWL71, NP71, Pad74, Pal71, PN70, PB73, Per72, PTF73, RB74a, RAU71, RGR71, SS71, SP74, Sha72, Shi71, SSD^{+}74, SBKH74, TKTVJ73, WHCFB74, Wel74, WACBB71, WG73, GL71]. Electron-Dense [HL74b, MHB73]. Electron-Microscope [Len72b]. Electron-Opaque [LC72]. Electrophoretic [RKT71]. Electrophysiology [Bal70]. Electroreceptor [SB73b]. Electrotonic [AB71b, PAB71]. Element [Bry70a, PN70, Stu71, Szo70]. Elements [ED70]. Elicited [BG74a]. Elongating [War74]. Elongation [Blo70, BC71, CR70a, LW74b, PdWW72, WWC72]. Embedded [HWP73, Nic71, OM74, TR73]. Embryo [BL71a, BN71, BB72c, BTLN73, BDN72, GG71, IK71, MW71, NFR^{+}74, Per73c, PP74, RK73, SH71, TB70, Ten70, TBT74, TSE70, YY74, Yoo70, RL74].

Embryogenesis [CWMH71, WCH71]. Embryogeny [CBWS74]. Embryonic [AESK71, BT70b, BT71, BT72, Cor73, GM72b, HS70, HM74a, JM74, KDR74, KK70, MG71, OW74, PP73a, PdWW72, SMD73, STM^{+}74, SHGH72, SO72b, TH73b, TBS71, Tre71, THT74, WR72]. Embryos [BW72, BH72, HLI72, HG71, KB72, MBS72a, MBS72b, RGGG71, SRA74, WH70]. Emerita [VT72]. Emptying [CSH71]. Enamel [WL71a]. Encystment [GK73a, RS74b, Rud71, SP73, SW72b]. End [Ros74]. Endings [BBHH74, Nak71]. Endocrine [Boq70, OLK^{+}71, Rub72]. Endocytosis [AH73, WRE71]. Endogenous [Fah70a, HM72, SK70b, SWSK71]. Endometrial [CA74]. Endoplasmic [All73, All74a, BB72d, FB70, HB72b, Hig74, LSP71a, LSP71b, LKMB73, Nak72, OPL72, SB74]. Endosperm [GB70, LKMB73, Vig70]. Endothelial [GCF74, SS70a, SC72a, SRS73, SJ72, WCF72]. Endothelium [GWS70, HBM73, PBK74, SSP74]. Energized [HB71c]. Energized-Twisted [HB71c]. Energy [Hac72, KPH^{+}74, WT70]. Energy-Linked [Hac72]. Enhancement [BL71a, IF71, WHG73]. Enriched [HBL70]. Enrichment [MJ74]. Entamoeba [KWR71, RW70b]. Entrapped [MGS^{+}74]. Entry [HG71]. Enucleated [MR74]. Envelope [HBS70, HN70b, Hen71, HB72a, KJ70, OD74, SHNB72, Sch70, SHH73, SW70]. Envelope-Associated [SHH73]. Envelopes [Blo70, ISWS71, OCWS74]. environmental [OYC^{+}73]. Enzymatic [DRE70, GR71, HLP73, HC72, PN70, SL70]. Enzyme [AESK71, GS70, KK73b, KDC71, MC71a, RCDMM71, RS74a, TT71, ZHW73]. Enzyme-Labeled [KDC71]. Enzymes [BF70, CMO73, GBN73, HB73, KGK72, KS71, SDCS74, WM72, DDP73, DPD73]. Enzymic [GFK72]. Eosinophilic [BF70]. Eosinophils [HS71a]. Epidermal [CF73, KK70, Kra71, ML74, SM74b, SM74c]. Epidermis [FC73b, MM70b, Uge71]. Epiphyseal [KSR^{+}70, KESA71]. Epithelia [CG73b]. Epithelial [Aue72, BB72c, BBC72, CKS74, DGM^{+}74, EB74b, EWBP71, FC73b, FC72, HMT71, JB72, LSO72, Nak72, QP73, Rub72, SB74, SCB72a, STM^{+}74, THT74]. Epithelial-Mesenchymal [BB72c, SCB72a]. Epithelio [MHT72]. Epithelio-Mesenchymal [MHT72]. Epithelium [Ben70, BY74b, BH71b, BS73, Dir71, EB74c, FR74a, GS71a, HN70a, HD73a, HK70b, HR71, LG72, MPEB71, MHW71, Mul71, NFR^{+}74, OPT74, PdWW72, QWT73, SG70a, Tre70, Tre71, VSF70, WD71]. Epon [TR73]. Epon-Araldite [TR73]. Equatorial [BT70c]. Equilibrium [Ste73]. Eremosphaera [Hol71]. Erythrocyte [EB74a, Nic73b, NP73, PTLL74, Red71, VHB^{+}72, ZLK74, dS72]. Erythrocytes [FCH73, KLJ^{+}74, PTLL74, ZLK74]. Erythroid [CDMR71, PB73, RI71]. Erythromycin [ABBA72]. Erythropoiesis [CWMH71, HI71, MI71a, MI71b, MI71c, MW71, WCH71]. Erythropoietic [CI73, HLI72, MM70a]. Erythropoietin [CDMR71]. Escargot [PA70]. Escherichia [HBS70, HN70b, Hen71, HB72a, SHNB72, WFO70]. Established [MGDT74, MGT74, PP71a]. Establishment [KSN73]. Ester [ZTB74]. Esterase [VR74]. Estrogen [CA74, PW71]. Etch [SRS73]. Etched [CA71, TM70]. Etching [GS74, PFSB72, dSB70]. Ethidium [HH71, KGK72, KSN73, Lei71, MHB73, SN71, UB74]. Etiology [BH71b]. Étude [Pag71]. Eucaryote [GHP73]. Eucaryotic [OO72, OW73a]. Euglena [CJC70, GJDVC74, JC73, WE70]. Euplotes [Klo70]. Events [LA70a, Pog72a]. Evidence [BT70c, Con71, GP73b, GE70, HKF71, HR73, Jef72, KD74, KW74b, POB71, RMR71b, Stu71, TBB^{+}73, WYE71, WB70, Wri73]. Examination [HW73, WACBB71]. Exchange [CCA70, FL74, Hin70, WE73]. Excision [HL74a]. Excitable [Per74, RE74]. Excitation [DSS72]. Excitation-Contraction [DSS72]. Excretion [Lee70]. Excystment [MB71a, RS74b]. Exocrine [AJ74a, AJ74b, GP73a, JP71a, JP71b, KP72, OPL72]. Exocytosis [KD74, MHW72, SW72b]. Exogenous [CK73, DD70, Fah70a, SCPH73, SSP73]. Experimental [AB71b, FG70a, GMdlTFGGF74, MM74b, Ove73]. Experimentally [RCBS70]. Explants [MT70]. Exportable [ASO^{+}71]. Exposed [DGW74, MBCP71]. Expression [JM74, KK73b, SH71, SA73a]. Expulsion [OBJ72]. Extended [JEG73]. Extensions [LLP72]. External [Bos73, DKHL74, HH74, LCRM71]. Extracellular [BT70b, FF70a, GP73a, Gri70b, Hua74, JN74, OC72, WT73a, WM72]. Extract [NFR^{+}74]. Extracted [GG72]. Extraction [BB71]. Extractive [HF70]. Extraneuronal [GHN74].

F1 [Lak73, GKP74]. F1-Histone [Lak73]. F2A1 [GPKJ73]. Fab [AAG^{+}74]. Faces [CG73b, WK74a]. Factor [GS74, OM74]. Factors [BN71, HDM74, RBK74, TC70]. Faecalis [HDM74]. Fallax [Gri73a]. Fast [HL74a, LA71]. Fasted [BW73a]. Fat [FC72, KMO^{+}70, Lar70, LM71, MJ70, She71b]. Fate [LM71, TH72]. Fatty [BY74a, RW73]. Fe [Fed74, SBKH74]. Feather [KDR74, MB71b]. Features [KKCP73, MHT72, PTF73]. Feeding [Tuc74]. Fenestrae [OPL72]. Fenestrations [WSS73]. Ferritin [DE71, FCH73, Fed74, NMS71, NS74, SY71, SS72b, RFBB74]. Ferritin-Conjugated [NMS71, NS74]. Ferritin-Labeled [SY71]. Fertilization [Bry70a, Bry70b, Kan70, LA70a, LA70b]. Fertilized [Aro73, CWT^{+}73]. Fetal [BB73b, CDMR71, HMT71, KL74b, LP72, LS74, Lef74a, Lef74b, MBCP71, O’B72, OLK^{+}71, OL73, SWSK71, WT70]. Fetoprotein [LS74]. Fiber [CF72, Dan72, Dan73, HL74a]. Fibers [ABE71, ABE72, Bun73, CG73a, Dan73, DU72, EF73, GP72, Gol71, GPR73, HMS73, Kay70, KBH73, MHB70, Nag72, Ova71, Ova72, PG70a, Per70, PB74, PHH70, PW73, Row71, Row73, RC73, SHP70, WPB72, Woo71]. Fibrillar [KPF74]. Fibrils [BDW71, LC72, ZF70]. Fibroblast [KLJ^{+}74]. Fibroblasts [BPA74, BTLN73, Cul74, Esp71, EMJT74, HH74, Per73c, PP74, SSC74, TS74, TBT74, UBKB74, WP72a, WP74b]. Figures [CA71]. Filament [KWR71, PD72]. Filamentous [BB73a, SFMD72, WFO70]. Filaments [All74a, BRB^{+}72, DS71, GWS70, LDC73, NSM70, PI70, PK71, RMM^{+}70, Roh74, SD70a, Szo70, UCB71]. Filtration [KNY74]. Fine [All71, AS71a, Arm70, Bal70, Bun73, CUMB71, CHM72b, DHN73, Dan73, DC74, EI71, Fah70a, Fah74, GR70, Han70a, HB71b, HLP73, Kay70, KP74, KT71, LC71, LK72, MCE71, NO71, Nak71, Ova71, Ova72, PBK74, RVHWU72, RP72, Roh74, Ten70, VSF70, War70, ZR73]. Fine-Structural [LK72]. Fish [KFP70, Nag72, WPB72]. Fission [SL73b]. Fixation [Bod70, BBW74, HP70, MBW71, Nan71a, OW73a, PV70a, PO73, PM72a, WABG74]. Fixatives [BD71, BdN74, DC72]. Fixed [CAC^{+}71, Gho71, HDM74]. Flagella [GF72, Sum74, WCBR72, WCR72]. Flagellar [CR70a, FKS71, GG72, GG74a, Kay70, SG73, War70]. Flagellates [Mül73, SD71b, Tam72b]. Flagellum [Bou71, HP74]. Flesh [RM71]. Flesh-Fly [RM71]. Flight [CSGB70, SS72a]. Flocculent [KK73a]. Flow [DJ74, RKT71, WD71]. Fluid [FR74a, OL73, SK71]. Fluid-Transporting [FR74a]. Fluidity [EF73]. Fluorescence [Lat74]. Fly [RM71]. Focal [MHB70]. Follicle [AA74, AY70, AS71b]. Follicles [Esp71, Whi72, WT73b]. Follicular [HSH^{+}71, SK70b]. Following [EH70, Klo70]. Food [All74b]. Force [All74a, Aro71]. Form [Coh71, HC74, MR74, RW72]. Formation [AB71a, BL70, BDH^{+}74, BdL73, BDW71, BFK^{+}70, CAH^{+}74, Dan72, Dan73, DBP70, Fah70b, Fli71, HCAP74, HAT71, HF71, KJ70, KWR71, LM70b, LKB71, LKMB73, LM70c, MPL71, MMS^{+}72, MHBM73, NT71a, NT71b, OU70, Ove73, Rat72, RP73, RET70, SA73b, Szu71, WS74a, dBTL73, dW71, dLB74, AJ73, OYC^{+}73]. Formations [Gob71]. Formed [FS73, GG74a, GG71]. Forming [KS72, WABG74]. Forssman [WR74b]. Fowl [MB71b]. Fowlpox [HP70]. Fraction [CGP70, CMY70, CMPF71, GPKJ73, KK73a, MJ70, MHM^{+}70, O’B72, dD71]. Fraction-Past [dD71]. Fractionated [Bow73]. Fractionation [ABS73, MBP72, TBT74, ZC71]. Fractions [BHdD70, BESP73, BEG73, BS70c, CGMD74, CS70, EBSP73, FBBdD72, FBP74, GAGGS72, GDH73, GAGGS71, GJKL73, HBL70, IP73, MJ74, NFR^{+}74, SB70a, SF74b, WACBB71, Wid72]. Fractographic [HWP73]. Fracture [BM74, CB70b, CG73b, FA74, Nan71b, SS74b, WK74a]. Fractured [BRB^{+}72, FF74, LR74]. Fracturing [PB74]. Fragmentation [Mol71]. Fragments [AAG^{+}74, HBS70, HN70b, Hen71, HB72a, KDC71, SHNB72, DKC71]. Free [IF71, RKT71, SDSN71, Tam72b, Woo74]. Free-Flow [RKT71]. Freeze [BM74, BOW70, BRB^{+}72, CB70b, CA71, FA74, FF74, GS74, LR74, PB74, PFSB72, SS74b, SRS73, TM70, dSB70]. Freeze-Dry [BOW70]. Freeze-Etch [SRS73]. Freeze-Etched [CA71, TM70]. Freeze-Etching [GS74, dSB70]. Freeze-Fracture [BM74, CB70b, SS74b]. Freeze-Fractured [BRB^{+}72, FF74, LR74]. Freeze-Fracturing [PB74]. Freezing [BM73]. French [Pag71]. Frequency [HDM74]. Fresh [Chr71]. Frog [AY70, BY74b, CHM72a, CHM73, CHM72b, Eds74, FF70b, FR74a, GD73, HR73, How74, Mag73, MPEB71, Mas71, Mas73, MF71, NF71, PRP74, PM73]. Front [WL73a]. Frozen [BV71, BIS72, Chr71]. Fucose [BL70, BL71b, BLH74, HSH^{+}71, PP73b]. Fucose- [BL71b, HSH^{+}71]. Function [BT71, Cus70a, Cus70b, CR70b, Goo71, Gor70a, Gor70b, JGR^{+}71, Kur74, MB71b, SHP70, SWSK71, WR74a, YSW71, GL70, LP70, TL70b]. Functional [WPB72, YA74]. Fundulus [LT71]. Fungus [AW72, CA71, DE71, Hea74]. Furrow [SD70a]. Further [BHdD70, Lak73, MHW71, UI72]. Fusarium [AW72]. Fusion [Aro73, Bak72, BH70, ISWS71, KLJ^{+}74, Lag73, LK72, MYWS72, McC74, Nak72, OKY72, PTLL74, SSS73, Shi71, WHG73, ZLK74].

G [FR74c, PR71a, Pot71, SS72c]. GABA [OK71]. Gairdneri [Nag72]. Galactose [WL71a]. Galactose- [WL71a]. Galactosyltransferase [WR74a]. Gallbladder [FR74a]. Gametes [Fus73]. Gamma [HBK70, Lam72]. Gamma-Aminobutyric [Lam72]. Ganglia [NNQH71, PW74, SPAR73]. Ganglion [Pea72]. Ganglioside [NLB73a]. Gap [AA74, DFDF73, DF74, FG72b, GS71a, GR71, GS72b, Goo74, GG74b, HG72, HR71, HBM73, PHPO74, SB70b, UB70]. Garfish [CGMD74]. Gas [WWB71, WABG74]. Gastric [IS74, RGR71]. Gene [GMP74, JM74, MABL71, MF71, WLH74]. Generate [All74a]. Generated [Til71]. Generation [THIM73]. Genes [GM72a]. Genesis [Lar70]. Genetic [BHLT74, KL74a, KB72, WWBG74]. Genome [LB72, PC73, SA73a]. Gentle [Nic73a]. Geometry [Har73, dW71]. Gerbil [Boq70]. Gerl [NNQH71]. Germ [CMN71, KG74]. German [ZBSW70]. Germarium [MS70a]. Germinal [OTE73]. Germinal-Center [OTE73]. Germination [Yoo70]. Gerris [Tur72]. Ghosts [EB74a, PTLL74, dS72]. Giant [HL74b, HFK70, HMS73, LLP72, OHPW74, TH73a]. Giemsa [MCIJ74]. Gill [GS71a, Nak72, SG70a]. Glancing [Gho71]. Gland [ASO^{+}71, Bel71, GM72a, GN73, Han70a, Han70b, HM72, KMO^{+}70, LH73, MT70, NP71, OB73, PP73b, Pel74, PHDN73, RS70, RR71, SMS72, Str71b, TO70, DKC71, EDE72]. Glands [BR72, BB72c, NB72]. Glia [HBL70]. Glia-Enriched [HBL70]. Glial [She73]. Glioma [DH74]. Globin [CAH^{+}74, HCAP74]. Globule [KMO^{+}70]. Glomerular [CF74, MS70c, RK74b]. Glomus [YCD70]. Glucagon [Det71, EH70]. Glucagon-Induced [Det71]. Gluco [LdH71]. Gluco-Corticosteroid-Induced [LdH71]. Glucocorticoid [BT70a]. Glucocorticoid-Induced [BT70a]. Glucose [MLGML71, RW73, WR72]. Glucose-Induced [MLGML71]. Glucuronyltransferase [BDN72]. Glutamate [FS70]. Glutamine [BL71a, KTC72, LRL71]. glutamyl [RL74]. Glutaraldehyde [CAC^{+}71, HF74, MBW71, OW73a, PKT74, PM72a]. Glutaraldehyde-Hydrogen [PM72a]. Glutathione [KR70]. Glycerinated [Col71]. Glycerol [BY74b, NF71]. glycerophosphate [OS74]. Glycogen [AP70, CSGB70, MC70, SH72a, SAE73, WT71, WD72]. Glycogene [PA70]. Glycogens [SP71a]. Glycolipid [BW70b]. Glycolipids [WR74b]. Glycoprotein [Ben70, BLH74, BW70b, WL71a]. Glycoproteins [AGS74, AGTS74, Cul74, FF70b, PP73b, Pel74, RC72]. Glycosaminoglycan [BB72c, HM74a]. Glycosaminoglycans [THT74]. Glycosyltransferases [RMR71b]. Glyoxysomes [GB70, TBB74, Vig70]. Glyptotendipes [WLH74]. Gold [SG70b]. Goldfish [Lam72, ZR73]. Golgi [Ben70, BL71b, BLH74, BESP73, BEG73, CMY70, CMPF71, CMN71, CSR^{+}74, EBSP73, FBP74, Fli71, GE70, MJ74, MHM^{+}70, NNQH71, SJ71, SS70a, SdlI72, Tre70, WF70, You73]. Graded [CF74]. Gradient [WP72b]. Grain [BP73a]. Gram [BT70c]. Gram-Positive [BT70c]. Granular [LLP72, MM73, Meb72, WM72]. Granulation [IF71]. Granule [ASO^{+}71, BHdD70, BF70, Cla70, FBBdD72, JP71a, MMM72, SH70]. Granules [Bai73, BCGL74, BB73d, BB74a, BWC74, HT70, JB72, Kan70, Key73, LC72, MCR74, MM70b, MT74, NHB72, Str71a, Str71b, Wat72, You71b]. Granulocyte [GR74]. Granulocytes [SSGJH73]. Granulolysis [POB71]. Granulosa [MM72]. Grasshopper [MGD71]. Green [BB73a, CPH74, EG71, HW74, MP70b, PA73, Pul74, SFMD72, WABG74]. Greening [GBN73, DSP70]. Grooves [Gho71]. Grow [Kni73]. Growing [CWB71, ML70, PV72, Pla72]. Grown [CUMB71, LP71, UB74]. Growth [All74b, BL74, BTLN73, BT70c, Bun73, Cla70, Coo73, FR74c, GK73a, GK73b, HPHB72, HT73, HD70b, IBK74, KL74b, LS74, Lef74a, Lef74b, LT74, MP70a, OP74, PD70a, PC71, PB74, RS74b, SD71a, SP73, SAE73, TBL70, Ten70, TML73, VHG74, WWC72, WR72, YSS^{+}72, YSW71, dW71]. Gryllidae [Cav72]. Guerin [FW73]. Guerin-Induced [FW73]. Guinea [APW^{+}73, BB73b, Bol74, CKS74, EWBP71, GDH73, KK73a, KP74, MJP71a, MJP71b, MJP71c, MC71a, MC72, Mel74, RKT71, UB72]. Gut [OC72]. Gymnogama [MMLM74]. Gymnostomatida [HC73].

H [NHS71, NLB73b, PO73, STM^{+}74, BL70, BL71b, Dal71, GAGSP73, GDH73, HSH^{+}71, KH71a, Len72b, PP73b, RVHWU72, Raw73, SS72c, SJ72, WL71a, WL74]. H-2 [NHS71, PO73, STM^{+}74]. Half [And70]. Halothane [HT74]. Hamster [CWB71, Eli72, Gol72, Gri73b, HH74, KKCP73, KSN73, LT70, LM73, NLB73a, NDAT71, OCWS74, POB74, PPF73, RE73, SM74a, SM73, SS72c, TL70a, WR74b, Wri73, KS74a]. Hamster-Mouse [Eli72]. Handling [SG70b]. Hansen [MR71b]. Haploid [MF71]. Hartmannella [RM72]. Having [Bur70]. Head [Lun72]. Heads [Tok74, ZBSW70]. Healing [Kra71, RET70]. Heart [AM73b, BS71, Bal70, Blo70, FG70b, GHN74, Gri70b, HS70, HB71c, LF72, MPAB70, MCS71, NF71, OM73, RW73, SMD73, SM71, WR72, Web72]. Heat [JSF70, JFdBJ73, WKG72]. Heat-Synchronized [JSF70, JFdBJ73]. Heating [Nic73a]. Heavy [CG73a, Con71]. HeLa [BF71, BS72, BR73, FH70, Gal74, JSP74, KGK72, KSE73, KR70, Lei71, MMS^{+}72, MHBM73, NK71, OCWS74, PD70a, PV72, PR70a, PR70b, RPK70, SHH73, SF74a, SD70b, SB72, SA71, SA72, SA73a, SA73b, WP72b]. Helical [Bur70, DBP70, HC74, KWR71, RW70b]. Hematopoietic [ZFG74].

Hemidesmosome [DRE70]. Hemispheres [YY74]. Hemocyanin [Fah70b]. Hemoflagellate [RW70a, Str72]. Hemoglobin [GNAB70]. Hemolysis [SCI73]. Hemopoietic [MG73]. hemoprotein [IKY70]. Heparan [KT72]. Hepatic [Cla70, GR71, RASP74]. Hepaticae [Car72]. Hepatocyte [DG71, GS72b, Goo74, GG74b, MGHA72, Sea72]. Hepatocytes [BL71b, BW73b, GNAB70, HFG73, Hig74, KL74b, LS74, Lef74a, Lef74b, LSP71a, LSP71b]. Hepatoma [BT70a, KW74b, KTC72, Pla72, PE72a, RRYT70, TT71, vRBvWW74, WMG74]. Hepatopancreas [BCGL74, CGL74]. Heterocyst [WABG74]. Heterocyst-Forming [WABG74]. Heterogeneity [YA74]. Heterogeneous [GP71, PD70a, SD70b]. Heterokaryons [OCWS74]. Heterologous [FCH73]. Heterophil [BHdD70]. Heterophils [HS71a]. Heteroploid [MF71]. Heterotypic [Arm70]. Hexagonal [HK70b]. High [BD72, Gla74, HN70a, KZ70, LKM71, Pul74, SP74]. Higher [LC72]. Highly [HDB73]. Hispid [HP74]. Histamine [LW74a]. Histochemical [FF70b, HKF71, TT71, Uge71]. Histochemistry [BS71]. Histocompatibility [NHS71, STM^{+}74]. Histological [Pay70]. Histolytica [KWR71, RW70b]. Histone [GPKJ73, GKP74, GWT74, Lak73, PR70a]. Histones [JEG73, MM70a]. Holes [SCI73]. Homo [OCWS74]. Homo- [OCWS74]. Homogenates [BESP73, BBHH74, EBSP73, GR73, TBT74]. Honeybee [Per72]. Hormone [Gro71, HRS74, HF73, MHW72, TBL70, TKTVJ73, VT74, WW72]. Hormone-Dependent [VT74]. Hormone-Stimulated [MHW72]. Hormones [MHSW71]. Horn [PP72]. Horny [LM70b]. Horseradish [AAG^{+}74, BTPT71, Kes70, MGS^{+}74, OL73, SC72b, SC72c, Str71a, WD71]. Horseshoe [LS71]. Hot [LK71]. House [Cav72]. Hrp [SC72b]. Human [Bak72, BT71, BT72, BB74a, BDW71, BWC74, DGW74, FTCM74, GR74, GCF74, HC70, KL74a, Lun72, Mag72, ML71, MHW71, MBW71, NO71, NMS71, Nic73b, NP73, PTLL74, Pog72b, RWB70, RE72, Rub72, RBK74, SCDS73, SDCS74, SH74, ST70, SMS72, SHM74, Ste74, VHB^{+}72, WL73b, WM72, ZF70, ZFG74, ZHW73, dS72]. Human-Mouse [KL74a]. Hyalin [SK70a]. Hyaline [Kan70]. Hybrid [Eli72, RFBB74]. Hybridization [GP71, MABL71, SD70b]. Hybridizing [LWD^{+}72]. Hybrids [DH74, KL74a, KCR70, WWvdN74]. Hydei [DBW73, HB71a]. Hydra [HG72, WYE71, Woo74]. hydrocarbon [WMG74]. Hydrocortisone [MT70]. Hydrogen [PM72a]. Hydrolases [Mül72]. Hydroosmosis [MHSW71]. Hydrophobic [DW70]. Hydrostatic [KZ70]. Hydroxylase [CB72, WMG74]. Hydroxytryptamine [GN73, SJ72, TdPP72]. Hyperplasia [FG70a]. Hypertonic [GG74b, RPK70]. Hypertrophying [JHK73]. Hypophysis [TVG70]. Hypothalamo [KW74a]. Hypothalamo-Neurohypophyseal [KW74a]. Hypothalamus [MR71a]. Hypothetical [BP73a].

Identification [BK73, GS74, Pad74, PHH70, RGR71]. Identified [AGS74, AGTS74, PM72b]. II [OW73b]. Ilyanassa [CWT^{+}73]. Image [SRA74]. Immobilization [PO73]. Immune [ISNC73, SC72b]. Immunis [GK72]. Immunoglobulin [SU70, ZUJP70]. Immunoglobulins [AAG^{+}74]. Immunohistochemical [LDdV72]. Immunological [DHN73]. Improved [ABS73, BV71, BWC74, CW74, HLP73]. Improving [PR71a]. Inability [DB74]. Incidental [HN72]. Incisor [WL71a, WL73a]. Including [WR74b]. Inclusion [WL71b]. Inclusions [Gri70a, SS70a]. Incompatibility [Jef72]. Incorporation [Dal71, GDH73, HSH^{+}71, PE72a, Pot71, RG72, Raw73, SWD74, SS72c]. Increase [PR71b, WE70]. Increased [LDP74, PAB71, SSK74]. Increases [Gri73b]. Increasing [TR73]. Incubated [FW72]. Incubation [GP73a]. Independence [RP73]. Independent [McC74]. Indication [MABL71]. Individual [Bra73, Fus73, Szu71]. Induced [AT70, BT70a, BS74, BW73b, BS73, CD73, DGM^{+}74, Det71, FW73, HW73, HB71a, HS72a, IWS72, KD74, KKCP73, Kis74, KH71a, KH71b, LdH71, LG72, LT74, MLGML71, MS70b, MHSW71, MHB73, MGK74, NLB73a, RASP74, RPK70, RAU71, SDCS74, SS72c, SSS74, SS73a, SS73b, Stu71, Sum74, WFO70]. inducing [RL74]. Induction [CA74, ISWS71, MYWS72, OCWS74, Pal73, SO72a, WMG74]. Inductors [HM74a]. Infected [DBP70, PP74, RO71]. Influence [BHLT74, DKHL74, Gro71, HP70, PKT74, SK71]. Influencing [HDM74]. Infradian [SAE73]. Infusionum [MR70]. Inherited [BH71b, LSO72]. Inhibited [CGB71, MCB71, MCB73, SP73]. Inhibition [AR74, BD72, BH70, CSH71, Dan72, Dan73, FR74c, HPHB72, HF71, HG71, HH71, KB72, NTKM73, Pas72, Pog72b, RI71, RK73, SG72, SWD74, SSS74, TML73, WR74c, WL70]. Inhibitors [Fli71, Goo71, HYC^{+}74, LT74, MHBM73, OM73, PC71, RU70, RO74, GL71]. Initiated [NB73]. Initiation [LT70, TL70a]. Injected [DW73, Str71a]. Injection [BLH74, GAGSP73, WL71a]. Injury [Bal70, Szu71]. Inner [BB73b, CGP70, Chl71, NMS71, SS73a, SS73b, TR73, CS74a]. Inorganic [BDM73, KLT71, KTT72, TLS70, TK71, TKT72]. Insect [Ash70, CSGB70, FS70, Lar70, LM71, Phi70, Ros73]. Insects [Rob72]. Insulin [AH73, BPA74, BEG73, BHLT74, EMJT74, HT70, MLGML71, MT70, OPT74]. Insulin-Stimulated [AH73]. Intact [HRWL71, KHTH70, PTLL74, Red71]. Integrity [Bod71]. Interaction [ASB73, LG70, MGS^{+}74, PW71, RW73, SC72b, SC72c, UP72, UP73]. Interactions [BH71c, CG71a, EB74b, MCIJ74, SMD73, SCB72a]. Intercalated [Woo74, dS72]. Intercellular [AA74, DBM74, DC70, FR74a, GR70, MS70a, RMR71a, WD71, WT73b]. Interface [BB72c, MHT72]. Interference [GW71, SSS74]. Interferon [Kni73, MCBBTG71]. Interfission [SSK74]. Intermedia [Hop72]. Intermediate [PG70a, Row73]. Intermediates [UB74]. Intermembrane [SSF74]. Intermicrotubule [HMC70]. Intermittent [KK73b]. Internalized [AAG^{+}74]. Interphase [BF71, CF73, ISWS71, KT71, Lat74, MYWS72, MPL71, MM72, OCWS74, OYC^{+}73]. Interpretation [CB70b, VDC71]. Interstitial [KAI^{+}74, LS73, NO71]. Interstitium [FL74]. Intertubule [Tuc74]. Intestinal [BSG71, BW70a, EB74c, FC72, HRW72, OL73, OW72, PW73, Rod73, RHW71, SJ71, SSP72]. Intestine [BL70, CKS74, Cor73, EWBP71, NN72, Rod70, Sti72, TC70, VSF70]. Intimate [Rub72]. Intoxication [TH73a]. Intracellular [BW72, EDE^{+}71, GK72, GE70, Hop72, HM74b, How74, JP71b, KDC71, LS73, LL72, Mül72, Pel74, SY71, TLS70, Tre70, VR74, WT73a]. Intracristal [HC71]. Intrafusal [KBH73, Ova71, Ova72]. Intramembrane [EB74a]. Intramembranous [MGK74]. Intramitochondrial [JB72, Mas71]. Intraneural [DD70]. Intranuclear [FSF72, Hin70, MBKM70, Tan73]. Intranucleolar [CP70]. Intraocular [GAGSP73]. Intrarenal [BOW70]. Intravascular [SK71]. Intrinsic [Col71]. Inulin [HM74b]. Invaginations [Mas73]. Invertebrate [AP70, HR71]. Investigation [DC74]. Investigations [MM74b]. Iodide [Nac74]. Iodination [HC72]. Ion [PKT74, PM73, SB73a]. Ionic [FL74]. Ionophores [KD74]. Ions [BS74, WHG73]. Iridophores [RP72, Roh74]. Iris [ZS72]. Iron [DG71, Fed74, HS71a, PB73]. Iron-Reactive [HS71a]. Irradians [SL70]. Irradiation [Pul74]. Irregular [FH70]. Islet [BHLT74]. Islets [APW^{+}73, EH70]. Isoenzyme [SF74a]. Isoenzymes [PV70a]. Isolated [AAWD74, ACBW^{+}74, ACWTS^{+}74, AT70, APW^{+}73, ABE71, ABE72, AYB^{+}72, BACF^{+}74, BACTS^{+}74, BB71, BDH^{+}74, BCG72, Bun73, BJ70, CB72, CGB71, CSR^{+}74, DW72, EWBP71, GR71, Hac72, HBL70, HHS70, HT70, HHH72, HLP73, HB71c, IBK74, KBC73, MMT^{+}74, MJ70, MCB71, MCB73, NHB72, NPD73, OO72, OW73a, PD70b, SWD74, SL72a, SM74b, Uge71, WT71, Wid72]. Isolating [BWC74]. Isolation [And74, BS70a, BS70b, BH71a, Bou71, Bry70a, CGMD74, CGM73, CMY70, CMPF71, CSGB70, CBS73, CT72, CBCT74, Cum72, CMN71, GR73, GS72b, Goo74, HS71b, HSWD72, HK70b, HDB73, Hol71, IK71, KK73a, Kan70, KMS73, KLKK73, MT73, MHM^{+}70, PFGG73, SM74c, TADH70, VZRE74, Whi72, ZZS71, ZC71]. Isoleucine [LM73]. Isopropyl [CPH74]. Isoproterenol [BS73, NB72]. Isoproterenol-Induced [BS73]. Isopycnic [BBHH74, CS74b]. Isotope [BKK73]. Isozymes [PV70b].

Junction [CHM73, CB70b, CHM72b, DFDF73, Fri71, HR73, KK70, Len70, Len72a, PRP74, Ros73, SG70a]. Junction-Ribosome [DFDF73]. Junctional [Bod71, FS74, LT71, MEW72, SMS72, WS74b]. Junctions [AA74, Arm70, BT70b, CHM72a, DF74, FC73b, FG72b, GS71a, Gob71, GR70, GR71, GS72b, Goo74, GG74b, HG72, HR71, HBM73, MM72, Nea73, PG70a, Per73a, Per73b, PHPO74, Ros72, SSF74, She70, SB70b, SL72b, UB70, UB72].

Kb [CGM73, RO71]. Keratin [KDR74]. Keratohyalin [MM70b, Uge71, UI72]. Ketogenesis [CMO73]. Kidney [AM70, BTPT71, BS70b, BPHBK74, Gol72, Gri73b, HKKSH72, KKCP73, MG71, Str71a, SG70b, TK71]. Kidneys [You71b]. Killing [Sto74]. Kinase [KLJ^{+}74, KL74a]. Kinase-Deficient [KLJ^{+}74]. Kinetic [OS72, VT72]. Kinetics [GWE74, LaV73]. Kinetochores [OD74]. Kinetodesmal [RC73]. Kinetoplast [RW70a, RW71b, RW72, Stu71]. kinetoplastic [RW71b]. Kinetosomes [Gri73a, Tam72b]. Klebsiella [Har70b]. Knife [WPB72]. Kupffer [Fah70a, WCF72].

L [BL70, KGK72, PH71, SSC73, SN71, UP72, UP73, ZBSW70, ZBSW70]. L-Fucose- [BL70]. L5178Y [BW70b]. Label [BL71b]. Labeled [BLH74, GAGSP73, GHN74, KDC71, PS71, PS70, SY71, TKTVJ73, WHCFB74, WABG74, DKC71, RFBB74]. Labeling [DHN73, OTE73, PS71, SIS74, SH72b, EDE72]. Lability [IBK74]. Labyrinthine [FS73]. Lacis [PHPO74]. Lack [GG71, SSS74]. Lacrimal [HM72]. Lactating [NP71]. Lactic [BS71]. Lacunae [Sea72]. Laevis [MK70, BW72, BdL73, CG72, KG74, Mil73, dLB74]. Lamella [Chl71, LC71]. Lamellae [FSF72, Mau70, NS70]. Lamellar [GR73, MP70b, WS74b]. Lamellibranch [SG70a]. Lamina [BDW71, VB70, VB72]. Lamprey [SJB70]. Langerhans [EH70]. Lanthanum [KO70, LF72, MPBA73, Nea73, She71a]. Lapse [Kri72]. Large [CS73, HHP70, MP70b, VSF70, ZC71]. Large-Scale [ZC71]. Larvae [ZC71]. Larval [BR72]. Larval-Pupal [BR72]. Laser [RB74a]. Late [Coo73, Pot71, SHH73]. Lateral [AS72, Dec74, FR74a]. Latex [SW72b]. Lattice [ABE71, ABE72, Eri74, HK70b, MP70b]. Layer [Kan70, Mab73]. Leaky [CG73b]. Leaves [GBN73]. Lecithin [LKMB73]. Lectin [GA73, NB73]. Lectin-Initiated [NB73]. Lectins [EB74c, GWE74]. Leech [PM72b]. Leishmania [Str72]. Length [CF72, Hag70]. Length-Tension [CF72]. Lengths [HN72]. Lens [HM74a, PdWW72, ZS72]. Lepisosteus [CGMD74]. Lesions [ISNC73]. Lethal [TGW73]. Leucine [Len72b]. Leucine- [Len72b]. Leukemia [YA74]. Leukocyte [Bai73, JB73]. Leukocytes [BHdD70, BF70, BB73d, BB74a, DGW74, FBBdD72, NHB72, RBK74, SCDS73, SDCS74, WT71, WM72, ZHW73]. Levels [CWB71, KR70]. Lewisi [RW70a]. Liees [PA70]. Life [Gri73a, KR70, PR70a, SH74]. Life-Span [SH74]. Ligation [Nea73]. Light [AI70, CG70, Gib70, HC74, NSG71, RB74a, SO72a, SP74, TKTVJ73]. Lignofibrils [LCRM71]. Like [CG73a, CD73, FGO71, Fri71, Mab73, Mul71, Ros71a, Ros71b, SS74b, SP71b]. Lily [Spi70]. Limulus [BH71c, DLC73, Fah70b, LDdV72, LDC73, LS71]. Line [DBP70, FA73a, HS70, LM73, NLB73a, Row71, SHDH74]. Linear [ZBSW70]. Lineardichroitische [ZBSW70]. Lines [CGB71, KFP70, MCB71, MCB73, MGDT74, MGT74, PP71a, SG74, Wri73, BSGS72]. Linked [Hac72]. Linking [All73]. Links [Har70a, Tuc74]. Linnaeus [CD73]. Lipase [BMS71, BMS73]. Lipid [BW70b, GDH73, NH74, SJ71, TAKS74, Wak74]. Lipid-Depleted [NH74]. Lipids [CB71a, CB71b]. Lipophosphonoglycan [BK74]. Lipoprotein [BMS71, BMS73, Cla70, SSS74]. Liposomes [MGS^{+}74]. Liquid [HD70b]. Lived [Pot71]. Liver [ABS73, ACBW^{+}74, ACWTS^{+}74, AT70, AYB^{+}72, BACF^{+}74, BACTS^{+}74, BN71, BESP73, BEG73, BT71, BHS72, BT72, BHM73, BS70c, BDN72, BJ70, CB72, CB70b, CMO73, CJMB73, CMY70, CMPF71, Chr71, DOP74, Det71, EBSP73, Fah70a, FBP74, FDE^{+}70, GE70, GD70, GR70, GS70, GFK72, Hac72, HHP70, HSH70, HLP73, HYC^{+}74, IP73, IK71, Ivá71, JEG73, KL72, Kle70, KHS74, LdD73a, LdD73b, LP72, LK70, MMT^{+}74, MHM^{+}70, MC71b, NPD73, O’B72, PHdD70, SB70a, SLWZ70, SG74, SF74b, SMH70, SSS74, Str71a, SdlI72, Tid70, TADH70, WRE71, Wak74, Wid72, WKG72, WL70, RFBB74]. Liver-Derived [SG74]. Livers [HB72b]. Lividus [EG71]. Living [Nic73a, SHM74]. Lizard [FS74, RP72]. Lobe [CWT^{+}73]. Lobes [FW72]. Lobster [Atw71, HBK70, HFK70, OK71]. Lobule [SLWZ70]. Local [TBW74]. Localisation [PA70]. Localization [AP70, BG72, BHB73, BSG71, BK74, Cav72, CG73a, CMO73, CAH^{+}74, CT74, DDP73, DPD73, Ett72, EB74c, Fah70a, FBBdD72, GHN74, HS71a, HCAP74, HM72, HB71b, HB73, IP73, JHK73, JR74, KDC71, LN74, LDdV72, MPEB71, Meb72, NMS71, OK71, PV70b, PP73b, QWT73, RVHWU72, RS74a, STM^{+}74, SV73, Sti72, SK70b, TLS70, TKTVJ73, TBB74, VSF70, WWvdN74, Wid72, WG73, WABG74, DKC71, RFBB74]. Location [Nan71b]. Locomotion [LW74b, SYW71]. Locus [DBW73, PH70]. Locus-Specific [DBW73]. Locusta [KNM74, ZBSW70]. Long [MCS71, Pot71]. Long-Lived [Pot71]. Long-Term [MCS71]. Loss [Fed74]. Low [AT70, CF73, EDE^{+}71, FC73b, HS72b, Per73a, Per73b, She70, Spi70]. Low-Amplitude [AT70]. Low-Resistance [CF73, FC73b, She70]. Lucidum [ABN70]. Lumbricoides [RS74a]. Lumenal [CBS72, SCB72b]. Lumina [Kay70]. Luminal [HK70b, VZRE74]. Lung [DHWHW71, GR73]. Lungs [SK71]. Luteinizing [TKTVJ73]. Lymph [AAG^{+}74, LdH71]. Lymphatic [Lea71]. Lymphocyte [BS74, Pog72a]. Lymphocytes [Bow73, Coo73, GWE74, MMS^{+}72, MGK74, OU70, Pog72b, PG72, WL73b]. Lymphoid [RE72, SWJ^{+}71]. Lysine [SSC73, UP72, UP73]. Lysis [ISNC73]. Lysosomal [Dec74, FW73, RB73, RB74b, SDCS74, ZHW73]. Lysosome [GKB74, Mul71]. Lysosome-Like [Mul71]. Lysosomes [Bow73, CB70a, GNAB70, NNQH71, SMH70]. Lysozyme [EAK70, KSR^{+}70, KESA71]. Lytechinus [Aro73].

M [EP72, KP71]. M4Rs [Pay70]. Macaca [PBGCP74]. Macro [Cum72, GPKJ73]. Macro- [Cum72, GPKJ73]. Macromolecular [CJC70, CBWS74, GdB73, GJDVC74, HB70, JC73, PG70b, WCDI72, PH71]. Macromolecule [MB71a]. Macronuclear [SW70]. Macronuclei [GKP74]. Macronucleus [Klo70]. Macrophage [AR74, Lut73]. Macrophages [FCH73, Fed74, FW73, HC70, RA73, SSC73, SSGJH73, SS72b, SC72b, SC72c]. Maintaining [CG70]. Major [Bry70a, Hoo72, HS73]. Malate [TBB74]. Male [Rat72]. Malignant [MHW71, WMKJ73]. Malnutrition [Dal71].

Malpighian [Kes70]. Mammalian [AM73a, BC71, CF72, CS73, DK73, DSS72, DMF74, EKP74, FDE^{+}70, FG72b, FF74, GSH71, GAP70, HT73, KFP70, Kle70, KH71a, MPBA73, MP70a, MSGE73, NS74, NF71, NN72, PG70a, Phi72a, RMM^{+}70, Row73, SSF74, SBRC70, SSHS74, SWR70, TCK72]. mammals [IKY70]. Mammary [KMO^{+}70, MT70, NP71, OPT74, PHDN73, SD70a, SB74]. Manganese [RdS73]. Manipulation [Ove73]. Mantis [POB71]. Many [GP73b]. Marchantia [MK70]. Marinus [SJB70]. Marker [GS70, WE73]. Markers [GP73a]. Marssonii [PA73]. Mass [GMP74]. Mast [Com71, KD74, Lag73, LW74a, Pad70, Pad74, RAU71, SS72b]. Mastigonemes [Bou71]. Mastocytoma [PS71]. Material [BL70, EI71, KK73a, SP74]. Mathematical [DJ74]. Mating [WE73]. Matrix [CGP70, DKHL74, EAK70, NU74, TPG72, WL71a, You71b]. Maturation [Coo73, CP70, GJ72, PG72, SW71]. Mature [Fus73, Mol71]. Maturing [WH73]. Means [Bod70, HN70a, Pad74, PM72a, RFBB74]. Measurement [CS73, GW71]. Measuring [JN74]. Mechanical [PKT74]. Mechanism [AT70, DBM74, OBJ72, RK74a, RO74, SM71, WM73b]. Mechanisms [BDN72, ZHW73]. Medial [Nak71]. Mediated [MHSW71, Nam72, SFK74, WHG73]. Mediates [WP74b]. Mediating [Cul74]. Medulla [AH73, Kis74, ZR73]. Medullary [LS73]. Megakaryocytes [TdPP72]. Meiosis [CG72, McL71, MR71b, Rat72, Wei72]. Meiotic [KT74a, KT74b, PGR72]. Melanocyte [Hop72, Mal71]. Melanocyte-Stimulating [Hop72]. Melanogaster [MS70a, PFGG73, Tok74, ZC71, OS74]. Melanogenesis [MB71b]. Melanoma [WOSE73]. Melanophores [MT74]. Melanosome [Mag73]. Melanosomes [ABN70, TH72]. Membrane [ABS73, ASB73, All74b, ASO^{+}71, AAG^{+}74, Aro73, AR74, BT70a, BL74, BKG73, BdL73, BK74, BB73d, Bry70a, Bry70b, CGMD74, CB70b, CGP70, CGM73, Chl71, CBS72, CD73, EMJT74, FKZ^{+}71, FF74, GP73a, GA73, GS70, HS71b, HBS70, HN70b, Hen71, HB72a, HHH72, HR73, HK70b, HC72, HYC^{+}74, ISNC73, Iwa71, KMO^{+}70, Key73, KBC73, KDS73, KS74b, KNY74, KL73, Lag73, McK74, MJ70, MJ74, MF71, MD72, MS70c, MP70c, Nan71b, NLB73b, NT71a, NT71b, PO73, PWT71, PKT74, Per74, RE72, Ros73, SSF74, SSS73, SS74a, SHNB72, SCI73, SRS73, SL72b, SF74b, SCB72b, Ste74, SS73a, SS73b, Str74, Szu71, TAKS74, TR73, WFO70, Woo74, WBSW74, dSB70, dS72, dLB74, vIG71, CS74a, GL71, OYC^{+}73]. Membrane-Associated [KL73]. Membrane-Bound [ABS73]. Membrane-Envelope [HBS70, HN70b, Hen71, HB72a, SHNB72]. Membrane-Like [CD73]. Membrane-To-Membrane [FKZ^{+}71]. Membranes [All73, ACBW^{+}74, ACWTS^{+}74, Bak72, BS70a, BACF^{+}74, BACTS^{+}74, BY74a, BW73b, CW71, CLBW72, CB71a, CB71b, Cla70, DHN73, DOP74, DK73, FDE^{+}70, GO70a, GO70b, GS71b, HSH70, HK70a, HK70b, HB72b, Hig74, Hin70, Hoo72, HS73, JSP74, Kle70, KLR74, KHS74, LR74, LKB71, MM74a, MPBA73, MJP71a, MJP71b, MJP71c, MC71a, MC72, Mel74, NHB72, Nak72, NMS71, NHS71, Nic73b, NP73, NPD73, NH74, OW72, OHPW74, PR71b, RE74, Red71, Saa74, She73, TM70, TADH70, UWWK71, VZRE74, DSP70]. Membranous [All71]. Mensuration [SHM74]. Meromyosin [CG73a, NSG71]. Mesangial [PHPO74]. Mesenchymal [BB72c, EB74b, MHT72, SCB72a, STM^{+}74]. Mesosome [Nan71a]. Mesosomes [HDM74]. Messenger [BH72, CAH^{+}74, GP71, HCAP74, HG71, SD70b]. Metabolic [AY70, JP71a, MHBM73, OM73]. Metabolism [AM70, BY71, HBK70, KPH^{+}74, LB72, OP74, Red71, RRYT70, SA71, SA72, VT72, WD72, WT70, You73]. Metabolisme [PA70]. Metachronal [AS72]. Metamorphosis [BW70a, MI71a, MI71b, MI71c]. Metaphase [ISWS71, Lak73, MYWS72, MGD71, OCWS74, SHM74, Ste73, OYC^{+}73]. Metaphase-Arrested [Lak73]. Metaplastic [ZS72]. Methane [SO72a]. Method [BS70a, BP73a, BWC74, BKK73, CL72, CKS74, DC73, HS71b, LKM71, PR71a, SP74, SY71, SMH70, TR73, TT71, TCK72]. Methods [Nic73a]. Methyl [AI70, GHP73, SO72a]. Methylation [Tid70]. Mg [Eds74]. Mice [BHLT74, Lan73, PC73, RU70, RASP74, SFS71]. Michaelis [WT73a]. Microbodies [GBN73, HB73, LW70, LM71, RCBS70, Vig70, WL70]. Microbody [RASP74]. Microcylinders [Kay70]. Microfibrillar [All71]. Microfibrils [Whi72]. Microfilaments [Aue72, Com73, CWT^{+}73, Gol72, LaF74, ML74, Per73c, PM73, RA73, SYW71]. Microfluorometry [DJ74]. Microirradiated [RB74a]. Micronuclear [MP70d]. Micronuclei [Cum72, GPKJ73, IWS72]. Microorganisms [Bai73]. Micropapillae [CG71b]. Microperoxidase [Fed71]. Microperoxisomes [CKS74]. Microscope [BP73a, BD71, Cha73, Coh71, Com71, GHN74, Gla74, HKF71, HN70a, HFK70, KD74, Kis74, KNM74, Len72b, MC70, MC73, MM74b, Nac72, NWL71, NP71, Pad74, PB73, Per72, PTF73, RAU71, RGR71, SS71, Sha72, Shi71, TKTVJ73, WACBB71, WG73]. Microscopic [Bod70, EAK70, HHP70, MBCP71]. Microscopy [AP74, BH71c, BJ70, Chr71, CW74, GW71, HWP73, Lat74, MM70a, Pal71, PN70, RB74a, SP74, SSD^{+}74, SBKH74, WHCFB74, Wel74]. Microsomal [KDS73, KHS74, KS74b, MC71a, NPD73, SB74]. Microsomes [ACBW^{+}74, ACWTS^{+}74, Bas71, BACF^{+}74, BACTS^{+}74, GD70, RFBB74]. Microsporidia [WT73a]. Microtubular [HC73, HA70, MS70b, Rob72, War74]. Microtubule [All74b, BB73c, BB73e, BB74b, Dal72, Eri74, FDBS71, RGGG71, RW71a, SD74, Tan73, Til71, Tuc74]. Microtubule-Associated [All74b]. Microtubules [BA74, BF71, BM74, Bur70, FL70, Fri71, FKS71, Gri70c, Har70a, HS72a, KZ70, KHTH70, ML71, McI74, McL71, MT74, OC72, Phi74, RA73, Sum74, TA70, TG70, TBB^{+}73, Tur72, WF73, WN73]. Microvilli [BG74a, IS74, LG72]. Microvillous [TC70]. Middle [LC71]. Midgut [RM71]. Migration [Ben70, BLH74, Kra71, MS70a, Pel74, WL74]. Migratoria [KNM74, ZBSW70]. Migratory [LW74b]. Milk [KMO^{+}70]. Mincing [HL74a]. Mineralization [WL73a]. Minutum [HHC71]. Mitochondria [ABBA72, AT70, BDM73, BWC74, BJ70, CMO73, CGL74, DMF74, FB70, GB72, GS70, HRWL71, Hac72, HKF71, HBL70, HC74, HSH70, HK70a, HDB73, HB71c, KL72, KSE73, KL73, Lag71, Lan73, Lar70, LK70, MHB73, MGHA72, MC71b, RS74a, SS72a, SSF74, SS70b, SN71, SPAR73, SG70b, TH73a, TC72, WHL70, Web72, WMKJ73, YSS^{+}72, You71b]. Mitochondrial [BMD^{+}72, CW71, CLBW72, CGP70, CLSA74, Dal71, FG70a, Gal74, GMP74, Gro71, HH71, HZM71, JGR^{+}71, Key73, KGK72, KL74a, KNM74, KW74b, Kur74, Lei71, LP71, LM73, LL72, MKPV74, MLD70, O’B72, PC73, PFGG73, SGL73, SJG72, SSD^{+}74, SS73a, SS73b, SA73a, SA73b, TR73, TPG72, UB74, WE73, Wat72, WKG72, WKCF73, ZKH73, CS74a]. Mitochondrial-Specific [LP71]. Mitochondriale [PA70]. Mitoplasts [Hac72]. Mitoribosomes [VSHA72]. Mitosis [AW72, AI70, BS73, Cos74, FH70, Hea74, HW74, ML71, OD74, OO71, RYR71, SHH73, WH70]. Mitotic [BB71, BC71, BR73, CF73, Cos74, DHN73, FKS71, HMC70, IBK74, RK74a, RPK70, Ste73, WR74b]. Mitotically [PW73]. Mixed [MG71]. Mo [SBKH74]. Mo-Fe [SBKH74]. Mobility [AR74, dS72]. Mobilization [WM72]. Mode [SA73b]. Model [Mal71, PV70a, SSS73]. Modification [SDCS74]. Modifications [WFO70]. Modified [BOW70, Red71, Tur72]. Modulating [McK74]. Mold [CS70, Ett72, GK73a, GK73b, HHC71, OKY72, Tan73]. Molecular [EDE^{+}71, TAKS74]. Molecules [Bos73, BH72, GP73b, KW74b, MWR72, WMKJ73, WKCF73]. Molluscan [GS71a]. Mongolian [Boq70]. Monkey [AB71a, And72, KL74a, You71a]. Monkey-Mouse [KL74a]. Monoamine [GS70]. Monocytes [HC70, NBF71]. Monolayer [BT71, BT72, BHM73, CJMB73, KL74b, LS74, Lef74a, Lef74b, PBK74, SM74a]. Mononuclear [WCF72]. Monophosphate [AM73a, FR72, FR74c, GS73, MHSW71, RM72, RF73, SSGJH73, Sto73a, Sto74, TML73]. Mori [TO70]. Morphogenesis [Aue72, BBC72, Dir71, EB74b, Phi70, TC71]. Morphogenetic [SW72a]. Morphological [BHdD70, BB74a, DW70, GAGGS71, Gri73a, GJKL73, KMO^{+}70, LP71, MB71a, OLK^{+}71, PAB71, Uge71, WT71, WPB72]. Morphology [BBC72, BT71, BHLT74, BG74a, CSR^{+}74, KGK72, MGHA72, MC71b, PPF73, PG72, RE73]. Morphometric [BW73b, HFG73, Lag71, Mat73, SSP74, Pag71]. Morphométrique [Pag71]. Morphometry [SdlI72]. Mosquitoes [AS71b]. Moth [WT73b]. Motile [BM74]. Motility [Gol71, Phi72a, SG73]. Motion [Tam72a]. Motive [All74a]. Motor [AP74, Dec74, PM72b, Ros74, VHG74]. Mounts [HK70b]. Mouse [BB72c, CDMR71, Com71, Cul74, Dir71, Dir74, Eli72, FCH73, Fed74, FS73, GJ72, GR70, GR71, GS72b, Goo74, Gri70b, GJKL73, HN70a, HT74, HD73b, IS74, JEG73, KLT71, KTT72, KT74a, KT74b, KLJ^{+}74, KL74a, Kni73, LaV73, LRL71, LKB71, LA71, MSM74, MGK74, MBCP71, MLMPF74, NHS71, NB72, PC71, PHDN73, PTF73, RdS73, RVHWU72, RK74b, RGR71, SW71, SRA74, SK71, SD70a, She71b, STM^{+}74, SF74b, SD74, SC72b, SC72c, TGW73, VHG74, WWvdN74, WP74b, WMKJ73, Woo71, WOSE73]. Movement [BR73, DB74, GG72, Mag73, MT74, TCMA73]. Movements [FL70, SW72a]. Mucopolysaccharide [BB72c, BBC72]. Mucopolysaccharide-Protein [BBC72]. Mucopolysaccharides [SC72a]. Mucosa [HRW72, MHT72, RHW71, RGR71]. Mucosal [JB72]. Mucosubstance [HS71a]. Mulatta [PBGCP74]. Müler [MC73]. Multinucleate [HW74]. Multiple [WLH74]. multiplication [Rub71]. Multivesicular [BL71b, Esp71, KJ70, Wak74]. Murine [BS73, PV70b, PS71, TS74, TML73]. Musca [CD73]. Muscle [Bal70, Bas71, Blo70, CUMB71, CB70a, CSGB70, CSH71, CF72, DS71, DLC73, DC70, EF73, EKP74, FS70, FC73a, FA74, GP72, Hag70, HSWD72, HL74a, How74, HMS73, IWM^{+}74, Iwa71, JHK73, JSJ71, JLS73, Kel71, KBH73, LDdV72, LDC73, LS71, Mat73, ML70, Nag72, NSG71, NF71, Ova71, Ova72, PG70a, PH70, PP73a, PHH70, PLRK72, RMM^{+}70, Ros71a, Ros72, RK71, Ros71c, Row71, SS72a, SHP70, SHDH74, She73, SSP73, SDSN71, SSD^{+}74, SPAR73, SV73, SWS74, SO72b, UB70, UCB71, UB72, WH70, BSGS72, OK71]. Muscles [Atw71, BS71, DSS72, LA71, Row73, SH72a]. Mushroom [McL71]. Mutant [JGR^{+}71, Lan73, SFS71, DSP70, GL70, LP70, TL70b]. Mutants [HN73, Mil73]. Mutation [BHLT74]. Myelin [CA71, GR73, Raw73, dW71]. Myelination [BHB73, RU70]. Myeloma [SW71]. Myoblast [BH70]. Myoblasts [GG71, Shi71]. Myocarde [Pag71]. Myocardial [FS74, FL74, Lag71]. Myocardium [FS73, Pag71]. Myofibrillar [LG73]. Myofibrils [Col71]. Myofilament [ABE71, ABE72]. Myofilaments [FC73a]. Myogenesis [Nam72, OS72, Prz71]. Myogenic [LK72, War74]. Myoglobin [SSP73]. Myoinositol [GDH73]. Myonemes [All73]. Myoneural [Ros72, Ros73]. Myosin [BRB^{+}72, Mab73, Nac72, PH70, PD72, Ros71a, Ros71b, SHGH72, Nac74]. Myosin-Like [Mab73, Ros71a, Ros71b]. Myosins [NSG71]. Myotome [TBS71]. Mytilus [AS72].

N [GHP73, TML73, AI70, CPH74, WABG74]. N-Desacetyl [AI70]. N-Labeled [WABG74]. N-Methyl [AI70]. N-Phenyl [CPH74]. Naegleria [FD71, SD71b]. Nafenopin [RASP74]. Nafenopin-Induced [RASP74]. Nasal [AS71a]. Nascent [BS70c, GE70, SB70a]. Nature [KW74a, KPF74, MBKM70, MM70b, ML70, Rae70]. Necklace [GS72a]. Necklaces [GG71]. Necturus [BTPT71]. Negative [CSR^{+}74, GS74]. Negative-Staining [GS74]. Negatively [HK70b]. Nematosomes [Pea72]. Neonatal [GAGGS71, MGHA72, Rod70, SM74a]. Nephrotoma [LaF74]. Nerve [BHB73, BBHH74, Bun73, CGMD74, Dan72, Dan73, FS70, FBS71, GAGSP73, Han70b, Kis74, LW74b, MG71, MHB70, Per70, PB74, RU70, RVHWU72, Raw73, TA70, WYE71, YSW71, YCD70, OK71]. Nerve-Acinar [Han70b]. Nerves [Eds74, dW71]. Nervous [DD70, HBK70, Lan73, PM72b, SJB70]. Net [WYE71]. Network [MM74b]. Neural [HM74a, JM74]. Neurites [MR74]. Neuroblast [TB70, Ten70]. Neuroblastoma [CG73a, DH74, HT74, LDP74, MR74, NB74]. Neuroblasts [MGD71]. Neurofilamentous [MM74b]. Neurogenic [WPB72].

Neurohypophyseal [KW74a, MHSW71]. Neuromuscular [CHM72a, CHM73, CHM72b, HR73, JA74, Len70, Len72a, PG70a, PRP74, She73, UB72]. Neuron [AGS74, AGTS74, AP74, HBL70, PM72b, VHG74]. Neuron-Enriched [HBL70]. Neuronal [GHN74, GD73, KLKK73, PBGCP74, SPAR73]. Neurons [Bra73, Bun73, Dan73, Dec74, Gri70a, HHS70, HHH72, LLP72, Len72b, MP73a, MP73b, MP73c, MBKM70, MM74b, NNQH71, PP72, SO72a, SL72b, TH73b]. Neurosecretion [KW74a]. Neurospora [CW71, CLBW72, HZM71, KL73, LL72, WL71b]. Neurulation [DF74]. Neutrophilic [BB74a, JB73]. Newborn [Rod73, SH72a, SK71]. Newly [AGS74]. Newt [BG72, OO71, SY74, She71b]. Nexus [Mat73, MW70, MM72]. Nexuses [Iwa71, MHW71]. Nicotian [GC74]. Nicotine [LA70b]. Nil [WR74b]. Nitella [All74a, CG70, Tur70]. Nitrogen [WABG74]. Nitrogenase [SBKH74]. Nitrous [BR73]. Node [AAG^{+}74, LdH71]. Nodes [WPB72]. Non [SF74a]. Non-Regan [SF74a]. Nongrowing [Coo73]. Nonhistone [HB71a]. Nonkinetochore [BA74]. Nonlethal [TBW74]. Nonparenchymal [CB72]. Nonproliferating [BHM73, CJMB73]. Nonsecreting [FF70a]. Nonuniform [LG73]. Norepinephrine [DD70, GHN74]. Normal [BdL73, BTLN73, Cul74, FC73b, FW73, GS73, LRL71, NLB73a, TS74, UBKB74, WWvdN74, WM73a, WMKJ73, dLB74]. Nosema [WT73a]. Notochord [HM74a]. Novikoff [KW74b, Pla72, PE72a]. Nuclear [AB74, Aro73, Blo70, Bos73, BH72, CS70, Dal71, ED70, FDE^{+}70, Gib70, Gor70a, Gor70b, GHP73, GP71, GMP74, Hin70, HM74b, ISWS71, KJ70, Kle70, KMS73, LWD^{+}72, MPL71, Mau71, MMS^{+}72, MHBM73, OD74, OCWS74, PD70a, Phi74, RR71, RYR71, Sch70, SHH73, SD70b, WG73, WR74c, WW71, WBSW74, EDE72, OYC^{+}73]. Nuclear-Cytoplasmic [Hin70, RYR71]. Nucleating [TG70]. Nuclei [AYB^{+}72, BH71a, DBW73, HB71a, IK71, KLKK73, KT71, Lat74, MSM74, MYWS72, MP70a, McC74, ML70, OCWS74, OO72, OW73a, Tan73]. Nucleic [LG70, Rud71]. Nucleocytoplasmic [CG71a, WBSW74]. Nucleofilum [Har70a]. Nucleohistone [GPR73]. Nucleolar [GMdlTFGGF74, GG71, KT74a, PG72, RR71, SD71a]. Nucleolar-Derived [RR71]. Nucleolate [Mil73]. Nucleoli [Gri70a, Ivá71, LLP72, MMT^{+}74, PP71a, PP71b, PP73c, RWB70, ZZS71]. Nucleolus [GLG70, GG70, NDAT71, Phi72b, SP71b]. Nucleolus-Like [SP71b]. Nucleoprotein [MCIJ74]. Nucleoside [LT74]. Nucleotidase [Wid72]. Nucleotide [Har70b, Pla72, SL73a, SS73a, SS73b]. Nucleotide-Induced [SS73a, SS73b]. Nucleotides [CSH71, CK73, HT70, WHCFB74]. Nucleus [Aro71, Hor72, KS71, MBP72, Nak71, TKT72, WG73]. Number [Ivá71]. Nutritionally [WFO70].

O [TML73]. Oblongata [Kis74, ZR73]. Observation [CB70b, GCF71, HN70a, NSM70]. Observations [Coh71, EI71, Eri74, HHP70, HN72, KG74, KP74, LW70, ME71, MHW71, MHB70, Nac72, Rat72, RAU71, RP72, Sea72, Shi71, SAE73, War70]. Observed [HDM74, KFP70, SBRC70]. Obsoleta [CWT^{+}73]. Obtained [MS70c]. Obtaining [LKM71]. Occludens [WK74b]. Occludentes [CG73b, GG74b, WK74a]. Occur [SB70b]. Occurrence [ABN70, HHC71, HK70b, MHW71, O’B72, Prz71]. Ochromonas [Bou71, BB73c, BB73e, BB74b, Gib70, GP73b, HP74, SJG72]. Octagonality [Mau71]. Octahedral [OW73b]. Odontoblasts [WL74]. Odontogenesis [SCB72a]. Oedogonium [CPH74]. Oenocytes [CD73]. Oleracea [MWR72]. Olfactory [CGMD74, GD73, Mul71]. Oligochaeta [DC73]. Oligomeric [WMKJ73]. Oligoribonucleoproteins [Uge71]. One [BH70]. Onset [TSE70]. Ontogenetic [BB72d]. Ontogeny [AA74]. Oocystis [PA73]. Oocyte [HF70, Hor72, HM74b, IBK74, MLMPF74, SS70b]. Oocytes [Cav72, DW72, EI71, GK72, Hin70, KFJ70, Mas71, Mas73, MKPV74, Sch70, SWD74]. Oogenesis [CG72]. Opaque [LC72]. Optic [GAGSP73]. Optical [Bur70]. Oral [RW71a]. Orconectes [GK72]. Ordered [Roh74]. Orderly [VB72]. Organ [BN71, Cor73, Far72, HRS74, MCS71, OLK^{+}71, TVG70, WPB72]. Organelle [Gib70]. Organelles [Bod71, BdN74, GN73, Vig70]. Organic [BDM73]. Organic-Inorganic [BDM73]. Organisation [ZBSW70]. Organization [GPR73, HG72, HA70, ME71, PA73, RMM^{+}70, SJB70, Ste74, UBKB74, War74, Wat72, Wei72, GL71, ZBSW70]. Organizing [Tan73]. Organotypic [SPAR73]. Organs [SWJ^{+}71, SB73b, ZC71]. Orientation [DU72]. Origin [Bou71, BDW71, LM71, ME71, MBKM70, SS70a]. Ors [Pul74]. Orthophosphate [AT70]. Orthophosphate-Induced [AT70]. Orthoptera [Cav72]. Oscillations [Har70b]. Osmiophilia [PR71b]. Osmium [FB70, PKT74, WS74c]. Osmolality [HP70]. Osmotic [AJ73, BD71, Coh71, HOB71, KBC73, SCI73, WD71]. Osmotically [WK74a]. Osseus [CGMD74]. Osteoclasts [HRS74]. Other [RW71a, Rub72, ZLK74, DDP73, DPD73, RW71b]. Outer [BY74a, CGP70, Coh71, GS70, HOB71, Key73, KBC73, LaV73, TM70, You71a]. Outgrowths [FC73b]. Outward [GAP70]. Ovarian [AA74, AY70, AS71b, Esp71, SFK74, WT73b]. Ovary [GM72a, HA70, MM72, PPF73, RE73, Wri73, KS74a]. Oviduct [AB71a, And72, And74, Dir71, Dir74, KMS73, PW71, SD74]. Oviducts [Har73]. Ovine [TKTVJ73]. Ovulation [Esp71]. Oxalate [PHH70, dMHM74]. Oxidase [GS70]. Oxidation [RW73, WR72]. Oxidative [HRWL71]. Oxide [BR73, DC73]. Oxygen [Har70b, Pul74]. Oxyntic [FF70a]. Oxysporum [AW72]. Oxytricha [Gri73a].

P [Wri73, ED70, IKY70]. P-450 [IKY70]. Packaging [BF70, Dec74, GK72, Hop72, KK71]. Packed [Woo74]. Pair [ST70]. Palestinensis [CB71a, CB71b, Chl71]. Pallidum [GK73a, GK73b]. Pancreas [BW73a, Bol74, Boq70, Chr71, GDH73, GP73a, HDB73, KP72, MJP71a, MJP71b, MJP71c, MC71a, MC72, OLK^{+}71, SM74a, SWR70]. Pancreatic [AJ74a, AJ74b, APW^{+}73, BY71, JP71a, JP71b, Mel74, OPL72, RCDMM71]. Pancreozymin [RCDMM71]. Para [BOW70]. Para-Aminohippuric [BOW70]. Paracrystalline [WL71b]. Paracrystals [NSG71]. Parafollicular [GN73, DKC71]. Paraformaldehyde [PO73]. Parallel [Per70]. Paramecium [ABBA72, All71, All74b, Cum72, FLM73, SSK74, Tam72a]. Parasite [WT73a]. Parasitic [RK74a]. Parathyroid [HRS74, NWL71]. Parenchymal [BHM73, CB72, CJMB73, HLP73]. Parietal [IS74, WS74c]. Parotid [ASO^{+}71, CJP72, CW74, Han70b, LH73, NB72, RS70, BG74b]. Pars [Hop72]. Part [Nic73a]. Parthenogenesis [SA70]. Partial [Cum72, HKF71, HL74a, Hol71, Mil73, RC73]. Partially [Ran74]. Participation [GE70]. Particle [Det71, EB74a]. Particles [Boq70, Chl71, LR74, MMT^{+}74, MGK74, O’B72, RKT71, SKMH70, SS72b, SS74b, WT71, Woo74, dS72]. Particular [AA74, RW70b]. Particulate [SP71a, SC72b]. Passage [BL71b]. Past [dD71]. Pathways [CF73, RC72]. Pattern [GLG70, GG70, HC73, Kra71, SSK74, WWC72, WS74c]. Patterns [Bra73, NLB73a, OTE73, Rob72, Rud71, Til71]. Patula [GdB73]. Pea [Yoo70]. Pecten [SL70]. Pectoralis [BS71]. Peptides [Hop72]. Performed [DRE70]. Perfused [BMS71, BCG72]. Perfusion [KAI^{+}74]. Perichromosomal [KT74a]. Perikarya [PBGCP74]. Perinuclear [EI71]. Periodic [CH73, KK73b, WE70]. Periodicity [SD71b]. Periods [PD70a, SS72c]. Peripheral [PG70b, RU70, Raw73, SWS74, Woo71, ZFG74, dW71]. Periphery [GKB74]. Peritoneal [FCH73, Fed74, SC72b, SC72c]. Permanent [SG74]. Permeability [AS71b, CF74, GR71, Hor72, HM74b, Kes70, Lea71, MPEB71, McK74, Nea73, NPD73, PKT74, SK71, SP71a, SSP72, SSP73]. Permeable [MEW72]. Peroxidase [AAG^{+}74, BTPT71, CA74, Fah70a, GA73, HM72, Kes70, MGS^{+}74, OL73, SC72b, SC72c, Str71a, SK70b, SWSK71, Str71b, TKTVJ73, VSF70, WD71, WL70]. Peroxidase-Labeled [TKTVJ73]. Peroxidase-Positive [Str71b]. Peroxide [PM72a]. Peroxisomal [GBN73]. Peroxisome [RASP74, SFMD72]. Peroxisomes [BB73b, Fah74, LdD73a, LdD73b, NN72, PHdD70]. Persistence [MG73]. Petromyzon [SJB70]. pH [HP70, JB73, OYC^{+}73, PH71, Rub71]. Phagocytes [WCF72]. Phagocytic [DGW74, JB73]. Phagocytizing [RA73, SSGJH73]. Phagocytosis [AR74, Bai73, CK73, GK73b, Lut73, Sto73b, WM72]. Phagosome [UWWK71]. Phase [BS73, FH70, FR74c, MHB70, PR71a]. Phases [Det71, MP70a]. Phenobarbital [BW73b, BDN72, CD73, HB72b, Hig74]. Phenobarbital-Induced [BW73b, CD73]. Phenobarbital-Sensitive [BDN72]. Phenobarbital-Treated [HB72b, Hig74]. Phenomena [RPK70]. Phenotypic [NLB73a]. Phenyl [CPH74]. Phloem [GC74]. Phorbol [ZTB74]. Phosphatase [BK73, BB73d, Det71, FBBdD72, Meb72].

Phosphatase-Bearing [Det71]. Phosphate [BCGL74, dMHM74]. Phosphatidic [Meb72]. Phosphine [DC73]. Phosphodiesterase [SL73a]. Phosphoenolpyruvate [vRBvWW74]. Phosphokinase [Lak73]. Phospholipase [NTKM73]. Phospholipases [FW73]. Phospholipid [BHB73, Red71]. Phospholipids [SSS74]. Phosphoprotein [WL73a]. Phosphorylation [GWT74, HRWL71]. Phosphotungstate [CSR^{+}74, RC73]. Phosphotungstic [LG70, LK71]. Photooxidation [NS70]. Photoreceptor [Coh71, HOB71, LSO72, NH74, Saa74]. Photoreceptor-Pigment [LSO72]. Photosynthetic [CG70, Nic71, OM74, GL71]. Photosynthetically [SS74b]. Phototaxis [SH73]. Phycobilisomes [EG71, GL72]. Phycomyces [DE71, OW73b]. Physarum [Ett72, HH71, Kur74, McC74, Nac72, Nac74, Tan73]. Physical [LG70, McC74, Nan71a, OO72]. Physicochemical [VHB^{+}72]. Physiological [SW70]. Physiologically [PM72b]. Physiology [HB70, RB73, RB74b]. Phytohemagglutinin [FTCM74, MMS^{+}72, Pog72a, Pog72b, Wri73]. Phytohemagglutinin-P [Wri73]. Phytohemagglutinin-Stimulated [MMS^{+}72]. Pig [Bol74, CKS74, EWBP71, GDH73, KK73a, KP74, MJP71a, MJP71b, MJP71c, MC71a, MC72, Mel74, RKT71, UB72, DKC71]. Pigeon [HL74a]. Pigeons [BW73a]. Pigment [BY74b, BH71b, LSO72, MMM72, MT74]. Pigmentation [WOSE73]. Pigmented [NFR^{+}74]. Pigs [APW^{+}73, BB73b]. Pine [WH73]. Pinocytic [BPHBK74]. Pinocytosis [BO72, MHSW71, SSC74]. Pituitaries [HF73]. Pituitary [AY70, PP73b, Pel74, TBL70]. Plain [KK73a]. Plane [Nan71b]. Plant [BC72, CBWS74, FSF72, KH70, LC71, LCRM71, MMLM74, NS74, WWC72]. Plants [LC72]. Plaques [DRE70, MR71b, PGR72]. Plasma [AAG^{+}74, AR74, BK74, CGMD74, CGM73, Chl71, CBS72, DK73, GA73, HS71b, HHH72, HK70b, JSP74, KMO^{+}70, MJ70, MJ74, Nan71b, PS71, RE72, SRS73, SCB72b, TADH70, UWWK71]. Plasmalemma [Gho71]. Plasmodia [Ett72, McC74]. Plasmodium [Tan73]. Plastic [HWP73]. Plastic-Embedded [HWP73]. Plastids [HB70, WB70]. Plate [Ros74]. Platelet [MCR74]. Platelets [BWC74, FTCM74, Ros71b, ZF70, ZTB74]. Pleiotropic [Wri73]. Plexus [QWT73]. Ploidy [SHM74]. Pneumocytes [MM73]. Pneumonocytes [Meb72]. Point [Gri70c]. Polar [CWT^{+}73, JB72]. Polarity [HKKSH72]. Polarized [WYE71, WT73b]. Poliovirus [DBP70, HW73]. Poliovirus-Induced [HW73]. Poliovirus-Infected [DBP70]. Poly [SSC73, UP72, UP73]. Poly-L-Lysine [SSC73, UP72, UP73]. Polyanions [AYB^{+}72]. Polycephalum [Ett72, HH71, Kur74, McC74, Nac72, Tan73]. Polylysine [Pad70, YY74]. Polylysine-Coated [YY74]. Polymerase [SCPH73]. Polymerization [THIM73]. Polymorpha [MK70]. Polymorphonuclear [Bai73, BB73d, BB74a, FBBdD72, JB73, NHB72, RBK74, SCDS73, SDCS74, WT71]. Polyoma [KKCP73]. Polypeptide [Hoo72, HS73]. Polypeptides [BL74, BS70c, SB70a]. Polyphemus [BH71c, Fah70b, LS71]. Polyploid [FH70]. polyribosome [AJ73]. Polyribosomes [Bos70, DBP70, HD70a, HF71]. Polysaccharide [RG72]. Polysaccharides [DC73]. Polysiphonia [TPG72]. Polysomal [SD70b]. Polysome [EMJT74, IP73, LKB71]. Polysome-Bound [IP73]. Polysphondylium [GK73a, GK73b]. Polytene [BC72, DBW73, GL74, HB71a]. Polyvalent [CSH71]. Pool [CWB71, Pot71]. Pools [Pla72]. Population [FH70, HRW72, KH70, RHW71, Rub71]. Populations [BBHH74, CS73, Det71, SSHS74]. Porcine [IBK74, MJ74, Sea72]. Pore [AB74, Mau70, MPL71, Mau71, MMS^{+}72, MHBM73]. Porous [RK74b]. Porphyridium [GL72, RG72]. Position [Tre70]. Positive [BT70c, Str71b]. Possible [Hin70, KAI^{+}74, Lee70, MB71b]. Posterior [TO70]. Postmitotic [Phi72b, PP73c, SP71b]. Postnatal [GJKL73, HM72, LA71]. Postsynaptic [BCC74, CBCT74, UB72]. Potassium [DKHL74, DGW74, HLP73, Nac74, TLS70]. Potent [ZTB74]. Potential [KPH^{+}74]. Precipitates [PHH70]. Precipitation [BKK73]. Precursor [AM70, MMT^{+}74, Pot71, WL74]. Precursors [CP70]. Preformed [Phi72b, PP73c]. Preimplantation [SRA74]. Premeiotic [KG74, KFJ70]. Preparation [ABS73, BV71, BJ70, CL72, CKS74, FS70, HHH72, MJ70, SMH70, CS74a]. Preparations [MS70c, SCPH73, OK71]. Prepared [BESP73, EBSP73, FBP74, HLP73, NU74]. Presence [AP72, BL71a, Fli71, IBK74, Pot71]. Present [dD71]. Presentation [HK70a]. Preservation [BdN74, CW74]. Pressure [KZ70]. Presynaptic [Bod71, HKF71]. Primary [KL74b, LP72, LS74, Lef74a, Lef74b, MP73a, MP73b, MP73c]. Primate [Har73, VR74]. Primate-Specific [VR74]. Primates [Bel71]. Primitive [CWMH71, WCH71]. Principal [SB74]. Prismatic [You71b]. Probe [DW70, JN74, SSD^{+}74]. Procedure [ABS73, Nac74]. Procedures [KMS73]. Process [BPHBK74, CJP72, HP73, MC71b, THIM73]. Processing [Rae70]. Procion [Pay70]. Procurement [Fus73]. Produced [BT72]. Produces [SF74a, Tre71]. Product [DBW73]. Production [DBW73, HYC^{+}74, KCR70, RP72, TBL70]. Products [PM72a]. Profiles [EMJT74]. Progesterone [PW71]. Progression [MGD71]. Prolactin [MT70, TBL70]. Prolamellar [FGO71]. Proliferating [Esp71, KCR70]. Proliferation [FR72, GWT74, HMT71, RASP74]. Proliferative [KH70, RS70]. Proline [WL74, SIS74]. Prolixus [HA70]. Prolonged [CHM72a]. Prominent [Bur70]. Promoter [ZTB74]. Propagating [SY74]. Properties [CGP70, CGM73, GKB74, GF72, GG74a, GP71, LK70, Nac74, Pay70, POB74, RKT71, SV73, SHGH72, SL70, SK70a]. Prophase [KFJ70, KT74a, KT74b, MYWS72, Tan73]. Prophasing [OCWS74, OYC^{+}73]. Prostaglandin [KK70]. Prostaglandin-B [KK70]. Prostaglandins [CK73]. Protection [PV70a]. Protein [ASO^{+}71, BBC72, BW73a, BW70b, CP70, CS73, Dav73, EP72, ED70, FW72, FDBS71, Fli71, Gal74, GAGGS72, GJ72, Goo71, HBS70, HN70b, Hen71, HB72a, HG71, Hol71, HRW72, HYC^{+}74, IWM^{+}74, KK71, Kan70, Kes70, KP72, LP71, LH73, Mab73, MC73, MMLM74, NWL71, Nic71, OP74, OB73, PGHK72, Per72, PdWW72, RO74, RHW71, SHNB72, SK71, SWD74, SB74, SD74, SBKH74, SSS74, VT72, WP74a, WW72, GL71]. Protein-Depleted [HRW72, RHW71]. Protein-Embedded [Nic71]. Proteinaceous [GK72]. Proteins [BCC74, BR72, BB72d, BS72, CG71a, Eds74, FLM73, HAT71, HB71a, HH74, JP71b, KHS74, LDdV72, LL72, LG73, PV72, RGGG71, RW71a, Ste74, SB72, TVG70, WW71]. Proteoglycan [BBC72]. Proteolysis [BS70c, SB70a, Zob73]. Proteus [PI70, PK71, PD70b, Szu71, WF70]. Protocols [TCK72]. Protofilaments [TBB^{+}73]. Protonation [MP70c]. Protozoan [RK74a]. Proximal [BTPT71, HKKSH72, Rod70, SB70b]. Pseudobranch [HC74, DDP73, DPD73]. Pulmonary [KAI^{+}74, MM73, SRS73, SJ72]. Pulverization [IWS72]. Pump [Sti72]. Pumps [QWT73]. Punctulata [And70, LA70a, LA70b, SA70, TG70]. Pupal [BR72]. Pupillae [UB72]. Purification [KHTH70, RC73]. Purified [HDB73, Nac74]. Purkinje [Lan73, MSM74, WMB74]. Puromycin [FC72, Kis74, MF71, TML73]. Puromycin-Induced [Kis74]. Purpureum [CCA70]. Pyknosis [LdH71]. pyrene [WMG74]. Pyrenoid [Hol71]. Pyridine [Har70b]. Pyriformis [SW70, CLSA74, Gor70a, Gor70b, GHP73, GPKJ73, KZ70, Kum70, LA74, Mül72, MP70d, NT71a, NT71b, OBJ72, RW71a, RC73, SS74a, SAE73, UB74]. Pyroantimonate [TLS70].

Qualitative [GJKL73]. Quantitation [MBW71]. Quantitative [BK70, Det71, GAGSP73, GJKL73, PV70a, Sto73b, VHG74]. Quantity [WB70]. Quarter [And70]. Quasi [MP70b]. Quasi-Crystalline [MP70b].

Rabbit [AA74, And74, BHdD70, BS70b, BB73d, CJP72, Esp71, FBBdD72, Fed74, GAGSP73, Gul72, HBL70, HS71a, KFJ70, MC70, MC73, NHB72, PBK74, Sti72, TB70, Ten70, TBS71, WT71, BSGS72]. Radicles [Yoo70]. Radioactive [PB73]. Radioautographic [BT70c, CJP72, EI71, FW72, HSH^{+}71, HF70, MC70, OU70, RWB70, RGR71, Sti72, SJ72, WL73a]. Radioautography [BL70, BLH74, BOW70, DHWHW71, Len72b, NWL71, Per72, SS71, WL71a, WL74]. Radioimmunoassay [BT72]. Rana [VBD70]. Ranvier [WPB72]. Rapid [BS70a, CS73, Eds74, LT74, SB73b]. Rapidly [BD72]. Rat [AH73, ABS73, ACBW^{+}74, ACWTS^{+}74, ASO^{+}71, AT70, AM73b, AYB^{+}72, BS71, BACF^{+}74, BACTS^{+}74, BHB73, BL70, BL71b, BESP73, BEG73, BHM73, BS70c, BPHBK74, BH71b, Bow73, BTLN73, BBHH74, BJ70, CB72, CMO73, CJMB73, CMY70, CMPF71, CT72, CBCT74, CMN71, DOP74, DD70, Det71, EBSP73, EB74c, Fah70a, FBP74, FG70b, FC72, FB70, FG72a, GP73a, GR73, GD70, GNAB70, GAGGS71, GS70, GFK72, HSH^{+}71, HHP70, Han70a, Han70b, HSH70, HKKSH72, HFG73, HM72, HK70b, HDB73, HF73, HLP73, HYC^{+}74, IP73, Kel71, KW74a, KL72, KLKK73, KL74b, KW74b, KP72, KHS74, LdH71, Lag71, LdD73a, LdD73b, LS74, Lef74a, Lef74b, LN74, LH73, LK70, LG73, MHT72, MMT^{+}74, Mat73, MM72, MHM^{+}70, MC71b, NWL71, NPD73, NNQH71, OLK^{+}71, OL73, OM73, Ova71, Ova72]. Rat [Pad70, Pag71, Pea72, PP73b, Pel74, Pla72, PHdD70, RS70, RC72, RCDMM71, Rod70, Rod73, RK74b, RAU71, RRYT70, SB70a, SHP70, SH72a, SM71, SF74b, SMH70, SSS74, SWSK71, Str71b, TAKS74, TK71, TBL70, TKTVJ73, TADH70, TBT74, VSF70, Wak74, WWvdN74, WL71a, WL73a, Whi72, Wid72, WKG72, WMKJ73, WL70, WMB74, Woo71, YSS^{+}72, YM73, YY74, YA74, dW71, BG74b, RFBB74]. Rate [CH73, HT73, LDP74]. Rates [LG73, SSHS74, SS72c]. Ratio [KAI^{+}74]. Rats [BLH74, Gri70a, HMT71, HB72b, Hig74, HRW72, LSO72, ML70, NSM70, RASP74, RHW71]. Ray [Pad74]. Rc [DBP70]. Reaction [DC74, LA70a, MM70a, Pad70, PM72a]. Reactions [Sto73a]. Reactivation [MT73]. Reactive [HS71a]. Reaggregating [KO70]. Rearrangements [SW72a]. Reassembly [Eri74, TC70]. Receptor [FTCM74, PP73a]. Receptors [CT74, Lut73, WWvdN74]. Reconstitution [Bry70b]. Reconstruction [ST70, WR74c]. Recover [BD72]. Recovery [HF71, PW74, SG72, TH73a, GL71]. Recycling [HR73]. Red [Bak72, Cha73, How74, HC72, MBW71, Nag72, NSG71, NMS71, NHS71, PG70a, RG72, Row73, Ste74, TA70, TS74, TM70, VBD70]. Redistribution [MGK74, RBK74]. Reduced [SA71, SA72]. Reducing [PR71b]. Reference [AA74, Nak71, RW70b]. Reformation [SP71b]. Regan [SF74a]. Regenerating [Len72b, MB71b, RVHWU72, YSS^{+}72]. Regeneration [GD73, HL74a, Ran74, TBW74, Tid70]. Region [Nag72, SM73]. Regions [HK70a]. Regular [KLR74]. Regularly [Chl71]. Regulation [AM73a, LT70, OB73, SG74, TL70a, WF73, WN73, vRBvWW74, AJ73, Rub71]. Regulator [KAI^{+}74]. Regulatory [GR74, GS73]. Reinforcement [Bur70]. Reinhardi [FGO71, Goo71, Hoo72, HS73, DSP70, GL70, GL71, LP70, TL70b]. Reinhardtii [BL74, CBS73, Cos74, HN73, MM74a, RF73, SH73]. Reinke [NO71]. Reinvestigation [NO71, Sch70]. Related [GdB73, Vig70]. Relating [BKG73, RP72]. Relation [IS74, JGR^{+}71, KWR71, LA70a, PC71, Rob72]. Relations [RYR71, SHP70]. Relationship [BDM73, CF72, FSH71, JSF70, NB73, NB72, PE72a, Rub72, SFK74, SG73]. Relationships [Cus70a, Cus70b, CR70b, Han70b, LSO72, Rud71]. Relative [PV72]. Release [BK70, BHLT74, DBW73, HR73, KDS73, KS74b, MLGML71, SDCS74, WL74, WM72, ZHW73, BG74b]. Released [Kan70, PC71]. Relief [Pas72]. Remigis [Tur72]. Removal [AYB^{+}72, BW73b, GG74a, BSGS72]. Renal [BCG72, CL72, Dav73, LS73, Lee70, SB70b]. Renewal [BY74a, BY74b, LaV73, TVG70, You71a]. Renografin [CS74b]. Reorganization [GMdlTFGGF74]. Reovirus [SSC73]. Repair [SO72b]. Repetitive [BG72, Fus73, SW71]. Replacement [VB72]. Replicas [CB70b]. Replicates [Stu71]. Replicating [WKCF73]. Replication [GL74, HT73, HF71, JSF70, Jef72, JFdBJ73, KW74b, LA74, MKPV74, MR70, WL73b]. Replicative [UB74]. Repopulation [Phi72b, PP73c]. Repositories [Ett72]. Representing [MP70a]. Reproductive [GdB73]. Require [SW72a]. Requirement [WL73b]. Requirements [KFP70, MB71a]. Resembling [Gri70a]. Resistance [AB71b, CF73, FC73b, KSN73, MF71, PAB71, Per73a, Per73b, Pul74, She70, Spi70, Str72]. Resistant [ABBA72, KSE73, POB74, Wri73]. Resolution [BBHH74, SS71, SP74]. Resorption [GP73a]. Respiration [HBS70, HN70b, Hen71, HB72a, SHNB72]. Respiratory [HLP73, KGK72, LB72]. Response [AY70, Coh71, DGS^{+}74, GCF71, PP72, SCDS73]. Responsiveness [OPT74]. Resting [RA73]. Restriction [AYB^{+}72]. Result [PR71b]. Resumption [SHH73]. Retardation [RU70]. Retention [Stu71, ZLK74]. Reticulum [All73, BB72d, DSS72, FS74, FB70, HSWD72, HB72b, Hig74, Kel71, LSP71a, LSP71b, LKMB73, LA71, Nak72, OPL72, SB74, SDSN71, TH73b, WD72, WSS73, WS74b, dMHM74]. Retina [BY74b, JR74, JM74, LaV73, Lam72, LRL71, MC70, MC73, RL74]. Retinal [BH71b, LSO72, NFR^{+}74]. Retinas [BL71a]. Retinula [POB71]. Retraction [SBRC70]. Revealed [CG73a, NWL71, TA70, WL71a]. Reversal [AI70, Coo73]. Reversible [FC73a, Nam72, WBSW74]. Reversion [KKCP73]. Revertant [CGB71, MCB71, MCB73]. Rhesus [AB71a, And72, You71a]. Rhizoplast [SD71b]. Rhodnius [HA70]. Rhodophyta [TPG72]. Rhodopsin [JR74, Saa74]. Rhynchosciara [GM72a, GL74, MABL71, Sha72, dBTL73]. Rhythm [BS73]. Ribonucleic [AM70, MP70d, Rae70, RR71, Tid70]. Ribonucleoprotein [DBW73, KWR71]. Ribonucleoprotein-Filament [KWR71]. Ribosomal [AM70, BC72, Cav72, CCA70, Coo73, CP70, Eli72, GM72a, GL74, KG74, Key73, KH71b, LL72, MMT^{+}74, NK71, PV72, Rae70, Wat72]. Ribosome [ASB73, DFDF73, HHP70, HBG74, Kum70, MBS72a, MBS72b, SB74]. Ribosome-Membrane [ASB73]. Ribosomes [ABS73, CBS73, CLSA74, KNM74, KL73, LKB71, MM74a, O’B72, PBGCP74, GL71]. Rich [APW^{+}73, CMY70, CMPF71, MHM^{+}70, Wak74]. Ricinus [GB70]. Rifampicin [Pog72b]. Rigor [GG74a]. Ring [Sch72, Szo70]. RNA [BPA74, BD72, BH72, CS70, CAH^{+}74, Con71, Coo73, EDE^{+}71, EDE72, Eli72, Fli71, GM72a, GAGSP73, Gib70, Goo71, GL71, GP71, HCAP74, HG71, JHK73, KS74a, KT74a, LWD^{+}72, MABL71, Mil73, MLD70, MLMPF74, NK71, PD70a, PR70b, Phi72b, PP73c, RO71, Sch73, SCPH73, SSC73, SHH73, SD70b, SA71, SP71b, SA72, SP73, WLH74, WG73, WBSW74]. RNA-DNA [MABL71, SD70b]. Rod [HOB71, KBC73, LaV73, ME71, NH74, You71a]. Rod-Bodies [ME71]. Rodlets [MBKM70]. Rods [HC71]. Role [AM73a, BH71b, FL74, GR74, Gol71, Hin70, Lee70, MT74, PC73, PM73, QP73, SC72a, SG70b, THIM73, VT74, WD72, You73]. Root [NNQH71, TB70, Ten70]. Rotational [All74a]. Rotor [CKS74]. Rough [OPL72, SB74]. Round [BH70]. Rounding [Sto73a]. Routine [LK71, LKM71]. Rubinellus [McL71]. Ruthenium [Cha73, How74, TA70, TS74].

S [BS72, SB72, MHBM73, O’B72, PV72, SS72c]. S-3 [BS72, SB72]. Sac [MW71]. Saccharide [NS74]. Saccharides [NS74]. Saccharomyces [Gho71, HPHB72, KPF74, LM70c, MR71b, PGR72, Wat72]. Saccinobaculus [MOL73, McI73]. Saccules [Coh71]. Sacs [ABN70]. Salamander [MCS71, TC70]. Salina [SGL73]. Salivary [BR72, BB72c, BBC72, GM72a, RR71, dBTL73, EDE72]. Salmo [Nag72]. Salmonid [KFP70]. Saltatory [FL70]. Salts [SPAR73]. Same [ABN70, Iwa71, SP74]. Sand [VT72]. Sap [LWD^{+}72, RR71]. Sapidus [BCGL74, CGL74]. Saponin [SCI73]. Sarcolemmal [HN72]. Sarcoma [LKB71, RdS73]. Sarcomere [Hag70, HN72]. Sarcophaga [RM71]. Sarcoplasmic [DSS72, FS74, HSWD72, Kel71, LA71, SDSN71, WD72, WSS73, WS74b, dMHM74]. Sartorius [PRP74]. satellite [RW71b]. Saturation [Kni73]. Say [Tur72]. Scaffold [VB72]. Scale [BFK^{+}70, MP70b, ZC71]. Scalloping [HN72]. Scanning [BH71c, HWP73, NP71, Pad74, PTF73, Wel74]. Schedule [JFdBJ73]. Schizophyllum [WE73]. Schwann [HFK70]. Sciara [BR72]. Sciatic [Eds74, RVHWU72, dW71]. Scorpion [Phi74]. Scrolls [CD73]. Sea [And70, BH72, Bry70a, Bry70b, Con71, Fus73, GG72, GF72, GG74a, HG71, Kan70, KB72, LA70a, LA70b, Mab73, MKPV74, RGGG71, RYR71, SA70]. Sebaceous [Bel71]. Secretagogues [GDH73]. Secreting [FF70a, FF70b]. Secretion [AH73, CA74, ERB74, EH70, FW72, GP73a, HD73a, HF73, IS74, KP72, Lag73, LW74a, MLGML71, MC71a, Mül72, SU70, SSS74, WW72, ZUJP70, dBTL73]. Secretory [ASO^{+}71, BR72, CJP72, EB74c, HT70, JP71b, KK71, LH73, RC72, SB74, WL71a]. Section [BOW70, NNQH71]. Sectioned [HK70b, HS72b]. Sections [BV71, BIS72, Chr71, DRE70, Gho71, GW71, LK71, LKM71, PH70, SY71, TR73, VDC71]. Sedimentation [LK70]. sediments [CS74a]. Seeds [AGMT71, MT71a, MT71b, MT71c]. Seen [Gho71]. Segment [HOB71, LaV73]. Segments [BY74a, Coh71, CSH71, KBC73, You71a]. Segregating [GP73b]. Segregation [BF70]. Selected [SP74]. Selection [PW73]. Selective [BB71, HK70a, KB72, KDS73, KS74b, KNY74, MPBA73, Rod70]. Seminiferous [CMN71]. Sendai [WHG73]. Sensilla [MCE71]. Sensitive [ABBA72, Aro71, BDN72, KSE73, POB74, TPG72]. Sensitivity [SMD73, SL73b]. Separated [ZR73]. Separation [CW71, CLBW72, MP70a, McC74, RKT71, SWJ^{+}71, WP72b]. Sephadex [NFR^{+}74]. Septa [KAI^{+}74, Per70]. Septate [GS71a, Gob71, HG72, HR71]. Sequence [MMS^{+}72, Raw73]. Sequences [GP71]. Sequential [Bai73, PN70, SP74]. Sequestered [DW73]. Series [TGW73]. Serine [IP73]. Serological [FKS71]. Serotonin [RGR71]. Serotonin-Synthesizing [RGR71]. Sertoli [Nea73]. Serum [BT72, BCG72, EMJT74, RC72, RK73, SSS74, Str74]. Serum-Starved [EMJT74]. Several [DMF74, KKCP73, KS71]. Sexual [Pal73]. Shape [BB73c, BB73e, BB74b]. Shaping [Phi74]. Sheaths [dW71]. Shifts [BS73]. Short [Hag70, HN72, MHB73]. Short-Term [MHB73]. Shortening [CR70a, LG72]. Shown [BLH74]. Shrimp [Nak72]. Shuttle [WG73]. Sialic [MS70c, RE72]. Sides [BK74, CB70b]. Sieve [ED70, PN70]. Significance [SIS74]. Silica [WP72b]. Silk [TO70]. Silkworm [TO70]. Silver [MM70a]. Simian [KSN73, NLB73a, WWvdN74]. Similarity [FKS71]. Simple [CL72, FA73a, PE72a]. Simultaneous [CS73]. Single [AGS74, AGTS74, CS73, Nic73a, PBGCP74, SFMD72, WPB72, WKCF73]. Sinus [YCD70]. Sinusoidal [BHS72]. Site [KS71, LL72, LKMB73, RB71]. Sites [BKG73, BMS71, FTCM74, KT74b, KP74, KWR71, MC73, Nic73b, NP73, PRP74, QP73, Sti72, TG70, TKTVJ73, Wel74, WLH74, WMB74, OK71]. Situ [CAH^{+}74, GR71, HCAP74, Hor72, LWD^{+}72]. Size [Ivá71]. Sizing [GB72]. Skeletal [CB70a, CSH71, EKP74, FA74, HSWD72, How74, JHK73, Kel71, LDC73, LA71, ML70, NF71, PG70a, Prz71, Row71, SHP70, SH72a, SWS74, BSGS72]. Skin [BDW71, KK70, Mag73, MPEB71, Pot71, PM73]. Slices [Gil71, BG74b]. Sliding [Sum74]. Slime [CS70, Ett72, GK73a, GK73b, HHC71, OKY72, Tan73]. Slit [RK74b]. Slow [HL74a, LA71]. Small [BL70, CKS74, EWBP71, NN72, Rod70, TC70]. Smooth [CUMB71, CF72, DS71, DSS72, DC70, FC73a, GD70, HB72b, Hig74, Iwa71, PH70, RMM^{+}70, Ros71a, RK71, Ros71c, SDSN71, SSD^{+}74, UB70, UCB71, UB72]. Snakes [OB73]. Snell [KPH^{+}74]. Sodium [HF70, QWT73, Sti72]. Soft [HD70b]. Solid [She70]. Solidissima [Wei72]. Solubilization [RC73]. Solubilizing [HT70]. Soluble [GF72, KR70, KS71, SC72c]. Solution [GB72, IF71, Zob73]. Solutions [BBW74, CS74b, RPK70, SDSN71]. Somata [WMB74]. Somatic [DH74, KL74a, KCR70, Sim70, SHM74, WWvdN74]. Some [GF72, HYC^{+}74, MT70, POB74, SHGH72, SL70, SK70a]. Somite [Min73]. Sorokiniana [Pul74]. Sorting [AP72]. Source [Mül72]. Sp [RK74a, Sha72]. Space [BT70b, FF70a, GP73a, SSF74]. Spaces [DC70, FR74a, WD71]. Span [SH74]. Spatial [KHS74]. Special [Nak71]. Specialization [Ros73, UB72]. Specializations [DC74]. Specialized [BT70b, SL72b]. Species [AP70]. Specific [BB73d, BB74a, DBW73, GM72b, GWT74, HAT71, HH71, LG70, LP71, NHB72, NS74, PGHK72, PW73, Pog72b, Pot71, SH72b, VR74, BSGS72]. Specificity [IWM^{+}74, RMR71a]. Spectrin [NMS71]. Spectrometry [Pad74]. Sperm [BH71c, FR74b, FF74, GG72, GF72, GG74a, Lun72, Phi70, Phi72a, Sum74, THIM73, Tok74, ZBSW70]. Sperm-Egg [BH71c]. Spermatids [Kay70, Phi74]. Spermatocytes [KT74b, LaF74, ST70, SM73]. Spermatogenesis [Car72, Tur70, VT72]. Spermatozoa [AP70, BB72a, BB72b, KP74, MSGE73, Rob72, Sha72, Woo71].

Spermatozoide [PA70]. Spermatozoon [BDR70]. Spermienköpfen [ZBSW70]. Spermiogenesis [KTT72]. Spheres [TR73]. Spheroidal [FS74]. Sphincter [UB72]. Spider [CHM72b, FA74, Hin70, She73]. Spinach [MWR72, OM74]. Spinacia [MWR72]. Spinal [OW74, VHG74]. Spindle [BC71, HMC70, ML71, MR71b, PGR72, Ste73]. Spindles [IBK74, MR71b]. Spiny [Atw71]. Spisula [Wei72]. Spleen [KS72]. Splitting [GG74b, dSB70]. Spoke [Tok74]. Spontaneous [PTF73]. Spore [LM70c]. Spores [CA71]. Spread [KT74b]. Spreading [Lut73, RdS73]. Squid [HL74b, OHPW74]. Squilla [POB71]. Stabilities [BH72]. Stability [CAC^{+}71, MSM74, PKT74]. Stacked [GS71b]. Staeger [WLH74]. Stages [KFJ70, OP74]. Stain [EAK70]. Stained [HK70b]. Staining [CSR^{+}74, FF70b, FB70, GS74, KO70, LKM71, Pay70, SY71, She71a, TS74, WF70]. Stains [LK71, NS74]. Starch [PN70]. Starfish [TSE70]. Stars [Fus73]. Starved [EMJT74, HRW72, RHW71]. State [PH70]. States [SW70]. Steady [JN74]. Stellate [Wak74]. Stem [MG73]. Stentor [HP73]. Stereological [Bol74, EKP74]. Stereoscope [HN70a]. Sternarchus [WPB72]. Steroid [SWD74, RL74]. Steroidogenic [KPH^{+}74]. Stiffness [SY74]. Stimulated [AH73, FS70, JP71b, LH73, MHW72, MMS^{+}72, NB72, Pog72b, PW74, WL73b]. Stimulates [RdS73]. Stimulating [Hop72]. Stimulation [AY70, BPA74, CHM72a, Coo73, EH70, GP73a, GWE74, MC71a, PR71b, SFK74, YCD70]. Stimuli [DGW74]. Stimulus [MLGML71]. Stimulus-Secretion [MLGML71]. Stingaree [ABN70]. Stomach [Atw71, FF70b, Rub72, WS74c]. Stomatopod [POB71]. Storage [Fah74, GN73, IF71, RGGG71, SAE73, TdPP72]. Strain [Pul74, RRYT70, Stu71, GL70, GL71, LP70, TL70b]. Strains [ABBA72, KSE73, PP71b, SH74, TBL70, RW71b]. Strand [WKCF73]. Strand-Containing [WKCF73]. Stranded [LG70, SSC73]. Strands [PLRK72]. Streaming [All74a, Tuc74]. Strength [Gri73b]. Streptococcus [HDM74]. Streptonigrin [Pul74]. Stress [Coh71]. Stretching [Tuc74]. Striated [DLC73, GWS70, LDdV72, LDC73, PHH70, Row73]. Striations [Woo71]. Strider [Tur72]. Stroma [TC71]. Structural [Arm70, AYB^{+}72, Blo70, Bry70a, CT72, Dan73, DC74, EI71, Fah70a, GR70, GS71b, HG72, HB71b, LK72, MHT72, PWT71, RVHWU72, RP72, Roh74, Szo70, VSF70, WS74a, YA74]. Structure [AA74, All71, AS71a, And72, Bal70, BMS73, Bou71, Bun73, Bur70, CUMB71, CB70b, CHM72b, Cus70a, Cus70b, CR70b, DHN73, DLC73, Eri74, Fah74, FS73, FL74, FA73a, FGO71, FC72, GEP71, Goo71, Gor70a, Gor70b, HHS70, Han70a, HLP73, Kay70, KP74, Kur74, Lei71, LC71, MS70b, MCE71, MK70, MP70c, NO71, Nak71, Nic71, Ova71, Ova72, PR72, PBK74, Phi70, RW72, SHP70, SCI73, SM73, Ten70, TPG72, WK74b, War70, WBSW74, ZR73, GL70, LP70, TL70b]. Structure-Function [Cus70a, Cus70b, CR70b]. Structures [All73, Esp71, HL74b, KLR74, KT71, MHB73, Mul71, SG73, Wat72]. Studied [AAG^{+}74, GP73a, GR71, NNQH71, PV70a, Per72]. Studies [AGMT71, AJ74a, AJ74b, ABE71, ABE72, BHdD70, BSG71, Car72, CT72, EAK70, EB72, EWBP71, FA70, FA71, FA73b, GSH71, GB70, Gor70a, Gor70b, GPKJ73, GMS73, HBL70, HB72b, Hig74, HWP73, IS74, JP71a, KFHF74, Kis74, Kri72, KS71, KNY74, Kur74, LG70, Lea71, LP72, LP71, MMT^{+}74, MGHA72, MT71a, MT71b, MT71c, NHB72, NSG71, NT71a, NT71b, OO72, OLK^{+}71, Pel74, RE74, Sha72, SS74b, SRS73, Sto73b, TC72, TO70, TT71, Uge71, VBD70, Web72, WKG72]. Study [ACBW^{+}74, ACWTS^{+}74, And70, Arm70, BACF^{+}74, BACTS^{+}74, Bel71, BW73b, Cha73, Com71, DG71, DBM74, FG70b, GAGSP73, GHN74, GC74, HSH^{+}71, HFK70, KR70, KNM74, LdH71, Lag71, LA70a, MC70, Mat73, MBCP71, NP71, OU70, Pag71, PN70, PTF73, Raw73, RWB70, SA70, VT72, VHG74, WE73]. Studying [SP71a, TBW74]. Stylonychia [Klo70]. Subcellular [ACBW^{+}74, ACWTS^{+}74, BACF^{+}74, BACTS^{+}74, BdN74, DOP74, Ett72, GN73, GJKL73, HFG73, MJ74, O’B72, RKT71, WACBB71, Wid72]. Subfractionation [GD70]. Subhuman [Bel71]. Sublines [SF74a]. Submaxillary [Str71b]. Submembranous [ZF70]. Submicrosomal [TAKS74]. Subplasmalemmal [RA73]. Subsequent [LA70a]. Substrata [EB72, NU74]. Substrate [Cul74, PV70a]. Substrate-Attached [Cul74]. Substratum [Gri73b]. Substructure [RK74b, Ros74, TA70]. Subtilis [MD72, Nan71a, Nan71b, vIG71]. Subunit [CCA70, Eri74, HHP70, PV72]. Subunits [CBS73, HK70b, TKTVJ73]. Suctorian [MR70, Tuc74]. Sugar [GS73]. Sulfate [BY71, Hua74, HA70, KT72, RG72, You73]. Sulfhydryl [KR70]. Sulfonate [SO72a]. Sulfoxide [KKCP73]. Sulfoxide-Induced [KKCP73]. Supergranal [SS74b]. Support [DB74]. Suppression [Nam72, WOSE73]. Surf [Wei72]. Surface [All73, AAWD74, BT70a, BS70a, BL70, BLH74, BBC72, Bos73, DGM^{+}74, DC74, DB74, EF73, Eri74, EB74c, EMJT74, ER74a, ER74b, FLM73, GKB74, GM72b, HHS70, HSH70, HK70b, Iwa71, KT72, LT71, LCRM71, Lut73, Mas73, NMS71, NS74, NLB73b, PWT71, PP74, PPF73, PTF73, RMR71b, RE73, RKT71, RBK74, SY74, She71a, TM70, UBKB74, VDC71, WR74a, Woo74, YY74, YA74, ZS72]. Surface-Bound [RBK74, UBKB74]. Surfaces [STM^{+}74]. Surrounding [HFK70]. Survival [WT73a]. Suspension [HD70b, LC71, Pla72]. Suspensions [Tok73]. Suturalis [LaF74]. Sv40 [CGB71, MCB71, MCB73]. Sv40-Transformed [CGB71, MCB71, MCB73]. Swallowed [OL73]. Sweat [SMS72]. Swelling [AT70, MHB70, MC71b]. Swimming [GG74a]. Swine [MGHA72]. Sympathetic [Bra73, Bun73, EH70, Gri70a, MP73a, MP73b, MP73c, MBKM70, PW74, TH73b]. Synapse [AB71b, HL74b, PAB71]. Synapses [MBCP71, PM72b, She73, VHG74, WYE71]. Synaptic [Bod70, BBW74, CHM73, CT72, CT74, Gri70b, HR73, KK73a, Nak71, PRP74, PW74, SJB70, TH73b, ZR73]. Synaptonemal [HHC71, SM73]. Synaptosomal [GAGGS72]. Synaptosome [GAGGS71, KK73a]. Synchronization [KH70, LM73]. Synchronized [BW70b, JSF70, JFdBJ73, SS72c, WE70]. Synchronizing [TCK72]. Synchronous [BL74]. Synchronously [CWB71]. Synchrony [PR71a]. Syndinium [RK74a]. Synechococcus [EG71]. Synthase [TBB74].

Syntheses [GdB73]. Synthesis [AGTS74, ASO^{+}71, BPA74, BL74, BD72, BHB73, BH70, BW73a, BS72, BW70b, CA74, Cla70, CBWS74, CG72, CP70, DKHL74, FLM73, Fli71, Gal74, GAGGS72, GK72, Gib70, Goo71, GP71, HM74a, HBS70, HN70b, Hen71, HB72a, HMT71, HG71, HS73, HH71, HYC^{+}74, IWM^{+}74, JHK73, JP71b, KDR74, KFJ70, KT74a, KK73b, KS71, Kum70, LDP74, Lat74, LdD73a, LdD73b, Lei71, LP71, LT70, LM73, LH73, LL72, MC70, MC73, MB71a, MC71a, Mil73, MD72, MLMPF74, NB72, OB73, PGHK72, PV72, PR70a, PG70b, PR70b, Pel74, PBK74, Per72, PdWW72, PHdD70, RGGG71, RW71a, RP73, RASP74, RO74, RK73, SO72a, SHNB72, SU70, Sch73, SLWZ70, SCPH73, SSHS74, SHH73, SSK74, SB72, SP71b, SP73, SO72b, SAE73, TVG70, TL70a, WH70, WL74]. Synthesis [WR74b, WW71, ZUJP70, AJ73, GL71, KS74a, PH71]. Synthesized [AGS74]. Synthesizing [RGR71]. Synthetase [BL71a, KTC72, LRL71]. Synthetic [GLG70, GG70, PLRK72, Rud71]. System [DD70, FG70b, GD73, HBK70, LA71, MPAB70, Pag71, PV70a, Per70, PM72b, SSS73, SJB70]. Système [Pag71]. Systems [All71, EB74b].

T [Atw71, FG70b, HN72, Kel71]. Tabacum [GC74]. Tactoids [Ros71b]. Tadpole [EI71]. Tadpoles [VBD70]. Tail [Tok74]. Tapetum [ABN70]. Tapo [DC72]. Tarantula [FA74]. Tarentolae [Str72]. Taste [Far72]. Tcrc [SF74a]. Tcrc-1 [SF74a]. Tcrc-2 [SF74a]. Technique [BOW70, Tok73, TKTVJ73]. Techniques [BV71, GS74, HF70, HLP73]. Teleost [MT74, DDP73, DPD73]. Temperature [HS72b, LW74a, SDSN71, SA71, SA72]. Temperature-Dependent [SDSN71]. Template [AYB^{+}72, NB72]. Temporal [Jef72, JB73]. Tension [CF72, Har70b, Pul74]. Tentacle [Tuc74]. Tentacles [HC73]. Tentans [EDE^{+}71, LWD^{+}72, RR71, EDE72]. Term [MHB73, MCS71]. Terminals [HKF71]. Testes [CMN71]. Testicular [NO71]. Testis [KLT71, KT74a, Tur72]. Testololactone [Sto74]. Testosterone [Sto73a]. Tetanic [CHM72a]. Tetradecanoyl [ZTB74]. Tetrahymena [CLSA74, GdB73, Gor70a, Gor70b, GHP73, GPKJ73, GKP74, HD70a, JSF70, Jef72, JFdBJ73, KZ70, Kum70, LA74, Mül72, MP70d, NT71a, NT71b, OBJ72, RW71a, Ran74, RO74, RB73, RB74b, RC73, SD71a, SS74a, SW70, SAE73, TBW74, UB74, WF73, WN73, WBSW74]. Tetrodotoxin [SMD73]. Tetroxide [PKT74]. Thallium [SPAR73]. Their [BD72, CBS73, Coh71, HSH70, NB72, PC71, Phi70, Sim70, SG73, dW71]. Theophylline [CK73]. Thermodynamic [Ste73]. Thermophilic [EG71]. Thermotropic [WBSW74]. Thick [DS71, LDC73, LK71, NNQH71]. Thickening [VB70]. Thickness [GW71]. Thin [Chr71, HS72b, LK71, NNQH71, SY71, TR73]. Thin-Sectioned [HS72b]. Thioglycolate [Lun72]. Thoracic [Bow73]. Thraustotheca [Hea74]. Three [And72, BBHH74, Gol71, JA74, ST70]. Three-Dimensional [And72, JA74, ST70]. Thrombin [FTCM74]. Throughout [SLWZ70]. thylakoid [DSP70]. Thylakoids [PA73]. Thymic [PG72]. Thymidine [Dal71, IWS72, KLJ^{+}74, KL74a, PE72a, Pot71, PS71]. Thymidine- [Dal71]. Thymidine-Labeled [PS71]. Thymine [OU70]. Thymus [LdH71]. Thyroglobulin [HSH^{+}71]. Thyroid [FW72, GN73, Gro71, HSH^{+}71, SH71, SK70b, SWSK71, WW72, DKC71]. Thyroidectomy [PP73b]. Thyroxin [VBD70]. Thyroxin-Treated [VBD70]. Tight [CG73b, FG72b, Nea73]. Tilia [ED70]. Time [AW72, Kri72, MMS^{+}72, Raw73, SSK74]. Time-Lapse [Kri72]. Time-Sequence [Raw73]. Timing [CG72]. Tissue [APW^{+}73, BS70a, BDH^{+}74, BMS71, CUMB71, CB70a, Chr71, CW74, GJ72, Gil71, GM72b, HS72b, KH70, KH71a, KNY74, NFR^{+}74, NU74, PH70, RWB70, SB73a, SS74b, SW72a, TR73, TT71, dD71, BG74b]. Tissue-Specific [GM72b]. Tissues [FG72b, PV70b, SSF74, Tok73, VR74, WMKJ73]. Toad [CG72, DGS^{+}74, DGM^{+}74, DC70, GCF71, KS72, MHW72, MR71a, WD71]. Tokophrya [Tuc74, MR70]. Tool [TBW74]. Tooth [STM^{+}74]. Topographic [NHS71]. Topographical [UBKB74]. Total [HKF71]. Traced [AP74]. Tracer [Kes70]. Tracers [GR71, SK71, SP71a]. Tracheal [HN70a]. Tracheids [WH73]. Transaminase [vRBvWW74]. Transcription [KT74b, KB72, Pog72b, SG72]. Transection [PP72]. Transfer [Tid70]. Transferase [SKMH70, RL74]. Transformants [PTF73]. Transformation [BR72, BS74, HRWL71, HS72a, Pog72a, RM72, ZS72]. Transformations [Hac72]. Transformed [BTLN73, CGB71, Cul74, GS73, KKCP73, KSN73, MCB71, MCB73, NLB73a, TS74, UBKB74, WWvdN74, WM73a]. Transient [LG72, MHB70]. Transitional [MPAB70]. Translational [Bur70, PR70a, dS72]. Translocation [MMM72]. Transmitter [BK70, CHM73, HR73]. Transport [All74b, AGS74, Bas71, BTPT71, CGL74, DGW74, Eds74, FBS71, GK72, GE70, GS73, GAP70, Hop72, HF70, JP71b, Len72b, LN74, LT74, PC71, PE72a, PM73, RO71, Rod70, Rod73, SC72a, Str74, WBSW74, GL71]. Transporting [FR74a]. Transverse [FS73, LA71, Pag71]. Trapezoid [Nak71]. Traverse [TCK72]. Treated [BdL73, HB72b, Hig74, KD74, KGK72, LA70a, MR74, NSM70, RU70, Ros71a, Ros71b, VBD70, WP72a, dLB74]. Treatment [BW73b, Lun72, MHB73, MS70c, NF71, PR71b, SD71c]. Triad [FA70, FA71, FA73b]. Triazole [WL70]. Tricarboxylic [HKF71]. Trichomonad [Mül73]. Trigeminal [BHB73, Pea72]. Triphosphatase [GG72, GC74, LN74, Nag72]. Triphosphate [GG74a, WP72a, WT73a]. Triplets [Har73]. Tris [DC73]. Tritiated [BIS72, BOW70, GHN74, IWS72, OTE73, Pot71, PS71]. Triton [GG72]. Tritosome [GKB74]. Triturus [BG72]. Trophic [RM72]. Trophozoites [KWR71, RW70b]. Trypanosoma [RW70a, Stu71, RW71b]. Trypsin [Ros71a, Ros71b, SG73]. Trypsin-Treated [Ros71a, Ros71b]. Tube [HM74a]. Tubulaire [Pag71]. Tubular [Bou71, GR73, GP72, LA71, Pag71, SS70a]. Tubule [BTPT71, FS73, HKKSH72, Kes70, SB70b]. Tubules [Atw71, BCG72, CMN71, HN72, Kel71, LS73]. Tubulin [BKK73, IBK74, Wei72]. Tubulovesicular [IS74]. Tumor [CSFC72, HRWL71, NLB73a, OP74, PKT74, PP74, TBL70, TML73, ZTB74]. Tumor-Promoter [ZTB74]. Tumors [She70]. Turnover [CHM73, GJ72, Gro71, LdD73a, LdD73b, LG73, NK71, PHdD70, SP73]. Twisted [HB71c, TPG72]. Two [Bai73, Mag73, McC74, Nak71, NHS71, PP71a, Ros72, SG74, SF74a, TKTVJ73]. Two-Dimensional [NHS71]. Type [Goo71, JGR^{+}71, MR71a, RO71]. Types [Bai73, BLH74, HL74a, Nak71, Ros72, Rub72]. Tyrosine [SG74, TT71, vRBvWW74, WMG74].

Ultracryotomy [Tok73]. Ultracytochemical [SAE73]. Ultramicroincineration [HS72b]. Ultrasensitive [JN74, KR70]. Ultrasonic [MS70c]. Ultrastructural [AY70, BHB73, BSG71, BC71, CGP70, CGL74, DG71, DKC71, DC72, HRWL71, Hac72, HS71a, JR74, KDC71, Lan73, MMT^{+}74, MT70, PV70b, PA73, PP74, RFBB74, SK71, TGW73, VBD70, WWvdN74, Web72, WYE71, Yoo70, ZFG74]. Ultrastructurally [Ros72]. Ultrastructure [ABBA72, AW72, Bas71, Bel71, CSGB70, CF72, DH74, EH70, Gri70c, HRS74, HMS73, HW74, JA74, KPH^{+}74, KSE73, Lun72, Mau70, MW70, Nag72, NH74, NDAT71, PG70a, Per74, Per73c, RW70b, Row71, Row73, Sch70, SJG72, SN71, YSW71]. Ultrathin [BV71, BIS72, DRE70, GW71]. Ultraviolet [Pul74, SO72a]. Unagitated [Rud71]. Unciliated [And74]. Undamped [Har70b]. Under-Replication [GL74]. Unfixed [CA71, HDM74]. Unique [SHDH74, She73]. Uniqueness [Nan71b]. Unit [Nak72]. Universality [Dal72]. Unmyelinated [MHB70]. Unstacked [GS71b]. Unstimulated [KP72]. Unsynchronized [MP70a]. Untersuchungen [ZBSW70]. Unusual [Fri71, LDC73, MR71a, RK74a, Rub72, Sea72, SS70b, Str71b, ZKH73]. Upon [BBC72, LB72, Szu71]. Upper [DB74]. Uptake [BCG72, CH73, FS70, FCH73, Fed74, GNAB70, PE72a, SSC73, SS72c, SJ72, OK71]. Uranyl [LK71]. Urchin [And70, BH72, Bry70a, Bry70b, Con71, GG72, GF72, GG74a, HG71, Kan70, KB72, LA70a, LA70b, Mab73, MKPV74, RGGG71, RYR71, SA70]. Urchins [Fus73]. Urethane [LA70a]. Urethane-Treated [LA70a]. Uridine [BDN72, GAGSP73, SS72c]. Uridine- [SS72c]. Urinary [CBS72, DGS^{+}74, DGM^{+}74, DC70, MHW72, SCB72b, VZRE74]. Urine [Str71a]. Use [Cha73, EB74c, GS70]. Using [BP73a, EAK70, Kes70, KDC71, PN70, PB73, TKTVJ73, WHCFB74]. Utilis [VSHA72]. Utility [TBW74]. Utilization [OU70]. Utilizing [NS74].

Vacuole [All74b, BK73, Det71, JP71a, JB73]. Vacuoles [PD70b, Tre71, WWB71]. Valine [PV72]. Valine-Deprived [PV72]. Variability [HC73]. Variable [SD71b]. Variant [NLB73a]. Variation [SH74, WB70]. Variations [FG72b, MS70c]. Variegatus [Aro73, HC74]. Various [AP70, BLH74, OP74, She70]. Varying [MGD71]. Vascular [DS71, GCF74]. Vasopressin [DGS^{+}74, DGM^{+}74, GCF71]. Vasopressin-Induced [DGM^{+}74]. Vegetative [HN73]. Venom [CHM72b, OB73]. Ventilation [KAI^{+}74]. Ventilation/ [KAI^{+}74]. Ventral [PP72, SWR70]. Ventricle [OO71]. Ventriculaire [Pag71]. Ventricular [FS73, Pag71]. Versus [LW74b]. Vertebrate [AP70, FC73a, JR74, PH70, Roh74, SB70b, SL72b, SWS74, UCB71]. Vesicle [HR73, OBJ72, PW74, SC72a]. Vesicles [Bod70, BBW74, CHM72a, CHM73, Gri70b, IF71, KK73a, KDS73, KS74b, Lee70, MB71b, PRP74, TH73b, dMHM74]. Vesicular [Boq70]. Viable [RM72]. Vibrating [JN74]. Vibrissal [Whi72]. Villous [HMT71]. Vinblastine [BKK73, DW72, HA70, KH71a, KH71b, MS70b]. Vinblastine- [KH71a]. Vinblastine-Induced [KH71b]. Vincristine [KH71a]. Vincristine-Induced [KH71a]. Vinegar [ZKH73]. Viperid [OB73]. Viral [PTF73, RO71]. Viridis [Hol71]. Virus [Cul74, HP70, KSN73, NLB73a, PP74, SS74b, WHG73, WWvdN74, WM73a]. Virus-Infected [PP74]. Virus-Like [SS74b]. Virus-Mediated [WHG73]. Virus-Transformed [Cul74, WM73a]. Visual [BY74a, BY74b, PBGCP74, Per72]. Visualization [BM73, DC72, DC73, OM74, WHCFB74, WL73a, Wel74, WMB74]. Visualized [BL70, WL74]. Vital [EAK70]. Vitamin [CJC70, GJDVC74, JC73, Wak74]. Vitellogenin [DW73]. Vitro [BC71, CF73, Com71, DKHL74, Dan72, Dan73, Eds74, FCH73, Fed74, FW72, FC73b, GP73a, GAP70, HSH^{+}71, HAT71, Kni73, MS70b, MG71, MG73, MC71a, OS72, PS71, SH70, SSC73, Shi71, SWR70, SC72b, SC72c, TS74, TML73, EDE72]. Vivo [Com71, GDH73, GP73a, HSH^{+}71, OP74, PW74, Raw73, SWR70, SSS74, SH72b]. Voltage [Gla74, HN70a]. Volume [CS73, PKT74, SK71]. Volumes [Sim70]. Volvox [Pal73]. Vorticella [All73].

Wall [BT70c, Bur70, DC72, KPF74, LCRM71, LM70c, MMLM74, Ros72]. Walls [LC72]. washout [RL74]. Water [OBJ72, Tur72]. Wave [AS72, SY74]. Waves [GG74a]. Way [BT70b]. Weaver [HD73b]. Weight [EDE^{+}71]. Which [Kni73, Stu71, Tre71]. White [Nag72, NSG71, PG70a, Row73]. Whole [And70, BL70, HK70b, KK73a]. Whorls [HYC^{+}74]. Wi [KK73b]. Wi-38 [KK73b]. Widow [CHM72b]. Wild [Goo71, JGR^{+}71]. Wild-Type [Goo71]. Will [Kni73]. Within [ABN70, Cav72, DU72, GNAB70, HB73, JB73, Len72b, WD71, Wat72]. Without [KDS73, KS74b, LKM71, PdWW72, PFSB72]. Wound [Kra71, RET70].

X [KD74, GG72, Pad74]. X-100 [GG72]. X-537A [KD74]. X-Ray [Pad74]. Xenopus [BW72, BdL73, BW70a, CG72, DW72, KG74, Mil73, dLB74]. XY [ST70]. Xylocaine [MBCP71].

y-1 [DSP70]. Yeast [BH71a, CH73, GPR73, Key73, KPF74, MR71b, SL73b].

Yeasts [JGR^{+}71]. Yellow [Pay70]. Yolk [GK72, HAT71, Mas71, MW71]. Yolk-Crystals [Mas71]. Young [PB74].

Z [ME71, Atw71, BSGS72, FA73a, HS70, Row71, Row73]. Z-Disk [ME71]. Zinc [BS74, RK73, WL73b]. Zonal [Bow73, BBHH74, CKS74]. Zonation [IKY70]. Zonula [WK74b]. Zonulae [CG73b, GG74b, WK74a]. zur [ZBSW70]. Zygnema [MP70b]. Zygote [Gul72]. Zymogen [JP71a]. Zymosan [SCDS73, SDCS74]. Zymosan-Induced [SDCS74].

References

Albertini:1974:ASI

[AA74] David F. Albertini and Everett Anderson. The appearance and structure of intercellular connections during the ontogeny of the rabbit ovarian follicle with particular reference to gap junctions. Journal of Cell Biology, 63(1):234–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/234.

Antoine:1974:PMI

[AAG^{+}74] Jean-Claude Antoine, Stratis Avrameas, Nicholas K. Gonatas, Anna Stieber, and Jacqueline O. Gonatas. Plasma membrane and internalized immunoglobulins of lymph node cells studied with conjugates of antibody or its fab fragments with horseradish peroxidase. Journal of Cell Biology, 63(1): 12–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/1/12.

Allen:1974:CCI

[AAWD74] Howard J. Allen, Charles Ault, Richard J. Winzler, and James F. Danielli. Chemical characterization of the isolated cell surface of Amoeba. Journal of Cell Biology, 60 (1):26–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/1/26.

Anderson:1971:FBB

[AB71a] Richard G. W. Anderson and Robert M. Brenner. The formation of basal bodies (centrioles) in the Rhesus monkey oviduct. Journal of Cell Biology, 50(1):10–??, July 1971.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/50/1/10.

Asada:1971:EAC

[AB71b] Y. Asada and M. V. L. Bennett. Experimental alteration of coupling resistance at an electrotonic synapse. Journal of Cell Biology, 49(1):159–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/1/159.

Aaronson:1974:ANP

[AB74] Robert Peter Aaronson and Günter Blobel. On the attachment of the nuclear pore complex. Journal of Cell Biology, 62(3):746–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/746.

Adoutte:1972:EEC

[ABBA72] A. Adoutte, M. Balmefrézol, J. Beisson, and J. André. The effects of erythromycin and chloramphenicol on the ultrastructure of mitochondria in sensitive and resistant strains of paramecium. Journal of Cell Biology, 54(1):8–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/54/1/8.

April:1971:MLS

[ABE71] Ernest W. April, Philip W. Brandt, and Gerald F. Elliott. The myofilament lattice: Studies on isolated fibers. Journal of Cell Biology, 51(1):72–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/1/72.

April:1972:MLS

[ABE72] Ernest W. April, Philip W. Brandt, and Gerald F. Elliott. The myofilament lattice: Studies on isolated fibers. Journal of Cell Biology, 53(1):53–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/1/53.

Arnott:1970:OMC

[ABN70] H. J. Arnott, A. C. G. Best, and J. A. C. Nicol. Occurrence of melanosomes and of crystal sacs within the same cell in the Tapetum Lucidum of the stingaree. Journal of Cell Biology, 46(2):426–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/2/426.

Adelman:1973:ICF

[ABS73] M. R. Adelman, Gunter Blobel, and David D. Sabatini. An improved cell fractionation procedure for the preparation of rat liver membrane-bound ribosomes. Journal of Cell Biology, 56(1):191–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/1/191.

Amar-Costesec:1974:ASMa

[ACBW^{+}74] Alain Amar-Costesec, Henri Beaufay, Maurice Wibo, Denise Thinès-Sempoux, Ernest Feytmans, Mariette Robbi, and Jacques Berthet. Analytical study of microsomes and isolated subcellular membranes from rat liver. Journal of Cell Biology, 61(1):201–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/1/201.

Amar-Costesec:1974:ASMb

[ACWTS^{+}74] Alain Amar-Costesec, Maurice Wibo, Denise Thinès-Sempoux, Henri Beaufay, and Jacques Berthet. Analytical study of microsomes and isolated subcellular membranes from rat liver. Journal of Cell Biology, 62(3):717–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
62/3/717.

Arsenis:1971:EAC

[AESK71] C. Arsenis, R. Eisenstein, L. W. Soble, and K. E. Kuettner. Enzyme activities in chick embryonic cartilage. Journal of Cell Biology, 49(2):459–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/2/459.

Allen:1971:SS

[AGMT71] C. Freeman Allen, Pearl Good, Hilton H. Mollenhauer, and Clara Totten. Studies on seeds. Journal of Cell Biology, 48 (3):542–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/3/542.

Ambron:1974:ATN

[AGS74] Richard T. Ambron, James E. Goldman, and James H. Schwartz. Axonal transport of newly synthesized glycoproteins in a single identified neuron of Aplysia Californica. Journal of Cell Biology, 61(3):665–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/3/665.

Ambron:1974:SGS

[AGTS74] Richard T. Ambron, James E. Goldman, Elizabeth Barnes Thompson, and James H. Schwartz. Synthesis of glycoproteins in a single identified neuron of Aplysia Californica. Journal of Cell Biology, 61(3):649–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/3/649.

Abrahams:1973:SEI

[AH73] Susan J. Abrahams and Eric Holtzman. Secretion and endocytosis in insulin-stimulated rat adrenal medulla cells. Journal of Cell Biology, 56(2):540–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/2/540.

Aronson:1970:RLA

[AI70] John Aronson and Shinya Inoué. Reversal by light of the action of N-methyl N-desacetyl colchicine on mitosis. Journal of Cell Biology, 45(2):470–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/470.

Armstrong:1973:ORA

[AJ73] Judith E. Armstrong and Russell L. Jones. Osmotic regulation of α-amylase synthesis and polyribosome formation in aleurone cells of barley. Journal of Cell Biology, 59(2): 444–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/2/444.

Amsterdam:1974:SDPa

[AJ74a] A. Amsterdam and J. D. Jamieson. Studies on dispersed pancreatic exocrine cells. Journal of Cell Biology, 63(3): 1037–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/1037.

Amsterdam:1974:SDPb

[AJ74b] A. Amsterdam and J. D. Jamieson. Studies on dispersed pancreatic exocrine cells. Journal of Cell Biology, 63(3): 1057–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/1057.

Allen:1971:FSM

[All71] Richard D. Allen. Fine structure of membranous and microfibrillar systems in the cortex of paramecium caudatum. Journal of Cell Biology, 49(1):1–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/1/1.

Allen:1973:SLM

[All73] Richard D. Allen. Structures linking the myonemes, endoplasmic reticulum, and surface membranes in the contractile ciliate vorticella. Journal of Cell Biology, 56(2):559– ??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/559.

Allen:1974:EFG

[All74a] Nina Strömgren Allen. Endoplasmic filaments generate the motive force for rotational streaming in nitella. Journal of Cell Biology, 63(1):270–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/1/270.

Allen:1974:FVM

[All74b] Richard D. Allen. Food vacuole membrane growth with microtubule-associated membrane transport in paramecium. Journal of Cell Biology, 63(3):904–??, December 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/63/3/

Ab:1970:MRP

[AM70] Geert Ab and Ronald A. Malt. Metabolism of ribosomal precursor ribonucleic acid in kidney. Journal of Cell Biology, 46 (2):362–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/2/362.

Abell:1973:RAC

[AM73a] C. W. Abell and T. M. Monahan. The role of adenosine 3’,5’-cyclic monophosphate in the regulation of mammalian cell division. Journal of Cell Biology, 59(3):549–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/549.

Anastasia:1973:ECC

[AM73b] J. V. Anastasia and R. L. McCarl. Effects of cortisol on cultured rat heart cells. Journal of Cell Biology, 57(1):109–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/1/109.

Anderson:1970:CSC

[And70] Everett Anderson. A cytological study of the centrifuged whole, half, and quarter eggs of the sea urchin Arbacia Punctulata. Journal of Cell Biology, 47(3):711–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
3/711.

Anderson:1972:TDS

[And72] Richard G. W. Anderson. The three-dimensional structure of the basal body from the Rhesus monkey oviduct. Journal of Cell Biology, 54(2):246–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/2/246.

Anderson:1974:ICU

[And74] Richard G. W. Anderson. Isolation of ciliated or unciliated basal bodies from the rabbit oviduct. Journal of Cell Biology, 60(2):393–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/2/393.

Anderson:1970:LGS

[AP70] Winston A. Anderson and Paul Personne. The localization of glycogen in the spermatozoa of various invertebrate and vertebrate species. Journal of Cell Biology, 44(1):29–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
1/29.

Armstrong:1972:CSP

[AP72] Peter B. Armstrong and David Parenti. Cell sorting in the presence of cytochalasin B. Journal of Cell Biology, 55(3): 542–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/3/542.

Atwood:1974:CMN

[AP74] H. L. Atwood and B. Pomeranz. Crustacean motor neuron connections traced by backfilling for electron microscopy. Journal of Cell Biology, 63(1):329–??, October 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/63/1/
329.

Andersson:1973:TCC

[APW^{+}73] Arne Andersson, Birger Petersson, Jan Westman, John Edwards, Gudmar Lundqvist, and Claes Hellerström. Tissue culture of A_{2} cell-rich pancreatic islets isolated from guinea pigs. Journal of Cell Biology, 57(1):241–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/241.

Axline:1974:IPP

[AR74] Stanton G. Axline and Eve P. Reaven. Inhibition of phagocytosis and plasma membrane mobility of the cultivated macrophage by cytochalasin B. Journal of Cell Biology, 62(3):647–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/647.

Armstrong:1970:FSS

[Arm70] Peter B. Armstrong. A fine structural study of adhesive cell junctions in heterotypic cell aggregates. Journal of Cell Biology, 47(1):197–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/1/197.

Aronson:1971:DCS

[Aro71] John F. Aronson. Demonstration of a colcemid-sensitive attractive force acting between the nucleus and a center. Journal of Cell Biology, 51(2):579–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/2/579.

Aronson:1973:NMF

[Aro73] John F. Aronson. Nuclear membrane fusion in fertilized lytechinus variegatus eggs. Journal of Cell Biology, 58(1): 126–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/1/126.

Anderson:1971:FSB

[AS71a] H. Clarke Anderson and Stanley W. Sajdera. The fine structure of bovine nasal cartilage. Journal of Cell Biology, 49 (3):650–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/650.

Anderson:1971:POF

[AS71b] Winston A. Anderson and Andrew Spielman. Permeability of the ovarian follicle of aedes aegypti mosquitoes. Journal of Cell Biology, 50(1):201–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/1/201.

Aiello:1972:MWL

[AS72] Edward Aiello and Michael A. Sleigh. The metachronal wave of lateral cilia of mytilus edulis. Journal of Cell Biology, 54(3):493–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/3/493.

Adelman:1973:RMI

[ASB73] M. R. Adelman, David D. Sabatini, and Günter Blobel. Ribosome-membrane interaction. Journal of Cell Biology, 56 (1):206–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/1/206.

Ashhurst:1970:ID

[Ash70] Doreen E. Ashhurst. An insect desmosome. Journal of Cell Biology, 46(2):421–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/2/421.

Amsterdam:1971:CSM

[ASO^{+}71] Abraham Amsterdam, Michael Schramm, Itzhak Ohad, Yoram Salomon, and Zvi Selinger. Concomitant synthesis of membrane protein and exportable protein of the secretory granule in rat parotid gland. Journal of Cell Biology, 50 (1):187–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/1/187.

Anagnosti:1970:MLA

[AT70] Evelyn Anagnosti and Henry Tedeschi. The mechanism of low-amplitude orthophosphate-induced swelling in isolated rat liver mitochondria. Journal of Cell Biology, 47(2):520– ??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/2/520.

Atwood:1971:ZTS

[Atw71] H. L. Atwood. Z and T tubules in stomach muscles of the spiny lobster. Journal of Cell Biology, 50(1):264–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
1/264.

Auerspebg:1972:MEM

[Aue72] Nelly Auerspebg. Microfilaments in epithelial morphogenesis. Journal of Cell Biology, 52(1):206–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/52/1/
206.

Aist:1972:UTC

[AW72] James R. Aist and P. H. Williams. Ultrastructure and time course of mitosis in the fungus fusarium oxysporum. Journal of Cell Biology, 55(2):368–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/2/368.

Anderson:1970:MUC

[AY70] John Walberg Anderson and Milton B. Yatvin. Metabolic and ultrastructural changes in the frog ovarian follicle in response to pituitary stimulation. Journal of Cell Biology, 46(3):491–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/3/491.

Arnold:1972:SAI

[AYB^{+}72] Eugene A. Arnold, David H. Yawn, David G. Brown, Robert C. Wyllie, and Donald S. Coffey. Structural alteration in isolated rat liver nuclei after removal of template restriction by polyanions. Journal of Cell Biology, 53(3):737–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/3/737.

Bergstrom:1974:NAC

[BA74] Beverly H. Bergstrom and John M. Arnold. Nonkinetochore association of chromatin and microtubules. Journal of Cell Biology, 62(3):917–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/3/917.

Beaufay:1974:ASMa

[BACF^{+}74] Henri Beaufay, Alain Amar-Costesec, Ernest Feytmans, Denise Thinès-Sempoux, Maurice Wibo, Mariette Robbi, and Jacques Berthet. Analytical study of microsomes and isolated subcellular membranes from rat liver. Journal of Cell Biology, 61(1):188–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/1/188.

Beaufay:1974:ASMb

[BACTS^{+}74] Henri Beaufay, Alain Amar-Costesec, Denise Thinès-Sempoux, Maurice Wibo, Mariette Robbi, and Jacques Berthet. Analytical study of microsomes and isolated subcellular membranes from rat liver. Journal of Cell Biology, 61 (1):213–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/1/213.

Bainton:1973:SDT

[Bai73] Dorothy Ford Bainton. Sequential degranulation of the two types of polymorphonuclear leukocyte granules during phagocytosis of microorganisms. Journal of Cell Biology, 58(2): 249–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/2/249.

Baker:1972:FHR

[Bak72] Richard F. Baker. Fusion of human red blood cell membranes. Journal of Cell Biology, 53(1):244–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/1/244.

Baldwin:1970:FSE

[Bal70] Kate M. Baldwin. The fine structure and electrophysiology of heart muscle cell injury. Journal of Cell Biology, 46(3): 455–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/455.

Baskin:1971:UCT

[Bas71] R. J. Baskin. Ultrastructure and calcium transport in crustacean muscle microsomes. Journal of Cell Biology, 48(1): 49–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/49.

Bibring:1971:SEI

[BB71] Thomas Bibring and Jane Baxandall. Selective extraction of isolated mitotic apparatus. Journal of Cell Biology, 48(2): 324–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/2/324.

Bearden:1972:BS

[BB72a] James Bearden and Irwin J. Bendet. Birefringence of spermatozoa. Journal of Cell Biology, 55(2):489–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/489.

Bendet:1972:BS

[BB72b] Irwin J. Bendet and James Bearden. Birefringence of spermatozoa. Journal of Cell Biology, 55(2):501–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/501.

Bernfield:1972:AMG

[BB72c] Merton R. Bernfield and Shib D. Banerjee. Acid mucopolysaccharide (glycosaminoglycan) at the epithelial-Mesenchymal interface of mouse embryo salivary glands. Journal of Cell Biology, 52(3):664–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/3/664.

Black:1972:OCP

[BB72d] Owen Black and Edward Bresnick. Ontogenetic changes of proteins of endoplasmic reticulum. Journal of Cell Biology, 52(3):733–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/733.

Bennett:1973:CCA

[BB73a] Allen Bennett and Lawrence Bogorad. Complementary chromatic adaptation in a filamentous blue-green alga. Journal of Cell Biology, 58(2):419–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/2/419.

Black:1973:PIA

[BB73b] Virginia H. Black and Bruce I. Bogart. Peroxisomes in inner adrenocortical cells of fetal and adult guinea pigs. Journal of Cell Biology, 57(2):345–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/2/345.

Bouck:1973:MBC

[BB73c] G. Benjamin Bouck and David L. Brown. Microtubule biogenesis and cell shape in ochromonas. Journal of Cell Biology, 56(2):340–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/340.

Bretz:1973:AAP

[BB73d] Ursula Bretz and Marco Baggiolini. Association of the alkaline phosphatase of rabbit polymorphonuclear leukocytes with the membrane of the specific granules. Journal of Cell Biology, 59(3):696–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/3/696.

Brown:1973:MBC

[BB73e] David L. Brown and G. Benjamin Bouck. Microtubule biogenesis and cell shape in ochromonas. Journal of Cell Biology, 56(2):360–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/360.

Bretz:1974:BMC

[BB74a] Ursula Bretz and Marco Baggiolini. Biochemical and morphological characterization of azurophil and specific granules of human neutrophilic polymorphonuclear leukocytes. Journal of Cell Biology, 63(1):251–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/1/251.

Brown:1974:MBC

[BB74b] David L. Brown and G. Benjamin Bouck. Microtubule biogenesis and cell shape in ochromonas. Journal of Cell Biology, 61(2):514–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/2/514.

Bernfield:1972:DSE

[BBC72] Merton R. Bernfield, Shib D. Banerjee, and Ronald H. Cohn. Dependence of salivary epithelial morphology and branching morphogenesis upon acid mucopolysaccharide-protein (proteoglycan) at the epithelial surface. Journal of Cell Biology, 52(3):674–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/674.

Bretz:1974:RTD

[BBHH74] Ursula Bretz, Marco Baggiolini, Rolf Hauser, and Christian Hodel. Resolution of three distinct populations of nerve endings from rat brain homogenates by zonal isopycnic centrifugation. Journal of Cell Biology, 61(2):466–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/466.

Boyne:1974:ECC

[BBW74] Alan F. Boyne, Timothy P. Bohan, and Terence H. Williams. Effects of calcium-containing fixation solutions on cholinergic synaptic vesicles. Journal of Cell Biology, 63(3):780– ??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/780.

Brinkley:1971:UAM

[BC71] B. R. Brinkley and Joiner Cartwright. Ultrastructural analysis of mitotic spindle elongation in mammalian cells in vitro. Journal of Cell Biology, 50(2):416–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/416.

Brady:1972:CRC

[BC72] Tom Brady and Mary E. Clutter. Cytolocalization of ribosomal cistrons in plant polytene chromosomes. Journal of Cell Biology, 53(3):827–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/3/827.

Banker:1974:PPD

[BCC74] G. Banker, L. Churchill, and C. W. Cotman. Proteins of the postsynaptic density. Journal of Cell Biology, 63(2): 456–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/2/456.

Bourdeau:1972:SAU

[BCG72] James E. Bourdeau, Frank A. Carone, and Charles E. Ganote. Serum albumin uptake in isolated perfused renal tubules. Journal of Cell Biology, 54(2):382–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/2/382.

Becker:1974:CPG

[BCGL74] Gerald L. Becker, Chung-Ho Chen, John W. Greenawalt, and Albert L. Lehninger. Calcium phosphate granules in the hepatopancreas of the blue crab callinectes sapidus. Journal of Cell Biology, 61(2):316–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/2/316.

Bone:1971:OEE

[BD71] Q. Bone and E. J. Denton. The osmotic effects of electron microscope fixatives. Journal of Cell Biology, 49(3):571–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/571.

Benedetto:1972:CRR

[BD72] Arrigo Benedetto and Wiktor Djaczenko. 37rc cells rapidly recover their RNA synthesis after inhibition with high doses of actinomycin D. Journal of Cell Biology, 52(1):171–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
52/1/171.

Binderman:1974:FBT

[BDH^{+}74] Itzhak Binderman, Dan Duksin, Arieh Harell, Ephraim Katzir (Katchalski), and Leo Sachs. Formation of bone tissue in culture from isolated bone cells. Journal of Cell Biology, 61 (2):427–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/2/427.

Bluemink:1973:NMF

[BdL73] John G. Bluemink and Siegfried W. de Laat. New membrane formation during cytokinesis in normal and cytochalasin Btreated eggs of Xenopus Laevis. Journal of Cell Biology, 59 (1):89–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/1/89.

Bonucci:1973:OIR

[BDM73] E. Bonucci, M. Derenzini, and V. Marinozzi. The organicinorganic relationship in calcified mitochondria. Journal of Cell Biology, 59(1):185–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/1/185.

Burchell:1972:DPS

[BDN72] Brian Burchell, Geoffrey J. Dutton, and Andrew M. Nemeth. Development of phenobarbital-sensitive control mechanisms for uridine Diphosphate glucuronyltransferase activity in chick embryo liver. Journal of Cell Biology, 55(2):448– ??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/448.

Borgers:1974:PSO

[BdN74] M. Borgers and S. de Nollin. The preservation of subcellular organelles of Candida albicans with conventional fixatives. Journal of Cell Biology, 62(2):574–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/2/574.

Baccetti:1970:SA

[BDR70] Baccio Baccetti, Romano Dallai, and Floriana Rosati. The spermatozoon of arthropoda. Journal of Cell Biology, 44(3): 681–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/3/681.

Briggaman:1971:FOB

[BDW71] Robert A. Briggaman, Frederic G. Dalldorf, and Clayton E. Wheeler. Formation and origin of basal lamina and anchoring fibrils in adult human skin. Journal of Cell Biology, 51(2): 384–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/384.

Bergeron:1973:IBR

[BEG73] J. J. M. Bergeron, W. H. Evans, and I. I. Geschwind. Insulin binding to rat liver Golgi fractions. Journal of Cell Biology, 59(3):771–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/3/771.

Bell:1971:CSS

[Bel71] Mary Bell. A comparative study of sebaceous gland ultrastructure in subhuman primates. Journal of Cell Biology, 49 (3):932–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/932.

Bennett:1970:MGG

[Ben70] Gary Bennett. Migration of glycoprotein from Golgi apparatus to cell coat in the columnar cells of the duodenal epithelium. Journal of Cell Biology, 45(3):668–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/668.

Bergeron:1973:GFP

[BESP73] J. J. M. Bergeron, J. H. Ehrenreich, P. Siekevitz, and G. E. Palade. Golgi fractions prepared from rat liver homogenates. Journal of Cell Biology, 59(1):73–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/1/73.

Bainton:1970:SPG

[BF70] Dorothy Ford Bainton and Marilyn G. Farquhar. Segregation and packaging of granule enzymes in eosinophilic leukocytes. Journal of Cell Biology, 45(1):54–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/45/1/54.

Bhisey:1971:CBM

[BF71] Avinash N. Bhisey and Jerome J. Freed. Cross-bridges on the microtubules of cooled interphase HeLa cells. Journal of Cell Biology, 50(2):557–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/2/557.

Brown:1970:SFC

[BFK^{+}70] R. Malcolm Brown, Werner W. Franke, Hans Kleinig, Heinz Falk, and Peter Sitte. Scale formation in chrysophycean algae. Journal of Cell Biology, 45(2):246–??, May 1970. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/45/2/
246.

Barsacchi:1972:CLR

[BG72] Giuseppina Barsacchi and Joseph G. Gall. Chromosomal localization of repetitive DNA in the newt, triturus. Journal of Cell Biology, 54(3):580–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/3/580.

Burgess:1974:AMD

[BG74a] David R. Burgess and Robert D. Grey. Alterations in morphology of developing microvilli elicited by cytochalasin B. Journal of Cell Biology, 62(2):566–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/2/566.

Butcher:1974:ECB

[BG74b] Fred R. Butcher and Rose H. Goldman. Effect of cytochalasin B and colchicine on α-amylase release from rat parotid tissue slices. Journal of Cell Biology, 60(2):519–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
2/519.

Bischoff:1970:IMF

[BH70] Richard Bischoff and Howard Holtzer. Inhibition of myoblast fusion after one round of DNA synthesis in 5bromodeoxyuridine. Journal of Cell Biology, 44(1):134–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/44/1/134.

Bhargava:1971:INY

[BH71a] Madhu M. Bhargava and Harlyn O. Halvorson. Isolation of nuclei from yeast. Journal of Cell Biology, 49(2):423–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
2/423.

Bok:1971:RPE

[BH71b] Dean Bok and Michael O. Hall. The role of the pigment epithelium in the etiology of inherited retinal dystrophy in the rat. Journal of Cell Biology, 49(3):664–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/664.

Brown:1971:SEI

[BH71c] George Gordon Brown and Walter J. Humphreys. Sperm-egg interactions of limulus polyphemus with scanning electron microscopy. Journal of Cell Biology, 51(3):904–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/904.

Brandhorst:1972:SNM

[BH72] Bruce P. Brandhorst and Tom Humphreys. Stabilities of nuclear and messenger RNA molecules in sea urchin embryos. Journal of Cell Biology, 53(2):474–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/474.

Benes:1973:ULP

[BHB73] Francine Benes, Joan A. Higgins, and Russell J. Barrnett. Ultrastructural localization of phospholipid synthesis in the rat trigeminal nerve during myelination. Journal of Cell Biology, 57(3):613–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/3/613.

Baggiolini:1970:FBM

[BHdD70] Marco Baggiolini, James G. Hirsch, and Christian de Duve. Further biochemical and morphological studies of granule fractions from rabbit heterophil leukocytes. Journal of Cell Biology, 45(3):586–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/3/586.

Boquist:1974:IMD

[BHLT74] L. Boquist, B. Hellman, Å. Lernmark, and I.-B. Täljedal. Influence of the mutation “diabetes” on insulin release and islet morphology in mice of different genetic backgrounds. Journal of Cell Biology, 62(1):77–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/1/77.

Bissell:1973:PCA

[BHM73] D. Montgomery Bissell, Lydia E. Hammaker, and Urs A. Meyer. Parenchymal cells from adult rat liver in nonproliferating monolayer culture. Journal of Cell Biology, 59(3): 722–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/3/722.

Bissell:1972:LSC

[BHS72] D. Montgomery Bissell, Lydia Hammaker, and Rudi Schmid. Liver sinusoidal cells. Journal of Cell Biology, 54(1):107–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
54/1/107.

Bernier:1972:DDT

[BIS72] Roch Bernier, Roberto Iglesias, and René Simard. Detection of DNA by tritiated actinomycin D on ultrathin frozen sections. Journal of Cell Biology, 53(3):798–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
3/798.

Butler:1970:PIR

[BJ70] W. H. Butler and J. D. Judah. Preparation of isolated rat liver mitochondria for electron microscopy. Journal of Cell Biology, 44(2):278–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/44/2/278.

Bittner:1970:QAT

[BK70] George D. Bittner and Donald Kennedy. Quantitative aspects of transmitter release. Journal of Cell Biology, 47(3): 585–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/585.

Bowers:1973:CIP

[BK73] Blair Bowers and Edward D. Korn. Cytochemical identification of phosphatase activity in the contractile vacuole of Acanthamoeba Castellanii. Journal of Cell Biology, 59(3): 784–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/3/784.

Bowers:1974:LLB

[BK74] Blair Bowers and Edward D. Korn. Localization of lipophosphonoglycan on both sides of Acanthamoeba plasma membrane. Journal of Cell Biology, 62(2):533–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/62/2/
533.

Beug:1973:DAM

[BKG73] H. Beug, F. E. Katz, and G. Gerisch. Dynamics of antigenic membrane sites relating to cell aggregation in Dictyostelium Discoideum. Journal of Cell Biology, 56(3):647–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
3/647.

Burnside:1973:TDI

[BKK73] Beth Burnside, Chris Kozak, and F. C. Kafatos. Tubulin determination by an isotope dilution-vinblastine precipitation method. Journal of Cell Biology, 59(3):755–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/755.

Bennett:1970:FCC

[BL70] G. Bennett and C. P. Leblond. Formation of cell coat material for the whole surface of columnar cells in the rat small intestine, as visualized by radioautography with L-fucose-^{3}H. Journal of Cell Biology, 46(2):409–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/2/409.

Beif-Lehrer:1971:ADE

[BL71a] Liane Beif-Lehrer. Actinomycin-d enhancement of glutamine synthetase activity in chick embryo retinas cultured in the presence of cortisol. Journal of Cell Biology, 51(1):303– ??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/303.

Bennett:1971:PFL

[BL71b] G. Bennett and C. P. Leblond. Passage of fucose-^{3}H label from the Golgi apparatus into dense and multivesicular bodies in the duodenal columnar cells and hepatocytes of the rat. Journal of Cell Biology, 51(3):875–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/875.

Beck:1974:SCM

[BL74] David P. Beck and R. P. Levine. Synthesis of chloroplast membrane polypeptides during synchronous growth of chlamydomonas reinhardtii. Journal of Cell Biology, 63(3): 759–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/759.

Bennett:1974:MGG

[BLH74] Gary Bennett, C. P. Leblond, and Antonio Haddad. Migration of glycoprotein from the Golgi apparatus to the surface of various cell types as shown by radioautography after labeled fucose injection into rats. Journal of Cell Biology, 60(1): 258–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/1/258.

Bloom:1970:SCN

[Blo70] Sherman Bloom. Structural changes in nuclear envelopes during elongation of heart muscle cells. Journal of Cell Biology, 44(1):218–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/1/218.

Bank:1973:VFD

[BM73] Harvey Bank and Peter Mazur. Visualization of freezing damage. Journal of Cell Biology, 57(3):729–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
3/729.

Bloodgood:1974:FFM

[BM74] Robert A. Bloodgood and Kenneth R. Miller. Freeze-fracture of microtubules and bridges in motile axostyles. Journal of Cell Biology, 62(3):660–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/3/660.

Bosmann:1972:MA

[BMD^{+}72] H. Bruce Bosmann, Marjorie W. Myers, Delena Dehond, Richard Ball, and Kenneth R. Case. Mitochondrial autonomy. Journal of Cell Biology, 55(1):147–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
1/147.

Blanchette-Mackie:1971:SLL

[BMS71] E. Joan Blanchette-Mackie and Robert O. Scow. Sites of lipoprotein lipase activity in adipose tissue perfused with chylomicrons. Journal of Cell Biology, 51(1):1–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/51/1/1.

Blanchette-Mackie:1973:ELL

[BMS73] E. Joan Blanchette-Mackie and Robert O. Scow. Effects of lipoprotein lipase on the structure of chylomicrons. Journal of Cell Biology, 58(3):689–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/3/689.

Benzo:1971:FCD

[BN71] Camillo A. Benzo and Andrew M. Nemeth. Factors controlling development of chick embryo liver cells during organ culture. Journal of Cell Biology, 48(2):235–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
2/235.

Bowers:1972:PAC

[BO72] Blair Bowers and Thomas E. Olszewski. Pinocytosis in Acanthamoeba Castellanii. Journal of Cell Biology, 53(3):681–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/3/681.

Bodian:1970:EMC

[Bod70] David Bodian. An electron microscopic characterization of classes of synaptic vesicles by means of controlled aldehyde fixation. Journal of Cell Biology, 44(1):115–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
1/115.

Bodian:1971:POJ

[Bod71] David Bodian. Presynaptic organelles and junctional integrity. Journal of Cell Biology, 48(3):707–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/48/3/
707.

Bolender:1974:SAG

[Bol74] Robert P. Bolender. Stereological analysis of the guinea pig pancreas. Journal of Cell Biology, 61(2):269–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/269.

Boquist:1970:CVP

[Boq70] Lennart Boquist. Cilia and vesicular particles in the endocrine pancreas of the Mongolian gerbil. Journal of Cell Biology, 45(3):532–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/3/532.

Boshes:1970:DP

[Bos70] R. A. Boshes. Drosophila polyribosomes. Journal of Cell Biology, 46(3):477–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/46/3/477.

Bosmann:1973:MEN

[Bos73] H. Bruce Bosmann. Molecules at the external nuclear surface. Journal of Cell Biology, 59(3):601–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/601.

Bouck:1971:SOI

[Bou71] G. Benjamin Bouck. The structure, origin, isolation, and composition of the tubular mastigonemes of the ochromonas flagellum. Journal of Cell Biology, 50(2):362–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
2/362.

Bordier:1970:IDT

[BOW70] Betrand Bordier, Leonard Ornstein, and Richard P. Wedeen. The intrarenal distribution of tritiated Para-aminohippuric acid determined by a modified technique of section freezedry radioautography. Journal of Cell Biology, 46(3):518– ??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/518.

Bowers:1973:LRT

[Bow73] William E. Bowers. Lysosomes in rat thoracic Duct lymphocytes fractionated by zonal centrifugation. Journal of Cell Biology, 59(1):177–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/1/177.

Blackett:1973:NMA

[BP73a] N. M. Blackett and D. M. Parry. A new method for analyzing electron microscope autoradiographs using hypothetical grain distributions. Journal of Cell Biology, 57(1):9–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/9.

Butcher:1973:CB

[BP73b] Fred R. Butcher and James F. Perdue. Cytochalasin B. Journal of Cell Biology, 56(3):857–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/3/857.

Baseman:1974:SIR

[BPA74] Joel B. Baseman, Domenic Paolini, and Harold Amos. Stimulation by insulin of RNA synthesis in chick fibroblasts. Journal of Cell Biology, 60(1):54–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/1/54.

Bode:1974:APP

[BPHBK74] Folkert Bode, Hannelore Pockrandt-Hemstedt, Karl Baumann, and Rolf Kinne. Analysis of the pinocytic process in rat kidney. Journal of Cell Biology, 63(3):998–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
3/998.

Been:1972:CSP

[BR72] Anita C. Been and Ellen M. Rasch. Cellular and secretory proteins of the salivary glands of sciara coprophila during the larval-pupal transformation. Journal of Cell Biology, 55(2): 420–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/420.

Brinkley:1973:NOE

[BR73] B. R. Brinkley and Potu N. Rao. Nitrous oxide: Effects on the mitotic apparatus and chromosome movement in HeLa cells. Journal of Cell Biology, 58(1):96–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/1/96.

Bray:1973:BPI

[Bra73] D. Bray. Branching patterns of individual sympathetic neurons in culture. Journal of Cell Biology, 56(3):702–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
3/702.

Bullivant:1972:FFM

[BRB^{+}72] S. Bullivant, D. G. Rayns, W. S. Bertaud, J. P. Chalcroft, and G. F. Grayston. Freeze-fractured myosin filaments. Journal of Cell Biology, 55(2):520–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/2/520.

Bryan:1970:IMS

[Bry70a] Joseph Bryan. The isolation of a major structural element of the sea urchin fertilization membrane. Journal of Cell Biology, 44(3):635–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/3/635.

Bryan:1970:RCC

[Bry70b] Joseph Bryan. On the reconstitution of the crystalline components of the sea urchin fertilization membrane. Journal of Cell Biology, 45(3):606–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/3/606.

Barland:1970:NRM

[BS70a] Peter Barland and Eleanor A. Schroeder. A new rapid method for the isolation of surface membranes from tissue culture cells. Journal of Cell Biology, 45(3):662–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/662.

Berger:1970:IBC

[BS70b] Sosamma J. Berger and Bertram Sacktor. Isolation and biochemical characterization of brush borders from rabbit kidney. Journal of Cell Biology, 47(3):637–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
3/637.

Blobel:1970:CPN

[BS70c] G. Blobel and D. D. Sabatini. Controlled proteolysis of nascent polypeptides in rat liver cell fractions. Journal of Cell Biology, 45(1):130–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/1/130.

Baba:1971:HLD

[BS71] Nobuhisa Baba and Hari M. Sharma. Histochemistry of lactic dehydrogenase in heart and pectoralis muscles of rat. Journal of Cell Biology, 51(3):621–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/3/621.

Borun:1972:SAC

[BS72] T. W. Borun and G. S. Stein. The synthesis of acidic chromosomal proteins during the cell cycle of HeLa s-3 cells. Journal of Cell Biology, 52(2):308–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/2/308.

Burns:1973:IIP

[BS73] E. Robert Burns and Lawrence E. Scheving. Isoproterenolinduced phase shifts in circadian rhythm of mitosis in murine corneal epithelium. Journal of Cell Biology, 56(2):605– ??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/605.

Berger:1974:CLT

[BS74] Nathan A. Berger and Sister Ann Marie Skinner. Characterization of lymphocyte transformation induced by zinc ions. Journal of Cell Biology, 61(1):45–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/1/45.

Benson:1971:SUL

[BSG71] Robert L. Benson, Bertram Sacktor, and John W. Greenawalt. Studies on the ultrastructural localization of intestinal disaccharidases. Journal of Cell Biology, 48(3):711–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
3/711.

Busch:1972:CSR

[BSGS72] Wayne A. Busch, M. H. Stromer, Darrel E. Goll, and A. Suzuki. Ca^{2+}-specific removal of Z lines from rabbit skeletal muscle. Journal of Cell Biology, 52(2):367–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
2/367.

Ballard:1970:GIA

[BT70a] Philip L. Ballard and Gordon M. Tomkins. Glucocorticoidinduced alteration of the surface membrane of cultured hepatoma cells. Journal of Cell Biology, 47(1):222–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/222.

Bennett:1970:ECB

[BT70b] M. V. L. Bennett and J. P. Trinkaus. Electrical coupling between embryonic cells by way of extracellular space and specialized junctions. Journal of Cell Biology, 44(3):592–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/44/3/592.

Briles:1970:REE

[BT70c] Eve Barak Briles and Alexander Tomasz. Radioautographic evidence for equatorial wall growth in a Gram-positive bacterium. Journal of Cell Biology, 47(3):786–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
3/786.

Bissell:1971:MFC

[BT71] D. Montgomery Bissell and Jeremiah G. Tilles. Morphology and function of cells of human embryonic liver in monolayer culture. Journal of Cell Biology, 50(1):222–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
1/222.

Bissell:1972:RSA

[BT72] D. Montgomery Bissell and Jeremiah G. Tilles. Radioimmunoassay of serum albumin produced by human embryonic liver in monolayer culture. Journal of Cell Biology, 53 (3):819–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/3/819.

Brailovsky:1973:GNT

[BTLN73] C. Brailovsky, M. Trudel, R. Lallier, and V. N. Nigam. Growth of normal and transformed rat embryo fibroblasts. Journal of Cell Biology, 57(1):124–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/1/124.

Bentzel:1971:BTH

[BTPT71] C. J. Bentzel, D. R. Tourville, B. Parsa, and T. B. Tomasi. Bidirectional transport of horseradish peroxidase in proximal tubule of necturus kidney. Journal of Cell Biology, 48(1): 197–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/197.

Bunge:1973:FSN

[Bun73] Mary Bartlett Bunge. Fine structure of nerve fibers and growth cones of isolated sympathetic neurons in culture. Journal of Cell Biology, 56(3):713–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/3/713.

Burton:1970:ODT

[Bur70] Paul R. Burton. Optical diffraction and translational reinforcement of microtubules having a prominent helical wall structure. Journal of Cell Biology, 44(3):693–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/693.

Bernhard:1971:ITP

[BV71] W. Bernhard and Annie Viron. Improved techniques for the preparation of ultrathin frozen sections. Journal of Cell Biology, 49(3):731–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/731.

Bonneville:1970:BBD

[BW70a] Mary A. Bonneville and Melvyn Weinstock. Brush border development in the intestinal absorptive cells of Xenopus during metamorphosis. Journal of Cell Biology, 44(1):151–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
44/1/151.

Bosmann:1970:SGG

[BW70b] H. Bruce Bosmann and R. Alan Winston. Synthesis of glycoprotein, glycolipid, protein, and lipid in synchronized l5178y cells. Journal of Cell Biology, 45(1):23–??, April 1970. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/45/1/23.

Baker:1972:ICC

[BW72] P. F. Baker and Anne E. Warner. Intracellular calcium and cell cleavage in early embryos of Xenopus Laevis. Journal of Cell Biology, 53(2):579–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/2/579.

Black:1973:PSP

[BW73a] Owen Black and Paul D. Webster. Protein synthesis in pancreas of fasted pigeons. Journal of Cell Biology, 57(1):1–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/1/1.

Bolender:1973:MSR

[BW73b] Robert P. Bolender and Ewald R. Weibel. A morphometric study of the removal of phenobarbital-induced membranes from hepatocytes after cessation of treatment. Journal of Cell Biology, 56(3):746–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/3/746.

Broekman:1974:IMI

[BWC74] M. Johan Broekman, Nelson P. Westmoreland, and Phin Cohen. An improved method for isolating alpha granules and mitochondria from human platelets. Journal of Cell Biology, 60(2):507–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/2/507.

Berg:1971:SMP

[BY71] Norton B. Berg and Richard W. Young. Sulfate metabolism in pancreatic acinar cells. Journal of Cell Biology, 50(2): 469–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/2/469.

Bibb:1974:RFA

[BY74a] Carol Bibb and Richard W. Young. Renewal of fatty acids in the membranes of visual cell outer segments. Journal of Cell Biology, 61(2):327–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/2/327.

Bibb:1974:RGV

[BY74b] Carol Bibb and Richard W. Young. Renewal of glycerol in the visual cells and pigment epithelium of the frog retina. Journal of Cell Biology, 62(2):378–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/2/378.

Cole:1971:DMF

[CA71] Garry T. Cole and Henry C. Aldrich. Demonstration of myelin figures in unfixed, freeze-etched fungus spores. Journal of Cell Biology, 51(3):873–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/3/873.

Churg:1974:IEP

[CA74] Andrew Churg and Winston A. Anderson. Induction of endometrial peroxidase synthesis and secretion by estrogen and estrogen antagonist. Journal of Cell Biology, 62(2): 449–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/449.

Carstensen:1971:SCF

[CAC^{+}71] E. L. Carstensen, William G. Aldridge, S. Z. Child, P. Sullivan, and H. H. Brown. Stability of cells fixed with glutaraldehyde and acrolein. Journal of Cell Biology, 50(2): 529–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/2/529.

Conkie:1974:SLG

[CAH^{+}74] D. Conkie, N. Affara, P. R. Harrison, J. Paul, and K. Jones. In situ localization of globin messenger RNA formation. Journal of Cell Biology, 63(2):414–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/414.

Carothers:1972:SSH

[Car72] Zane B. Carothers. Studies of spermatogenesis in the hepaticae. Journal of Cell Biology, 52(2):273–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
2/273.

Cave:1972:LRD

[Cav72] Mac Donald Cave. Localization of ribosomal DNA within oocytes of the house Cricket, acheta domesticus (orthoptera: Gryllidae). Journal of Cell Biology, 55(2):310–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/310.

Canonico:1970:LSM

[CB70a] Peter G. Canonico and John W. C. Bird. Lysosomes in skeletal muscle tissue. Journal of Cell Biology, 45(2):321–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/321.

Chalcroft:1970:ILC

[CB70b] J. P. Chalcroft and S. Bullivant. An interpretation of liver cell membrane and junction structure based on observation of freeze-fracture replicas of both sides of the fracture. Journal of Cell Biology, 47(1):49–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/1/49.

Chlapowski:1971:ALMa

[CB71a] Francis J. Chlapowski and R. Neal Band. Assembly of lipids into membranes in Acanthamoeba Palestinensis. Journal of Cell Biology, 50(3):625–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/3/625.

Chlapowski:1971:ALMb

[CB71b] Francis J. Chlapowski and R. Neal Band. Assembly of lipids into membranes in Acanthamoeba Palestinensis. Journal of Cell Biology, 50(3):634–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/3/634.

Cantrell:1972:BHA

[CB72] Elroy Cantrell and Edward Bresnick. Benzpyrene hydroxylase activity in isolated parenchymal and nonparenchymal cells of rat liver. Journal of Cell Biology, 52(2):316–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
52/2/316.

Cotman:1974:IPD

[CBCT74] C. W. Cotman, G. Banker, L. Churchill, and D. Taylor. Isolation of postsynaptic densities from rat brain. Journal of Cell Biology, 63(2):441–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/2/441.

Chlapowski:1972:LPM

[CBS72] Francis J. Chlapowski, Mary A. Bonneville, and L. Andrew Staehelin. Lumenal plasma membrane of the urinary bladder. Journal of Cell Biology, 53(1):92–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/1/92.

Chua:1973:ICC

[CBS73] Nam-Hai Chua, G. Blobel, and P. Siekevitz. Isolation of cytoplasmic and chloroplast ribosomes and their dissociation into active subunits from chlamydomonas reinhardtii. Journal of Cell Biology, 57(3):798–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/3/798.

Clutter:1974:MSD

[CBWS74] M. Clutter, T. Brady, V. Walbot, and I. Sussex. Macromolecular synthesis during plant embryogeny. Journal of Cell Biology, 63(3):1097–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/3/1097.

Ceccarini:1970:RSE

[CCA70] Costante Ceccarini, Maria S. Campo, and Franca Andronico. Ribosomal subunit exchange in Dictyostelium Purpureum. Journal of Cell Biology, 46(2):428–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/2/428.

Clark:1973:PIM

[CD73] Margaret K. Clark and Paul A. Dahm. Phenobarbitalinduced, membrane-like scrolls in the oenocytes of Musca Domestica Linnaeus. Journal of Cell Biology, 56(3):870–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/56/3/870.

Chui:1971:EEF

[CDMR71] David H. K. Chui, Meir Djaldetti, Paul A. Marks, and Richard A. Rifkind. Erythropoietin effects on fetal mouse erythroid cells. Journal of Cell Biology, 51(3):585–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
51/3/585.

Cooke:1972:CBF

[CF72] Peter H. Cooke and Fredric S. Fay. Correlation between fiber length, ultrastructure, and the length-tension relationship of mammalian smooth muscle. Journal of Cell Biology, 52(1): 105–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/1/105.

Cavoto:1973:LRP

[CF73] Francis V. Cavoto and B. Allen Flaxman. Low-resistance pathways between mitotic and interphase epidermal cells in vitro. Journal of Cell Biology, 58(1):223–??, July 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/58/1/
223.

Caulfield:1974:PGC

[CF74] John P. Caulfield and Marilyn G. Farquhar. The permeability of glomerular capillaries to graded dextrans. Journal of Cell Biology, 63(3):883–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/3/883.

Craig:1970:DLE

[CG70] I. W. Craig and A. Gibor. A direct light effect on maintaining photosynthetic activity of nitella chloroplasts. Journal of Cell Biology, 44(2):305–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/44/2/305.

Chatterjee:1971:PNI

[CG71a] Samar Chatterjee and Lester Goldstein. Proteins in nucleocytoplasmic interactions. Journal of Cell Biology, 48(1): 202–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/202.

Clawson:1971:M

[CG71b] C. C. Clawson and Robert A. Good. Micropapillae. Journal of Cell Biology, 48(1):207–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/1/207.

Craig:1971:CDC

[CG71c] I. W. Craig and A. Gibor. Chloroplast development in charophyceae. Journal of Cell Biology, 49(3):950–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/950.

Coggins:1972:TMD

[CG72] Lesley Watson Coggins and Joseph G. Gall. The timing of meiosis and DNA synthesis during early oogenesis in the toad, Xenopus Laevis. Journal of Cell Biology, 52(3): 569–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/3/569.

Chang:1973:LAL

[CG73a] Cheng-Ming Chang and Robert D. Goldman. The localization of actin-like fibers in cultured neuroblastoma cells as revealed by heavy meromyosin binding. Journal of Cell Biology, 57(3):867–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/3/867.

Claude:1973:FFZ

[CG73b] Philippa Claude and Daniel A. Goodenough. Fracture faces of zonulae occludentes from “tight” and “leaky” epithelia. Journal of Cell Biology, 58(2):390–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/2/390.

Culp:1971:CIR

[CGB71] Lloyd A. Culp, William J. Grimes, and Paul H. Black. Contact-inhibited revertant cell lines isolated from sv40transformed cells. Journal of Cell Biology, 50(3):682–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/50/3/682.

Chen:1974:BUA

[CGL74] Chung-Ho Chen, John W. Greenawalt, and Albert L. Lehninger. Biochemical and ultrastructural aspects of Ca^{2+} transport by mitochondria of the hepatopancreas of the blue crab Callinectes Sapidus. Journal of Cell Biology, 61(2): 301–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/2/301.

Charalampous:1973:IPP

[CGM73] F. C. Charalampous, N. K. Gonatas, and A. D. Melbourne. Isolation and properties of the plasma membrane of Kb cells. Journal of Cell Biology, 59(2):421–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
2/421.

Chacko:1974:ICP

[CGMD74] George K. Chacko, David E. Goldman, Harish C. Malhotra, and Maynard M. Dewey. Isolation and characterization of plasma membrane fractions from garfish lepisosteus osseus olfactory nerve. Journal of Cell Biology, 62(3):831– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/831.

Chan:1970:BUP

[CGP70] T. L. Chan, John W. Greenawalt, and Peter L. Pedersen. Biochemical and ultrastructural properties of a mitochondrial inner membrane fraction deficient in outer membrane and matrix activities. Journal of Cell Biology, 45(2):291–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/291.

Carter:1973:PCR

[CH73] B. L. A. Carter and H. O. Halvorson. Periodic changes in rate of amino acid uptake during yeast cell cycle. Journal of Cell Biology, 58(2):401–??, August 1973. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/2/401.

Chambers:1973:URR

[Cha73] Velma C. Chambers. The use of ruthenium red in an electron microscope study of cytophagocytosis. Journal of Cell Biology, 57(3):872–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/3/872.

Chlapowski:1971:RAP

[Chl71] Francis J. Chlapowski. Regularly arranged particles on the inner lamella of the plasma membrane of Acanthamoeba Palestinensis. Journal of Cell Biology, 51(3):893–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
51/3/893.

Ceccarelli:1972:DVF

[CHM72a] B. Ceccarelli, W. P. Hurlbut, and A. Mauro. Depletion of vesicles from frog neuromuscular junctions by prolonged tetanic stimulation. Journal of Cell Biology, 54(1):30–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
54/1/30.

Clark:1972:CFS

[CHM72b] Allen W. Clark, William P. Hurlbut, and Alexander Mauro. Changes in the fine structure of the neuromuscular junction of the frog caused by black widow spider venom. Journal of Cell Biology, 52(1):1–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/1/1.

Ceccarelli:1973:TTS

[CHM73] B. Ceccarelli, W. P. Hurlbut, and A. Mauro. Turnover of transmitter and synaptic vesicles at the frog neuromuscular junction. Journal of Cell Biology, 57(2):499–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
2/499.

Christensen:1971:FTS

[Chr71] A. Kent Christensen. Frozen thin sections of fresh tissue for electron microscopy, with a description of pancreas and liver. Journal of Cell Biology, 51(3):772–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/772.

Chan:1973:CEC

[CI73] Lee-Nien L. Chan and V. M. Ingram. Culture of erythropoietic cells from chick blastoderms. Journal of Cell Biology, 56 (3):861–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/861.

Carell:1970:VBM

[CJC70] Edgar F. Carell, Pamela Leban Johnston, and Alan R. Christopher. Vitamin B_{12} and the macromolecular composition of euglena. Journal of Cell Biology, 47(2):525– ??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/2/525.

Chapman:1973:PCA

[CJMB73] Geoffrey S. Chapman, Albert L. Jones, Urs A. Meyer, and D. Montgomery Bissell. Parenchymal cells from adult rat liver in nonproliferating monolayer culture. Journal of Cell Biology, 59(3):735–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/3/735.

Castle:1972:RAS

[CJP72] J. David Castle, James D. Jamieson, and George E. Palade. Radioautographic analysis of the secretory process in the parotid acinar cell of the rabbit. Journal of Cell Biology, 53(2):290–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/290.

Cox:1973:DPE

[CK73] Joyce P. Cox and Manfred L. Karnovsky. The depression of phagocytosis by exogenous cyclic nucleotides, prostaglandins, and theophylline. Journal of Cell Biology, 59(2):480–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/59/2/480.

Connock:1974:ZRM

[CKS74] M. J. Connock, P. R. Kirk, and A. P. Sturdee. A zonal rotor method for the preparation of microperoxisomes from epithelial cells of guinea pig small intestine. Journal of Cell Biology, 61(1):123–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/1/123.

Chertok:1972:SMP

[CL72] Robert J. Chertok and Suzanne Lake. A simple method for the preparation of renal brush borders. Journal of Cell Biology, 54(2):426–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/2/426.

Claude:1970:GDC

[Cla70] Albert Claude. Growth and differentiation of cytoplasmic membranes in the course of lipoprotein granule synthesis in the hepatic cell. Journal of Cell Biology, 47(3):745– ??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/745.

Cassady:1972:SMM

[CLBW72] W. E. Cassady, E. H. Leiter, A. Bergquist, and R. P. Wagner. Separation of mitochondrial membranes of neurospora crassa. Journal of Cell Biology, 53(1):66–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/1/66.

Curgy:1974:MCR

[CLSA74] Jean-Jacques Curgy, Gérard Ledoigt, Barbara J. Stevens, and Jean André. Mitochondrial and cytoplasmic ribosomes from Tetrahymena Pyriformis. Journal of Cell Biology, 60 (3):628–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/3/628.

Cunningham:1971:IGC

[CMN71] W. P. Cunningham, H. H. Mollenhauer, and S. E. Nyquist. Isolation of germ cell Golgi apparatus from seminiferous tubules of rat testes. Journal of Cell Biology, 51(1):273– ??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/273.

Chapman:1973:LEK

[CMO73] M. John Chapman, Leonard R. Miller, and Joseph A. Ontko. Localization of the enzymes of ketogenesis in rat liver mitochondria. Journal of Cell Biology, 58(2):284–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
2/284.

Cheetham:1971:IGA

[CMPF71] R. D. Cheetham, D. James Morré, Carol Pannek, and Daniel S. Friend. Isolation of a Golgi apparatus-rich fraction from rat liver. Journal of Cell Biology, 49(3):899–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/899.

Cheetham:1970:IGA

[CMY70] R. D. Cheetham, D. James Morré, and Wayne N. Yunghans. Isolation of a Golgi apparatus-rich fraction from rat liver. Journal of Cell Biology, 44(3):492–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/44/3/492.

Cohen:1971:EMO

[Coh71] Adolph I. Cohen. Electron microscope observations on form changes in photoreceptor outer segments and their saccules in response to osmotic stress. Journal of Cell Biology, 48 (3):547–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/3/547.

Colby:1971:IBG

[Col71] Richard H. Colby. Intrinsic birefringence of glycerinated myofibrils. Journal of Cell Biology, 51(3):763–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/763.

Combs:1971:EMS

[Com71] John W. Combs. An electron microscope study of mouse mast cells arising in vivo and in vitro. Journal of Cell Biology, 48(3):676–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/3/676.

Comly:1973:MCC

[Com73] Lucy T. Comly. Microfilaments in chaos carolinensis. Journal of Cell Biology, 58(1):230–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/1/230.

Conway:1971:ERH

[Con71] Carolyn M. Conway. Evidence for RNA in the heavy bodies of sea urchin eggs. Journal of Cell Biology, 51(3):889– ??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/889.

Cooper:1973:DRL

[Coo73] Herbert L. Cooper. Degradation of 28s RNA late in ribosomal RNA maturation in nongrowing lymphocytes and its reversal after growth stimulation. Journal of Cell Biology, 59(1): 250–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/1/250.

Corradino:1973:ECI

[Cor73] R. A. Corradino. Embryonic chick intestine in organ culture. Journal of Cell Biology, 58(1):64–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/1/64.

Coss:1974:MCR

[Cos74] Ronald A. Coss. Mitosis in chlamydomonas reinhardtii basal bodies and the mitotic apparatus. Journal of Cell Biology, 63(1):325–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/1/325.

Craig:1970:AIM

[CP70] Nessly C. Craig and Robert P. Perry. Aberrant intranucleolar maturation of ribosomal precursors in the absence of protein synthesis. Journal of Cell Biology, 45(3):554–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/554.

Coss:1974:EIP

[CPH74] Ronald A. Coss and Jeremy D. Pickett-Heaps. The effects of isopropyl N-phenyl carbamate on the green alga Oedogonium Cardiacum. Journal of Cell Biology, 63(1):84–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/84.

Coyne:1970:FES

[CR70a] Brian Coyne and Joel L. Rosenbaum. Flagellar elongation and shortening in chlamydomonas. Journal of Cell Biology, 47(3):777–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/3/777.

Cushman:1970:SFRc

[CR70b] Samuel W. Cushman and Martin A. Rizack. Structurefunction relationships in the adipose cell. Journal of Cell Biology, 46(2):354–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/2/354.

Cocucci:1970:RCN

[CS70] S. M. Cocucci and M. Sussman. RNA in cytoplasmic and nuclear fractions of cellular slime mold amebas. Journal of Cell Biology, 45(2):399–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/399.

Crissman:1973:RSM

[CS73] Harry A. Crissman and John A. Steinkamp. Rapid, simultaneous measurement of DNA, protein, and cell volume in single cells from large mammalian cell populations. Journal of Cell Biology, 59(3):766–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/3/766.

Carabez:1974:MIM

[CS74a] A. Carabez and F. Sandoval. A mitochondrial inner membrane preparation that sediments at 100 g. Journal of Cell Biology, 62(3):877–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/3/877.

Chan:1974:ICC

[CS74b] R. T. L. Chan and I. E. Scheffler. Isopycnic centrifugation of chromatin in renografin solutions. Journal of Cell Biology, 61(3):780–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/3/780.

Cavallo:1972:TA

[CSFC72] Tito Cavallo, Robert Sade, Judah Folkman, and Ramzi S. Cotran. Tumor angiogenesis. Journal of Cell Biology, 54(2): 408–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/2/408.

Childress:1970:IUB

[CSGB70] Charles C. Childress, Bertram Sacktor, I. William Grossman, and Ernest Bueding. Isolation, ultrastructure, and biochemical characterization of glycogen in insect flight muscle. Journal of Cell Biology, 45(1):83–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/45/1/83.

Comissiong:1971:IES

[CSH71] E. A. Comissiong, D. W. Stanley, and H. O. Hultin. Inhibition of emptying of skeletal muscle cell segments by adenine nucleotides and polyvalent cations. Journal of Cell Biology, 49(2):430–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/430.

Cunningham:1974:EPN

[CSR^{+}74] William P. Cunningham, L. Andrew Staehelin, Robert W. Rubin, Ross Wilkins, and Mary Bonneville. Effects of phosphotungstate negative staining on the morphology of the isolated Golgi apparatus. Journal of Cell Biology, 62(2): 491–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/491.

Cotman:1972:ISS

[CT72] Carl W. Cotman and Dwan Taylor. Isolation and structural studies on synaptic complexes from rat brain. Journal of Cell Biology, 55(3):696–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/3/696.

Cotman:1974:LCC

[CT74] C. W. Cotman and D. Taylor. Localization and characterization of concanavalin a receptors in the synaptic cleft. Journal of Cell Biology, 62(1):236–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/1/236.

Culp:1974:SAG

[Cul74] Lloyd A. Culp. Substrate-attached glycoproteins mediating adhesion of normal and virus-transformed mouse fibroblasts. Journal of Cell Biology, 63(1):71–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/1/71.

Cummings:1972:IPC

[Cum72] Donald J. Cummings. Isolation and partial characterization of macro- and micronuclei from paramecium aurelia. Journal of Cell Biology, 53(1):105–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/1/105.

Campbell:1971:FSS

[CUMB71] Gordon R. Campbell, Yasuo Uehara, Gerda Mark, and Geoffrey Burnstock. Fine structure of smooth muscle cells grown in tissue culture. Journal of Cell Biology, 49(1):21–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
1/21.

Cushman:1970:SFRa

[Cus70a] Samuel W. Cushman. Structure-function relationships in the adipose cell. Journal of Cell Biology, 46(2):326–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
2/326.

Cushman:1970:SFRb

[Cus70b] Samuel W. Cushman. Structure-function relationships in the adipose cell. Journal of Cell Biology, 46(2):342–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
2/342.

Cassady:1971:SMM

[CW71] William E. Cassady and Robert P. Wagner. Separation of mitochondrial membranes of neurospora crassa. Journal of Cell Biology, 49(2):536–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/2/536.

Cope:1974:IPP

[CW74] G. H. Cope and M. A. Williams. Improved preservation of parotid tissue for electron microscopy. Journal of Cell Biology, 60(1):292–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/1/292.

Chapman:1971:APL

[CWB71] J. D. Chapman, R. G. Webb, and J. Borsa. ATP pool levels in synchronously growing Chinese hamster cells. Journal of Cell Biology, 49(1):229–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/1/229.

Campbell:1971:PEE

[CWMH71] G. Le M. Campbell, H. Weintraub, B. H. Mayall, and H. Holtzer. Primitive erythropoiesis in early chick embryogenesis. Journal of Cell Biology, 50(3):669–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/669.

Conrad:1973:MPL

[CWT^{+}73] Gary W. Conrad, Daniel C. Williams, F. Rudolf Turner, Kenneth M. Newrock, and Rudolf A. Raff. Microfilaments in the polar lobe constriction of fertilized eggs of ilyanassa obsoleta. Journal of Cell Biology, 59(1):228–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/228.

Dallman:1971:MIT

[Dal71] Peter R. Dallman. Malnutrition: Incorporation of thymidine-3H into nuclear and mitochondrial DNA. Journal of Cell Biology, 51(2):549–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/2/549.

Dales:1972:CUM

[Dal72] Samuel Dales. Concerning the universality of a microtubule antigen in animal cells. Journal of Cell Biology, 52(3): 748–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/3/748.

Daniels:1972:CIN

[Dan72] Mathew P. Daniels. Colchicine inhibition of nerve fiber formation in vitro. Journal of Cell Biology, 53(1):164–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
1/164.

Daniels:1973:FSC

[Dan73] Mathew P. Daniels. Fine structural changes in neurons and nerve fibers associated with colchicine inhibition of nerve fiber formation in vitro. Journal of Cell Biology, 58(2): 463–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/2/463.

Davidson:1973:PAR

[Dav73] Samuel James Davidson. Protein absorption by renal cells. Journal of Cell Biology, 59(1):213–??, October 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/59/1/
213.

DiPasquale:1974:UCS

[DB74] Albert DiPasquale and Paul B. Bell. The upper cell surface: Its inability to support active cell movement in culture. Journal of Cell Biology, 62(1):198–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/1/198.

Deman:1974:SMI

[DBM74] Joris J. Deman, Erik A. Bruyneel, and Marc M. Mareel. A study on the mechanism of intercellular adhesion. Journal of Cell Biology, 60(3):641–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/60/3/641.

Djaczenko:1970:FHP

[DBP70] Wiktor Djaczenko, Arrigo Benedetto, and Renato Pezzi. Formation of helical polyribosomes in poliovirus-infected cells of the 37 Rc line. Journal of Cell Biology, 45(1):173–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
1/173.

deBianchi:1973:FSS

[dBTL73] A. G. de Bianchi, W. R. Terra, and F. J. S. Lara. Formation of salivary secretion in Rhynchosciara Americana. Journal of Cell Biology, 58(2):470–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/2/470.

Derksen:1973:PRL

[DBW73] J. Derksen, H. D. Berendes, and E. Willart. Production and release of a locus-specific ribonucleoprotein product in polytene nuclei of Drosophila Hydei. Journal of Cell Biology, 59(3):661–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/3/661.

DiBona:1970:ESM

[DC70] Donald R. DiBona and Mortimer M. Civan. The effect of smooth muscle on the intercellular spaces in toad urinary bladder. Journal of Cell Biology, 46(2):235–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
2/235.

Djaczenko:1972:VNU

[DC72] Wiktor Djaczenko and Antonio Cassone. Visualization of new ultrastructural components in the cell wall of Candida albicans with fixatives containing tapo. Journal of Cell Biology, 52(1):186–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/1/186.

Djaczenko:1973:VPC

[DC73] Wiktor Djaczenko and Carmen Calenda. Visualization of polysaccharides in the cuticle of oligochaeta by the tris 1aziridinyl phosphine oxide method. Journal of Cell Biology, 57(3):859–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/3/859.

Dewel:1974:FSI

[DC74] William C. Dewel and Wallis H. Clark. A fine structural investigation of surface specializations and the cortical reaction in eggs of the cnidarian bunodosoma cavernata. Journal of Cell Biology, 60(1):78–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/60/1/78.

Descarries:1970:IDE

[DD70] Laurent Descarries and Bernard Droz. Intraneural distribution of exogenous norepinephrine in the central nervous system of the rat. Journal of Cell Biology, 44(2):385–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
44/2/385.

deDuve:1971:TFP

[dD71] Christian de Duve. Tissue fraction-past and present. Journal of Cell Biology, 50(1):??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/1/20d.

Dendy:1973:LKAa

[DDP73] Leslie A. Dendy, Russell L. Deter, and Charles W. Philpott. Localization of Na^{+}, K^{+}-ATPase and other enzymes in teleost pseudobranch. Journal of Cell Biology, 57(3):675–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/3/675.

David:1971:FFP

[DE71] Charles N. David and Kenneth Easterbrook. Ferritin in the fungus phycomyces. Journal of Cell Biology, 48(1): 15–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/15.

Decker:1974:LPD

[Dec74] Robert S. Decker. Lysosomal packaging in differentiating and degenerating anuran lateral motor column neurons. Journal of Cell Biology, 61(3):599–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/3/599.

Deter:1971:QCD

[Det71] Russell L. Deter. Quantitative characterization of dense body, autophagic vacuole, and acid phosphatase-bearing particle populations during the early phases of glucagon-induced autophagy in rat liver. Journal of Cell Biology, 48(3): 473–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/3/473.

Decker:1974:AGJ

[DF74] Robert S. Decker and Daniel S. Friend. Assembly of gap junctions during amphibian neurulation. Journal of Cell Biology, 62(1):32–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/1/32.

David-Ferreira:1973:GJR

[DFDF73] J. F. David-Ferreira and K. L. David-Ferreira. Gap junctionribosome association after autolysis. Journal of Cell Biology, 58(1):226–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/1/226.

Dallman:1971:EID

[DG71] Peter R. Dallman and Joseph R. Goodman. The effects of iron deficiency on the hepatocyte: a biochemical and ultrastructural study. Journal of Cell Biology, 48(1):79–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
48/1/79.

Davis:1974:VIC

[DGM^{+}74] Walter L. Davis, David B. P. Goodman, James H. Martin, James L. Matthews, and Howard Rasmussen. Vasopressininduced changes in the toad urinary bladder epithelial surface. Journal of Cell Biology, 61(2):544–??, May 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/61/2/
544.

Davis:1974:ECB

[DGS^{+}74] Walter L. Davis, David B. P. Goodman, Richard J. Schuster, Howard Rasmussen, and James H. Martin. Effects of cytochalasin B on the response of toad urinary bladder to Vasopressin. Journal of Cell Biology, 63(3):986–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
3/986.

Dunham:1974:PAA

[DGW74] Philip B. Dunham, Ira M. Goldstein, and Gerald Weissmann. Potassium and amino acid transport in human leukocytes exposed to phagocytic stimuli. Journal of Cell Biology, 63(1): 215–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/1/215.

Daniels:1974:UNG

[DH74] Mathew P. Daniels and Bernd Hamprecht. The ultrastructure of neuroblastoma glioma somatic cell hybrids. Journal of Cell Biology, 63(2):691–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/691.

Dales:1973:FSI

[DHN73] Samuel Dales, Konrad C. Hsu, and Ariaki Nagayama. The fine structure and immunological labeling of the achromatic mitotic apparatus after disruption of cell membranes. Journal of Cell Biology, 59(3):643–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/3/643.

Darrah:1971:RCL

[DHWHW71] Hilary K. Darrah, John Hedley-Whyte, and E. Tessa Hedley-Whyte. Radioautography of cholesterol in lung. Journal of Cell Biology, 49(2):345–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/2/345.

Dirksen:1971:CMD

[Dir71] Ellen Roter Dirksen. Centriole morphogenesis in developing ciliated epithelium of the mouse oviduct. Journal of Cell Biology, 51(1):286–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/1/286.

Dirksen:1974:CMO

[Dir74] Ellen Roter Dirksen. Ciliogenesis in the mouse oviduct. Journal of Cell Biology, 62(3):899–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/3/899.

Dean:1974:MAD

[DJ74] Phillip N. Dean and James H. Jett. Mathematical analysis of DNA distributions derived from flow microfluorometry. Journal of Cell Biology, 60(2):523–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/2/523.

DePierre:1973:PMM

[DK73] Joseph W. DePierre and Manfred L. Karnovsky. Plasma membranes of mammalian cells. Journal of Cell Biology, 56(2):275–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/275.

DeGrandi:1971:ULC

[DKC71] P. B. De Grandi, J. P. Kraehenbuhl, and M. A. Campiche. Ultrastructural localization of calcitonin in the parafollicular cells of pig thyroid gland with cytochrome c-labeled antibody fragments. Journal of Cell Biology, 50(2):446–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
2/446.

Daniel:1974:IEP

[DKHL74] Jon C. Daniel, Robert A. Kosher, James E. Hamos, and James W. Lash. Influence of external potassium on the synthesis and deposition of matrix components by chondrocytes in vitro. Journal of Cell Biology, 63(3):843–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
3/843.

deLaat:1974:NMF

[dLB74] Siegfried W. de Laat and John G. Bluemink. New membrane formation during cytokinesis in normal and cytochalasin Btreated eggs of Xenopus Laevis. Journal of Cell Biology, 60 (3):529–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/3/529.

Dewey:1973:SLS

[DLC73] Maynard M. Dewey, Rhea J. C. Levine, and David E. Colflesh. Structure of limulus striated muscle. Journal of Cell Biology, 58(3):574–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/3/574.

Durphy:1974:DSD

[DMF74] Michael Durphy, Paul N. Manley, and Errol C. Friedberg. A demonstration of several deoxyribonuclease activities in mammalian cell mitochondria. Journal of Cell Biology, 62(3): 695–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/695.

deMeis:1974:CCO

[dMHM74] L. de Meis, W. Hasselbach, and R. D. Machado. Characterization of calcium oxalate and calcium phosphate deposits in sarcoplasmic reticulum vesicles. Journal of Cell Biology, 62 (2):505–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/505.

DeHeer:1974:CRL

[DOP74] David H. DeHeer, Merle S. Olson, and R. Neal Pinckard. Characterization of rat liver subcellular membranes. Journal of Cell Biology, 60(2):460–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/2/460.

Dendy:1973:LKAb

[DPD73] Leslie A. Dendy, Charles W. Philpott, and Russell L. Deter. Localization of Na^{+}, K^{+}-ATPase and other enzymes in teleost pseudobranch. Journal of Cell Biology, 57(3):689–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/3/689.

Douglas:1970:EDD

[DRE70] William H. J. Douglas, Robert C. Ripley, and Richard A. Ellis. Enzymatic digestion of desmosome and hemidesmosome plaques performed on ultrathin sections. Journal of Cell Biology, 44(1):211–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/1/211.

Devine:1971:TFV

[DS71] Carrick E. Devine and Andrew P. Somlyo. Thick filaments in vascular smooth muscle. Journal of Cell Biology, 49(3): 636–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/636.

daSilva:1972:TMM

[dS72] Pedro Pinto da Silva. Translational mobility of the membrane intercalated particles of human erythrocyte ghosts. Journal of Cell Biology, 53(3):777–??, June 1972. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/3/777.

daSilva:1970:MSF

[dSB70] Pedro Pinto da Silva and Daniel Branton. Membrane splitting in freeze-etching. Journal of Cell Biology, 45(3):598–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
45/3/598.

DePetrocellis:1970:CCC

[DSP70] Benita De Petrocellis, Philip Siekevitz, and George E. Palade. Changes in chemical composition of thylakoid membranes during greening of the y-1 mutant of Chlamydomonas reinhardi. Journal of Cell Biology, 44(3):618–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/3/
618.

Devine:1972:SRE

[DSS72] Carrick E. Devine, Avril V. Somlyo, and Andrew P. Somlyo. Sarcoplasmic reticulum and excitation-contraction coupling in mammalian smooth muscles. Journal of Cell Biology, 52 (3):690–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/3/690.

Dusenbery:1972:ODW

[DU72] David B. Dusenbery and Robert B. Uretz. The orientation of DNA within 80-angstrom chromatin fibers. Journal of Cell Biology, 52(3):639–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/639.

Dyckman:1970:MAB

[DW70] Jacob Dyckman and Joel K. Weltman. A morphological analysis of binding of a hydrophobic probe to cells. Journal of Cell Biology, 45(1):192–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/1/192.

deFWebster:1971:GPM

[dW71] Henry deF. Webster. The geometry of peripheral myelin sheaths during their formation and growth in rat sciatic nerves. Journal of Cell Biology, 48(2):348–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
2/348.

Dumont:1972:EVI

[DW72] James N. Dumont and R. A. Wallace. The effects of vinblastine on isolated Xenopus oocytes. Journal of Cell Biology, 53(2):605–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/605.

Dehn:1973:SIV

[DW73] Paula F. Dehn and Robin A. Wallace. Sequestered and injected vitellogenin. Journal of Cell Biology, 58(3):721– ??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/3/721.

Eisenstein:1970:EMS

[EAK70] Reuben Eisenstein, Charalampos Arsenis, and Klaus E. Kuettner. Electron microscopic studies of cartilage matrix using lysozyme as a vital stain. Journal of Cell Biology, 46(3): 626–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/626.

Elsdale:1972:CSS

[EB72] Tom Elsdale and Jonathan Bard. Collagen substrata for studies on cell behavior. Journal of Cell Biology, 54(3):626– ??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/3/626.

Elgsaeter:1974:IPA

[EB74a] Arnljot Elgsaeter and Daniel Branton. Intramembrane particle aggregation in erythrocyte ghosts. Journal of Cell Biology, 63(3):1018–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/3/1018.

Elsdale:1974:CIM

[EB74b] Tom Elsdale and Jonathan Bard. Cellular interactions in morphogenesis of epithelial Mesenchymal systems. Journal of Cell Biology, 63(1):343–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/1/343.

Etzler:1974:DLC

[EB74c] Marilynn E. Etzler and Margaret L. Branstrator. Differential localization of cell surface and secretory components in rat intestinal epithelium by use of lectins. Journal of Cell Biology, 62(2):329–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/2/329.

Ehrenreich:1973:GFP

[EBSP73] J. H. Ehrenreich, J. J. M. Bergeron, P. Siekevitz, and G. E. Palade. Golgi fractions prepared from rat liver homogenates. Journal of Cell Biology, 59(1):45–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/1/45.

Everet:1970:NPP

[ED70] Ray F. Everet and Bharati P. Deshpande. Nuclear P protein in sieve elements of Tilia Americana. Journal of Cell Biology, 44(2):462–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/2/462.

Egyhazi:1971:IDL

[EDE^{+}71] E. Egyházi, B. Daneholt, J.-E. Edström, B. Lambert, and U. Ringborg. Intracellular distribution of low molecular weight RNA in chironomus tentans. Journal of Cell Biology, 48(1):120–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/1/120.

Egyhazi:1972:EAV

[EDE72] E. Egyházi, B. D’Monte, and J.-E. Edström. Effects of αamanitin on in vitro labeling of RNA from defined nuclear components in salivary gland cells from Chironomus tentans. Journal of Cell Biology, 53(2):523–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/523.

Edstrom:1974:ECM

[Eds74] Anders Edström. Effects of Ca^{2+} and Mg^{2+} on rapid axonal transport of proteins in vitro in frog sciatic nerves. Journal of Cell Biology, 61(3):812–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/3/812.

Edidin:1973:FSC

[EF73] Michael Edidin and Douglas Fambrough. Fluidity of the surface of cultured muscle fibers. Journal of Cell Biology, 57 (1):27–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/27.

Edwards:1971:PTB

[EG71] Mercedes R. Edwards and Elisabeth Gantt. Phycobilisomes of the thermophilic blue-green alga synechococcus lividus. Journal of Cell Biology, 50(3):896–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/896.

Esterhuizen:1970:UCC

[EH70] A. C. Esterhuizen and S. L. Howell. Ultrastructure of the a-cells of cat islets of langerhans following sympathetic stimulation of glucagon secretion. Journal of Cell Biology, 46(3): 593–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/593.

Eddy:1971:FSR

[EI71] E. M. Eddy and Susumu Ito. Fine structural and radioautographic observations on dense perinuclear cytoplasmic material in tadpole oocytes. Journal of Cell Biology, 49(1):90–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/1/90.

Eisenberg:1974:SAM

[EKP74] Brenda R. Eisenberg, Aileen M. Kuda, and James B. Peter. Stereological analysis of mammalian skeletal muscle. Journal of Cell Biology, 60(3):732–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/3/732.

Eliceiri:1972:RRH

[Eli72] George L. Eliceiri. The ribosomal RNA of hamster-mouse hybrid cells. Journal of Cell Biology, 53(1):177–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
1/177.

Evans:1974:CEI

[EMJT74] R. Blair Evans, Vera Morhenn, Albert L. Jones, and Gordon M. Tomkins. Concomitant effects of insulin on surface membrane conformation and polysome profiles of serumstarved Balb/c 3T3 fibroblasts. Journal of Cell Biology, 61 (1):95–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/1/95.

Eaton:1972:MBP

[EP72] Barbra L. Eaton and Frank A. Pepe. M band protein. Journal of Cell Biology, 55(3):681–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/681.

Everhart:1974:CCCa

[ER74a] Leighton P. Everhart and Robert W. Rubin. Cyclic changes in the cell surface. Journal of Cell Biology, 60(2):434– ??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/434.

Everhart:1974:CCCb

[ER74b] Leighton P. Everhart and Robert W. Rubin. Cyclic changes in the cell surface. Journal of Cell Biology, 60(2):442– ??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/442.

Ehrlich:1974:EAA

[ERB74] H. Paul Ehrlich, Russell Ross, and Paul Bornstein. Effects of antimicrotubular agents on the secretion of collagen. Journal of Cell Biology, 62(2):390–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/2/390.

Erickson:1974:MSL

[Eri74] Harold P. Erickson. Microtubule surface lattice and subunit structure and observations on reassembly. Journal of Cell Biology, 60(1):153–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/1/153.

Espey:1971:MSP

[Esp71] Lawrence L. Espey. Multivesicular structures in proliferating fibroblasts of rabbit ovarian follicles during ovulation. Journal of Cell Biology, 48(2):437–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/2/437.

Ettienne:1972:SLC

[Ett72] Earl Ettienne. Subcellular localization of calcium repositories in plasmodia of the acellular slime mold Physarum Polycephalum. Journal of Cell Biology, 54(1):179–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
1/179.

Evans:1971:SEC

[EWBP71] E. M. Evans, J. M. Wrigglesworth, K. Burdett, and W. F. R. Pover. Studies on epithelial cells isolated from guinea pig small intestine. Journal of Cell Biology, 51(2):452– ??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/452.

Franzini-Armstrong:1970:ST

[FA70] Clara Franzini-Armstrong. Studies of the triad. Journal of Cell Biology, 47(2):488–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/2/488.

Franzini-Armstrong:1971:ST

[FA71] Clara Franzini-Armstrong. Studies of the triad. Journal of Cell Biology, 49(1):196–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/1/196.

Franzini-Armstrong:1973:SSZ

[FA73a] Clara Franzini-Armstrong. The structure of a simple Z line. Journal of Cell Biology, 58(3):630–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
3/630.

Franzini-Armstrong:1973:ST

[FA73b] Clara Franzini-Armstrong. Studies of the triad. Journal of Cell Biology, 56(1):120–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/1/120.

Franzini-Armstrong:1974:FFS

[FA74] Clara Franzini-Armstrong. Freeze fracture of skeletal muscle from the tarantula spider. Journal of Cell Biology, 61 (2):501–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/2/501.

Fahimi:1970:FSL

[Fah70a] H. Dariush Fahimi. The fine structural localization of endogenous and exogenous peroxidase activity in kupffer cells of rat liver. Journal of Cell Biology, 47(1):247–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/247.

Fahrenbach:1970:CHF

[Fah70b] Wolf H. Fahrenbach. The cyanoblast: Hemocyanin formation in limulus polyphemus. Journal of Cell Biology, 44(2): 445–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/2/445.

Fahimi:1974:EBS

[Fah74] H. Dariush Fahimi. Effect of buffer storage on fine structure and catalase cytochemistry of peroxisomes. Journal of Cell Biology, 63(2):675–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/2/675.

Farbman:1972:DTB

[Far72] Albert I. Farbman. Differentiation of taste buds in organ culture. Journal of Cell Biology, 52(2):489–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
2/489.

Friend:1970:OSE

[FB70] Daniel S. Friend and Gail E. Brassil. Osmium staining of endoplasmic reticulum and mitochondria in the rat adrenal cortex. Journal of Cell Biology, 46(2):252–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/46/2/
252.

Farquhar:1972:CLA

[FBBdD72] Marilyn G. Farquhar, Dorothy F. Bainton, Marco Baggiolini, and Christian de Duve. Cytochemical localization of acid phosphatase activity in granule fractions from rabbit polymorphonuclear leukocytes. Journal of Cell Biology, 54 (1):141–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/1/141.

Farquhar:1974:CGF

[FBP74] Marilyn G. Farquhar, J. J. M. Bergeron, and George E. Palade. Cytochemistry of Golgi fractions prepared from rat liver. Journal of Cell Biology, 60(1):8–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/60/1/8.

Fernandez:1971:ATC

[FBS71] H. L. Fernandez, P. R. Burton, and F. E. Samson. Axoplasmic transport in the crayfish nerve cord. Journal of Cell Biology, 51(1):176–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/1/176.

Friedman:1972:EPS

[FC72] Harold I. Friedman and Robert R. Cardell. Effects of puromycin on the structure of rat intestinal epithelial cells during fat absorption. Journal of Cell Biology, 52(1):15– ??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/1/15.

Fay:1973:RDM

[FC73a] Fredric S. Fay and P. H. Cooke. Reversible disaggregation of myofilaments in vertebrate smooth muscle. Journal of Cell Biology, 56(2):399–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/2/399.

Flaxman:1973:LRJ

[FC73b] B. Allen Flaxman and Francis V. Cavoto. Low-resistance junctions in epithelial outgrowths from normal and cancerous epidermis in vitro. Journal of Cell Biology, 58(1):219–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
1/219.

Fedorko:1973:ADF

[FCH73] Martha E. Fedorko, Nicholas L. Cross, and James G. Hirsch. Appearance and distribution of ferritin in mouse peritoneal macrophages in vitro after uptake of heterologous erythrocytes. Journal of Cell Biology, 57(2):289–??, May 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/57/2/
289.

Fulton:1971:BBC

[FD71] Chandler Fulton and Allan D. Dingle. Basal bodies, but not centrioles, in naegleria. Journal of Cell Biology, 51(3): 826–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/826.

Feit:1971:MP

[FDBS71] Howard Feit, Gary R. Dutton, Samuel H. Barondes, and Michael L. Shelanski. Microtubule protein. Journal of Cell Biology, 51(1):138–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/1/138.

Franke:1970:NMM

[FDE^{+}70] Werner W. Franke, Barbara Deumling, Baerbel Ermen, Ernst-Dieter Jarasch, and Hans Kleinig. Nuclear membranes from mammalian liver. Journal of Cell Biology, 46 (2):379–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/2/379.

Feder:1971:M

[Fed71] Ned Feder. Microperoxidase. Journal of Cell Biology, 51(1): 339–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/339.

Fedorko:1974:LIM

[Fed74] Martha E. Fedorko. Loss of iron from mouse peritoneal macrophages in vitro after uptake of [^{55} fe]ferritin and [^{55} fe]ferritin rabbit antiferritin complexes. Journal of Cell Biology, 62(3):802–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/802.

Forte:1970:DES

[FF70a] Trudy M. Forte and John G. Forte. Definition of the extracellular space in secreting and nonsecreting oxyntic cells. Journal of Cell Biology, 47(3):782–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/3/782.

Forte:1970:HSC

[FF70b] Trudy M. Forte and John G. Forte. Histochemical staining and characterization of glycoproteins in acid-secreting cells of frog stomach. Journal of Cell Biology, 47(2):437–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
47/2/437.

Friend:1974:MDF

[FF74] Daniel S. Friend and Don W. Fawcett. Membrane differentiations in freeze-fractured mammalian sperm. Journal of Cell Biology, 63(2):641–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/2/641.

Finegold:1970:MDE

[FG70a] Milton J. Finegold and Laura E. Greene. Mitochondrial damage in experimental congenital adrenal hyperplasia. Journal of Cell Biology, 45(2):455–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/455.

Forssmann:1970:SSR

[FG70b] W. G. Forssmann and L. Girardier. A study of the T system in rat heart. Journal of Cell Biology, 44(1):1–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
1/1.

Friend:1972:DCC

[FG72a] Daniel S. Friend and Norton B. Gilula. A distinctive cell contact in the rat adrenal cortex. Journal of Cell Biology, 53 (1):148–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/148.

Friend:1972:VTG

[FG72b] Daniel S. Friend and Norton B. Gilula. Variations in tight and gap junctions in mammalian tissues. Journal of Cell Biology, 53(3):758–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/3/758.

Friedberg:1971:PBL

[FGO71] I. Friedberg, I. Goldberg, and I. Ohad. A prolamellar bodylike structure in chlamydomonas reinhardi. Journal of Cell Biology, 50(1):268–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/1/268.

Firket:1970:ACP

[FH70] H. Firket and A. F. Hopper. Abnormal cycle phase durations in polyploid cells or cells with irregular mitosis in a HeLa cell population. Journal of Cell Biology, 44(3):678–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/678.

Fulton:1971:SSF

[FKS71] Chandler Fulton, R. E. Kane, and R. E. Stephens. Serological similarity of flagellar and mitotic microtubules. Journal of Cell Biology, 50(3):762–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/3/762.

Franke:1971:MMC

[FKZ^{+}71] Werner W. Franke, Jürgen Kartenbeck, Hanswalter Zentgraf, Ulrich Scheer, and Heinz Falk. Membrane-to-membrane cross-bridges. Journal of Cell Biology, 51(3):881–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
51/3/881.

Freed:1970:ACS

[FL70] Jerome J. Freed and Marcia M. Lebowitz. The association of a class of saltatory movements with microtubules in cultured cells. Journal of Cell Biology, 45(2):334–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/334.

Frank:1974:MIS

[FL74] J. S. Frank and G. A. Langer. The myocardial interstitium: Its structure and its role in ionic exchange. Journal of Cell Biology, 60(3):586–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/3/586.

Flickinger:1971:DFG

[Fli71] Charles J. Flickinger. Decreased formation of Golgi bodies in amebae in the presence of RNA and protein synthesis inhibitors. Journal of Cell Biology, 49(1):221–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
1/221.

Finger:1973:SPS

[FLM73] Irving Finger, Patricia Lavanchy, and Ann Meany. Synthesis of paramecium surface proteins. Journal of Cell Biology, 56(2):434–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/434.

Froehlich:1972:EAC

[FR72] Jeffrey E. Froehlich and Martin Rachmeler. Effect of adenosine 3’-5’-cyclic monophosphate on cell proliferation. Journal of Cell Biology, 55(1):19–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/1/19.

Frederiksen:1974:ADL

[FR74a] Ole Frederiksen and Jørgen Rostgaard. Absence of dilated lateral intercellular spaces in fluid-transporting frog gallbladder epithelium. Journal of Cell Biology, 61(3):830–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
3/830.

Friend:1974:ADS

[FR74b] Daniel S. Friend and Irene Rudolf. Acrosomal disruption in sperm. Journal of Cell Biology, 63(2):466–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/466.

Froehlich:1974:ICG

[FR74c] Jeffrey E. Froehlich and Martin Rachmeler. Inhibition of cell growth in the G_{1} phase by adenosine 3’,5’-cyclic monophosphate. Journal of Cell Biology, 60(1):249–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
60/1/249.

Friedman:1971:ABM

[Fri71] Morton H. Friedman. Arm-bearing microtubules associated with an unusual desmosome-like junction. Journal of Cell Biology, 49(3):916–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/916.

Faeder:1970:GUS

[FS70] Isabelle R. Faeder and Miriam M. Salpeter. Glutamate uptake by a stimulated insect nerve muscle preparation. Journal of Cell Biology, 46(2):300–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/2/300.

Forbes:1973:LSF

[FS73] M. S. Forbes and N. Sperelakis. A labyrinthine structure formed from a transverse tubule of mouse ventricular myocardium. Journal of Cell Biology, 56(3):865–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
3/865.

Forbes:1974:SBJ

[FS74] M. S. Forbes and N. Sperelakis. Spheroidal bodies in the junctional sarcoplasmic reticulum of lizard myocardial cells. Journal of Cell Biology, 60(3):602–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/3/602.

Franke:1972:ICA

[FSF72] Werner W. Franke, Ulrich Scheer, and Hansjörg Fritsch. Intranuclear and cytoplasmic annulate lamellae in plant cells. Journal of Cell Biology, 53(3):823–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/3/823.

Fonte:1971:RCC

[FSH71] Virginia G. Fonte, Robert L. Searls, and S. Robert Hilfer. The relationship of cilia with cell division and differentiation. Journal of Cell Biology, 49(1):226–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/1/226.

Feagler:1974:ETP

[FTCM74] John R. Feagler, Thomas W. Tillack, David D. Chaplin, and Philip W. Majerus. The effects of thrombin on phytohemagglutinin receptor sites in human platelets. Journal of Cell Biology, 60(3):541–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/3/541.

Fuseler:1973:RPM

[Fus73] John W. Fuseler. Repetitive procurement of mature gametes from individual sea stars and sea urchins. Journal of Cell Biology, 57(3):879–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/3/879.

Feeney:1972:BRA

[FW72] Lynette Feeney and Steven L. Wissig. A biochemical and radioautographic analysis of protein secretion by thyroid lobes incubated in vitro. Journal of Cell Biology, 53(2):510–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/2/510.

Franson:1973:LPN

[FW73] Richard C. Franson and Moseley Waite. Lysosomal phospholipases A_{1} and A_{2} of normal and Bacillus Calmette Guerininduced alveolar macrophages. Journal of Cell Biology, 56 (3):621–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/621.

Gonatas:1973:DPM

[GA73] Nicholas K. Gonatas and Stratis Avrameas. Detection of plasma membrane carbohydrates with lectin peroxidase conjugates. Journal of Cell Biology, 59(2):436–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
2/436.

Gonatas:1971:MBC

[GAGGS71] Nicholas K. Gonatas, Lucila Autilio-Gambetti, Pierluigi Gambetti, and Brenda Shafer. Morphological and biochemical changes in rat synaptosome fractions during neonatal development. Journal of Cell Biology, 51(2):484–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
2/484.

Gambetti:1972:PSS

[GAGGS72] P. Gambetti, L. A. Autilio-Gambetti, N. K. Gonatas, and B. Shafer. Protein synthesis in synaptosomal fractions. Journal of Cell Biology, 52(3):526–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/3/526.

Gambetti:1973:QAS

[GAGSP73] Pierluigi Gambetti, Lucila Autilio-Gambetti, Brenda Shafer, and LaDorna Pfaff. Quantitative autoradiographic study of labeled RNA in rabbit optic nerve after intraocular injection of [^{3}H]uridine. Journal of Cell Biology, 59(3):677–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
59/3/677.

Galper:1974:MPS

[Gal74] Jonas B. Galper. Mitochondrial protein synthesis in HeLa cells. Journal of Cell Biology, 60(3):755–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/60/3/
755.

Gross:1970:OTC

[GAP70] Stephen R. Gross, Lewis Aronow, and William B. Pratt. The outward transport of cortisol by mammalian cells in vitro. Journal of Cell Biology, 44(1):103–??, January 1970. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/1/
103.

Gerhardt:1970:DSG

[GB70] B. P. Gerhardt and Harry Beevers. Developmental studies on glyoxysomes in ricinus endosperm. Journal of Cell Biology, 44 (1):94–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/1/94.

Gear:1972:DCS

[GB72] Adrian R. L. Gear and Jana M. Bednarek. Direct counting and sizing of mitochondria in solution. Journal of Cell Biology, 54(2):325–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/2/325.

Gruber:1973:DMP

[GBN73] Peter J. Gruber, Wayne M. Becker, and Eldon H. Newcomb. The development of microbodies and peroxisomal enzymes in greening bean leaves. Journal of Cell Biology, 56(2): 500–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/500.

Gilder:1974:BCS

[GC74] Jamison Gilder and James Cronshaw. A biochemical and cytochemical study of adenosine triphosphatase activity in the phloem of nicotian a tabacum. Journal of Cell Biology, 60(1): 221–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/1/221.

Grantham:1971:DOT

[GCF71] Jared Grantham, F. E. Cuppage, and Darrell Fanestil. Direct observation of toad bladder response to Vasopressin. Journal of Cell Biology, 48(3):695–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/3/695.

Gimbrone:1974:HVE

[GCF74] Michael A. Gimbrone, Ramzi S. Cotran, and Judah Folkman. Human vascular endothelial cells in culture. Journal of Cell Biology, 60(3):673–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/3/673.

Glaumann:1970:SSM

[GD70] Hans Glaumann and Gustav Dallner. Subfractionation of smooth microsomes from rat liver. Journal of Cell Biology, 47 (1):34–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/1/34.

Graziadei:1973:NRF

[GD73] P. P. C. Graziadei and R. S. DeHan. Neuronal regeneration in frog olfactory system. Journal of Cell Biology, 59(2): 525–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/2/525.

Gabe:1973:MSR

[GdB73] P. R. Gabe and L. E. de Bault. Macromolecular syntheses related to the reproductive cyst of Tetrahymena patula. Journal of Cell Biology, 59(3):615–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/3/615.

Gerber:1973:ESI

[GDH73] Dorothy Gerber, Margaret Davies, and Lowell E. Hokin. The effects of secretagogues on the incorporation of [2-3H]myoinositol into lipid in cytological fractions in the pancreas of the guinea pig in vivo. Journal of Cell Biology, 56 (3):736–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/736.

Glaumann:1970:EPG

[GE70] Hans Glaumann and Jan L. E. Ericsson. Evidence for the participation of the Golgi apparatus in the intracellular transport of nascent albumin in the liver cell. Journal of Cell Biology, 47(3):555–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/47/3/555.

Gantt:1971:CSC

[GEP71] E. Gantt, M. R. Edwards, and L. Provasoli. Chloroplast structure of the cryptophyceae. Journal of Cell Biology, 48 (2):280–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/2/280.

Gibbons:1972:SPB

[GF72] I. R. Gibbons and E. Fronk. Some properties of bound and soluble dynein from sea urchin sperm flagella. Journal of Cell Biology, 54(2):365–??, August 1972. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/2/365.

Greengard:1972:CDR

[GFK72] Olga Greengard, Micheline Federman, and W. Eugene Knox. Cytomorphometry of developing rat liver and its application to enzymic differentiation. Journal of Cell Biology, 52(2): 261–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/261.

Ghosh:1970:DSPb

[GG70] Sibdas Ghosh and Ira Ghosh. DNA synthetic pattern in the nucleolus. Journal of Cell Biology, 45(1):181–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
1/181.

Granick:1971:NNC

[GG71] S. Granick and Donna Granick. Nucleolar necklaces in chick embryo myoblasts formed by lack of arginine. Journal of Cell Biology, 51(3):636–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/3/636.

Gibbons:1972:FMA

[GG72] Barbara H. Gibbons and I. R. Gibbons. Flagellar movement and adenosine triphosphatase activity in sea urchin sperm extracted with Triton X-100. Journal of Cell Biology, 54 (1):75–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/1/75.

Gibbons:1974:PFR

[GG74a] Barbara H. Gibbons and I. R. Gibbons. Properties of flagellar “rigor waves” formed by abrupt removal of adenosine triphosphate from actively swimming sea urchin sperm. Journal of Cell Biology, 63(3):970–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/3/970.

Goodenough:1974:SHG

[GG74b] Daniel A. Goodenough and Norton B. Gilula. The splitting of hepatocyte gap junctions and zonulae occludentes with hypertonic disaccharides. Journal of Cell Biology, 61(3):575–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
61/3/575.

Gershon:1974:EMA

[GHN74] Michael D. Gershon, Martin Hagopian, and Eladio A. Nunez. An electron microscope autoradiographic study of the neuronal and extraneuronal localization of labeled amine in the heart of the bat after administration of tritiated norepinephrine. Journal of Cell Biology, 62(3):610–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
3/610.

Ghosh:1971:GPS

[Gho71] B. K. Ghosh. Grooves in the plasmalemma of Saccharomyces Cerevisiae seen in glancing sections of double aldehydefixed cells. Journal of Cell Biology, 48(1):192–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
1/192.

Gorovsky:1973:NAN

[GHP73] Martin A. Gorovsky, Stanley Hattman, and Gloria Lorick Pleger. [^{6} n]methyl adenine in the nuclear DNA of a eucaryote, Tetrahymena Pyriformis. Journal of Cell Biology, 56 (3):697–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/697.

Gibbs:1970:NOR

[Gib70] Sarah P. Gibbs. Nuclear and organelle RNA synthesis in ochromonas: the effects of light. Journal of Cell Biology, 46(3):599–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/3/599.

Gillespie:1971:CBT

[Gil71] Elizabeth Gillespie. Colchicine binding in tissue slices. Journal of Cell Biology, 50(2):544–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/544.

Gilbert:1972:PTD

[GJ72] Brian E. Gilbert and Terry C. Johnson. Protein turnover during maturation of mouse brain tissue. Journal of Cell Biology, 53(1):143–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/1/143.

Goetz:1974:VBM

[GJDVC74] George H. Goetz, Pamela L. Johnston, Kathy Dobrosielski-Vergona, and Edgar F. Carell. Vitamin B_{12} and the macromolecular composition of euglena. Journal of Cell Biology, 62(3):672–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/672.

Grove:1973:QQM

[GJKL73] William E. Grove, Terry C. Johnson, Paul Kelly, and Marvin Luttges. Qualitative and quantitative morphological changes in subcellular fractions from mouse cerebral cortex during postnatal development. Journal of Cell Biology, 58(3):676– ??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/3/676.

Ganion:1972:IST

[GK72] L. R. Ganion and R. G. Kessel. Intracellular synthesis, transport, and packaging of proteinaceous yolk in oocytes of orconectes immunis. Journal of Cell Biology, 52(2):420– ??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/420.

Githens:1973:BCD

[GK73a] S. Githens and M. L. Karnovsky. Biochemical changes during growth and encystment of the cellular slime mold polysphondylium pallidum. Journal of Cell Biology, 58(3):

522–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/3/522.

Githens:1973:PCS

[GK73b] S. Githens and M. L. Karnovsky. Phagocytosis by the cellular slime mold polysphondylium pallidum during growth and development. Journal of Cell Biology, 58(3):536–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
3/536.

Gersten:1974:LPB

[GKB74] Douglas M. Gersten, Thomas W. Kimmerer, and H. Bruce Bosmann. The lysosome periphery: Biochemical and electrokinetic properties of the tritosome surface. Journal of Cell Biology, 60(3):764–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/3/764.

Gorovsky:1974:HFT

[GKP74] Martin A. Gorovsky, Josephine Bowen Keevert, and Gloria Lorick Pleger. Histone F1 of Tetrahymena macronuclei. Journal of Cell Biology, 61(1):134–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/1/134.

Goodenough:1970:CSF

[GL70] Ursula W. Goodenough and R. P. Levine. Chloroplast structure and function in ac-20 , a mutant strain of Chlamydomonas reinhardi. Journal of Cell Biology, 44(3):547–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/44/3/547.

Goodenough:1971:EIRb

[GL71] Ursula W. Goodenough and R. P. Levine. The effects of inhibitors of RNA and protein synthesis on the recovery of chloroplast ribosomes, membrane organization, and photosynthetic electron transport in the ac-20 strain of Chlamydomonas reinhardi. Journal of Cell Biology, 50(1):50–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
50/1/50.

Gantt:1972:PPC

[GL72] E. Gantt and C. A. Lipschultz. Phycobilisomes of porphyridium cruentum. Journal of Cell Biology, 54(2):313–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
2/313.

Gambarini:1974:URR

[GL74] A. G. Gambarini and F. J. S. Lara. Under-replication of ribosomal cistrons in polytene chromosomes of rhynchosciara. Journal of Cell Biology, 62(1):215–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/1/215.

Glauert:1974:HVE

[Gla74] Audrey M. Glauert. The high voltage electron microscope in biology. Journal of Cell Biology, 63(3):717–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
3/717.

Ghosh:1970:DSPa

[GLG70] Sibdas Ghosh, Renate Lettré, and Ira Ghosh. DNA synthetic pattern in the nucleolus. Journal of Cell Biology, 45 (1):177–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/1/177.

Gambarini:1972:RRG

[GM72a] A. G. Gambarini and R. Meneghini. Ribosomal RNA genes in salivary gland and ovary of rhynchosciara angelae. Journal of Cell Biology, 54(2):421–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/2/421.

Goldschneider:1972:TSC

[GM72b] Irving Goldschneider and A. A. Moscona. Tissue-specific cellsurface antigens in embryonic cells. Journal of Cell Biology, 53(2):435–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/435.

Gimenez-Martin:1974:EAN

[GMdlTFGGF74] G. Giménez-Martín, C. de la Torre, M. E. Fernández-Gómez, and A. González-Fernández. Experimental analysis of nucleolar reorganization. Journal of Cell Biology, 60(2):502– ??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/502.

Grimes:1974:NGD

[GMP74] Gary W. Grimes, Henry R. Mahler, and Philip S. Perlman. Nuclear gene dosage effects on mitochondrial mass and DNA. Journal of Cell Biology, 61(3):565–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/3/565.

Grinnell:1973:SCA

[GMS73] Frederick Grinnell, Mary Milam, and Paul A. Srere. Studies on cell adhesion. Journal of Cell Biology, 56(3):659–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/56/3/659.

Gershon:1973:SSO

[GN73] Michael D. Gershon and Eladio A. Nunez. Subcellular storage organelles for 5-hydroxytryptamine in parafollicular cells of the thyroid gland. Journal of Cell Biology, 56(3): 676–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/676.

Goldfischer:1970:HUR

[GNAB70] Sidney Goldfischer, Alex B. Novikoff, Arline Albala, and Luis Biempica. Hemoglobin uptake by rat hepatocytes and its breakdown within lysosomes. Journal of Cell Biology, 44(3): 513–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/3/513.

Goldberg:1970:BCMa

[GO70a] I. Goldberg and I. Ohad. Biogenesis of chloroplast membranes. Journal of Cell Biology, 44(3):563–??, March 1970.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/3/
563.

Goldberg:1970:BCMb

[GO70b] I. Goldberg and I. Ohad. Biogenesis of chloroplast membranes. Journal of Cell Biology, 44(3):572–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/3/
572.

Gobel:1971:AAS

[Gob71] Stephen Gobel. Axo-axonic septate junctions in the basket formations of the cat cerebellar cortex. Journal of Cell Biology, 51(1):328–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/1/328.

Goldman:1971:RTC

[Gol71] Robert D. Goldman. The role of three cytoplasmic fibers in bhk-21 cell motility. Journal of Cell Biology, 51(3):752– ??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/752.

Goldman:1972:ECB

[Gol72] Robert D. Goldman. The effects of cytochalasin B on the microfilaments of baby hamster kidney (bhk-21) cells. Journal of Cell Biology, 52(2):246–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/2/246.

Goodenough:1971:EIRa

[Goo71] Ursula W. Goodenough. The effects of inhibitors of RNA and protein synthesis on chloroplast structure and function in wild-type chlamydomonas reinhardi. Journal of Cell Biology, 50(1):35–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/1/35.

Goodenough:1974:BIM

[Goo74] Daniel A. Goodenough. Bulk isolation of mouse hepatocyte gap junctions. Journal of Cell Biology, 61(2):557–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/557.

Gorovsky:1970:SNSa

[Gor70a] Martin A. Gorovsky. Studies on nuclear structure and function in Tetrahymena Pyriformis. Journal of Cell Biology, 47(3):619–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/3/619.

Gorovsky:1970:SNSb

[Gor70b] Martin A. Gorovsky. Studies on nuclear structure and function in Tetrahymena Pyriformis. Journal of Cell Biology, 47(3):631–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/3/631.

Greenberg:1971:HPD

[GP71] Jay R. Greenberg and Robert P. Perry. Hybridization properties of DNA sequences directing the synthesis of messenger RNA and heterogeneous nuclear RNA. Journal of Cell Biology, 50(3):774–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/3/774.

Gilai:1972:ECT

[GP72] Arieh Gilai and I. Parnas. Electromechanical coupling in tubular muscle fibers. Journal of Cell Biology, 52(3):626–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/52/3/626.

Geuze:1973:CMR

[GP73a] J. J. Geuze and C. Poort. Cell membrane resorption in the rat exocrine pancreas cell after in vivo stimulation of the secretion, as studied by in vitro incubation with extracellular space markers. Journal of Cell Biology, 57(1):159–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/159.

Gibbs:1973:AEM

[GP73b] Sarah P. Gibbs and Ronald J. Poole. Autoradiographic evidence for many segregating DNA molecules in the chloroplast of ochromonas danica. Journal of Cell Biology, 59(2): 318–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/2/318.

Gorovsky:1973:SHF

[GPKJ73] Martin A. Gorovsky, Gloria Lorick Pleger, Josephine Bowen Keevert, and Carol A. Johmann. Studies on histone fraction F2A1 in macro- and micronuclei of Tetrahymena Pyriformis. Journal of Cell Biology, 57(3):773–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/3/773.

Gray:1973:OYN

[GPR73] Robert H. Gray, Joan B. Peterson, and Hans Ris. The organization of yeast nucleohistone fibers. Journal of Cell Biology, 58(1):244–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/1/244.

Goodenough:1970:FSA

[GR70] Daniel A. Goodenough and Jean Paul Revel. A fine structural analysis of intercellular junctions in the mouse liver. Journal of Cell Biology, 45(2):272–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/272.

Goodenough:1971:PIS

[GR71] Daniel A. Goodenough and Jean-Paul Revel. The permeability of isolated and in situ mouse hepatic gap junctions studied with enzymatic tracers. Journal of Cell Biology, 50 (1):81–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/1/81.

Gil:1973:ICL

[GR73] Joan Gil and Oscar K. Reiss. Isolation and characterization of lamellar bodies and tubular myelin from rat lung homogenates. Journal of Cell Biology, 58(1):152–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
1/152.

Gallin:1974:RRD

[GR74] John I. Gallin and Alan S. Rosenthal. The regulatory role of divalent cations in human granulocyte chemotaxis. Journal of Cell Biology, 62(3):594–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/3/594.

Grillo:1970:CIR

[Gri70a] Mary A. Grillo. Cytoplasmic inclusions resembling nucleoli in sympathetic neurons of adult rats. Journal of Cell Biology, 45(1):100–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/1/100.

Grillo:1970:ESV

[Gri70b] Mary A. Grillo. Extracellular synaptic vesicles in the mouse heart. Journal of Cell Biology, 47(2):547–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
2/547.

Grim:1970:UCM

[Gri70c] J. Norman Grim. Ultrastructure of ciliary microtubules after critical point drying. Journal of Cell Biology, 45(2):466–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
45/2/466.

Grimes:1973:MDK

[Gri73a] Gary W. Grimes. Morphological discontinuity of kinetosomes during the life cycle of oxytricha fallax. Journal of Cell Biology, 57(1):229–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/1/229.

Grinnell:1973:CIS

[Gri73b] Frederick Grinnell. Concanavalin a increases the strength of baby hamster kidney cell attachment to substratum. Journal of Cell Biology, 58(3):602–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/3/602.

Gross:1971:CMT

[Gro71] Nicholas J. Gross. Control of mitochondrial turnover under the influence of thyroid hormone. Journal of Cell Biology, 48 (1):29–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/29.

Greenawalt:1970:AUM

[GS70] John W. Greenawalt and Carl Schnaitman. An appraisal of the use of monoamine oxidase as an enzyme marker for the outer membrane of rat liver mitochondria. Journal of Cell Biology, 46(1):173–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/1/173.

Gilula:1971:SGJ

[GS71a] Norton B. Gilula and Peter Satir. Septate and gap junctions in molluscan Gill epithelium. Journal of Cell Biology, 51(3): 869–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/869.

Goodenough:1971:SDS

[GS71b] Ursula W. Goodenough and L. Andrew Staehelin. Structural differentiation of stacked and unstacked chloroplast membranes. Journal of Cell Biology, 48(3):594–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
3/594.

Gilula:1972:CN

[GS72a] Norton B. Gilula and Peter Satir. The ciliary necklace. Journal of Cell Biology, 53(2):494–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/494.

Goodenough:1972:IMH

[GS72b] Daniel A. Goodenough and Walther Stoeckenius. The isolation of mouse hepatocyte gap junctions. Journal of Cell Biology, 54(3):646–??, September 1972. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/3/646.

Grimes:1973:DCA

[GS73] William J. Grimes and Judith L. Schroeder. Dibutyryl cyclic adenosine 3’ 5’ monophosphate, sugar transport, and regulatory control of cell division in normal and transformed cells. Journal of Cell Biology, 56(2):487–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
2/487.

Garber:1974:ICC

[GS74] Melvin P. Garber and Peter L. Steponkus. Identification of chloroplast coupling factor by freeze-etching and negativestaining techniques. Journal of Cell Biology, 63(1):24–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
63/1/24.

Gautschi:1971:SDC

[GSH71] J. R. Gautschi, R. Schindler, and C. Hürni. Studies on the division cycle of mammalian cells. Journal of Cell Biology, 51(3):653–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/3/653.

Gulyas:1972:RZ

[Gul72] Bela J. Gulyas. The rabbit zygote. Journal of Cell Biology, 55(3):533–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/55/3/533.

Gillis:1971:AMT

[GW71] Jean-Marie Gillis and Maurice Wibo. Accurate measurement of the thickness of ultrathin sections by interference microscopy. Journal of Cell Biology, 49(3):947–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/947.

Gunther:1974:KCC

[GWE74] Gary R. Gunther, John L. Wang, and Gerald M. Edelman. The kinetics of cellular commitment during stimulation of lymphocytes by lectins. Journal of Cell Biology, 62 (2):366–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/366.

Giacomelli:1970:CSA

[GWS70] Filiberto Giacomelli, Joseph Wiener, and David Spiro. Crossstriated arrays of filaments in endothelium. Journal of Cell Biology, 45(1):188–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/1/188.

Gurley:1974:CCS

[GWT74] Lawrence R. Gurley, Ronald A. Walters, and Robert A. Tobey. Cell cycle-specific changes in histone phosphorylation associated with cell proliferation and chromosome condensation. Journal of Cell Biology, 60(2):356–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
2/356.

Huebner:1970:EVS

[HA70] Erwin Huebner and Everett Anderson. The effects of vinblastine sulfate on the microtubular organization of the ovary of rhodnius prolixus. Journal of Cell Biology, 46(1):191–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
1/191.

Hackenbrock:1972:ELU

[Hac72] Charles R. Hackenbrock. Energy-linked ultrastructural transformations in isolated liver mitochondria and mitoplasts. Journal of Cell Biology, 53(2):450–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/450.

Hagopian:1970:CBS

[Hag70] Martin Hagopian. Contraction bands at short sarcomere length in chick muscle. Journal of Cell Biology, 47(3):790– ??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/790.

Hand:1970:FSE

[Han70a] Arthur R. Hand. The fine structure of Von Ebner’s gland of the rat. Journal of Cell Biology, 44(2):340–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
2/340.

Hand:1970:NAC

[Han70b] Arthur R. Hand. Nerve-acinar cell relationships in the rat parotid gland. Journal of Cell Biology, 47(2):540–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
47/2/540.

Harris:1970:AAM

[Har70a] W. F. Harris. The arrangement of the axonemal microtubules and links of echinosphaerium nucleofilum. Journal of Cell Biology, 46(1):183–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/1/183.

Harrison:1970:UOP

[Har70b] David E. F. Harrison. Undamped oscillations of pyridine nucleotide and oxygen tension in chemostat cultures of klebsiella aerogenes. Journal of Cell Biology, 45(3):514–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/514.

Harris:1973:GTC

[Har73] W. F. Harris. The geometry of triplets in ciliary basal bodies from primate oviducts. Journal of Cell Biology, 58(1):237–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
58/1/237.

Hausman:1971:DCY

[HAT71] Steven J. Hausman, Lucy M. Anderson, and William H. Telfer. The dependence of Cecropia yolk formation in vitro on specific blood proteins. Journal of Cell Biology, 48(2): 303–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/2/303.

Henningsen:1970:MPP

[HB70] K. W. Henningsen and J. E. Boynton. Macromolecular physiology of plastids. Journal of Cell Biology, 44(2):290– ??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/2/290.

Helmsing:1971:IAN

[HB71a] Pieter J. Helmsing and Hans D. Berendes. Induced accumulation of nonhistone proteins in polytene nuclei of Drosophila Hydei. Journal of Cell Biology, 50(3):893–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/893.

Higgins:1971:FSL

[HB71b] Joan A. Higgins and Russell J. Barrnett. Fine structural localization of acyltransferases. Journal of Cell Biology, 50 (1):102–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/1/102.

Hunter:1971:ETC

[HB71c] George R. Hunter and Gerald P. Brierley. On the “energizedtwisted” configuration of isolated beef heart mitochondria. Journal of Cell Biology, 50(1):250–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/1/250.

Hendler:1972:RPS

[HB72a] Richard W. Hendler and Amelia H. Burgess. Respiration and protein synthesis in Escherichia coli membrane-envelope fragments. Journal of Cell Biology, 55(2):266–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/266.

Higgins:1972:SBS

[HB72b] Joan A. Higgins and Russell J. Barrnett. Studies on the biogenesis of smooth endoplasmic reticulum membranes in livers of phenobarbital-treated rats. Journal of Cell Biology, 55(2): 282–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/282.

Huang:1973:LEW

[HB73] A. H. C. Huang and H. Beevers. Localization of enzymes within microbodies. Journal of Cell Biology, 58(2):379–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/58/2/379.

Harris:1974:CRB

[HBG74] Elizabeth H. Harris, John E. Boynton, and Nicholas W. Gillham. Chloroplast ribosome biogenesis in chlamydomonas. Journal of Cell Biology, 63(1):160–??, October 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/63/1/
160.

Hall:1970:MGA

[HBK70] Z. W. Hall, M. D. Bownds, and E. A. Kravitz. The metabolism of gamma aminobutyric acid in the lobster nervous system. Journal of Cell Biology, 46(2):290–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
2/290.

Hamberger:1970:CSM

[HBL70] Anders Hamberger, Christian Blomstrand, and Albert L. Lehninger. Comparative studies on mitochondria isolated from neuron-enriched and glia-enriched fractions of rabbit and beef brain. Journal of Cell Biology, 45(2):221–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/221.

Huttner:1973:GJA

[HBM73] Istvan Hüttner, Michel Boutet, and Robert H. More. Gap junctions in arterial endothelium. Journal of Cell Biology, 57 (1):247–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/247.

Hendler:1970:RPSa

[HBS70] Richard W. Hendler, Amelia H. Burgess, and Raymond Scharff. Respiration and protein synthesis in Escherichia coli membrane-envelope fragments. Journal of Cell Biology, 44(2):376–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/2/376.

Hanifin:1970:HMM

[HC70] J. M. Hanifin and M. J. Cline. Human monocytes and macrophages. Journal of Cell Biology, 46(1):97–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
1/97.

Hall:1971:IR

[HC71] J. D. Hall and F. L. Crane. Intracristal rods. Journal of Cell Biology, 48(2):420–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/48/2/420.

Hubbard:1972:EIR

[HC72] Ann L. Hubbard and Zanvil A. Cohn. The enzymatic iodination of the red cell membrane. Journal of Cell Biology, 55(2): 390–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/390.

Holt:1973:PVM

[HC73] Portia A. Holt and John O. Corliss. Pattern variability in microtubular arrays associated with the tentacles of actinobolina (ciliatea: Gymnostomatida). Journal of Cell Biology, 58(1):213–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/58/1/213.

Harb:1974:HFC

[HC74] Joseph M. Harb and D. Eugene Copeland. Helical form cristae in the mitochondria of the pseudobranch of cyprinodon variegatus adapted to conditions of constant light and darkness. Journal of Cell Biology, 62(3):886–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
3/886.

Harrison:1974:SLG

[HCAP74] P. R. Harrison, D. Conkie, N. Affara, and J. Paul. In situ localization of globin messenger RNA formation. Journal of Cell Biology, 63(2):402–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/402.

Hartman:1970:PT

[HD70a] H. Hartman and R. M. Dowben. Polyribosomes from Tetrahymena. Journal of Cell Biology, 44(3):676–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/676.

Horwitz:1970:GCC

[HD70b] Allen L. Horwitz and Albert Dorfman. The growth of cartilage cells in soft agar and liquid suspension. Journal of Cell Biology, 45(2):434–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/2/434.

Hay:1973:SCC

[HD73a] Elizabeth D. Hay and James W. Dodson. Secretion of collagen by corneal epithelium. Journal of Cell Biology, 57 (1):190–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/190.

Hirano:1973:CAW

[HD73b] Asao Hirano and Herbert M. Dembitzer. Cerebellar alterations in the weaver mouse. Journal of Cell Biology, 56(2): 478–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/478.

Hodarnau:1973:IHP

[HDB73] Adriana Hodârnâu, Silvia Dancea, and Octavian Bârzu. Isolation of highly purified mitochondria from rat pancreas.

Journal of Cell Biology, 59(1):222–??, October 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/59/1/
222.

Higgins:1974:FIF

[HDM74] Michael L. Higgins and Lolita Daneo-Moore. Factors influencing the frequency of mesosomes observed in fixed and unfixed cells of streptococcus faecalis. Journal of Cell Biology, 61(2):288–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/2/288.

Heath:1974:MFT

[Hea74] I. Brent Heath. Mitosis in the fungus thraustotheca clavata. Journal of Cell Biology, 60(1):204–??, January 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/60/1/
204.

Hendler:1971:RPS

[Hen71] Richard W. Hendler. Respiration and protein synthesis in Escherichia coli membrane-envelope fragments. Journal of Cell Biology, 51(3):664–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/3/664.

Horowitz:1970:AST

[HF70] Samuel B. Horowitz and I. Robert Fenichel. Analysis of sodium transport in the amphibian oocyte by extractive and radioautographic techniques. Journal of Cell Biology, 47(1): 120–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/1/120.

Hodgson:1971:FPD

[HF71] John R. Hodgson and Harold W. Fisher. Formation of polyribosomes during recovery from contact inhibition of replication. Journal of Cell Biology, 49(3):945–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/945.

Hopkins:1973:HSC

[HF73] Colin R. Hopkins and Marilyn G. Farquhar. Hormone secretion by cells dissociated from rat anterior pituitaries. Journal of Cell Biology, 59(2):276–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/2/276.

Herzog:1974:EGC

[HF74] Volloer Herzog and H. Dariush Fahimi. The effect of glutaraldehyde on catalase. Journal of Cell Biology, 60(1): 303–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/1/303.

Herzfeld:1973:SMB

[HFG73] Annemarie Herzfeld, Micheline Federman, and Olga Greengard. Subcellular morphometric and biochemical analyses of developing rat hepatocytes. Journal of Cell Biology, 57 (2):475–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/2/475.

Holtzman:1970:CEM

[HFK70] Eric Holtzman, Alan R. Freeman, and Leonard A. Kashner. A cytochemical and electron microscope study of channels in the Schwann cells surrounding lobster giant axons. Journal of Cell Biology, 44(2):438–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/44/2/438.

Hogan:1971:EPS

[HG71] Brigid Hogan and Paul R. Gross. The effect of protein synthesis inhibition on the entry of messenger RNA into the cytoplasm of sea urchin embryos. Journal of Cell Biology, 49 (3):692–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/692.

Hand:1972:SOS

[HG72] Arthur R. Hand and Stephen Gobel. The structural organization of the septate and gap junctions of hydra. Journal of Cell Biology, 52(2):397–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/2/397.

Horwitz:1971:SIE

[HH71] Henry B. Horwitz and Charles E. Holt. Specific inhibition by ethidium bromide of mitochondrial DNA synthesis in Physarum Polycephalum. Journal of Cell Biology, 49 (2):546–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/546.

Hynes:1974:CEP

[HH74] Richard O. Hynes and Kenneth C. Humphryes. Characterization of the external proteins of hamster fibroblasts. Journal of Cell Biology, 62(2):438–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/2/438.

Haskins:1971:OSC

[HHC71] Edward F. Haskins, A. Arlan Hinchee, and Richard A. Cloney. The occurrence of synaptonemal complexes in the slime mold echinostelium minutum De bary. Journal of Cell Biology, 51(3):898–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/3/898.

Henn:1972:PPM

[HHH72] Fritz A. Henn, Hans-Arne Hansson, and Anders Hamberger. Preparation of plasma membrane from isolated neurons. Journal of Cell Biology, 53(3):654–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/3/654.

Haga:1970:EMO

[HHP70] J. Y. Haga, M. G. Hamilton, and M. L. Petermann. Electron microscopic observations on the large subunit of the rat liver ribosome. Journal of Cell Biology, 47(1):211–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/211.

Hamberger:1970:SSI

[HHS70] Anders Hamberger, Hans-Arne Hansson, and Johan Sjöstrand. Surface structure of isolated neurons. Journal of Cell Biology, 47(2):319–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/2/319.

Hagopian:1971:DCE

[HI71] Helen K. Hagopian and Vernon M. Ingram. Developmental changes of erythropoiesis in cultured chick blastoderms. Journal of Cell Biology, 51(2):440–??, November 1971. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/51/2/
440.

Higgins:1974:SBS

[Hig74] Joan A. Higgins. Studies on the biogenesis of smooth endoplasmic reticulum membranes in hepatocytes of phenobarbitaltreated rats. Journal of Cell Biology, 62(3):635–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
62/3/635.

Hinsch:1970:PRI

[Hin70] Gertrude W. Hinsch. Possible role of intranuclear membranes in nuclear-cytoplasmic exchange in spider crab oocytes. Journal of Cell Biology, 47(2):531–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/2/531.

Herrmann:1970:SPD

[HK70a] R. G. Herrmann and K. V. Kowallik. Selective presentation of DNA-regions and membranes in chloroplasts and mitochondria. Journal of Cell Biology, 45(1):198–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
1/198.

Hicks:1970:IPM

[HK70b] R. M. Hicks and B. Ketterer. Isolation of the plasma membrane of the luminal surface of rat bladder epithelium, and the occurrence of a hexagonal lattice of subunits both in negatively stained whole mounts and in sectioned membranes.

Journal of Cell Biology, 45(3):542–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/45/3/542.

Hajos:1971:EMH

[HKF71] Ferenc Hajós and Sándor Kerpel-Fronius. Electron microscope histochemical evidence for a partial or total block of the tricarboxylic acid cycle in the mitochondria of presynaptic axon terminals. Journal of Cell Biology, 51(1):216–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
51/1/216.

Heidrich:1972:PPT

[HKKSH72] Hans-G Heidrich, Rolf Kinne, Eva Kinne-Saffran, and Kurt Hannig. The polarity of the proximal tubule cell in rat kidney. Journal of Cell Biology, 54(2):232–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/2/232.

Hikida:1974:RPF

[HL74a] Robert S. Hikida and Joseph A. Lombardo. Regeneration of pigeon fast and slow muscle fiber types after partial excision and mincing. Journal of Cell Biology, 61(2):414–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/414.

Hillman:1974:CCE

[HL74b] D. E. Hillman and R. Llins. Calcium-containing electrondense structures in the axons of the squid giant synapse. Journal of Cell Biology, 61(1):146–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/1/146.

Hagopian:1972:ECC

[HLI72] Helen K. Hagopian, Judith A. Lippke, and Vernon M. Ingram. Erythropoietic cell cultures from chick embryos. Journal of Cell Biology, 54(1):98–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/1/98.

Howard:1973:FSP

[HLP73] Roger B. Howard, Joseph C. Lee, and Leroy A. Pesch. The fine structure, potassium content, and respiratory activity of isolated rat liver parenchymal cells prepared by improved enzymatic techniques. Journal of Cell Biology, 57(3):642–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/3/642.

Herzog:1972:LEP

[HM72] V. Herzog and F. Miller. The localization of endogenous peroxidase in the lacrimal gland of the rat during postnatal development. Journal of Cell Biology, 53(3):662–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
3/662.

Hay:1974:GSE

[HM74a] Elizabeth D. Hay and Stephen Meier. Glycosaminoglycan synthesis by embryonic inductors: Neural tube, notochord, and lens. Journal of Cell Biology, 62(3):889–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
3/889.

Horowitz:1974:NPI

[HM74b] Samuel B. Horowitz and Leonard C. Moore. The nuclear permeability, intracellular distribution, and diffusion of inulin in the amphibian oocyte. Journal of Cell Biology, 60(2): 405–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/405.

Hepler:1970:IBM

[HMC70] Peter K. Hepler, J. Richard McIntosh, and Story Cleland. Intermicrotubule bridges in mitotic spindle apparatus. Journal of Cell Biology, 45(2):438–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/438.

Hoyle:1973:UBG

[HMS73] Graham Hoyle, Patricia A. McNeill, and Allen I. Selverston. Ultrastructure of barnacle giant muscle fibers. Journal of Cell Biology, 56(1):74–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/1/74.

Hermos:1971:DSP

[HMT71] John A. Hermos, Minnie Mathan, and Jerry S. Trier. DNA synthesis and proliferation by villous epithelial cells in fetal rats. Journal of Cell Biology, 50(1):255–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/50/1/
255.

Hama:1970:SOT

[HN70a] Kiyoshi Hama and Fumio Nagata. A stereoscope observation of tracheal epithelium of mouse by means of the high voltage electron microscope. Journal of Cell Biology, 45(3):654–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
45/3/654.

Hendler:1970:RPSb

[HN70b] Richard W. Hendler and N. Nanninga. Respiration and protein synthesis in Escherichia coli membrane-envelope fragments. Journal of Cell Biology, 46(1):114–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/46/1/
114.

Hagopian:1972:SSS

[HN72] Martin Hagopian and Eladio A. Nunez. Sarcolemmal scalloping at short sarcomere lengths with incidental observations on the T tubules. Journal of Cell Biology, 53(1):252–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
1/252.

Howell:1973:CMC

[HN73] Stephen H. Howell and Jay A. Naliboff. Conditional mutants in chlamydomonas reinhardtii blocked in the vegetative cell cycle. Journal of Cell Biology, 57(3):760–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
3/760.

Heller:1971:OBR

[HOB71] Joram Heller, Thomas J. Ostwald, and Dean Bok. The osmotic behavior of rod photoreceptor outer segment discs. Journal of Cell Biology, 48(3):633–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/3/633.

Holdsworth:1971:IPC

[Hol71] Robert H. Holdsworth. The isolation and partial characterization of the pyrenoid protein of eremosphaera viridis. Journal of Cell Biology, 51(2):499–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/2/499.

Hoober:1972:MPC

[Hoo72] J. Kenneth Hoober. A major polypeptide of chloroplast membranes of chlamydomonas reinhardi. Journal of Cell Biology, 52(1):84–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/1/84.

Hopkins:1972:BIT

[Hop72] C. R. Hopkins. The biosynthesis, intracellular transport, and packaging of melanocyte-stimulating peptides in the amphibian pars intermedia. Journal of Cell Biology, 53(3):642–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/3/642.

Horowitz:1972:PAO

[Hor72] Samuel B. Horowitz. The permeability of the amphibian oocyte nucleus, in situ. Journal of Cell Biology, 54(3):609– ??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/3/609.

Howell:1974:IBR

[How74] John N. Howell. Intracellular binding of ruthenium red in frog skeletal muscle. Journal of Cell Biology, 62(1):242–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
1/242.

Hyde:1970:IPO

[HP70] James M. Hyde and Dietrich Peters. The influence of pH and osmolality on fixation of the fowlpox virus core. Journal of Cell Biology, 46(1):179–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/46/1/179.

Huang:1973:CPC

[HP73] B. Huang and D. R. Pitelka. The contractile process in the ciliate, Stentor Coeruleus. Journal of Cell Biology, 57 (3):704–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/3/704.

Holwill:1974:DHF

[HP74] M. E. J. Holwill and Patricia D. Peters. Dynamics of the hispid flagellum of ochromonas danica. Journal of Cell Biology, 62(2):322–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/2/322.

Haber:1972:CIC

[HPHB72] James E. Haber, John G. Peloquin, Harlyn O. Halvorson, and Gary G. Borisy. Colcemid inhibition of cell growth and the characterization of a colcemid-binding activity in Saccharomyces Cerevisiae. Journal of Cell Biology, 55(2):355– ??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/355.

Hudspeth:1971:CGS

[HR71] A. J. Hudspeth and J. P. Revel. Coexistence of gap and septate junctions in an invertebrate epithelium. Journal of Cell Biology, 50(1):92–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/1/92.

Heuser:1973:ERS

[HR73] J. E. Heuser and T. S. Reese. Evidence for recycling of synaptic vesicle membrane during transmitter release at the frog neuromuscular junction. Journal of Cell Biology, 57 (2):315–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/2/315.

Holtrop:1974:EPH

[HRS74] Marijke E. Holtrop, Lawrence G. Raisz, and Hollis A. Simmons. The effects of parathyroid hormone, colchicine, and calcitonin on the ultrastructure and the activity of osteoclasts in organ culture. Journal of Cell Biology, 60(2):346– ??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/346.

Hopper:1972:CPC

[HRW72] A. F. Hopper, P. M. Rose, and R. W. Wannemacher. Cell population changes in the intestinal mucosa of proteindepleted or starved rats. Journal of Cell Biology, 53(1): 225–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/225.

Hackenbrock:1971:OPU

[HRWL71] Charles R. Hackenbrock, Terry G. Rehn, Eugene C. Weinbach, and John J. Lemasters. Oxidative phosphorylation and ultrastructural transformation in mitochondria in the intact ascites tumor cell. Journal of Cell Biology, 51(1):123–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
51/1/123.

Hagopian:1970:DZL

[HS70] Martin Hagopian and David Spiro. Derivation of the Z line in the embryonic chick heart. Journal of Cell Biology, 44 (3):683–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/3/683.

Hardin:1971:ULD

[HS71a] J. H. Hardin and S. S. Spicer. Ultrastructural localization of dialyzed iron-reactive mucosubstance in rabbit heterophils, basophils, and eosinophils. Journal of Cell Biology, 48(2): 368–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/2/368.

Heine:1971:MIP

[HS71b] Jochen W. Heine and Carl A. Schnaitman. A method for the isolation of plasma membrane of animal cells. Journal of Cell Biology, 48(3):703–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/48/3/703.

Hinkley:1972:AIT

[HS72a] Robert E. Hinkley and Frederick E. Samson. Anestheticinduced transformation of axonal microtubules. Journal of Cell Biology, 53(1):258–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/1/258.

Hohman:1972:LTU

[HS72b] Wayne Hohman and Harald Schraer. Low temperature ultramicroincineration of thin-sectioned tissue. Journal of Cell Biology, 55(2):328–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/2/328.

Hoober:1973:CSM

[HS73] J. Kenneth Hoober and William J. Stegeman. Control of the synthesis of a major polypeptide of chloroplast membranes in chlamydomonas reinhardi. Journal of Cell Biology, 56 (1):1–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/1/1.

Heidrich:1970:SCR

[HSH70] Hans-G. Heidrich, Roland Stahn, and Kurt Hannig. The surface charge of rat liver mitochondria and their membranes. Journal of Cell Biology, 46(1):137–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/1/137.

Haddad:1971:RSV

[HSH^{+}71] A. Haddad, Meredith D. Smith, Annette Herscovics, N. J. Nadler, and C. P. Leblond. Radioautographic study of in vivo and in vitro incorporation of fucose-^{3}H into thyroglobulin by rat thyroid follicular cells. Journal of Cell Biology, 49 (3):856–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/856.

Heuson-Stiennon:1972:ICS

[HSWD72] Jeanine-Anne Heuson-Stiennon, Jean-Claude Wanson, and Pierre Drochmans. Isolation and characterization of the sarcoplasmic reticulum of skeletal muscle. Journal of Cell Biology, 55(2):471–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/55/2/471.

Hellman:1970:SEN

[HT70] B. Hellman and I.-B. Täljedal. Solubilizing effect of nucleotides on isolated insulin secretory granules. Journal of Cell Biology, 47(1):289–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/47/1/289.

Hand:1973:DRD

[HT73] Roger Hand and Igor Tamm. DNA replication: Direction and rate of chain growth in mammalian cells. Journal of Cell Biology, 58(2):410–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/58/2/410.

Hinkley:1974:EHC

[HT74] Robert E. Hinkley and Alvin G. Telser. The effects of halothane on cultured mouse neuroblastoma cells. Journal of Cell Biology, 63(2):531–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/531.

Huang:1974:EEC

[Hua74] D. Huang. Effect of extracellular chondroitin sulfate on cultured chondrocytes. Journal of Cell Biology, 62(3):881– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/881.

Heding:1973:CEP

[HW73] Linda D. Heding and David A. Wolff. The cytochemical examination of poliovirus-induced cell damage. Journal of Cell Biology, 59(2):530–??, November 1973. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/2/530.

Hudson:1974:UMC

[HW74] Peggy R. Hudson and J. Robert Waaland. Ultrastructure of mitosis and cytokinesis in the multinucleate green alga acrosiphonia. Journal of Cell Biology, 62(2):274–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
2/274.

Humphreys:1973:FSP

[HWP73] Walter J. Humphreys, Tomasz J. Wodzicki, and Jerome J. Paulin. Fractographic studies of plastic-embedded cells by scanning electron microscopy. Journal of Cell Biology, 56 (3):876–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/876.

Hwang:1974:PMW

[HYC^{+}74] Kou M. Hwang, Linda C. Yang, Christine K. Carrico, Rose A. Schulz, John B. Schenkman, and Alan C. Sartorelli. Production of membrane whorls in rat liver by some inhibitors of protein synthesis. Journal of Cell Biology, 62(1):20–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
1/20.

Howell:1971:MBN

[HZM71] Neil Howell, Carol A. Zuiches, and Kenneth D. Munkres. Mitochondrial biogenesis in neurospora crassa. Journal of Cell Biology, 50(3):721–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/3/721.

Inoue:1974:GLC

[IBK74] Shinya Inoué, Gary G. Borisy, and Daniel P. Kiehart. Growth and lability of Chaetopterus oocyte mitotic spindles isolated in the presence of porcine brain tubulin. Journal of Cell Biology, 62(1):175–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/1/175.

Iwayama:1971:EGA

[IF71] Takashi Iwayama and J. B. Furness. Enhancement of the granulation of adrenergic storage vesicles in drug-free solution. Journal of Cell Biology, 48(3):699–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/48/3/
699.

Incefy:1971:IBC

[IK71] Genevieve S. Incefy and Attallah Kappas. Isolation and biochemical characterization of nuclei from chick embryo liver. Journal of Cell Biology, 50(2):385–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/385.

Ichikawa:1970:ZHP

[IKY70] Yoshiyuki Ichikawa, Masao Kuroda, and Toshio Yamano. Zonation of hemoprotein P-450 and cytochrome b_{5} in the adrenocortex of mammals. Journal of Cell Biology, 45(3): 640–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/3/640.

Ikehara:1973:LPB

[IP73] Yukio Ikehara and Henry C. Pitot. Localization of polysomebound albumin and serine dehydratase in rat liver cell fractions. Journal of Cell Biology, 59(1):28–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/59/1/28.

Ito:1974:SDA

[IS74] S. Ito and G. C. Schofield. Studies on the depletion and accumulation of microvilli and changes in the tubulovesicular compartment of mouse parietal cells in relation to gastric acid secretion. Journal of Cell Biology, 63(2):364–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/364.

Iles:1973:MLI

[ISNC73] G. H. Iles, P. Seeman, D. Naylor, and B. Cinader. Membrane lesions in immune lysis. Journal of Cell Biology, 56(2):

528–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/528.

Ikeuchi:1971:INE

[ISWS71] Tatsuro Ikeuchi, Mitsuyoshi Sanbe, Herbert Weinfeld, and Avery A. Sandberg. Induction of nuclear envelopes around metaphase chromosomes after fusion with interphase cells. Journal of Cell Biology, 51(1):104–??, October 1971. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/51/1/
104.

Ivanyi:1971:NSN

[Ivá71] Dagmar Iványi. Number and size of nucleoli in binucleate liver cells. Journal of Cell Biology, 51(1):313–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/313.

Iwayama:1971:NBA

[Iwa71] Takashi Iwayama. Nexuses between areas of the surface membrane of the same arterial smooth muscle cell. Journal of Cell Biology, 49(2):521–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/2/521.

Ingwall:1974:SCC

[IWM^{+}74] Joanne S. Ingwall, Cynthia D. Weiner, Manuel F. Morales, Elaine Davis, and Frank E. Stockdale. Specificity of creatine in the control of muscle protein synthesis. Journal of Cell Biology, 62(1):145–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/1/145.

Ikeuchi:1972:CPM

[IWS72] Tatsuro Ikeuchi, Herbert Weinfeld, and Avery A. Sandberg. Chromosome pulverization in micronuclei induced by tritiated thymidine. Journal of Cell Biology, 52(1):97–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
52/1/97.

Jahromi:1974:TDU

[JA74] S. S. Jahromi and H. L. Atwood. Three-dimensional ultrastructure of the crayfish neuromuscular apparatus. Journal of Cell Biology, 63(2):599–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/599.

Jasper:1972:PDI

[JB72] D. K. Jasper and J. R. Bronk. Polar distribution of intramitochondrial granules of mucosal epithelial cells. Journal of Cell Biology, 54(1):190–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/1/190.

Jensen:1973:TCP

[JB73] Michael S. Jensen and Dorothy F. Bainton. Temporal changes in pH within the phagocytic vacuole of the polymorphonuclear neutrophilic leukocyte. Journal of Cell Biology, 56(2):379–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/379.

Johnston:1973:VBM

[JC73] Pamela Leban Johnston and Edgar F. Carell. Vitamin B_{12} and the macromolecular composition of euglena. Journal of Cell Biology, 57(3):668–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/3/668.

Jeffery:1972:ETI

[Jef72] William R. Jeffery. Evidence for a temporal incompatibility between DNA replication and division during the cell cycle of Tetrahymena. Journal of Cell Biology, 53(3):624–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
3/624.

Johmann:1973:HAC

[JEG73] Carol Johmann, Ronald A. Eckhardt, and Martin A. Gorovsky. The histones associated with condensed and extended chromatin of mouse liver. Journal of Cell Biology, 58 (1):119–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/1/119.

Jeffery:1973:ASD

[JFdBJ73] William R. Jeffery, Joseph Frankel, Lawrence E. de Bault, and Leslie M. Jenkins. Analysis of the schedule of DNA replication in heat-synchronized Tetrahymena. Journal of Cell Biology, 59(1):1–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/1/1.

Jakovcic:1971:CCW

[JGR^{+}71] S. Jakovcic, G. S. Getz, M. Rabinowitz, H. Jakob, and H. Swift. Cardiolipin content of wild type and mutant yeasts in relation to mitochondrial function and development. Journal of Cell Biology, 48(3):490–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/3/490.

Jablecki:1973:ALN

[JHK73] Charles K. Jablecki, John E. Heuser, and Seymour Kaufman. Autoradiographic localization of new RNA synthesis in hypertrophying skeletal muscle. Journal of Cell Biology, 57 (3):743–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/3/743.

Jewett:1973:CCM

[JLS73] Paul H. Jewett, Stephen D. Leonard, and Joachim R. Sommer. Chicken cardiac muscle. Journal of Cell Biology, 56(2): 595–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/595.

Jones:1974:ECA

[JM74] R. E. Jones and A. A. Moscona. Effects of cytosine arabinoside on differential gene expression in embryonic neural retina. Journal of Cell Biology, 61(3):688–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
3/688.

Jaffe:1974:UVP

[JN74] Lionel F. Jaffe and Richard Nuccitelli. An ultrasensitive vibrating probe for measuring steady extracellular currents. Journal of Cell Biology, 63(2):614–??, November 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/63/2/
614.

Jamieson:1971:CVC

[JP71a] James D. Jamieson and George E. Palade. Condensing vacuole conversion and zymogen granule discharge in pancreatic exocrine cells: Metabolic studies. Journal of Cell Biology, 48 (3):503–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/3/503.

Jamieson:1971:SIT

[JP71b] James D. Jamieson and George E. Palade. Synthesis, intracellular transport, and discharge of secretory proteins in stimulated pancreatic exocrine cells. Journal of Cell Biology, 50(1):135–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/1/135.

Jan:1974:ULR

[JR74] Lily Yeh Jan and Jean-Paul Revel. Ultrastructural localization of rhodopsin in the vertebrate retina. Journal of Cell Biology, 62(2):257–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/2/257.

Jeffery:1970:RBD

[JSF70] William R. Jeffery, Kenneth D. Stuart, and Joseph Frankel. The relationship between deoxyribonucleic acid replication and cell division in heat-synchronized Tetrahymena. Journal of Cell Biology, 46(3):533–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/3/533.

Jewett:1971:CM

[JSJ71] Paul H. Jewett, J. R. Sommer, and E. A. Johnson. Cardiac muscle. Journal of Cell Biology, 49(1):50–??, April 1971.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/49/1/50.

Johnsen:1974:HCP

[JSP74] Sven Johnsen, Torbjørn Stokke, and Hans Prydz. HeLa cell plasma membranes. Journal of Cell Biology, 63(2):357– ??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/2/357.

Kapanci:1974:CIC

[KAI^{+}74] Yusuf Kapanci, André Assimacopoulos, Claude Irle, André Zwahlen, and Giulio Gabbiani. ”contractile interstitial cells” in pulmonary alveolar septa: a possible regulator of Ventilation/ perfusion ratio? Journal of Cell Biology, 60(2): 375–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/375.

Kane:1970:DIH

[Kan70] R. E. Kane. Direct isolation of the hyaline layer protein released from the cortical granules of the sea urchin egg at fertilization. Journal of Cell Biology, 45(3):615–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/615.

Kaye:1970:FSA

[Kay70] Jerome S. Kaye. The fine structure and arrangement of microcylinders in the lumina of flagellar fibers in Cricket spermatids. Journal of Cell Biology, 45(2):416–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/416.

Kotzin:1972:SIG

[KB72] Brian L. Kotzin and Robert F. Baker. Selective inhibition of genetic transcription in sea urchin embryos. Journal of Cell Biology, 55(1):74–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/1/74.

Korenbrot:1973:MCO

[KBC73] Juan I. Korenbrot, Dennis T. Brown, and Richard A. Cone. Membrane characteristics and osmotic behavior of isolated rod outer segments. Journal of Cell Biology, 56(2):389– ??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/389.

Konishi:1973:CIM

[KBH73] A. Konishi, W. S. Beacham, and C. C. Hunt. Centrioles in intrafusal muscle fibers. Journal of Cell Biology, 59(3): 749–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/3/749.

Klebe:1970:CPP

[KCR70] Robert J. Klebe, Tchaw-Ren Chen, and Frank H. Ruddle. Controlled production of proliferating somatic cell hybrids. Journal of Cell Biology, 45(1):74–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/45/1/74.

Kagayama:1974:EME

[KD74] M. Kagayama and W. W. Douglas. Electron microscope evidence of calcium-induced exocytosis in mast cells treated with 48/80 or the ionophores a-23187 and x-537a. Journal of Cell Biology, 62(2):519–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/2/519.

Kraehenbuhl:1971:ULI

[KDC71] J. P. Kraehenbuhl, P. B. De Grandi, and M. A. Campiche. Ultrastructural localization of intracellular antigen using enzyme-labeled antibody fragments. Journal of Cell Biology, 50(2):432–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/2/432.

Kemp:1974:KSD

[KDR74] D. J. Kemp, P. Y. Dyer, and G. E. Rogers. Keratin synthesis during development of the embryonic chick feather. Journal of Cell Biology, 62(1):114–??, July 1974. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/1/114.

Kreibich:1973:SRC

[KDS73] Gert Kreibich, Pascale Debey, and David D. Sabatini. Selective release of content from microsomal vesicles without membrane disassembly. Journal of Cell Biology, 58(2):436–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
58/2/436.

Kelly:1971:SRT

[Kel71] A. M. Kelly. Sarcoplasmic reticulum and T tubules in differentiating rat skeletal muscle. Journal of Cell Biology, 49 (2):335–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/335.

Kessel:1970:PDM

[Kes70] R. G. Kessel. The permeability of dragonfly malpighian tubule cells to protein using horseradish peroxidase as a tracer. Journal of Cell Biology, 47(1):299–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/299.

Kuettner:1971:LEC

[KESA71] K. E. Kuettner, R. Eisenstein, L. W. Soble, and C. Arsenis. Lysozyme in epiphyseal cartilage. Journal of Cell Biology, 49(2):450–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/450.

Keyhani:1973:RGA

[Key73] Ezzatollah Keyhani. Ribosomal granules associated with outer mitochondrial membrane in aerobic yeast cells. Journal of Cell Biology, 58(2):480–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/2/480.

Kaye:1974:SC

[KFHF74] Gordon I. Kaye, Cecilia M. Fenoglio, Frank B. Hoefle, and Jorge Fischbarg. Studies on the cornea. Journal of Cell Biology, 61(2):537–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/2/537.

Kennelly:1970:DPD

[KFJ70] J. J. Kennelly, R. H. Foote, and R. C. Jones. Duration of premeiotic deoxyribonucleic acid synthesis and the stages of prophase I in rabbit oocytes. Journal of Cell Biology, 47(3): 577–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/577.

Kleeman:1970:ODC

[KFP70] Karl T. Kleeman, J. L. Fryer, and K. S. Pilcher. Observed differences in CO_{2} requirements between mammalian and salmonid fish cell lines. Journal of Cell Biology, 47(3):796– ??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/796.

Kalt:1974:OEG

[KG74] Marvin R. Kalt and Joseph G. Gall. Observations on early germ cell development and premeiotic ribosomal DNA amplification in Xenopus Laevis. Journal of Cell Biology, 62(2): 460–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/460.

King:1972:REM

[KGK72] Mary E. King, Gabriel C. Godman, and Donald W. King. Respiratory enzymes and mitochondrial morphology of HeLa and L cells treated with chloramphenicol and ethidium bromide. Journal of Cell Biology, 53(1):127–??, April 1972. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/53/1/
127.

Kovacs:1970:SPP

[KH70] Charles J. Kovacs and Jack Van’t Hof. Synchronization of a proliferative population in a cultured plant tissue. Journal of Cell Biology, 47(2):536–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/2/536.

Krishan:1971:BCV

[KH71a] Awtar Krishan and Dora Hsu. Binding of colchicine-^{3}H to vinblastine- and vincristine-induced crystals in mammalian tissue culture cells. Journal of Cell Biology, 48(2):407– ??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/2/407.

Krishan:1971:VIR

[KH71b] Awtar Krishan and Dora Hsu. Vinblastine-induced ribosomal complexes. Journal of Cell Biology, 49(3):927–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/927.

Kreibich:1974:SAP

[KHS74] Gert Kreibich, Ann L. Hubbard, and David D. Sabatini. On the spatial arrangement of proteins in microsomal membranes from rat liver. Journal of Cell Biology, 60(3):616–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/60/3/616.

Kirkpatrick:1970:PIM

[KHTH70] Joel B. Kirkpatrick, Lyon Hyams, Virginia L. Thomas, and Peter M. Howley. Purification of intact microtubules from brain. Journal of Cell Biology, 47(2):384–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
2/384.

Kishi:1974:EMS

[Kis74] Kiyoshi Kishi. Electron microscope studies on puromycininduced changes in nerve cells of the medulla oblongata. Journal of Cell Biology, 63(3):1091–??, December 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/63/3/
1091.

Kilarski:1970:FMB

[KJ70] Wincenty Kilarski and Andrzej Jasiński. The formation of multivesicular bodies from the nuclear envelope. Journal of Cell Biology, 45(2):205–??, May 1970. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/205.

Kischer:1970:PBE

[KK70] C. Ward Kischer and Joe S. Keeter. Prostaglandin-B_{1}, embryonic skin, and the dermo–epidermal junction. Journal of Cell Biology, 47(1):303–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/47/1/303.

Kafatos:1971:PSP

[KK71] Fotis C. Kafatos and Vassili Kiortsis. The packaging of a secretory protein. Journal of Cell Biology, 48(2):426–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
48/2/426.

Kadota:1973:ICV

[KK73a] Ken Kadota and Tomoko Kadota. Isolation of coated vesicles, plain synaptic vesicles, and flocculent material from a crude synaptosome fraction of guinea pig whole brain. Journal of Cell Biology, 58(1):135–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/1/135.

Klevecz:1973:IDS

[KK73b] Robert R. Klevecz and Leon N. Kapp. Intermittent DNA synthesis and periodic expression of enzyme activity in the cell cycle of wi-38. Journal of Cell Biology, 58(3):564–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
58/3/564.

Kisch:1973:DSI

[KKCP73] Alexander L. Kisch, Robert O. Kelley, Harry Crissman, and Linda Paxton. Dimethyl sulfoxide-induced reversion of several features of polyoma transformed baby hamster kidney cells (bhk-21). Journal of Cell Biology, 57(1):38–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/38.

Kimberg:1972:ECA

[KL72] Daniel V. Kimberg and John N. Loeb. Effects of cortisone administration on rat liver mitochondria. Journal of Cell Biology, 55(3):635–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/3/635.

Kuriyama:1973:MAR

[KL73] Yoshiaki Kuriyama and David J. L. Luck. Membraneassociated ribosomes in mitochondria of neurospora crassa. Journal of Cell Biology, 59(3):776–??, December 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/59/3/
776.

Kit:1974:GCM

[KL74a] Saul Kit and Wai-Choi Leung. Genetic control of mitochondrial thymidine kinase in human-mouse and monkey-mouse somatic cell hybrids. Journal of Cell Biology, 61(1):35–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
61/1/35.

Koch:1974:GCD

[KL74b] K. Koch and H. L. Leffert. Growth control of differentiated fetal rat hepatocytes in primary monolayer culture. Journal of Cell Biology, 62(3):780–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/3/780.

Kleinig:1970:NMM

[Kle70] Hans Kleinig. Nuclear membranes from mammalian liver. Journal of Cell Biology, 46(2):396–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/2/396.

Kit:1974:ACC

[KLJ^{+}74] Saul Kit, Wai-Choi Leung, George Jorgensen, David Trkula, and Del Rose Dubbs. Acquisition of chick cytosol thymidine kinase activity by thymidine kinase-deficient mouse fibroblast cells after fusion with chick erythrocytes. Journal of Cell Biology, 63(2):505–??, November 1974. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/2/505.

Knusel:1973:INN

[KLKK73] Alois Knüsel, Béatrice Lehner, Clive C. Kuenzle, and Gonzague S. Kistler. Isolation of neuronal nuclei from rat brain cortex. Journal of Cell Biology, 59(3):762–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/762.

Kloetzel:1970:CDM

[Klo70] John A. Kloetzel. Compartmentalization of the developing macronucleus following conjugation in stylonychia and euplotes. Journal of Cell Biology, 47(2):395–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
2/395.

Knutton:1974:RSM

[KLR74] S. Knutton, A. R. Limbrick, and J. D. Robertson. Regular structures in membranes. Journal of Cell Biology, 62(3):679– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/679.

Kierszenbaum:1971:DIC

[KLT71] Abraham L. Kierszenbaum, Cesar M. Libanati, and Carlos J. Tandler. The distribution of inorganic cations in mouse testis. Journal of Cell Biology, 48(2):314–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
2/314.

Keenan:1970:BMC

[KMO^{+}70] T. W. Keenan, D. James Morré, Diane E. Olson, W. N. Yunghans, and Stuart Patton. Biochemical and morphological comparison of plasma membrane and milk fat globule membrane from bovine mammary gland. Journal of Cell Biology, 44(1):80–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/1/80.

Knowler:1973:CCN

[KMS73] John T. Knowler, Harold L. Moses, and Thomas C. Spelsberg. Comparison and characterization of nuclear isolation procedures as applied to chick oviduct. Journal of Cell Biology, 59(3):685–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/3/685.

Knight:1973:IES

[Kni73] E. Knight. Interferon: Effect on the saturation density to which mouse cells will grow in vitro. Journal of Cell Biology, 56(3):846–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/3/846.

Kleinow:1974:EMS

[KNM74] W. Kleinow, W. Neupert, and F. Miller. Electron microscope study of mitochondrial 60s and cytoplasmic 80s ribosomes from locusta migratoria. Journal of Cell Biology, 62(3):860– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/860.

Kumegawa:1974:SES

[KNY74] M. Kumegawa, T. Nakamura, and T. Yamada. Studies on the effect of selective membrane filtration on the differentiation of tissue cultures. Journal of Cell Biology, 62(3):908– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/908.

Khan:1970:LSD

[KO70] Talat A. Khan and Jane Overton. Lanthanum staining of developing chick cartilage and reaggregating cartilage cells. Journal of Cell Biology, 44(2):433–??, February 1970. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/2/
433.

Kundrat:1971:MB

[KP71] Ellen Kundrat and Frank A. Pepe. The M band. Journal of Cell Biology, 48(2):340–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/2/340.

Kramer:1972:USP

[KP72] M. F. Kramer and C. Poort. Unstimulated secretion of protein from rat exocrine pancreas cells. Journal of Cell Biology, 52(1):147–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/1/147.

Koehler:1974:FSO

[KP74] James K. Koehler and William D. Perkins. Fine structure observations on the distribution of antigenic sites on guinea pig spermatozoa. Journal of Cell Biology, 60(3):789–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
3/789.

Kopecka:1974:DFC

[KPF74] Marie Kopecká, H. J. Phaff, and G. H. Fleet. Demonstration of a fibrillar component in the cell wall of the yeast Saccharomyces Cerevisiae and its chemical nature. Journal of Cell Biology, 62(1):66–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/1/66.

Kimmel:1974:USP

[KPH^{+}74] G. L. Kimmel, F. G. Péron, A. Haksar, E. Bedigian, W. F. Robidoux, and M. T. Lin. Ultrastructure, steroidogenic potential, and energy metabolism of the Snell adrenocortical carcinoma 494. Journal of Cell Biology, 62(1):152–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
1/152.

Klein:1970:UAS

[KR70] P. Klein and E. Robbins. An ultrasensitive assay for soluble sulfhydryl and its application to the study of glutathione levels during the HeLa life cycle. Journal of Cell Biology, 46 (1):165–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/1/165.

Krawczyk:1971:PEC

[Kra71] Walter S. Krawczyk. A pattern of epidermal cell migration during wound healing. Journal of Cell Biology, 49(2):247–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/2/247.

Krishan:1972:CBT

[Kri72] Awtar Krishan. Cytochalasin-b: Time-lapse cinematographic studies on its effects on cytokinesis. Journal of Cell Biology, 54(3):657–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/3/657.

Kuehl:1971:SSS

[KS71] LeRoy Kuehl and Earl N. Sumsion. Studies on the site of synthesis of several soluble enzymes of the cell nucleus. Journal of Cell Biology, 50(1):1–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/1/1.

Kraft:1972:DAF

[KS72] N. Kraft and Ken Shortman. Differentiation of antibodyforming cells in toad spleen. Journal of Cell Biology, 52(2): 438–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/438.

Kedinger:1974:AAR

[KS74a] Claude Kedinger and Rene Simard. The action of α-amanitin on RNA synthesis in Chinese hamster ovary cells. Journal of Cell Biology, 63(3):831–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/3/831.

Kreibich:1974:SRC

[KS74b] Gert Kreibich and David D. Sabatini. Selective release of content from microsomal vesicles without membrane disassembly. Journal of Cell Biology, 61(3):789–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/61/3/

Kislev:1973:ECU

[KSE73] N. Kislev, C. M. Spolsky, and J. M. Eisenstadt. Effect of chloramphenicol on the ultrastructure of mitochondria in sensitive and resistant strains of HeLa. Journal of Cell Biology, 57(2):571–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/2/571.

Klietmann:1973:ECE

[KSN73] Wolfgang Klietmann, Nobuhiro Sato, and Margit M. K. Nass. Establishment and characterization of ethidium bromide resistance in simian virus 40-transformed hamster cells. Journal of Cell Biology, 58(1):11–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/1/11.

Kuettner:1970:LEC

[KSR^{+}70] K. E. Kuettner, L. W. Soble, R. D. Ray, R. L. Croxen, M. Passovoy, and R. Eisenstein. Lysozyme in epiphyseal cartilage. Journal of Cell Biology, 44(2):329–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
2/329.

Kuroiwa:1971:FSI

[KT71] Tsuneyoshi Kuroiwa and Nobunori Tanaka. Fine structures of interphase nuclei. Journal of Cell Biology, 49(3):939–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/939.

Kraemer:1972:CCD

[KT72] Paul M. Kraemer and Robert A. Tobey. Cell-cycle dependent desquamation of heparan sulfate from the cell surface. Journal of Cell Biology, 55(3):713–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/713.

Kierszenbaum:1974:NPR

[KT74a] A. L. Kierszenbaum and Laura L. Tres. Nucleolar and perichromosomal RNA synthesis during meiotic prophase in the mouse testis. Journal of Cell Biology, 60(1):39–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
60/1/39.

Kierszenbaum:1974:TSS

[KT74b] A. L. Kierszenbaum and Laura L. Tres. Transcription sites in spread meiotic prophase chromosomes from mouse spermatocytes. Journal of Cell Biology, 63(3):923–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
3/923.

Kulka:1972:CDG

[KTC72] Richard G. Kulka, Gordon M. Tomkins, and Richard B. Crook. Clonal differences in glutamine synthetase activity of hepatoma cells. Journal of Cell Biology, 54(1):175–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
1/175.

Kierszenbaum:1972:DIC

[KTT72] A. L. Kierszenbaum, Laura L. Tres, and C. J. Tandler. The distribution of inorganic cations during spermiogenesis in the mouse. Journal of Cell Biology, 53(1):239–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
1/239.

Kumar:1970:RST

[Kum70] Ajit Kumar. Ribosome synthesis in Tetrahymena Pyriformis. Journal of Cell Biology, 45(3):623–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/45/3/623.

Kuroiwa:1974:SMS

[Kur74] Tsuneyoshi Kuroiwa. Studies on mitochondrial structure and function in Physarum Polycephalum. Journal of Cell Biology, 63(1):299–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/1/299.

Kent:1974:NHN

[KW74a] Christine Kent and M. A. Williams. The nature of hypothalamo-neurohypophyseal neurosecretion in the rat. Journal of Cell Biology, 60(3):554–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/3/554.

Koike:1974:EDR

[KW74b] Katsuro Koike and David R. Wolstenholme. Evidence for discontinuous replication of circular mitochondrial DNA molecules from Novikoff rat ascites hepatoma cells. Journal of Cell Biology, 61(1):14–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/1/14.

Kress:1971:SCR

[KWR71] Yvonne Kress, Murray Wittner, and Robert M. Rosenbaum. Sites of cytoplasmic ribonucleoprotein-filament assembly in relation to helical body formation in axenic trophozoites of entamoeba histolytica. Journal of Cell Biology, 49(3):773–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/773.

Kennedy:1970:EHH

[KZ70] John R. Kennedy and Arthur M. Zimmerman. The effects of high hydrostatic pressure on the microtubules of Tetrahymena Pyriformis. Journal of Cell Biology, 47(3): 568–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/568.

Longo:1970:CSR

[LA70a] Frank J. Longo and Everett Anderson. A cytological study of the relation of the cortical reaction to subsequent events of fertilization in urethane-treated eggs of the sea urchin, Arbacia Punctulata. Journal of Cell Biology, 47(3):646– ??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/646.

Longo:1970:ENF

[LA70b] Frank J. Longo and Everett Anderson. The effects of nicotine on fertilization in the sea urchin, Arbacia Punctulata. Journal of Cell Biology, 46(2):308–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/2/308.

Luff:1971:CSR

[LA71] A. R. Luff and H. L. Atwood. Changes in the sarcoplasmic reticulum and transverse tubular system of fast and slow skeletal muscles of the mouse during postnatal development. Journal of Cell Biology, 51(2):369–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
2/369.

Lykkesfeldt:1974:EBD

[LA74] Anne E. Lykkesfeldt and H. A. Andersen. The effect of 5-bromodeoxyuridine on DNA replication and cell division in Tetrahymena Pyriformis. Journal of Cell Biology, 62(2): 316–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/316.

LaFountain:1974:MSN

[LaF74] James R. LaFountain. Microfilaments in spermatocytes of nephrotoma suturalis. Journal of Cell Biology, 60(3): 784–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/3/784.

Laguens:1971:MSM

[Lag71] Ruben Laguens. Morphometric study of myocardial mitochondria in the rat. Journal of Cell Biology, 48(3):673–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/48/3/673.

Lagunoff:1973:MFD

[Lag73] David Lagunoff. Membrane fusion during mast cell secretion. Journal of Cell Biology, 57(1):252–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/1/252.

Lake:1973:FCF

[Lak73] Robert S. Lake. Further characterization of the f1-histone phosphokinase of metaphase-arrested animal cells. Journal of Cell Biology, 58(2):317–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/2/317.

Lam:1972:BCG

[Lam72] Dominic M. K. Lam. The biosynthesis and content of gammaaminobutyric acid in the goldfish retina. Journal of Cell Biology, 54(2):225–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/2/225.

Landis:1973:UCM

[Lan73] Story C. Landis. Ultrastructural changes in the mitochondria of cerebellar Purkinje cells of nervous mutant mice. Journal of Cell Biology, 57(3):782–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/3/782.

Larsen:1970:GMI

[Lar70] W. J. Larsen. Genesis of mitochondria in insect fat body. Journal of Cell Biology, 47(2):373–??, November 1970. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/47/2/
373.

Latt:1974:DDS

[Lat74] Samuel A. Latt. Detection of DNA synthesis in interphase nuclei by fluorescence microscopy. Journal of Cell Biology, 62 (2):546–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/546.

LaVail:1973:KRO

[LaV73] Matthew M. LaVail. Kinetics of rod outer segment renewal in the developing mouse retina. Journal of Cell Biology, 58(3): 650–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/3/650.

Leenders:1972:ECR

[LB72] Hans J. Leenders and Pieter J. A. Beckers. The effect of changes in the respiratory metabolism upon genome activity. Journal of Cell Biology, 55(2):257–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/257.

Leppard:1971:FSM

[LC71] Gary G. Leppard and J. Ross Colvin. Fine structure of the middle lamella of aggregates of plant cells in suspension culture. Journal of Cell Biology, 50(1):237–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
1/237.

Leppard:1972:EOF

[LC72] Gary G. Leppard and J. Ross Colvin. Electron-opaque fibrils and granules in and between the cell walls of higher plants. Journal of Cell Biology, 53(3):695–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/3/695.

Leppard:1971:LEC

[LCRM71] Gary G. Leppard, J. Ross Colvin, Dyson Rose, and S. M. Martin. Lignofibrils on the external cell wall surface of cultured plant cells. Journal of Cell Biology, 50(1):63–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
1/63.

Levine:1973:UST

[LDC73] Rhea J. C. Levine, Maynard M. Dewey, and David E. Colflesh. Unusual striated thick filaments in limulus skeletal muscle. Journal of Cell Biology, 57(2):591–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
2/591.

Lazarow:1973:STRa

[LdD73a] Paul B. Lazarow and Christian de Duve. The synthesis and turnover of rat liver peroxisomes. Journal of Cell Biology, 59(2):491–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/491.

Lazarow:1973:STRb

[LdD73b] Paul B. Lazarow and Christian de Duve. The synthesis and turnover of rat liver peroxisomes. Journal of Cell Biology, 59(2):507–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/507.

Levine:1972:ILC

[LDdV72] Rhea J. C. Levine, Maynard M. Dewey, and George W. de Villafranca. Immunohistochemical localization of contractile proteins in limulus striated muscle. Journal of Cell Biology, 55(1):221–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/55/1/221.

LaPushin:1971:SGC

[LdH71] Ramah W. La Pushin and Etienne de Harven. A study of gluco-corticosteroid-induced pyknosis in the thymus and lymph node of the adrenalectomized rat. Journal of Cell Biology, 50(3):583–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/3/583.

Lanks:1974:IRA

[LDP74] Karl W. Lanks, Jan M. Dorwin, and Bruno Papirmeister. Increased rate of acetylcholinesterase synthesis in differentiating neuroblastoma cells. Journal of Cell Biology, 63(3): 824–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/824.

Leak:1971:SPL

[Lea71] Lee V. Leak. Studies on the permeability of lymphatic capillaries. Journal of Cell Biology, 50(2):300–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
2/300.

Lee:1970:PRV

[Lee70] Sin Hang Lee. The possible role of the vesicles in renal ammonia excretion. Journal of Cell Biology, 45(3):644–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/644.

Leffert:1974:GCDa

[Lef74a] H. L. Leffert. Growth control of differentiated fetal rat hepatocytes in primary monolayer culture. Journal of Cell Biology, 62(3):767–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/767.

Leffert:1974:GCDb

[Lef74b] H. L. Leffert. Growth control of differentiated fetal rat hepatocytes in primary monolayer culture. Journal of Cell Biology, 62(3):792–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/792.

Leibowitz:1971:EEB

[Lei71] Robert D. Leibowitz. The effect of ethidium bromide on mitochondrial DNA synthesis and mitochondrial DNA structure in HeLa cells. Journal of Cell Biology, 51(1):116–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/116.

Lentz:1970:DNJ

[Len70] Thomas L. Lentz. Development of the neuromuscular junction. Journal of Cell Biology, 47(2):423–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
2/423.

Lentz:1972:DNJ

[Len72a] Thomas L. Lentz. Development of the neuromuscular junction. Journal of Cell Biology, 55(1):93–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/55/1/93.

Lentz:1972:DLD

[Len72b] Thomas L. Lentz. Distribution of leucine-^{3}H during axoplasmic transport within regenerating neurons as determined by electron-microscope radioautography. Journal of Cell Biology, 52(3):719–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/719.

Langer:1972:LHC

[LF72] G. A. Langer and J. S. Frank. Lanthanum in heart cell culture. Journal of Cell Biology, 54(3):441–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
3/441.

Ladoulis:1970:PCS

[LG70] Charles T. Ladoulis and Thomas J. Gill. Physical chemical studies on the specific interaction of an acriflavinephosphotungstic acid complex with double-stranded nucleic acids. Journal of Cell Biology, 47(2):500–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
2/500.

Lecount:1972:TSM

[LG72] Thomas S. Lecount and Robert D. Grey. Transient shortening of microvilli induced by cycloheximide in the duodenal epithelium of the chicken. Journal of Cell Biology, 53 (2):601–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/601.

Low:1973:NRT

[LG73] Robert B. Low and Alfred L. Goldberg. Nonuniform rates of turnover of myofibrillar proteins in rat diaphragm. Journal of Cell Biology, 56(2):590–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/2/590.

Lillie:1973:SPS

[LH73] John H. Lillie and Seong S. Han. Secretory protein synthesis in the stimulated rat parotid gland. Journal of Cell Biology, 59(3):708–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/3/708.

Loeb:1970:SPR

[LK70] John N. Loeb and Daniel V. Kimberg. Sedimentation properties of rat liver mitochondria. Journal of Cell Biology, 46 (1):17–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/1/17.

Locke:1971:HAP

[LK71] M. Locke and N. Krishnan. Hot alcoholic phosphotungstic acid and uranyl acetate as routine stains for thick and thin sections. Journal of Cell Biology, 50(2):550–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
2/550.

Lipton:1972:FSA

[LK72] Bruce H. Lipton and Irwin R. Konigsberg. A fine-structural analysis of the fusion of myogenic cells. Journal of Cell Biology, 53(2):348–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/2/348.

Lee:1971:ARM

[LKB71] Se Yong Lee, Velibor Krsmanovic, and George Brawerman. Attachment of ribosomes to membranes during polysome formation in mouse sarcoma 180 cells. Journal of Cell Biology, 49(3):683–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/683.

Locke:1971:RMO

[LKM71] M. Locke, N. Krishnan, and J. T. McMahon. A routine method for obtaining high contrast without staining sections. Journal of Cell Biology, 50(2):540–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/540.

Lord:1973:ERS

[LKMB73] J. M. Lord, T. Kagawa, T. S. Moore, and H. Beevers. Endoplasmic reticulum as the site of lecithin formation in castor bean endosperm. Journal of Cell Biology, 57(3):659–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
3/659.

Lizardi:1972:ISS

[LL72] Paul M. Lizardi and David J. L. Luck. The intracellular site of synthesis of mitochondrial ribosomal proteins in neurospora crassa. Journal of Cell Biology, 54(1):56–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
1/56.

Lasek:1972:GEN

[LLP72] Raymond J. Lasek, Carolyn K. Lee, and Ronald J. Przybylski. Granular extensions of the nucleoli in giant neurons of Aplysia Californica. Journal of Cell Biology, 55(1): 237–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/1/237.

Lakhotia:1970:CBD

[LM70a] S. C. Lakhotia and A. S. Mukherjee. Chromosomal basis of dosage compensation in Drosophila. Journal of Cell Biology, 47(1):18–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/1/18.

Lavker:1970:FHC

[LM70b] Robert M. Lavker and A. Gedeon Matoltsy. Formation of horny cells. Journal of Cell Biology, 44(3):501–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/501.

Lynn:1970:DSW

[LM70c] Richard R. Lynn and Paul T. Magee. Development of the spore wall during ascospore formation in Saccharomyces Cerevisiae. Journal of Cell Biology, 44(3):688–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/688.

Locke:1971:OFM

[LM71] M. Locke and J. T. McMahon. The origin and fate of microbodies in the fat body of an insect. Journal of Cell Biology, 48 (1):61–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/61.

Ley:1973:SMD

[LM73] Kenneth D. Ley and Marilyn M. Murphy. Synchronization of mitochondrial DNA synthesis in Chinese hamster cells (line cho) deprived of isoleucine. Journal of Cell Biology, 58(2): 340–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/2/340.

Leuenberger:1974:LTA

[LN74] Peter M. Leuenberger and Alex B. Novikoff. Localization of transport adenosine triphosphatase in rat cornea. Journal of Cell Biology, 60(3):721–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/60/3/721.

Levine:1970:CSF

[LP70] R. P. Levine and A. Paszewski. Chloroplast structure and function in ac-20 , a mutant strain of Chlamydomonas reinhardi. Journal of Cell Biology, 44(3):540–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/540.

Lenk:1971:MSC

[LP71] R. Lenk and S. Penman. Morphological studies of cells grown in the absence of mitochondrial-specific protein synthesis. Journal of Cell Biology, 49(2):541–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/2/541.

Leffert:1972:SPC

[LP72] H. L. Leffert and D. Paul. Studies on primary cultures of differentiated fetal liver cells. Journal of Cell Biology, 52 (3):559–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/3/559.

Landis:1974:APF

[LR74] Dennis M. D. Landis and T. S. Reese. Arrays of particles in freeze-fractured astrocytic membranes. Journal of Cell Biology, 60(1):316–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/1/316.

Lavail:1971:GSN

[LRL71] Matthew M. Lavail and Liane Reif-Lehrer. Glutamine synthetase in the normal and dystrophic mouse retina. Journal of Cell Biology, 51(1):348–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/1/348.

Leyton:1971:CMH

[LS71] R. A. Leyton and E. H. Sonnenblick. Cardiac muscle of the horseshoe crab, limulus polyphemus. Journal of Cell Biology, 48(1):101–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/1/101.

Ledingham:1973:BIT

[LS73] Janet M. Ledingham and F. O. Simpson. Bundles of intracellular tubules in renal medullary interstitial cells. Journal of Cell Biology, 57(2):594–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/2/594.

Leffert:1974:AFB

[LS74] H. L. Leffert and S. Sell. Alpha_{1}-fetoprotein biosynthesis during the growth cycle of differentiated fetal rat hepatocytes in primary monolayer culture. Journal of Cell Biology, 61 (3):823–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/3/823.

LaVail:1972:PPE

[LSO72] Matthew M. LaVail, Richard L. Sidman, and Deborah O’Neil. Photoreceptor-pigment epithelial cell relationships in rats with inherited retinal degeneration. Journal of Cell Biology, 53(1):185–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/1/185.

Leskes:1971:DERa

[LSP71a] A. Leskes, P. Siekevitz, and G. E. Palade. Differentiation of endoplasmic reticulum in hepatocytes. Journal of Cell Biology, 49(2):264–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/2/264.

Leskes:1971:DERb

[LSP71b] A. Leskes, P. Siekevitz, and G. E. Palade. Differentiation of endoplasmic reticulum in hepatocytes. Journal of Cell Biology, 49(2):288–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/2/288.

Ley:1970:RID

[LT70] K. D. Ley and R. A. Tobey. Regulation of initiation of DNA synthesis in Chinese hamster cells. Journal of Cell Biology, 47(2):453–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/2/453.

Lentz:1971:DJC

[LT71] Thomas L. Lentz and J. P. Trinkaus. Differentiation of the junctional complex of surface cells in the developing fundulus blastoderm. Journal of Cell Biology, 48(3):455–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
3/455.

Lingwood:1974:RCN

[LT74] Clifford A. Lingwood and D. Brian Thomas. Rapid changes in nucleoside transport induced by growth inhibitors. Journal of Cell Biology, 61(2):359–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/2/359.

Lung:1972:UCD

[Lun72] Ben Lung. Ultrastructure and chromatin disaggregation of human sperm head with thioglycolate treatment. Journal of Cell Biology, 52(1):179–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/1/179.

Lutton:1973:EPS

[Lut73] J. D. Lutton. The effect of phagocytosis and spreading on macrophage surface receptors for concanavalin A. Journal of Cell Biology, 56(2):611–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/2/611.

Legg:1970:NOM

[LW70] Peter G. Legg and Richard L. Wood. New observations on microbodies. Journal of Cell Biology, 45(1):118–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
1/118.

Lagunoff:1974:TDM

[LW74a] David Lagunoff and Helen Wan. Temperature dependence of mast cell histamine secretion. Journal of Cell Biology, 61 (3):809–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/3/809.

Letourneau:1974:MCL

[LW74b] Paul C. Letourneau and Norman K. Wessells. Migratory cell locomotion versus nerve axon elongation. Journal of Cell Biology, 61(1):56–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/1/56.

Lambert:1972:SDD

[LWD^{+}72] B. Lambert, L. Wieslander, B. Daneholt, E. Egyházi, and U. Ringborg. In situ demonstration of DNA hybridizing with chromosomal and nuclear sap RNA in chironomus tentans. Journal of Cell Biology, 53(2):407–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/407.

Mabuchi:1973:MLP

[Mab73] Issei Mabuchi. A myosin-like protein in the cortical layer of the sea urchin egg. Journal of Cell Biology, 59(2):542– ??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/2/542.

Meneghini:1971:IGA

[MABL71] R. Meneghini, H. A. Armelin, J. Balsamo, and F. J. S. Lara. Indication of gene amplification in rhynchosciara by RNA-DNA hybridization. Journal of Cell Biology, 49(3):913–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/913.

Magalhaes:1972:NCC

[Mag72] Maria C. Magalhães. A new crystal-containing cell in human adrenal cortex. Journal of Cell Biology, 55(1):126–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
55/1/126.

Magun:1973:TAC

[Mag73] Bruce Magun. Two actions of cyclic AMP on melanosome movement in frog skin. Journal of Cell Biology, 57(3):845–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/3/845.

Malawista:1971:MM

[Mal71] Stephen E. Malawista. The melanocyte model. Journal of Cell Biology, 49(3):848–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/3/848.

Massover:1971:IYC

[Mas71] William H. Massover. Intramitochondrial yolk-crystals of frog oocytes. Journal of Cell Biology, 48(2):266–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
2/266.

Massover:1973:CSI

[Mas73] William H. Massover. Complex surface invaginations in frog oocytes. Journal of Cell Biology, 58(2):485–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
2/485.

Matter:1973:MSN

[Mat73] Alex Matter. A morphometric study on the nexus of rat cardiac muscle. Journal of Cell Biology, 56(3):690–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
3/690.

Maul:1970:UPC

[Mau70] Gerd G. Maul. Ultrastructure of pore complexes of annulate lamellae. Journal of Cell Biology, 46(3):604–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
3/604.

Maul:1971:ONP

[Mau71] Gerd G. Maul. On the octagonality of the nuclear pore complex. Journal of Cell Biology, 51(2):558–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
2/558.

Mattar:1971:MCR

[MB71a] Fathy E. Mattar and Thomas J. Byers. Morphological changes and the requirements for macromolecule synthesis during excystment of Acanthamoeba Castellanii. Journal of Cell Biology, 49(2):507–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/2/507.

Maul:1971:PFC

[MB71b] Gerd G. Maul and J. A. Brumbaugh. On the possible function of coated vesicles in melanogenesis of the regenerating fowl feather. Journal of Cell Biology, 48(1):41–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
1/41.

Model:1971:EMS

[MBCP71] Pat G. Model, Murray B. Bornstein, Stanley M. Crain, and George D. Pappas. An electron microscopic study of the development of synapses in cultured fetal mouse cerebrum continuously exposed to xylocaine. Journal of Cell Biology, 49(2):362–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/362.

Masurovsky:1970:ODN

[MBKM70] Edmund B. Masurovsky, Helena H. Benitez, Seung U. Kim, and Margaret R. Murray. Origin, development, and nature of intranuclear rodlets and associated bodies in chicken sympathetic neurons. Journal of Cell Biology, 44(1):172–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
44/1/172.

Monneron:1972:FND

[MBP72] Ariane Monneron, Günter Blobel, and George E. Palade. Fractionation of the nucleus by divalent cations. Journal of Cell Biology, 55(1):104–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/1/104.

Morimoto:1972:RCCa

[MBS72a] T. Morimoto, G. Blobel, and D. D. Sabatini. Ribosome crystallization in chicken embryos. Journal of Cell Biology, 52(2): 338–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/338.

Morimoto:1972:RCCb

[MBS72b] T. Morimoto, G. Blobel, and D. D. Sabatini. Ribosome crystallization in chicken embryos. Journal of Cell Biology, 52(2): 355–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/355.

Morel:1971:QHR

[MBW71] François M. M. Morel, Richard F. Baker, and Harold Wayland. Quantitation of human red blood cell fixation by glutaraldehyde. Journal of Cell Biology, 48(1):91–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
1/91.

Magalhaes:1970:EMR

[MC70] M. M. Magalhães and Antonio Coimbra. Electron microscope radioautographic study of glycogen synthesis in the rabbit retina. Journal of Cell Biology, 47(1):263–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/263.

Meldolesi:1971:VSE

[MC71a] Jacopo Meldolesi and Dario Cova. In vitro stimulation of enzyme secretion and the synthesis of microsomal membranes in the pancreas of the guinea pig. Journal of Cell Biology, 51(2): 396–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/396.

Myron:1971:MSP

[MC71b] D. R. Myron and J. L. Connelly. The morphology of the swelling process in rat liver mitochondria. Journal of Cell Biology, 48(2):291–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/2/291.

Meldolesi:1972:CCM

[MC72] Jacopo Meldolesi and Dario Cova. Composition of cellular membranes in the pancreas of the guinea pig. Journal of Cell Biology, 55(1):1–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/1/1.

Magalhaes:1973:EMA

[MC73] M. M. Magalhães and Antonio Coimbra. Electron microscope autoradiographic detection of sites of protein synthesis in the rabbit retina Müler cells. Journal of Cell Biology, 57 (1):77–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/77.

McNutt:1971:CIR

[MCB71] N. Scott McNutt, Lloyd A. Culp, and Paul H. Black. Contact-inhibited revertant cell lines isolated from sv40transformed cells. Journal of Cell Biology, 50(3):691–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/50/3/691.

McNutt:1973:CIR

[MCB73] N. Scott McNutt, Lloyd A. Culp, and Paul H. Black. Contact-inhibited revertant cell lines isolated from sv40transformed cells. Journal of Cell Biology, 56(2):412–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/56/2/412.

Macieira-Coelho:1971:ICD

[MCBBTG71] A. Macieira-Coelho, D. Brouty-Boyé, M. T. Thomas, and I. Gresser. Interferon and cell division. Journal of Cell Biology, 48(2):415–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/2/415.

McCormick:1974:PSN

[McC74] J. J. McCormick. Physical separation of nuclei from two independent plasmodia of Physarum Polycephalum after fusion. Journal of Cell Biology, 62(1):227–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/1/227.

Moran:1971:FSC

[MCE71] David T. Moran, Kent M. Chapman, and Richard A. Ellis. The fine structure of cockroach campaniform sensilla. Journal of Cell Biology, 48(1):155–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/1/155.

McIntosh:1973:ASb

[McI73] J. Richard McIntosh. The axostyle of saccinobaculus. Journal of Cell Biology, 56(2):324–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/2/324.

McIntosh:1974:BBM

[McI74] J. R. McIntosh. Bridges between microtubules. Journal of Cell Biology, 61(1):166–??, April 1974. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/1/166.

Meisner:1974:DNI

[MCIJ74] Lorraine F. Meisner, Thomas W. Chuprevich, Stanley L. Inhorn, and Collin B. Johnson. Dye-nucleoprotein interactions in giemsa banding. Journal of Cell Biology, 61(1):248–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
61/1/248.

McKanna:1974:PMM

[McK74] James A. McKanna. Permeability modulating membrane coats. Journal of Cell Biology, 63(1):317–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/63/1/
317.

McLaughlin:1971:CMD

[McL71] David J. McLaughlin. Centrosomes and microtubules during meiosis in the mushroom boletus rubinellus. Journal of Cell Biology, 50(3):737–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/3/737.

Martin:1974:CCP

[MCR74] James H. Martin, Freida L. Carson, and George J. Race. Calcium-containing platelet granules. Journal of Cell Biology, 60(3):775–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/3/775.

Millhouse:1971:LTO

[MCS71] Edward W. Millhouse, John J. Chiakulas, and Lawrence E. Scheving. Long-term organ culture of the Salamander heart. Journal of Cell Biology, 48(1):1–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/1/1.

Mindich:1972:MSB

[MD72] Leonard Mindich and Samuel Dales. Membrane synthesis in bacillus subtilis. Journal of Cell Biology, 55(1):32–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
55/1/32.

MacDonald:1971:OOZ

[ME71] Ronald D. MacDonald and Andrew G. Engel. Observations on organization of z-disk components and on rod-bodies of z-disk origin. Journal of Cell Biology, 48(2):431–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
2/431.

Meban:1972:LPA

[Meb72] C. Meban. Localization of phosphatidic acid phosphatase activity in granular pneumonocytes. Journal of Cell Biology, 53(1):249–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/249.

Meldolesi:1974:DCM

[Mel74] J. Meldolesi. Dynamics of cytoplasmic membranes in guinea pig pancreatic acinar cells. Journal of Cell Biology, 61 (1):1–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/1/1.

Machen:1972:PJC

[MEW72] Terry E. Machen, David Erlij, and F. B. P. Wooding. Permeable junctional complexes. Journal of Cell Biology, 54(2): 302–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/2/302.

Mezger-Freed:1971:PRH

[MF71] Liselotte Mezger-Freed. Puromycin resistance in haploid and heteroploid frog cells: Gene or membrane determined? Journal of Cell Biology, 51(3):742–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/3/742.

Medoff:1971:VAM

[MG71] Judith Medoff and Jerome Gross. In vitro aggregation of mixed embryonic kidney and nerve cells. Journal of Cell Biology, 50(2):457–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/2/457.

Meints:1973:PHS

[MG73] Russell Meints and Eugene Goldwasser. The persistence of hemopoietic stem cells in vitro. Journal of Cell Biology, 56(2):429–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/429.

Mueller:1971:EVC

[MGD71] G. A. Mueller, Mary Esther Gaulden, and Wanzer Drane. The effects of varying concentrations of colchicine on the progression of grasshopper neuroblasts into metaphase. Journal of Cell Biology, 48(2):253–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/2/253.

Miranda:1974:ACDa

[MGDT74] Armand F. Miranda, Gabriel C. Godman, Arline D. Deitch, and Stuart W. Tanenbaum. Action of cytochalasin D on cells of established lines. Journal of Cell Biology, 61(2):481–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
61/2/481.

Mersmann:1972:SBM

[MGHA72] H. J. Mersmann, J. Goodman, J. M. Houk, and S. Anderson. Studies on the biochemistry of mitochondria and cell morphology in the neonatal swine hepatocyte. Journal of Cell Biology, 53(2):335–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/2/335.

McIntyre:1974:CIR

[MGK74] John A. McIntyre, Norton B. Gilula, and Morris J. Karnovsky. Cryoprotectant-induced redistribution of intramembranous particles in mouse lymphocytes. Journal of Cell Biology, 60(1):192–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/60/1/192.

Magee:1974:ICL

[MGS^{+}74] Wayne E. Magee, Charles W. Goff, Jean Schoknecht, M. Diane Smith, and K. Cherian. The interaction of cationic liposomes containing entrapped horseradish peroxidase with cells in culture. Journal of Cell Biology, 63(2):492–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/492.

Miranda:1974:ACDb

[MGT74] Armand F. Miranda, Gabriel C. Godman, and Stuart W. Tanenbaum. Action of cytochalasin D on cells of established lines. Journal of Cell Biology, 62(2):406–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
2/406.

Mire:1970:OTP

[MHB70] J. J. Mire, W. J. Hendelman, and R. P. Bunge. Observations on a transient phase of focal swelling in degenerating unmyelinated nerve fibers. Journal of Cell Biology, 45(1):9–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
1/9.

McGill:1973:EDS

[MHB73] Manley McGill, T. C. Hsu, and B. R. Brinkley. Electrondense structures in mitochondria induced by short-term ethidium bromide treatment. Journal of Cell Biology, 59(1): 260–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/1/260.

Maul:1973:EMI

[MHBM73] Helmut M. Maul, Betty Yee Li Hsu, Thaddeus M. Borun, and Gerd G. Maul. Effect of metabolic inhibitors on nuclear pore formation during the HeLa S_{3} cell cycle. Journal of Cell Biology, 59(3):669–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/3/669.

Morre:1970:IGA

[MHM^{+}70] D. James Morré, R. L. Hamilton, H. H. Mollenhauer, R. W. Mahley, W. P. Cunningham, R. D. Cheetham, and V. S. Lequire. Isolation of a Golgi apparatus-rich fraction from rat liver. Journal of Cell Biology, 44(3):484–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/484.

Masur:1971:CBP

[MHSW71] Sandra K. Masur, Eric Holtzman, Irving L. Schwartz, and Roderich Walter. Correlation between pinocytosis and hydroosmosis induced by neurohypophyseal hormones and mediated by adenosine 3’,5’-cyclic monophosphate. Journal of Cell Biology, 49(3):582–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/3/582.

Mathan:1972:SFE

[MHT72] Minnie Mathan, John A. Hermos, and Jerry S. Trier. Structural features of the epithelio-Mesenchymal interface of rat duodenal mucosa during development. Journal of Cell Biology, 52(3):577–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/577.

McNutt:1971:FOO

[MHW71] N. Scott McNutt, Robert A. Hershberg, and Ronald S. Weinstein. Further observations on the occurrence of nexuses in benign and malignant human cervical epithelium. Journal of Cell Biology, 51(3):805–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/3/805.

Masur:1972:HSE

[MHW72] Sandra K. Masur, Eric Holtzman, and Roderich Walter. Hormone-stimulated exocytosis in the toad urinary bladder. Journal of Cell Biology, 52(1):211–??, January 1972. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/52/1/

Maniatis:1971:EDAa

[MI71a] George M. Maniatis and Vernon M. Ingram. Erythropoiesis during amphibian metamorphosis. Journal of Cell Biology, 49(2):372–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/372.

Maniatis:1971:EDAb

[MI71b] George M. Maniatis and Vernon M. Ingram. Erythropoiesis during amphibian metamorphosis. Journal of Cell Biology, 49(2):380–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/380.

Maniatis:1971:EDAc

[MI71c] George M. Maniatis and Vernon M. Ingram. Erythropoiesis during amphibian metamorphosis. Journal of Cell Biology, 49(2):390–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/390.

Miller:1973:CRS

[Mil73] Leo Miller. Control of 5s RNA synthesis during early development of anucleolate and partial nucleolate mutants of Xenopus Laevis. Journal of Cell Biology, 59(3):624–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/624.

Minor:1973:SC

[Min73] Ronald R. Minor. Somite chondrogenesis. Journal of Cell Biology, 56(1):27–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/1/27.

McKeel:1970:PCP

[MJ70] Daniel W. McKeel and Leonard Jarett. Preparation and characterization of a plasma membrane fraction from isolated fat cells. Journal of Cell Biology, 44(2):417–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
2/417.

McKeel:1974:EAC

[MJ74] Daniel W. McKeel and Leonard Jarett. The enrichment of adenylate cyclase in the plasma membrane and Golgi subcellular fractions of porcine adenohypophysis. Journal of Cell Biology, 62(1):231–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/1/231.

Meldolesi:1971:CCMa

[MJP71a] J. Meldolesi, J. D. Jamieson, and G. E. Palade. Composition of cellular membranes in the pancreas of the guinea pig. Journal of Cell Biology, 49(1):109–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/1/109.

Meldolesi:1971:CCMb

[MJP71b] J. Meldolesi, J. D. Jamieson, and G. E. Palade. Composition of cellular membranes in the pancreas of the guinea pig. Journal of Cell Biology, 49(1):130–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/1/130.

Meldolesi:1971:CCMc

[MJP71c] J. Meldolesi, J. D. Jamieson, and G. E. Palade. Composition of cellular membranes in the pancreas of the guinea pig. Journal of Cell Biology, 49(1):150–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/1/150.

Moser:1970:CSA

[MK70] John W. Moser and Gerald L. Kreitner. Centrosome structure in anthoceros laevis and marchantia polymorpha. Journal of Cell Biology, 44(2):454–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/44/2/454.

Matsumoto:1974:MDR

[MKPV74] Lloyd Matsumoto, Harumi Kasamatsu, Lajos Pikó, and Jerome Vinograd. Mitochondrial DNA replication in sea urchin oocytes. Journal of Cell Biology, 63(1):146–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
63/1/146.

Moss:1970:NDN

[ML70] F. P. Moss and C. P. Leblond. Nature of dividing nuclei in skeletal muscle of growing rats. Journal of Cell Biology, 44(2):459–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/2/459.

McIntosh:1971:DSM

[ML71] J. Richard McIntosh and Story C. Landis. The distribution of spindle microtubules during mitosis in cultured human cells. Journal of Cell Biology, 49(2):468–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/49/2/468.

Malech:1974:MEC

[ML74] Harry L. Malech and Thomas L. Lentz. Microfilaments in epidermal cancer cells. Journal of Cell Biology, 60(2):473– ??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/473.

Montenecourt:1970:MRC

[MLD70] Bland S. Montenecourt, Margaret E. Langsam, and Donald T. Dubin. Mitochondrial RNA from cultured animal cells. Journal of Cell Biology, 46(2):245–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/2/245.

Malaisse-Lagae:1971:SSC

[MLGML71] F. Malaisse-Lagae, M. H. Greider, W. J. Malaisse, and P. E. Lacy. The stimulus-secretion coupling of glucose-induced insulin release. Journal of Cell Biology, 49(2):530–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
2/530.

Moore:1974:RSM

[MLMPF74] G. P. M. Moore, Sue Lintern-Moore, Hannah Peters, and M. Faber. RNA synthesis in the mouse oocyte. Journal of Cell Biology, 60(2):416–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/2/416.

MacRae:1970:EMA

[MM70a] Edith K. MacRae and Gerald D. Meetz. Electron microscopy of the ammoniacal Silver reaction for histones in the erythropoietic cells of the chick. Journal of Cell Biology, 45 (2):235–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/2/235.

Matoltsy:1970:CNK

[MM70b] A. Gedeon Matoltsy and Margit N. Matoltsy. The chemical nature of keratohyalin granules of the epidermis. Journal of Cell Biology, 47(3):593–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/3/593.

Merk:1972:NJB

[MM72] Frederick B. Merk and N. Scott McNutt. Nexus junctions between dividing and interphase granulosa cells of the rat ovary. Journal of Cell Biology, 55(2):511–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/511.

Massaro:1973:PGP

[MM73] Gloria D. Massaro and Donald Massaro. Pulmonary granular pneumocytes. Journal of Cell Biology, 59(1):246–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/246.

Margulies:1974:RBC

[MM74a] Maurice M. Margulies and Allan Michaels. Ribosomes bound to chloroplast membranes in chlamydomonas reinhardtii. Journal of Cell Biology, 60(1):65–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/1/65.

Metuzals:1974:EME

[MM74b] J. Metuzals and W. E. Mushynski. Electron microscope and experimental investigations of the neurofilamentous network in Deiters’ neurons. Journal of Cell Biology, 61(3):701–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
61/3/701.

Miller:1974:CCC

[MMLM74] David H. Miller, Ira Seth Mellman, Derek T. A. Lamport, and Maureen Miller. The chemical composition of the cell wall of chlamydomonas gymnogama and the concept of a plant cell wall protein. Journal of Cell Biology, 63(2):420– ??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/2/420.

McGuire:1972:CBP

[MMM72] Joseph McGuire, Gisela Moellmann, and Frank McKeon. Cytochalasin B and pigment granule translocation. Journal of Cell Biology, 52(3):754–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/3/754.

Maul:1972:TSN

[MMS^{+}72] Gerd G. Maul, Helmut M. Maul, Joseph E. Scogna, Michael W. Lieberman, Gary S. Stein, Betty Yee-Li Hsu, and Thaddeus W. Borun. Time sequence of nuclear pore formation in phytohemagglutinin-stimulated lymphocytes and in HeLa cells during the cell cycle. Journal of Cell Biology, 55(2):433–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/55/2/433.

Matsuura:1974:UBS

[MMT^{+}74] S. Matsuura, T. Morimoto, Y. Tashiro, T. Higashinakagawa, and M. Muramatsu. Ultrastructural and biochemical studies on the precursor ribosomal particles isolated from rat liver nucleoli. Journal of Cell Biology, 63(2):629–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/629.

Mollenhauer:1971:FMD

[Mol71] Hilton H. Mollenhauer. Fragmentation of mature Dictyosome Cisternae. Journal of Cell Biology, 49(1):212–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
1/212.

McIntosh:1973:ASa

[MOL73] J. Richard McIntosh, Edward S. Ogata, and Story C. Landis. The axostyle of saccinobaculus. Journal of Cell Biology, 56(2):304–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/304.

McBride:1970:SNR

[MP70a] O. Wesley McBride and Elbert A. Peterson. Separation of nuclei representing different phases of the growth cycle from unsynchronized mammalian cell cultures. Journal of Cell Biology, 47(1):132–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/1/132.

Mclean:1970:LSQ

[MP70b] Robert J. Mclean and George F. Pessoney. A large scale quasi-crystalline lamellar lattice in chloroplasts of the green alga zygnema. Journal of Cell Biology, 45(3):522–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/522.

Murakami:1970:PCM

[MP70c] Satoru Murakami and Lester Packer. Protonation and chloroplast membrane structure. Journal of Cell Biology, 47(2): 332–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/2/332.

Murti:1970:MRA

[MP70d] Kuruganti G. Murti and David M. Prescott. Micronuclear ribonucleic acid in Tetrahymena Pyriformis. Journal of Cell Biology, 47(2):460–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/47/2/460.

Mains:1973:PCDa

[MP73a] Richard E. Mains and Paul H. Patterson. Primary cultures of dissociated sympathetic neurons. Journal of Cell Biology, 59(2):329–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/329.

Mains:1973:PCDb

[MP73b] Richard E. Mains and Paul H. Patterson. Primary cultures of dissociated sympathetic neurons. Journal of Cell Biology, 59(2):346–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/346.

Mains:1973:PCDc

[MP73c] Richard E. Mains and Paul H. Patterson. Primary cultures of dissociated sympathetic neurons. Journal of Cell Biology, 59(2):361–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/361.

Martinez-Palomo:1970:TCC

[MPAB70] A. Martinez-Palomo, J. Alanis, and D. Benitez. Transitional cardiac cells of the conductive system of the dog heart. Journal of Cell Biology, 47(1):1–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/1/1.

Martinez-Palomo:1973:SDL

[MPBA73] A. Martinez-Palomo, D. Benitez, and J. Alanis. Selective deposition of lanthanum in mammalian cardiac cell membranes. Journal of Cell Biology, 58(1):1–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/1/1.

Martinez-Palomo:1971:LPB

[MPEB71] A. Martinez-Palomo, D. Erlij, and H. Bracho. Localization of permeability barriers in the frog skin epithelium. Journal of Cell Biology, 50(2):277–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/2/277.

Maul:1971:FDN

[MPL71] Gerd G. Maul, Joseph W. Price, and Michael W. Lieberman. Formation and distribution of nuclear pore complexes in interphase. Journal of Cell Biology, 51(2):405–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
2/405.

Millecchia:1970:BBR

[MR70] Lyndell L. Millecchia and Maria A. Rudzinska. Basal body replication and ciliogenesis in a suctorian, Tokophrya Infusionum. Journal of Cell Biology, 46(3):553–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
3/553.

McKenna:1971:CUC

[MR71a] Olivia C. McKenna and Jack Rosenbluth. Characterization of an unusual catecholamine-containing cell type in the toad hypothalamus. Journal of Cell Biology, 48(3):650–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
3/650.

Moens:1971:SSP

[MR71b] Peter B. Moens and Ellen Rapport. Spindles, spindle plaques, and meiosis in the yeast Saccharomyces Cerevisiae (Hansen). Journal of Cell Biology, 50(2):344–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/344.

Miller:1974:ENC

[MR74] Richard A. Miller and Frank H. Ruddle. Enucleated neuroblastoma cells form neurites when treated with dibutyryl cyclic amp. Journal of Cell Biology, 63(1):295–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/295.

Mahowald:1970:IMC

[MS70a] Anthony P. Mahowald and Joan M. Strassheim. Intercellular migration of centrioles in the germarium of Drosophila Melanogaster. Journal of Cell Biology, 45(2):306–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/306.

Marantz:1970:SMC

[MS70b] Robert Marantz and Michael L. Shelanski. Structure of microtubular crystals induced by vinblastine in vitro. Journal of Cell Biology, 44(1):234–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/44/1/234.

Mohos:1970:VSA

[MS70c] Steven C. Mohos and Lorant Skoza. Variations in the sialic acid concentration of glomerular basement membrane preparations obtained by ultrasonic treatment. Journal of Cell Biology, 45(2):450–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/2/450.

Millette:1973:CDM

[MSGE73] C. F. Millette, P. G. Spear, W. E. Gall, and G. M. Edelman. Chemical dissection of mammalian spermatozoa. Journal of Cell Biology, 58(3):662–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/3/662.

Mares:1974:SDP

[MSM74] V. Mareš, B. Schultze, and W. Maurer. Stability of DNA in Purkinje cell nuclei of the mouse. Journal of Cell Biology, 63(2):665–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/2/665.

Mills:1970:SUE

[MT70] Elinor S. Mills and Yale J. Topper. Some ultrastructural effects of insulin, hydrocortisone, and prolactin on mammary gland explants. Journal of Cell Biology, 44(2):310–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
44/2/310.

Mollenhauer:1971:SSa

[MT71a] Hilton H. Mollenhauer and Clara Totten. Studies on seeds. Journal of Cell Biology, 48(2):387–??, February 1971. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/48/2/

Mollenhauer:1971:SSb

[MT71b] Hilton H. Mollenhauer and Clara Totten. Studies on seeds. Journal of Cell Biology, 48(2):395–??, February 1971. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/48/2/
395.

Mollenhauer:1971:SSc

[MT71c] Hilton H. Mollenhauer and Clara Totten. Studies on seeds. Journal of Cell Biology, 48(3):533–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/3/533.

Mooseker:1973:IRA

[MT73] Mark S. Mooseker and Lewis G. Tilney. Isolation and reactivation of the axostyle. Journal of Cell Biology, 56(1): 13–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/1/13.

Murphy:1974:RMM

[MT74] Douglas B. Murphy and Lewis G. Tilney. The role of microtubules in the movement of pigment granules in teleost melanophores. Journal of Cell Biology, 61(3):757–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
3/757.

Mulvaney:1971:CLL

[Mul71] B. D. Mulvaney. Chemography of lysosome-like structures in olfactory epithelium. Journal of Cell Biology, 51(2):568– ??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/568.

Muller:1972:SAH

[Mül72] Miklós Müller. Secretion of acid hydrolases and its intracellular source in Tetrahymena Pyriformis. Journal of Cell Biology, 52(2):478–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/2/478.

Muller:1973:BCT

[Mül73] Miklós Müller. Biochemical cytology of trichomonad flagellates. Journal of Cell Biology, 57(2):453–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/57/2/
453.

McNutt:1970:UN

[MW70] N. Scott McNutt and Ronald S. Weinstein. The ultrastructure of the nexus. Journal of Cell Biology, 47(3):666– ??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/666.

Miura:1971:EBY

[MW71] Yasusada Miura and Fred H. Wilt. The effects of 5bromodeoxyuridine on yolk sac erythropoiesis in the chick embryo. Journal of Cell Biology, 48(3):523–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/48/3/
523.

Manning:1972:CDM

[MWR72] Jerry E. Manning, David R. Wolstenholme, and Oliver C. Richards. Circular DNA molecules associated with chloroplasts of spinach, spinacia oleracea. Journal of Cell Biology, 53(2):594–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/594.

Matsui:1972:IPI

[MYWS72] Sei-Ichi Matsui, Hiroshi Yoshida, Herbert Weinfeld, and Avery A. Sandberg. Induction of prophase in interphase nuclei by fusion with metaphase cells. Journal of Cell Biology, 54 (1):120–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/1/120.

Nachmias:1972:EMO

[Nac72] Vivianne T. Nachmias. Electron microscope observations on myosin from Physarum Polycephalum. Journal of Cell Biology, 52(3):648–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/648.

Nachmias:1974:PPM

[Nac74] V. T. Nachmias. Properties of Physarum Myosin purified by a potassium iodide procedure. Journal of Cell Biology, 62(1):54–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/1/54.

Nag:1972:UAT

[Nag72] Asish C. Nag. Ultrastructure and adenosine triphosphatase activity of red and white muscle fibers of the caudal region of a fish, salmo gairdneri. Journal of Cell Biology, 55(1): 42–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/1/42.

Nakajima:1971:FSM

[Nak71] Yasuko Nakajima. Fine structure of the medial nucleus of the trapezoid body of the bat with special reference to two types of synaptic endings. Journal of Cell Biology, 50(1):121–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
50/1/121.

Nakao:1972:FUM

[Nak72] Taisuke Nakao. Fusion of unit membranes of the endoplasmic reticulum in the Gill epithelial cell of shrimp, caridina. Journal of Cell Biology, 53(3):853–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/3/853.

Nameroff:1972:CMR

[Nam72] Mark Nameroff. Contact-mediated reversible suppression of myogenesis. Journal of Cell Biology, 54(1):166–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
1/166.

Nanninga:1971:MBS

[Nan71a] N. Nanninga. The mesosome of bacillus subtilis as affected by chemical and physical fixation. Journal of Cell Biology, 48(1):

219–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/219.

Nanninga:1971:ULF

[Nan71b] N. Nanninga. Uniqueness and location of the fracture plane in the plasma membrane of bacillus subtilis. Journal of Cell Biology, 49(2):564–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/2/564.

Novi:1972:CCT

[NB72] Anna M. Novi and Renato Baserga. Changes in chromatin template activity and their relationship to DNA synthesis in mouse parotid glands stimulated by isoproterenol. Journal of Cell Biology, 55(3):554–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/554.

Noonan:1973:RCB

[NB73] Kenneth D. Noonan and Max M. Burger. The relationship of concanavalin a binding to lectin-initiated cell agglutination. Journal of Cell Biology, 59(1):134–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/134.

Narotzky:1974:BAC

[NB74] Robert Narotzky and William Bondareff. Biogenic amines in cultured neuroblastoma and astrocytoma cells. Journal of Cell Biology, 63(1):64–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/1/64.

Nichols:1971:DM

[NBF71] Barbara A. Nichols, Dorothy Ford Bainton, and Marilyn G. Farquhar. Differentiation of monocytes. Journal of Cell Biology, 50(2):498–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/2/498.

Noel:1971:UND

[NDAT71] J. S. Noel, W. C. Dewey, J. H. Abel, and R. P. Thompson. Ultrastructure of the nucleolus during the Chinese hamster cell cycle. Journal of Cell Biology, 49(3):830–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/830.

Neaves:1973:PSC

[Nea73] William B. Neaves. Permeability of sertoli cell tight junctions to lanthanum after ligation of ductus deferens and ductuli efferentes. Journal of Cell Biology, 59(3):559–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/559.

Niemeyer:1971:CGT

[NF71] G. Niemeyer and W. G. Forssmann. Comparison of glycerol treatment in frog skeletal muscle and mammalian heart. Journal of Cell Biology, 50(2):288–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/288.

Newsome:1974:ECA

[NFR^{+}74] D. A. Newsome, R. T. Fletcher, W. G. Robison, K. R. Kenyon, and G. J. Chader. Effects of cyclic AMP and sephadex fractions of chick embryo extract on cloned retinal pigmented epithelium in tissue culture. Journal of Cell Biology, 61(2):369–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/2/369.

Nir:1974:ULD

[NH74] Izhak Nir and Michael O. Hall. The ultrastructure of lipiddepleted rod photoreceptor membranes. Journal of Cell Biology, 63(2):587–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/2/587.

Nachman:1972:SIM

[NHB72] Ralph Nachman, James G. Hirsch, and Marco Baggiolini. Studies on isolated membranes of azurophil and specific granules from rabbit polymorphonuclear leukocytes. Journal of Cell Biology, 54(1):133–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/1/133.

Nicolson:1971:TDT

[NHS71] Garth L. Nicolson, Robert Hyman, and S. J. Singer. The twodimensional topographic distribution of h-2 histocompatibility alloantigens on mouse red blood cell membranes. Journal of Cell Biology, 50(3):905–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/3/905.

Nicolson:1971:SPA

[Nic71] Garth L. Nicolson. Structure of the photosynthetic apparatus in protein-embedded chloroplasts. Journal of Cell Biology, 50 (1):258–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/1/258.

Nicklas:1973:MGD

[Nic73a] R. Bruce Nicklas. Methods for gentle, differential heating of part of a single living cell. Journal of Cell Biology, 59(3): 595–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/3/595.

Nicolson:1973:ASHa

[Nic73b] Garth L. Nicolson. Anionic sites of human erythrocyte membranes. Journal of Cell Biology, 57(2):373–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
2/373.

Nair:1971:THR

[NK71] C. N. Nair and E. Knight. Turnover of HeLa ribosomal RNA. Journal of Cell Biology, 50(3):787–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/787.

Nigam:1973:GPP

[NLB73a] Vijai N. Nigam, R. Lallier, and C. Brailovsky. Ganglioside patterns and phenotypic characteristics in a normal variant and a transformed back variant of a simian virus 40-induced hamster tumor cell line. Journal of Cell Biology, 58(2): 307–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/2/307.

Noonan:1973:CCD

[NLB73b] Kenneth D. Noonan, Arnold J. Levine, and Max M. Burger. Cell cycle-dependent changes in the surface membrane as detected with [^{3}h] concanavalin A. Journal of Cell Biology, 58 (2):491–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/2/491.

Nicolson:1971:LSI

[NMS71] Garth L. Nicolson, V. T. Marchesi, and S. J. Singer. The localization of spectrin on the inner surface of human red blood cell membranes by ferritin-conjugated antibodies. Journal of Cell Biology, 51(1):265–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/1/265.

Novikoff:1972:PAC

[NN72] Phyllis M. Novikoff and Alex B. Novikoff. Peroxisomes in absorptive cells of mammalian small intestine. Journal of Cell Biology, 53(2):532–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/2/532.

Novikoff:1971:GAG

[NNQH71] Phyllis M. Novikoff, Alex B. Novikoff, Nelson Quintana, and Jean-Jacques Hauw. Golgi apparatus, gerl, and lysosomes of neurons in rat dorsal root ganglia, studied by thick section and thin section cytochemistry. Journal of Cell Biology, 50(3):859–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/3/859.

Nagano:1971:RFS

[NO71] Toshio Nagano and Iwao Ohtsuki. Reinvestigation of the fine structure of Reinke’s crystal in the human testicular interstitial cell. Journal of Cell Biology, 51(1):148–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/148.

Nemanic:1971:SEM

[NP71] Michael K. Nemanic and Dorothy R. Pitelka. A scanning electron microscope study of the lactating mammary gland. Journal of Cell Biology, 48(2):410–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
2/410.

Nicolson:1973:ASHb

[NP73] Garth L. Nicolson and Richard G. Painter. Anionic sites of human erythrocyte membranes. Journal of Cell Biology, 59(2):395–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/395.

Nilsson:1973:PMM

[NPD73] Robert Nilsson, Elisabeth Peterson, and Gustav Dallner. Permeability of microsomal membranes isolated from rat liver. Journal of Cell Biology, 56(3):762–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/3/762.

Nir:1970:PDD

[NS70] Izhak Nir and Arnold M. Seligman. Photooxidation of diaminobenzidine (dab) by chloroplast lamellae. Journal of Cell Biology, 46(3):617–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/3/617.

Nicolson:1974:DAM

[NS74] Garth L. Nicolson and S. J. Singer. The distribution and asymmetry of mammalian cell surface saccharides utilizing ferritin-conjugated plant agglutinins as specific saccharide stains. Journal of Cell Biology, 60(1):236–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
1/236.

Nakamura:1971:CSL

[NSG71] A. Nakamura, F. Sreter, and J. Gergely. Comparative studies of light meromyosin paracrystals derived from red, white, and cardiac muscle myosins. Journal of Cell Biology, 49 (3):883–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/3/883.

Nickerson:1970:OFA

[NSM70] Peter A. Nickerson, Floyd R. Skelton, and Agostino Molteni. Observation of filaments in the adrenal of androgen-treated rats. Journal of Cell Biology, 47(1):277–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/47/1/
277.

Nozawa:1971:SMFa

[NT71a] Yoshinori Nozawa and Guy A. Thompson. Studies of membrane formation in Tetrahymena Pyriformis. Journal of Cell Biology, 49(3):712–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/712.

Nozawa:1971:SMFb

[NT71b] Yoshinori Nozawa and Guy A. Thompson. Studies of membrane formation in Tetrahymena Pyriformis. Journal of Cell Biology, 49(3):722–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/722.

Nameroff:1973:ICD

[NTKM73] Mark Nameroff, John A. Trotter, John M. Keller, and Erlynda Munar. Inhibition of cellular differentiation by phospholipase C. Journal of Cell Biology, 58(1):107–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
1/107.

Nogami:1974:SPB

[NU74] Hiroshi Nogami and Marshall R. Urist. Substrata prepared from bone matrix for chondrogenesis in tissue culture. Journal of Cell Biology, 62(2):510–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/2/510.

Nakagami:1971:EPC

[NWL71] K. Nakagami, H. Warshawsky, and C. P. Leblond. The elaboration of protein and carbohydrate by rat parathyroid cells as revealed by electron microscope radioautography. Journal of Cell Biology, 51(3):596–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/3/596.

OBrien:1972:OAR

[O’B72] Thomas W. O’Brien. Occurrence of active 80 S ribosomes in subcellular particles in the mitochondrial fraction of fetal bovine liver. Journal of Cell Biology, 53(2):590–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
2/590.

Oron:1973:RPS

[OB73] U. Oron and A. Bdolah. Regulation of protein synthesis in the venom gland of viperid snakes. Journal of Cell Biology, 56(1): 177–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/1/177.

Organ:1972:MWE

[OBJ72] Alan E. Organ, Eugene C. Bovee, and Theodore L. Jahn. The mechanism of the water expulsion vesicle of the ciliate Tetrahymena Pyriformis. Journal of Cell Biology, 55(3): 644–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/3/644.

OLoughlin:1972:EMA

[OC72] G. T. O’Loughlin and T. C. Chambers. Extracellular microtubules in the aphid gut. Journal of Cell Biology, 53 (2):575–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/575.

Obara:1974:PIN

[OCWS74] Yoshitaka Obara, Lee S. Chai, Herbert Weinfeld, and Avery A. Sandberg. Prophasing of interphase nuclei and induction of nuclear envelopes around metaphase chromosomes in HeLa and Chinese hamster homo- and heterokaryons. Journal of Cell Biology, 62(1):104–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/1/104.

Oakley:1974:KAN

[OD74] B. R. Oakley and J. D. Dodge. Kinetochores associated with the nuclear envelope in the mitosis of a dinoflagellate. Journal of Cell Biology, 63(1):322–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/1/322.

Oschman:1974:ACM

[OHPW74] J. L. Oschman, T. A. Hall, P. D. Peters, and B. J. Wall. Association of calcium with membranes of squid giant axon. Journal of Cell Biology, 61(1):156–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/1/156.

Orkand:1971:LSA

[OK71] Paula M. Orkand and Edward A. Kravitz. Localization of the sites of γ-aminobutyric acid (GABA) uptake in lobster nerve– muscle preparations. Journal of Cell Biology, 49(1):75–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/1/75.

Ono:1972:CFC

[OKY72] H. Ono, S. Kobayashi, and K. Yanagisawa. Cell fusion in the cellular slime mold, Dictyostelium Discoideum. Journal of Cell Biology, 54(3):665–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/3/665.

Orlic:1973:FRI

[OL73] Donald Orlic and Robert Lev. Fetal rat intestinal absorption of horseradish peroxidase from swallowed amniotic fluid. Journal of Cell Biology, 56(1):106–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
1/106.

Orci:1971:MBS

[OLK^{+}71] Lelio Orci, Andre E. Lambert, Yasunori Kanazawa, Mylene Amherdt, Charles Rouiller, and Albert E. Renold. Morphological and biochemical studies of B cells of fetal rat endocrine pancreas in organ culture. Journal of Cell Biology, 50(3): 565–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/3/565.

Orloff:1973:EMI

[OM73] Kenneth G. Orloff and Richard L. McCarl. The effects of metabolic inhibitors on cultured rat heart cells. Journal of Cell Biology, 57(1):225–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/1/225.

Oleszko:1974:VPC

[OM74] Susan Oleszko and Evangelos N. Moudrianakis. The visualization of the photosynthetic coupling factor in embedded spinach chloroplasts. Journal of Cell Biology, 63(3):936– ??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/936.

Oberpriller:1971:MAN

[OO71] John Oberpriller and Jean C. Oberpriller. Mitosis in adult newt ventricle. Journal of Cell Biology, 49(2):560–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
2/560.

Olins:1972:PSI

[OO72] Donald E. Olins and Ada L. Olins. Physical studies of isolated eucaryotic nuclei. Journal of Cell Biology, 53(3):715–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
3/715.

Olivotto:1974:PMT

[OP74] Massimo Olivotto and Francesco Paoletti. Protein metabolism in tumor cells at various stages of growth in vivo. Journal of Cell Biology, 62(3):585–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/3/585.

Orci:1972:FRE

[OPL72] Lelio Orci, Alain Perrelet, and Arthur A. Like. Fenestrae in the rough endoplasmic reticulum of the exocrine pancreatic cells. Journal of Cell Biology, 55(1):245–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
1/245.

Oka:1974:CIR

[OPT74] Takami Oka, John W. Perry, and Yale J. Topper. Changes in insulin responsiveness during development of mammary epithelium. Journal of Cell Biology, 62(2):550–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
2/550.

ONeill:1972:KAM

[OS72] Michael C. O’Neill and Frank E. Stockdale. A kinetic analysis of myogenesis in vitro. Journal of Cell Biology, 52(1): 52–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/1/52.

OBrien:1974:GCD

[OS74] Stephen J. O’Brien and Yoshio Shimada. The α-glycerophosphate cycle in Drosophila melanogaster. Journal of Cell Biology, 63(3):864–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/3/864.

Osogoe:1973:PLG

[OTE73] Bunsuke Osogoe, R. W. Tyler, and N. B. Everett. The patterns of labeling of germinal-center cells with tritiated deoxycytidine. Journal of Cell Biology, 57(1):215–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/215.

Osogoe:1970:RSU

[OU70] Bunsuke Osogoe and Ayako Ueki. A radioautographic study of the utilization of deoxycytidine for the formation of deoxyribonucleic acid-thymine in lymphocytes. Journal of Cell Biology, 46(2):403–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/2/403.

Ovalle:1971:FSR

[Ova71] William K. Ovalle. Fine structure of rat intrafusal muscle fibers. Journal of Cell Biology, 51(1):83–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/51/1/83.

Ovalle:1972:FSR

[Ova72] William K. Ovalle. Fine structure of rat intrafusal muscle fibers. Journal of Cell Biology, 52(2):382–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
2/382.

Overton:1973:EMD

[Ove73] Jane Overton. Experimental manipulation of desmosome formation. Journal of Cell Biology, 56(3):636–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
3/636.

Oschman:1972:CBI

[OW72] James L. Oschman and Betty J. Wall. Calcium binding to intestinal membranes. Journal of Cell Biology, 55(1): 58–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/1/58.

Olins:1973:GFI

[OW73a] Donald E. Olins and Everline B. Wright. Glutaraldehyde fixation of isolated eucaryotic nuclei. Journal of Cell Biology, 59(2):304–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/59/2/304.

Ootaki:1973:OCP

[OW73b] T. Ootaki and J. J. Wolken. Octahedral crystals in phycomyces. II. Journal of Cell Biology, 57(2):278–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
2/278.

OConnor:1974:CDE

[OW74] Thomas M. O’Connor and Charles R. Wyttenbach. Cell death in the embryonic chick spinal cord. Journal of Cell Biology, 60(2):448–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/2/448.

Obara:1973:CBE

[OYC^{+}73] Yoshitaka Obara, Hiroshi Yoshida, Lee S. Chai, Herbert Weinfeld, and Avery A. Sandberg. Contrast between the environmental pH dependencies of prophasing and nuclear membrane formation in interphase–metaphase cells. Journal of Cell Biology, 58(3):608–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/3/608.

Personne:1970:LME

[PA70] Paul Personne and Winston Anderson. Localisation mitochondriale d’enzymes liees Au metabolisme du glycogene dans le spermatozoide De l’escargot. Journal of Cell Biology, 44(1):20–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/1/20.

Pendland:1973:UOC

[PA73] Jacquelyn C. Pendland and H. C. Aldrich. Ultrastructural organization of chloroplast thylakoids of the green alga “oocystis marssonii”. Journal of Cell Biology, 57(2):306–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
2/306.

Pappas:1971:MCI

[PAB71] George D. Pappas, Y. Asada, and M. V. L. Bennett. Morphological correlates of increased coupling resistance at an electrotonic synapse. Journal of Cell Biology, 49(1):173–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
1/173.

Padawer:1970:RRM

[Pad70] Jacques Padawer. The reaction of rat mast cells to polylysine. Journal of Cell Biology, 47(2):352–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
2/352.

Padawer:1974:IMC

[Pad74] Jacques Padawer. Identification of mast cells in the scanning electron microscope by means of X-ray spectrometry. Journal of Cell Biology, 61(3):641–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/3/641.

Pager:1971:EMD

[Pag71] Jeanne Pager. Étude morphométrique du système tubulaire transverse du myocarde ventriculaire de rat. (French) [Morphometric study of the transverse tubular system of rat ventricular myocardium]. Journal of Cell Biology, 50(1):233–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
50/1/233.

Palade:1971:ACB

[Pal71] George E. Palade. Albert claude and the beginnings of biological electron microscopy. Journal of Cell Biology, 50(1):??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
50/1/5d.

Pall:1973:SIV

[Pal73] Martin L. Pall. Sexual induction in Volvox carteri. Journal of Cell Biology, 59(1):238–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/1/238.

Pasternak:1972:RCI

[Pas72] C. A. Pasternak. Relief from contact inhibition. Journal of Cell Biology, 53(1):231–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/1/231.

Payton:1970:HSP

[Pay70] B. W. Payton. Histological staining properties of procion yellow m4rs. Journal of Cell Biology, 45(3):659–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/659.

Parry:1973:EMA

[PB73] D. M. Parry and N. M. Blackett. Electron microscope autoradiography of erythroid cells using radioactive iron. Journal of Cell Biology, 57(1):16–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/1/16.

Pfenninger:1974:FFN

[PB74] Karl H. Pfenninger and Richard P. Bunge. Freeze-fracturing of nerve growth cones and Young fibers. Journal of Cell Biology, 63(1):180–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/1/180.

Palay:1974:NPD

[PBGCP74] S. L. Palay, S. Billings-Gagliardi, and V. Chan-Palay. Neuronal perikarya with dispersed, single ribosomes in the visual cortex of macaca mulatta. Journal of Cell Biology, 63(3): 1074–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/1074.

Perlman:1974:FSC

[PBK74] Marcia Perlman, Jules L. Baum, and Gordon I. Kaye. Fine structure and collagen synthesis activity of monolayer cultures of rabbit corneal endothelium. Journal of Cell Biology, 63(1):306–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/1/306.

Pariser:1971:TIR

[PC71] Robert J. Pariser and Dennis D. Cunningham. Transport inhibitors released by 3t3 mouse cells and their relation to growth control. Journal of Cell Biology, 49(2):525–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
2/525.

Piko:1973:RMG

[PC73] Lajos Pikó and David G. Chase. Role of the mitochondrial genome during early development in mice. Journal of Cell Biology, 58(2):357–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/58/2/357.

Pagoulatos:1970:CHN

[PD70a] Gerassimos N. Pagoulatos and James E. Darnell. A comparison of the heterogeneous nuclear RNA of HeLa cells in different periods of the cell growth cycle. Journal of Cell Biology, 44(3):476–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/3/476.

Prusch:1970:CIC

[PD70b] Robert D. Prusch and Philip B. Dunham. Contraction of isolated contractile vacuoles from Amoeba Proteus. Journal of Cell Biology, 46(2):431–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/46/2/431.

Pepe:1972:MF

[PD72] Frank A. Pepe and Barbara Drucker. The myosin filament. Journal of Cell Biology, 52(2):255–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
2/255.

Piatigorsky:1972:CEC

[PdWW72] Joram Piatigorsky, Henry deF. Webster, and Miriam Wollberg. Cell elongation in the cultured embryonic chick lens epithelium with and without protein synthesis. Journal of Cell Biology, 55(1):82–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/1/82.

Plagemann:1972:TTC

[PE72a] Peter G. W. Plagemann and John Erbe. Thymidine transport by cultured Novikoff hepatoma cells and uptake by simple diffusion and relationship to incorporation into deoxyribonucleic acid. Journal of Cell Biology, 55(1):161–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
1/161.

Plagemann:1972:CB

[PE72b] Peter G. W. Plagemann and Richard D. Estensen. Cytochalasin B. Journal of Cell Biology, 55(1):179–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
1/179.

Peach:1972:NRT

[Pea72] Roy Peach. Nematosomes in the rat trigeminal ganglion. Journal of Cell Biology, 55(3):718–??, December 1972. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/55/3/
718.

Pelletier:1974:ASS

[Pel74] Georges Pelletier. Autoradiographic studies of synthesis and intracellular migration of glycoproteins in the rat anterior pituitary gland. Journal of Cell Biology, 62(1):185–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
1/185.

Peracchia:1970:SPS

[Per70] Camillo Peracchia. A system of parallel septa in crayfish nerve fibers. Journal of Cell Biology, 44(1):125–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
1/125.

Perrelet:1972:PSV

[Per72] Alain Perrelet. Protein synthesis in the visual cells of the honeybee drone as studied with electron microscope radioautography. Journal of Cell Biology, 55(3):595–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
3/595.

Peracchia:1973:LRJa

[Per73a] Camillo Peracchia. Low resistance junctions in crayfish. Journal of Cell Biology, 57(1):54–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/1/54.

Peracchia:1973:LRJb

[Per73b] Camillo Peracchia. Low resistance junctions in crayfish. Journal of Cell Biology, 57(1):66–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/1/66.

Perdue:1973:DUC

[Per73c] J. F. Perdue. The distribution, ultrastructure, and chemistry of microfilaments in cultured chick embryo fibroblasts. Journal of Cell Biology, 58(2):265–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/2/265.

Peracchia:1974:EMU

[Per74] Camillo Peracchia. Excitable membrane ultrastructure. Journal of Cell Biology, 61(1):107–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/1/107.

Polan:1973:ICM

[PFGG73] Mary Lake Polan, Susan Friedman, Joseph G. Gall, and Walter Gehring. Isolation and characterization of mitochondrial DNA from Drosophila Melanogaster. Journal of Cell Biology, 56(2):580–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/580.

Plattner:1972:FEC

[PFSB72] Helmut Plattner, Walter M. Fischer, Werner W. Schmitt, and Luis Bachmann. Freeze etching of cells without cryoprotectants. Journal of Cell Biology, 53(1):116–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
1/116.

Padykula:1970:UNJ

[PG70a] Helen A. Padykula and Geraldine F. Gauthier. The ultrastructure of the neuromuscular junctions of mammalian red, white, and intermediate skeletal muscle fibers. Journal of Cell Biology, 46(1):27–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/1/27.

Pederson:1970:MSD

[PG70b] Thoru Pederson and Seymour Gelfant. Macromolecular synthesis in dogfish peripheral blood cells. Journal of Cell Biology, 45(1):183–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/1/183.

Potmesil:1972:NMM

[PG72] Milan Potmesil and Anna Goldfeder. Nucleolar morphology and maturation of thymic lymphocytes. Journal of Cell Biology, 53(3):832–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/3/832.

Paul:1972:SPS

[PGHK72] M. Paul, M. R. Goldsmith, J. R. Hunsley, and F. C. Kafatos. Specific protein synthesis in cellular differentiation. Journal of Cell Biology, 55(3):653–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/653.

Peterson:1972:MSP

[PGR72] Joan B. Peterson, Robert H. Gray, and Hans Ris. Meiotic spindle plaques in Saccharomyces Cerevisiae. Journal of Cell Biology, 53(3):837–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/3/837.

Panner:1970:LSA

[PH70] Bernard J. Panner and Carl R. Honig. Locus and state of aggregation of myosin in tissue sections of vertebrate smooth muscle. Journal of Cell Biology, 44(1):52–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/1/52.

Perlin:1971:EAP

[PH71] Mina Perlin and Jules V. Hallum. Effect of acid pH on macromolecular synthesis in L cells. Journal of Cell Biology, 49 (1):66–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/1/66.

Poole:1970:STR

[PHdD70] Brian Poole, Tokuhiko Higashi, and Christian de Duve. The synthesis and turnover of rat liver peroxisomes. Journal of Cell Biology, 45(2):408–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/2/408.

Pitelka:1973:CCM

[PHDN73] Dorothy R. Pitelka, Susan T. Hamamoto, Joan G. Duafala, and Michael K. Nemanic. Cell contacts in the mouse mammary gland. Journal of Cell Biology, 56(3):797–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
3/797.

Podolsky:1970:IOP

[PHH70] R. J. Podolsky, T. Hall, and S. L. Hatchett. Identification of oxalate precipitates in striated muscle fibers. Journal of Cell Biology, 44(3):699–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/44/3/699.

Phillips:1970:IST

[Phi70] David M. Phillips. Insect sperm: Their structure and morphogenesis. Journal of Cell Biology, 44(2):243–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
2/243.

Phillips:1972:CAM

[Phi72a] David M. Phillips. Comparative analysis of mammalian sperm motility. Journal of Cell Biology, 53(2):561–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
2/561.

Phillips:1972:RPN

[Phi72b] Stephanie Gordon Phillips. Repopulation of the postmitotic nucleolus by preformed RNA. Journal of Cell Biology, 53 (3):611–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/3/611.

Phillips:1974:NSA

[Phi74] David M. Phillips. Nuclear shaping in the absence of microtubules in scorpion spermatids. Journal of Cell Biology, 62(3):911–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/3/911.

Pricam:1974:GJM

[PHPO74] Claude Pricam, Fabienne Humbert, Alain Perrelet, and Lelio Orci. Gap junctions in mesangial and lacis cells. Journal of Cell Biology, 63(1):349–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/1/349.

Pollard:1970:CFA

[PI70] Thomas D. Pollard and Susumu Ito. Cytoplasmic filaments of Amoeba Proteus. Journal of Cell Biology, 46(2):267–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
46/2/267.

Pollard:1971:FAP

[PK71] Thomas D. Pollard and Edward D. Korn. Filaments of Amoeba Proteus. Journal of Cell Biology, 48(1):216–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
48/1/216.

Penttila:1974:IGO

[PKT74] Antti Penttila, Hannu Kalimo, and Benjamin F. Trump. Influence of glutaraldehyde and/or osmium tetroxide on cell volume, ion content, mechanical stability, and membrane permeability of Ehrlich ascites tumor cells. Journal of Cell Biology, 63(1):197–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/1/197.

Plagemann:1972:NPN

[Pla72] Peter G. W. Plagemann. Nucleotide pools in Novikoff rat hepatoma cells growing in suspension culture. Journal of Cell Biology, 52(1):131–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/1/131.

Purdy:1972:SSC

[PLRK72] Joyce E. Purdy, Melvyn Lieberman, Anne E. Roggeveen, and R. Gary Kirk. Synthetic strands of cardiac muscle. Journal of Cell Biology, 55(3):563–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/563.

Peracchia:1972:FMG

[PM72a] Camillo Peracchia and Brant S. Mittler. Fixation by means of glutaraldehyde-hydrogen peroxide reaction products. Journal of Cell Biology, 53(1):234–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/53/1/234.

Purves:1972:DSP

[PM72b] Dale Purves and Uel J. McMahan. The distribution of synapses on a physiologically identified motor neuron in the central nervous system of the leech. Journal of Cell Biology, 55(1):205–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/55/1/205.

Pratley:1973:RMF

[PM73] James N. Pratley and Nancy K. McQuillen. The role of microfilaments in frog skin ion transport. Journal of Cell Biology, 56(3):850–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/3/850.

Palevitz:1970:SSE

[PN70] Barry A. Palevitz and Eldon H. Newcomb. A study of sieve element starch using sequential enzymatic digestion and electron microscopy. Journal of Cell Biology, 45(2):383–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/383.

Parr:1973:IMH

[PO73] Earl L. Parr and Jane S. Oei. Immobilization of membrane h-2 antigens by paraformaldehyde fixation. Journal of Cell Biology, 59(2):537–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/2/537.

Perrelet:1971:EGR

[POB71] Alain Perrelet, Lelio Orci, and Fritz Baumann. Evidence for granulolysis in the retinula cells of a stomatopod crustacean, squilla mantis. Journal of Cell Biology, 48(3):684–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/48/
3/684.

Peterson:1974:SBP

[POB74] Robert H. F. Peterson, Judith A. O’Neil, and June L. Biedler. Some biochemical properties of Chinese hamster cells sensitive and resistant to actinomycin D. Journal of Cell Biology, 63(3):773–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/3/773.

Pogo:1972:EEL

[Pog72a] Beatriz G. T. Pogo. Early events in lymphocyte transformation by phytohemagglutinin. Journal of Cell Biology, 53 (3):635–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/3/635.

Pogo:1972:SIR

[Pog72b] Beatriz G. T. Pogo. Specific inhibition by rifampicin of transcription in human lymphocytes stimulated by phytohemagglutinin. Journal of Cell Biology, 55(2):515–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/515.

Potten:1971:ELI

[Pot71] Christopher S. Potten. Early and late incorporation of tritiated thymidine into skin cells and the presence of a long-lived G_{0}-specific precursor pool. Journal of Cell Biology, 51(3):

855–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/855.

Phillips:1971:DCN

[PP71a] David M. Phillips and Stephanie G. Phillips. Distinctive characteristics of nucleoli of two established cell lines. Journal of Cell Biology, 49(3):803–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/3/803.

Phillips:1971:NDC

[PP71b] Stephanie G. Phillips and David M. Phillips. Nucleoli of diploid cell strains. Journal of Cell Biology, 49(3):785–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/785.

Price:1972:RVH

[PP72] Donald L. Price and Keith R. Porter. The response of ventral Horn neurons to axonal transection. Journal of Cell Biology, 53(1):24–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/24.

Paterson:1973:AAR

[PP73a] Bruce Paterson and Joav Prives. Appearance of acetylcholine receptor in differentiating cultures of embryonic chick breast muscle. Journal of Cell Biology, 59(1):241–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/241.

Pelletier:1973:DGA

[PP73b] G. Pelletier and R. Puviani. Detection of glycoproteins and autoradiographic localization of [^{3}H]fucose in the thyroidectomy cells of rat anterior pituitary gland. Journal of Cell Biology, 56(2):600–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/600.

Phillips:1973:RPN

[PP73c] David M. Phillips and Stephanie Gordon Phillips. Repopulation of postmitotic nucleoli by preformed RNA. Journal of Cell Biology, 58(1):54–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/58/1/54.

Phillips:1974:UDC

[PP74] E. R. Phillips and J. F. Perdue. Ultrastructural distribution of cell surface antigens in avian tumor virus-infected chick embryo fibroblasts. Journal of Cell Biology, 61(3):743–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
61/3/743.

Porter:1973:CSM

[PPF73] Keith Porter, David Prescott, and Jearl Frye. Changes in surface morphology of Chinese hamster ovary cells during the cell cycle. Journal of Cell Biology, 57(3):815–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
3/815.

Pederson:1970:ATC

[PR70a] Thoru Pederson and Elliott Robbins. Absence of translational control of histone synthesis during the HeLa cell life cycle. Journal of Cell Biology, 45(3):509–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/509.

Pederson:1970:RSH

[PR70b] Thoru Pederson and Elliott Robbins. RNA synthesis in HeLa cells. Journal of Cell Biology, 47(3):734–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
3/734.

Pederson:1971:MIS

[PR71a] Thoru Pederson and Elliott Robbins. A method for improving synchrony in the G_{2} phase of the cell cycle. Journal of Cell Biology, 49(3):942–??, June 1971. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/3/942.

Peracchia:1971:IOA

[PR71b] Camillo Peracchia and J. David Robertson. Increase in osmiophilia of axonal membranes of crayfish as a result of electrical stimulation, asphyxia, or treatment with reducing agents. Journal of Cell Biology, 51(1):223–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/223.

Pederson:1972:CSC

[PR72] Thoru Pederson and Elliott Robbins. Chromatin structure and the cell division cycle. Journal of Cell Biology, 55(2): 322–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/322.

Politoff:1974:CBS

[PRP74] Alberto L. Politoff, Steven Rose, and George D. Pappas. The calcium binding sites of synaptic vesicles of the frog sartorius neuromuscular junction. Journal of Cell Biology, 61 (3):818–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/3/818.

Przybylski:1971:OCD

[Prz71] Ronald J. Przybylski. Occurrence of centrioles during skeletal and cardiac myogenesis. Journal of Cell Biology, 49(1): 214–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/1/214.

Priest:1970:DLC

[PS70] Jean H. Priest and Robert H. Shikes. Distribution of labeled chromatin. Journal of Cell Biology, 47(1):99–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/99.

Potten:1971:LMM

[PS71] Christopher S. Potten and Jean C. Schaer. Labeling of murine mastocytoma cells in vitro with plasma tritiated thymidine-labeled animals. Journal of Cell Biology, 51(3): 862–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/862.

Porter:1973:SEM

[PTF73] Keith R. Porter, George J. Todaro, and Virginia Fonte. A scanning electron microscope study of surface features of viral and spontaneous transformants of mouse Balb/ 3t3 cells. Journal of Cell Biology, 59(3):633–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
3/633.

Peretz:1974:FIH

[PTLL74] Hava Peretz, Zivia Toister, Yehudith Laster, and Abraham Loyter. Fusion of intact human erythrocytes and erythrocyte ghosts. Journal of Cell Biology, 63(1):1–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/1.

Pulich:1974:RHO

[Pul74] Warren M. Pulich. Resistance to high oxygen tension, streptonigrin, and ultraviolet irradiation in the green alga chlorella sorokiniana strain ors. Journal of Cell Biology, 62(3):904– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/904.

Papadimitriou:1970:EFS

[PV70a] J. M. Papadimitriou and P. Van Duijn. Effects of fixation and substrate protection on the isoenzymes of aspartate aminotransferase studied in a quantitative cytochemical model system. Journal of Cell Biology, 47(1):71–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/47/1/71.

Papadimitriou:1970:ULI

[PV70b] J. M. Papadimitriou and P. Van Duijn. The ultrastructural localization of the isozymes of aspartate aminotransferase in murine tissues. Journal of Cell Biology, 47(1):84–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/84.

Pawlowski:1972:CRS

[PV72] Philip J. Pawlowski and Maurice H. Vaughan. Comparison of the relative synthesis of the proteins of the 50 S ribosomal subunit in growing and valine-deprived HeLa cells. Journal of Cell Biology, 52(2):409–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/2/409.

Palmiter:1971:IEP

[PW71] Richard D. Palmiter and Joan T. Wrenn. Interaction of estrogen and progesterone in chick oviduct development. Journal of Cell Biology, 50(3):598–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/3/598.

Podolsky:1973:SSM

[PW73] Daniel K. Podolsky and Milton M. Weiser. Specific selection of mitotically active intestinal cells by concanavalin aderivatized fibers. Journal of Cell Biology, 58(2):497–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
58/2/497.

Pysh:1974:SVD

[PW74] Joseph J. Pysh and Ronald G. Wiley. Synaptic vesicle depletion and recovery in cat sympathetic ganglia electrically stimulated in vivo. Journal of Cell Biology, 60(2):365–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
60/2/365.

Pasternak:1971:SAS

[PWT71] C. A. Pasternak, A. M. H. Warmsley, and D. B. Thomas. Structural alterations in the surface membrane during the cell cycle. Journal of Cell Biology, 50(2):562–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
2/562.

Quinton:1973:RAS

[QP73] P. M. Quinton and C. W. Philpott. A role for anionic sites in epithelial architecture. Journal of Cell Biology, 56(3): 787–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/787.

Quinton:1973:LSP

[QWT73] P. M. Quinton, Ernest M. Wright, and John McD. Tormey. Localization of sodium pumps in the choroid Plexus epithelium. Journal of Cell Biology, 58(3):724–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
3/724.

Reaven:1973:SMM

[RA73] Eve P. Reaven and Stanton G. Axline. Subplasmalemmal microfilaments and microtubules in resting and phagocytizing cultivated macrophages. Journal of Cell Biology, 59(1): 12–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/1/12.

Rae:1970:NPR

[Rae70] Peter M. M. Rae. The nature and processing of ribosomal ribonucleic acid in a dinoflagellate. Journal of Cell Biology, 46(1):106–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/1/106.

Rannestad:1974:RCP

[Ran74] John Rannestad. The regeneration of cilia in partially deciliated Tetrahymena. Journal of Cell Biology, 63(3):1009– ??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/1009.

Reddy:1974:NIH

[RASP74] Jarnardan K. Reddy, Daniel L. Azarnoff, Donald J. Svoboda, and Jada D. Prasad. Nafenopin-induced hepatic microbody (peroxisome) proliferation and catalase synthesis in rats and mice. Journal of Cell Biology, 61(2):344–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/344.

Rattner:1972:OCF

[Rat72] J. B. Rattner. Observations of centriole formation in male meiosis. Journal of Cell Biology, 54(1):20–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/54/1/20.

Rohlich:1971:EMO

[RAU71] Pál Röhlich, Per Anderson, and Börje Uvnäs. Electron microscope observations on compound 48/80-induced degranulation in rat mast cells. Journal of Cell Biology, 51(2): 465–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/465.

Rawlins:1973:TSA

[Raw73] Frank A. Rawlins. A time-sequence autoradiographic study of the in vivo incorporation of [1, 2-^{3}H]cholesterol into peripheral nerve myelin. Journal of Cell Biology, 58(1):42–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
1/42.

Richardson:1971:CSA

[RB71] John B. Richardson and Aurèle Beaulnes. The cellular site of action of angiotensin. Journal of Cell Biology, 51(2):419– ??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/419.

Rothstein:1973:LPT

[RB73] Thomas L. Rothstein and J. J. Blum. Lysosomal physiology in Tetrahymena. Journal of Cell Biology, 57(3):630–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
3/630.

Rattner:1974:LEM

[RB74a] J. B. Rattner and Michael W. Berns. Light and electron microscopy of laser microirradiated chromosomes. Journal of Cell Biology, 62(2):526–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/2/526.

Rothstein:1974:LPT

[RB74b] Thomas L. Rothstein and J. J. Blum. Lysosomal physiology in Tetrahymena. Journal of Cell Biology, 62(3):844– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/844.

Ryan:1974:FAR

[RBK74] Graeme B. Ryan, Joan Z. Borysenko, and Morris J. Karnovsky. Factors affecting the redistribution of surfacebound concanavalin a on human polymorphonuclear leukocytes. Journal of Cell Biology, 62(2):351–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/62/2/
351.

Redman:1972:SPR

[RC72] Colvin M. Redman and M. George Cherian. The secretory pathways of rat serum glycoproteins and albumin. Journal of Cell Biology, 52(2):231–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/2/231.

Rubin:1973:PPP

[RC73] Robert W. Rubin and William P. Cunningham. Partial purification and phosphotungstate solubilization of basal bodies and kinetodesmal fibers from Tetrahymena Pyriformis. Journal of Cell Biology, 57(3):601–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/3/601.

Reddy:1970:MEA

[RCBS70] Janardan Reddy, Masahiro Chiga, Sukhum Bunyaratvej, and Donald Svoboda. Microbodies in experimentally altered cells. Journal of Cell Biology, 44(1):226–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
1/226.

Reggio:1971:EPR

[RCDMM71] H. Reggio, H. Cailla-Deckmyn, and G. Marchis-Mouren. Effect of pancreozymin on rat pancreatic enzyme biosynthesis. Journal of Cell Biology, 50(2):333–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/333.

Rabinovitch:1973:MSA

[RdS73] Michel Rabinovitch and Mary Jo de Stefano. Manganese stimulates adhesion and spreading of mouse sarcoma I ascites cells. Journal of Cell Biology, 59(1):165–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/165.

Rosenberg:1972:SAP

[RE72] Steven A. Rosenberg and Albert B. Einstein. Sialic acids on the plasma membrane of cultured human lymphoid cells. Journal of Cell Biology, 53(2):466–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/466.

Rubin:1973:ECC

[RE73] R. W. Rubin and L. P. Everhart. The effect of cell-to-cell contact on the surface morphology of Chinese hamster ovary cells. Journal of Cell Biology, 57(3):837–??, June 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
3/837.

Rash:1974:SEM

[RE74] John E. Rash and Mark H. Ellisman. Studies of excitable membranes. Journal of Cell Biology, 63(2):567–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/567.

Redman:1971:PMI

[Red71] Colvin M. Redman. Phospholipid metabolism in intact and modified erythrocyte membranes. Journal of Cell Biology, 49(1):35–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/1/35.

Ross:1970:WHC

[RET70] Russell Ross, Newton B. Everett, and Ruth Tyler. Wound healing and collagen formation. Journal of Cell Biology, 44 (3):645–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/3/645.

Rubin:1973:ACM

[RF73] Robert W. Rubin and Philip Filner. Adenosine 3 ’, 5 ’-cyclic monophosphate in chlamydomonas reinhardtii. Journal of Cell Biology, 56(3):628–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/3/628.

Remacle:1974:ULC

[RFBB74] José Remacle, Stanley Fowler, Henri Beaufay, and Jacques Berthet. Ultrastructural localization of cytochrome B_{5} on rat liver microsomes by means of hybrid antibodies labeled with ferritin. Journal of Cell Biology, 61(1):237–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
1/237.

Ramus:1972:ISC

[RG72] J. Ramus and S. T. Groves. Incorporation of sulfate into the capsular polysaccharide of the red alga porphyridium. Journal of Cell Biology, 54(2):399–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/2/399.

Raff:1971:SSM

[RGGG71] Rudolf A. Raff, Gerald Greenhouse, Kenneth W. Gross, and Paul R. Gross. Synthesis and storage of microtubule proteins by sea urchin embryos. Journal of Cell Biology, 50(2): 516–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/2/516.

Rubin:1971:EMR

[RGR71] Walter Rubin, Michael D. Gershon, and Leonard L. Ross. Electron microscope radioautographic identification of serotonin-synthesizing cells in the mouse gastric mucosa. Journal of Cell Biology, 50(2):399–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/399.

Rose:1971:CPC

[RHW71] P. M. Rose, A. F. Hopper, and R. W. Wannemacher. Cell population changes in the intestinal mucosa of proteindepleted or starved rats. Journal of Cell Biology, 50(3):887– ??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/3/887.

Reynolds:1971:ECD

[RI71] Linda W. Reynolds and Vernon M. Ingram. Erythroid cell differentiation and the inhibition of cytokinesis by cytochalasin. Journal of Cell Biology, 51(2):433–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
2/433.

Ross:1971:SMCa

[RK71] Russell Ross and Seymour J. Klebanoff. The smooth muscle cell. Journal of Cell Biology, 50(1):159–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/50/1/
159.

Rubin:1973:IDS

[RK73] H. Rubin and T. Koide. Inhibition of DNA synthesis in chick embryo cultures by deprivation of either serum or zinc. Journal of Cell Biology, 56(3):777–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/3/777.

Ris:1974:UMM

[RK74a] Hans Ris and Donna F. Kubai. An unusual mitotic mechanism in the parasitic protozoan syndinium Sp. Journal of Cell Biology, 60(3):702–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/60/3/702.

Rodewald:1974:PSG

[RK74b] Richard Rodewald and Morris J. Karnovsky. Porous substructure of the glomerular slit diaphragm in the rat and mouse. Journal of Cell Biology, 60(2):423–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
2/423.

Ryan:1971:FFE

[RKT71] K. J. Ryan, H. Kalant, and E. Llewellyn Thomas. Free-flow electrophoretic separation and electrical surface properties of subcellular particles from guinea pig brain. Journal of Cell Biology, 49(2):235–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/2/235.

Reif-Lehrer:1974:EWI

[RL74] Liane Reif-Lehrer. Effect of washout of inducing steroid on γ-glutamyl transferase activity in the cultured chick embryo retina. Journal of Cell Biology, 63(1):312–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/312.

Radford:1971:CEE

[RM71] Sharon V. Radford and Donald W. Misch. The cytological effect of ecdysterone on the midgut cells of the flesh-fly sarcophaga bullata. Journal of Cell Biology, 49(3):702–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/702.

Raizada:1972:TTH

[RM72] M. K. Raizada and C. R. Krishna Murti. Transformation of trophic hartmannella culbertsoni into viable cysts by cyclic 3’,5’-adenosine monophosphate. Journal of Cell Biology, 52 (3):743–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/3/743.

Rice:1970:OCF

[RMM^{+}70] Robert V. Rice, Joan A. Moses, G. M. McManus, Arlene C. Brady, and Lorraine M. Blasik. The organization of contractile filaments in a mammalian smooth muscle. Journal of Cell Biology, 47(1):183–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/47/1/183.

Roth:1971:AIA

[RMR71a] Stephen Roth, Edward J. McGuire, and Saul Roseman. An assay for intercellular adhesive specificity. Journal of Cell Biology, 51(2):525–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/2/525.

Roth:1971:ECS

[RMR71b] Stephen Roth, Edward J. McGuire, and Saul Roseman. Evidence for cell-surface glycosyltransferases. Journal of Cell Biology, 51(2):536–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/51/2/536.

Raskas:1971:TVR

[RO71] Heschel J. Raskas and Case K. Okubo. Transport of viral RNA in Kb cells infected with adenovirus type 2. Journal of Cell Biology, 49(2):438–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/49/2/438.

Roberts:1974:MAP

[RO74] Charles T. Roberts and Eduardo Orias. On the mechanism of adaptation to protein synthesis inhibitors by Tetrahymena. Journal of Cell Biology, 62(3):707–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
3/707.

Robison:1972:MPS

[Rob72] W. Gerald Robison. Microtubular patterns in spermatozoa of coccid insects in relation to bending. Journal of Cell Biology, 52(1):66–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/1/66.

Rodewald:1970:SAT

[Rod70] Richard Rodewald. Selective antibody transport in the proximal small intestine of the neonatal rat. Journal of Cell Biology, 45(3):635–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/3/635.

Rodewald:1973:ITA

[Rod73] Richard Rodewald. Intestinal transport of antibodies in the newborn rat. Journal of Cell Biology, 58(1):189–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
1/189.

Rohrlich:1974:FSD

[Roh74] Susannah T. Rohrlich. Fine structural demonstration of ordered arrays of cytoplasmic filaments in vertebrate iridophores. Journal of Cell Biology, 62(2):295–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
2/295.

Rosenbluth:1971:MLA

[Ros71a] Jack Rosenbluth. Myosin-like aggregates in trypsin-treated smooth muscle cells. Journal of Cell Biology, 48(1):174– ??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/1/174.

Rosenbluth:1971:MLT

[Ros71b] Jack Rosenbluth. Myosin-like tactoids in trypsin-treated blood platelets. Journal of Cell Biology, 50(3):900–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/50/3/900.

Ross:1971:SMCb

[Ros71c] Russell Ross. The smooth muscle cell. Journal of Cell Biology, 50(1):172–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/1/172.

Rosenbluth:1972:MJT

[Ros72] Jack Rosenbluth. Myoneural junctions of two ultrastructurally distinct types in earthworm body wall muscle. Journal of Cell Biology, 54(3):566–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/3/566.

Rosenbluth:1973:MSI

[Ros73] Jack Rosenbluth. Membrane specialization at an insect myoneural junction. Journal of Cell Biology, 59(1):143–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
59/1/143.

Rosenbluth:1974:SAM

[Ros74] Jack Rosenbluth. Substructure of amphibian motor end plate. Journal of Cell Biology, 62(3):755–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
3/755.

Rowe:1971:UZL

[Row71] R. W. Rowe. Ultrastructure of the Z line of skeletal muscle fibers. Journal of Cell Biology, 51(3):674–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/674.

Rowe:1973:UZD

[Row73] R. W. D. Rowe. The ultrastructure of Z disks from white, intermediate, and red fibers of mammalian striated muscles. Journal of Cell Biology, 57(2):261–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/2/261.

Rohrlich:1972:FSO

[RP72] Susannah T. Rohrlich and Keith R. Porter. Fine structural observations relating to the production of color by the iridophores of a lizard, anolis carolinensis. Journal of Cell Biology, 53(1):38–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/1/38.

Rattner:1973:ICF

[RP73] J. B. Rattner and Stephanie G. Phillips. Independence of centriole formation and DNA synthesis. Journal of Cell Biology, 57(2):359–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/2/359.

Robbins:1970:CMP

[RPK70] Elliott Robbins, Thoru Pederson, and Paul Klein. Comparison of mitotic phenomena and effects induced by hypertonic solutions in HeLa cells. Journal of Cell Biology, 44(2): 400–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/2/400.

Ringborg:1971:NDR

[RR71] U. Ringborg and L. Rydlander. Nucleolar-derived ribonucleic acid in chromosomes, nuclear sap, and cytoplasm of chironomus tentans salivary gland cells. Journal of Cell Biology, 51(2):355–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/2/355.

Rugstad:1970:MBC

[RRYT70] Hans E. Rugstad, Stephen H. Robinson, Claudine Yannoni, and Armen H. Tashjian. Metabolism of bilirubin by a clonal strain of rat hepatoma cells. Journal of Cell Biology, 47(3): 703–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/3/703.

Redman:1970:PBD

[RS70] Robert S. Redman and Leo M. Sreebny. Proliferative behavior of differentiating cells in the developing rat parotid gland. Journal of Cell Biology, 46(1):81–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/1/81.

Rew:1974:ELA

[RS74a] Robert S. Rew and Howard J. Saz. Enzyme localization in the anaerobic mitochondria of ascaris lumbricoides. Journal of Cell Biology, 63(1):125–??, October 1974. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/1/125.

Roti:1974:EBG

[RS74b] L. W. Roti Roti and A. R. Stevens. Effect of 5bromodeoxyuridine on growth, encystment, and excystment of Acanthamoeba Castellanii. Journal of Cell Biology, 61 (1):233–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/1/233.

Rawlins:1970:RPN

[RU70] Frank A. Rawlins and Betty G. Uzman. Retardation of peripheral nerve myelination in mice treated with inhibitors of cholesterol biosynthesis. Journal of Cell Biology, 46(3): 505–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/505.

Rubin:1971:PPD

[Rub71] H. Rubin. pH and population density in the regulation of animal cell multiplication. Journal of Cell Biology, 51(3): 686–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/686.

Rubin:1972:UIR

[Rub72] Walter Rubin. An unusual intimate relationship between endocrine cells and other types of epithelial cells in the human stomach. Journal of Cell Biology, 52(1):219–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
1/219.

Rudick:1971:RBN

[Rud71] V. L. Rudick. Relationships between nucleic acid synthetic patterns and encystment in aging unagitated cultures of Acanthamoeba Castellanii. Journal of Cell Biology, 49(2): 498–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/498.

Rawlins:1972:FSL

[RVHWU72] F. A. Rawlins, G. M. Villegas, E. T. Hedley-Whyte, and B. G. Uzman. Fine structural localization of cholesterol-1,2-^{3}H in degenerating and regenerating mouse sciatic nerve. Journal of Cell Biology, 52(3):615–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/3/615.

Renger:1970:KDA

[RW70a] Hartmut C. Renger and David R. Wolstenholme. Kinetoplast deoxyribonucleic acid of the hemoflagellate trypanosoma lewisi. Journal of Cell Biology, 47(3):689–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
3/689.

Rosenbaum:1970:UBA

[RW70b] Robert M. Rosenbaum and Murray Wittner. Ultrastructure of bacterized and axenic trophozoites of entamoeba histolytica with particular reference to helical bodies. Journal of Cell Biology, 45(2):367–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/2/367.

Rannestad:1971:SMO

[RW71a] John Rannestad and Norman E. Williams. The synthesis of microtubule and other proteins of the oral apparatus in Tetrahymena Pyriformis. Journal of Cell Biology, 50(3):709– ??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/3/709.

Renger:1971:KOS

[RW71b] Harmut C. Renger and David R. Wolstenholme. Kinetoplast and other satellite DNAs of kinetoplastic and dyskinetoplastic strains of Trypanosoma. Journal of Cell Biology, 50(2): 533–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/2/533.

Renger:1972:FSK

[RW72] Hartmut C. Renger and David R. Wolstenholme. The form and structure of kinetoplast DNA of crithidia. Journal of Cell Biology, 54(2):346–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/2/346.

Rosenthal:1973:IFA

[RW73] Miriam D. Rosenthal and Joseph B. Warshaw. Interaction of fatty acid and glucose oxidation by cultured heart cells. Journal of Cell Biology, 58(2):332–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/2/332.

Recher:1970:CRS

[RWB70] L. Recher, J. Whitescarver, and L. Briggs. A cytochemical and radioautographic study of human tissue culture cell nucleoli. Journal of Cell Biology, 45(3):479–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/479.

Rustad:1971:NCR

[RYR71] Ronald C. Rustad, Shuhei Yuyama, and Lynne C. Rustad. Nuclear-cytoplasmic relations in the mitosis of sea urchin eggs. Journal of Cell Biology, 49(3):906–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
3/906.

Sachs:1970:CSA

[SA70] Martin I. Sachs and Everett Anderson. A cytological study of artificial parthenogenesis in the sea urchin Arbacia Punctulata. Journal of Cell Biology, 47(1):140–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/140.

Stevens:1971:RMH

[SA71] Ronald H. Stevens and Harold Amos. RNA metabolism in HeLa cells at reduced temperature. Journal of Cell Biology, 50(3):818–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/3/818.

Stevens:1972:RMH

[SA72] Ronald H. Stevens and Harold Amos. RNA metabolism in HeLa cells at reduced temperature. Journal of Cell Biology, 54(1):1–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/1/1.

Storrie:1973:EMG

[SA73a] Brian Storrie and Giuseppe Attardi. Expression of the mitochondrial genome in HeLa cells. Journal of Cell Biology, 56 (3):819–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/819.

Storrie:1973:MMF

[SA73b] Brian Storrie and Giuseppe Attardi. Mode of mitochondrial formation in HeLa cells. Journal of Cell Biology, 56 (3):833–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/833.

Saari:1974:ABR

[Saa74] John C. Saari. The accessibility of bovine rhodopsin in photoreceptor membranes. Journal of Cell Biology, 63(2):480– ??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/2/480.

Sutherland:1973:UOG

[SAE73] Alexander Sutherland, Gregory A. Antipa, and Charles F. Ehret. Ultracytochemical observations of glycogen synthesis and storage during infradian growth of Tetrahymena Pyriformis. Journal of Cell Biology, 58(1):240–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
1/240.

Sabatini:1970:CPN

[SB70a] D. D. Sabatini and G. Blobel. Controlled proteolysis of nascent polypeptides in rat liver cell fractions. Journal of Cell Biology, 45(1):146–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/1/146.

Silverblatt:1970:GJO

[SB70b] Fredric J. Silverblatt and Ruth Ellen Bulger. Gap junctions occur in vertebrate renal proximal tubule cells. Journal of Cell Biology, 47(2):513–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/2/513.

Stein:1972:SAC

[SB72] G. S. Stein and T. W. Borun. The synthesis of acidic chromosomal proteins during the cell cycle of HeLa s-3 cells. Journal of Cell Biology, 52(2):292–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/2/292.

Schlaepfer:1973:ECI

[SB73a] W. W. Schlaepfer and R. P. Bunge. Effects of calcium ion concentration on the degeneration of amputated axons in tissue culture. Journal of Cell Biology, 59(2):456–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
2/456.

Szamier:1973:RDA

[SB73b] R. Bruce Szamier and Michael V. L. Bennett. Rapid degeneration of ampullary electroreceptor organs after denervation. Journal of Cell Biology, 56(2):466–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
2/466.

Slaby:1974:CRC

[SB74] Frank Slaby and Carolyn Brown. Changes in the ribosome content, principal microsomal protein composition, and secretory character of mammary epithelial rough endoplasmic reticulum during differentiation. Journal of Cell Biology, 61 (3):613–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/3/613.

Stasny:1974:EMM

[SBKH74] J. T. Stasny, R. C. Burns, B. D. Korant, and R. W. F. Hardy. Electron microscopy of the Mo-Fe protein from azotobacter nitrogenase. Journal of Cell Biology, 60(1):311–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
1/311.

Shepro:1970:AEC

[SBRC70] David Shepro, Frank A. Belamarich, Lois Robblee, and Francis C. Chao. Antimotility effect of cytochalasin B observed in mammalian clot retraction. Journal of Cell Biology, 47(2): 544–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/2/544.

Shirahama:1972:RMV

[SC72a] Tsuranobu Shirahama and Alan S. Cohen. The role of mucopolysaccharides in vesicle architecture and endothelial transport. Journal of Cell Biology, 52(1):198–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
1/198.

Steinman:1972:IPH

[SC72b] Ralph M. Steinman and Zanvil A. Cohn. The interaction of particulate horseradish peroxidase (hrp)-anti hrp immune complexes with mouse peritoneal macrophages in vitro. Journal of Cell Biology, 55(3):616–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/616.

Steinman:1972:ISH

[SC72c] Ralph M. Steinman and Zanvil A. Cohn. The interaction of soluble horseradish peroxidase with mouse peritoneal macrophages in vitro. Journal of Cell Biology, 55(1):186– ??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/1/186.

Slavkin:1972:EMI

[SCB72a] Harold C. Slavkin, Richard Croissant, and Pablo Bringas. Epithelial-Mesenchymal interactions during odontogenesis. Journal of Cell Biology, 53(3):841–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/3/841.

Staehelin:1972:LPM

[SCB72b] L. Andrew Staehelin, Francis J. Chlapowski, and Mary A. Bonneville. Lumenal plasma membrane of the urinary bladder. Journal of Cell Biology, 53(1):73–??, April 1972. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/53/1/73.

Skosey:1973:ECB

[SCDS73] John L. Skosey, Donald Chow, Evelyn Damgaard, and Leif B. Sorensen. Effect of cytochalasin B on response of human polymorphonuclear leukocytes to zymosan. Journal of Cell Biology, 57(1):237–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/1/237.

Scheer:1970:UNE

[Sch70] Ulrich Scheer. The ultrastructure of the nuclear envelope of amphibian oocytes: a reinvestigation. Journal of Cell Biology, 45(2):445–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/2/445.

Schroeder:1972:CR

[Sch72] Thomas E. Schroeder. The contractile ring. Journal of Cell Biology, 53(2):419–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/53/2/419.

Schneider:1973:RSC

[Sch73] Lawrence K. Schneider. RNA synthesis in chromosome connectives. Journal of Cell Biology, 56(2):608–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
2/608.

Seeman:1973:SMH

[SCI73] P. Seeman, D. Cheng, and G. H. Iles. Structure of membrane holes in osmotic and saponin hemolysis. Journal of Cell Biology, 56(2):519–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/2/519.

Sederoff:1973:RSE

[SCPH73] R. Sederoff, R. Clynes, M. Poncz, and S. Hachtel. RNA synthesis by exogenous RNA polymerase on cytological preparations of chromosomes. Journal of Cell Biology, 57(2):538–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/2/538.

Scott:1970:FDF

[SD70a] Dougald G. Scott and Charles W. Daniel. Filaments in the division furrow of mouse mammary cells. Journal of Cell Biology, 45(2):461–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/45/2/461.

Soeiro:1970:CBH

[SD70b] Ruy Soeiro and James E. Darnell. A comparison between heterogeneous nuclear RNA and polysomal messenger RNA in HeLa cells by RNA-DNA hybridization. Journal of Cell Biology, 44(3):467–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/3/467.

Satir:1971:NAT

[SD71a] Birgit Satir and Ellen Roter Dirksen. Nucleolar aging in Tetrahymena during the cultural growth cycle. Journal of Cell Biology, 48(1):143–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/48/1/143.

Simpson:1971:VPR

[SD71b] Peter A. Simpson and Allan D. Dingle. Variable periodicity in the rhizoplast of naegleria flagellates. Journal of Cell Biology, 51(1):323–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/1/323.

Socher:1971:AT

[SD71c] S. H. Socher and D. Davidson. 5-aminouracil treatment. Journal of Cell Biology, 48(2):248–??, February 1971. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/48/2/

Staprans:1974:MPD

[SD74] Ilona Staprans and Ellen Roter Dirksen. Microtubule protein during ciliogenesis in the mouse oviduct. Journal of Cell Biology, 62(1):164–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/1/164.

Skosey:1974:MZI

[SDCS74] John L. Skosey, Evelyn Damgaard, Donald Chow, and Leif B. Sorensen. Modification of zymosan-induced release of lysosomal enzymes from human polymorphonuclear leukocytes by cytochalasin B. Journal of Cell Biology, 62(3):625– ??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/3/625.

Sturgess:1972:MGA

[SdlI72] J. M. Sturgess and F. A. de la Iglesia. Morphometry of the Golgi apparatus in developing liver. Journal of Cell Biology, 55(2):524–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/55/2/524.

Somlyo:1971:SRT

[SDSN71] Andrew P. Somlyo, Carrick E. Devine, Avril V. Somlyo, and Stanley R. North. Sarcoplasmic reticulum and the temperature-dependent contraction of smooth muscle in calcium-free solutions. Journal of Cell Biology, 51(3):722– ??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/722.

Searle:1972:OUC

[Sea72] Roger F. Searle. Observations on the unusual cytoplasmic lacunae of the porcine hepatocyte. Journal of Cell Biology, 54 (2):434–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/2/434.

Singer:1974:CTH

[SF74a] Robert M. Singer and William H. Fishman. Characterization of two HeLa sublines: Tcrc-1 produces Regan isoenzyme and tcrc-2, non-Regan isoenzyme. Journal of Cell Biology, 60 (3):777–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/3/777.

Stadleb:1974:CCB

[SF74b] Joachim Stadleb and Werner W. Franke. Characterization of the colchicine binding of membrane fractions from rat and mouse liver. Journal of Cell Biology, 60(1):297–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
1/297.

Schlaeger:1974:EMS

[SFK74] Dorothy A. Schlaeger, Morton S. Fuchs, and Suk-Hee Kang. Ecdysone-mediated stimulation of dopa decarboxylase activity and its relationship to ovarian development in aedes aegypti. Journal of Cell Biology, 61(2):454–??, May 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/61/2/
454.

Stewart:1972:CDS

[SFMD72] Kenneth D. Stewart, Gary L. Floyd, Karl R. Mattox, and Melanie E. Davis. Cytochemical demonstration of a single peroxisome in a filamentous green alga. Journal of Cell Biology, 54(2):431–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/2/431.

Seegmiller:1971:NCM

[SFS71] R. Seegmiller, F. C. Fraser, and H. Sheldon. A new chondrodystrophic mutant in mice. Journal of Cell Biology, 48 (3):580–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/3/580.

Satir:1970:CJL

[SG70a] P. Satir and N. B. Gilula. The cell junction in a lamellibranch Gill ciliated epithelium. Journal of Cell Biology, 47(2):468– ??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/2/468.

Stuve:1970:RMH

[SG70b] J. Stuve and P. Galle. Role of mitochondria in the handling of gold by the kidney. Journal of Cell Biology, 44(3): 667–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/3/667.

Sawicki:1971:DCA

[SG71] Stanley G. Sawicki and Gabriel C. Godman. On the differential cytotoxicity of actinomycin D. Journal of Cell Biology, 50(3):746–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/3/746.

Sawicki:1972:RTA

[SG72] Stanley G. Sawicki and Gabriel C. Godman. On the recovery of transcription after inhibition by actinomycin D. Journal of Cell Biology, 55(2):299–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/2/299.

Summers:1973:ETD

[SG73] Keith E. Summers and I. R. Gibbons. Effects of trypsin digestion on flagellar structures and their relationship to motility. Journal of Cell Biology, 58(3):618–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
3/618.

Sellers:1974:RTA

[SG74] Larry Sellers and Daryl Granner. Regulation of tyrosine aminotransferase activity in two liver-derived permanent cell lines. Journal of Cell Biology, 60(2):337–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
2/337.

Schmitt:1973:MBD

[SGL73] H. Schmitt, H. Grossfeld, and U. Z. Littauer. Mitochondrial biogenesis during differentiation of artemia salina cysts. Journal of Cell Biology, 58(3):643–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/3/643.

Seil:1970:CGC

[SH70] Fredrick J. Seil and Robert M. Herndon. Cerebellar granule cells in vitro. Journal of Cell Biology, 45(2):212–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/212.

Spooner:1971:EDC

[SH71] Brian S. Spooner and S. Robert Hilfer. The expression of differentiation by chick embryo thyroid in cell culture. Journal of Cell Biology, 48(2):225–??, February 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/2/225.

Schiaffino:1972:ADG

[SH72a] S. Schiaffino and Věra Hanzlíková. Autophagic degradation of glycogen in skeletal muscles of the newborn rat. Journal of Cell Biology, 52(1):41–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/1/41.

Swinton:1972:VSL

[SH72b] David C. Swinton and Philip C. Hanawalt. In vivo specific labeling of chlamydomonas chloroplast DNA. Journal of Cell Biology, 54(3):592–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/3/592.

Stavis:1973:PCR

[SH73] Robert L. Stavis and Rona Hirschberg. Phototaxis in chlamydomonas reinhardtii. Journal of Cell Biology, 59(2):367– ??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/2/367.

Smith:1974:VLS

[SH74] J. R. Smith and L. Hayflick. Variation in the life-span of clones derived from human diploid cell strains. Journal of Cell Biology, 62(1):48–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/1/48.

Shay:1972:EMS

[Sha72] Jerry W. Shay. Electron microscope studies of spermatozoa of rhynchosciara Sp. Journal of Cell Biology, 54(3):598– ??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/3/598.

Schubert:1974:CUM

[SHDH74] David Schubert, A. John Harris, Carrick E. Devine, and Stephen Heinemann. Characterization of a unique muscle cell line. Journal of Cell Biology, 61(2):398–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/398.

Sheridan:1970:LRJ

[She70] Judson D. Sheridan. Low-resistance junctions between cancer cells in various solid tumors. Journal of Cell Biology, 45 (1):91–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/1/91.

Shea:1971:LSS

[She71a] Stephen M. Shea. Lanthanum staining of the surface coat of cells. Journal of Cell Biology, 51(3):611–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/611.

Sheridan:1971:ECB

[She71b] Judson D. Sheridan. Electrical coupling between fat cells in newt fat body and mouse Brown fat. Journal of Cell Biology, 50(3):795–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/3/795.

Sherman:1973:UAG

[She73] R. G. Sherman. Unique arrangement of glial membranes between adjacent neuromuscular synapses in a spider muscle. Journal of Cell Biology, 59(1):234–??, October 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/59/1/

Sreter:1972:SPE

[SHGH72] F. Sreter, S. Holtzer, J. Gergely, and H. Holtzer. Some properties of embryonic myosin. Journal of Cell Biology, 55(3): 586–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/3/586.

Simmons:1973:NEA

[SHH73] T. Simmons, P. Heywood, and L. Hodge. Nuclear envelopeassociated resumption of RNA synthesis in late mitosis of HeLa cells. Journal of Cell Biology, 59(1):150–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/150.

Shimada:1971:EMO

[Shi71] Yutaka Shimada. Electron microscope observations on the fusion of chick myoblasts in vitro. Journal of Cell Biology, 48(1):128–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/1/128.

Sprague:1974:PLC

[SHM74] C. A. Sprague, H. Hoehn, and G. M. Martin. Ploidy of living clones of human somatic cells determined by mensuration at metaphase. Journal of Cell Biology, 60(3):781–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
3/781.

Scharff:1972:RPS

[SHNB72] R. Scharff, R. W. Hendler, N. Nanninga, and A. H. Burgess. Respiration and protein synthesis in Escherichia coli membrane-envelope fragments. Journal of Cell Biology, 53(1):1–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/1.

Schiaffino:1970:RBS

[SHP70] S. Schiaffino, Věra Hanzlíková, and Sandra Pierobon. Relations between structure and function in rat skeletal muscle fibers. Journal of Cell Biology, 47(1):107–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/107.

Simons:1970:CSC

[Sim70] J. W. I. M. Simons. Characterization of somatic cells by determination of their volumes with a Coulter counter. Journal of Cell Biology, 46(3):610–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/3/610.

Steward:1974:LCC

[SIS74] F. C. Steward, H. W. Israel, and M. M. Salpeter. The labeling of cultured cells of Acer with [^{14}C]Proline and its significance. Journal of Cell Biology, 60(3):695–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/3/695.

Sage:1971:CDC

[SJ71] Jean A. Sage and Ralph A. Jersild. Comparative distribution of carbohydrates and lipid droplets in the Golgi apparatus of intestinal absorptive cells. Journal of Cell Biology, 51(1): 333–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/333.

Strum:1972:RDH

[SJ72] Judy M. Strum and Alain F. Junod. Radioautographic demonstration of 5-hydroxytryptamine-^{3}H uptake by pulmonary endothelial cells. Journal of Cell Biology, 54(3):456– ??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/3/456.

Smith:1970:OSA

[SJB70] David S. Smith, U. Järlfors, and R. Beránek. The organization of synaptic axoplasm in the lamprey (Petromyzon Marinus) central nervous system. Journal of Cell Biology, 46(2): 199–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/2/199.

Smith-Johannsen:1972:ECC

[SJG72] Heidi Smith-Johannsen and Sarah P. Gibbs. Effects of chloramphenicol on chloroplast and mitochondrial ultrastructure in ochromonas danica. Journal of Cell Biology, 52(3): 598–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/3/598.

Stephens:1970:SPH

[SK70a] R. E. Stephens and R. E. Kane. Some properties of hyalin. Journal of Cell Biology, 44(3):611–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/44/3/611.

Strum:1970:CLE

[SK70b] Judy M. Strum and Morris J. Karnovsky. Cytochemical localization of endogenous peroxidase in thyroid follicular cells. Journal of Cell Biology, 44(3):655–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/44/3/655.

Schneeberger:1971:IIF

[SK71] Eveline E. Schneeberger and Morris J. Karnovsky. The influence of intravascular fluid volume on the permeability of newborn and adult mouse lungs to ultrastructural protein tracers. Journal of Cell Biology, 49(2):319–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/49/
2/319.

Shelton:1970:CPA

[SKMH70] Emma Shelton, Edward L. Kuff, Elizabeth S. Maxwell, and J. T. Harrington. Cytoplasmic particles and aminoacyl transferase I activity. Journal of Cell Biology, 45(1):1–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
1/1.

Stephens:1970:SEP

[SL70] R. E. Stephens and E. E. Levine. Some enzymatic properties of axonemes from the cilia of pecten irradians. Journal of Cell Biology, 46(2):416–??, August 1970. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/46/2/416.

Shephard:1972:BIA

[SL72a] David C. Shephard and Wendy B. Levin. Biosynthesis in isolated acetabularia chloroplasts. Journal of Cell Biology, 54 (2):279–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/54/2/279.

Sotelo:1972:SMJ

[SL72b] Constantino Sotelo and Rodolfo Llinás. Specialized membrane junctions between neurons in the vertebrate cerebellar cortex. Journal of Cell Biology, 53(2):271–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
2/271.

Schmidt:1973:CNP

[SL73a] Susan Y. Schmidt and Richard N. Lolley. Cyclic-nucleotide phosphodiesterase. Journal of Cell Biology, 57(1):117–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/1/117.

Stetten:1973:CSF

[SL73b] Gail Stetten and Seymour Lederberg. Colcemid sensitivity of fission yeast. Journal of Cell Biology, 56(1):259–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
1/259.

Schreiber:1970:DAS

[SLWZ70] Gerhard Schreiber, Rainer Lesch, Ursula Weinssen, and Josef Zähringer. The distribution of albumin synthesis throughout the liver lobule. Journal of Cell Biology, 47(1):285–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
47/1/285.

Shaler:1971:MCA

[SM71] Robert C. Shaler and Richard L. McCarl. The mechanism of cortisol action in cultured rat heart cells. Journal of Cell Biology, 49(1):205–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/1/205.

Solari:1973:SCR

[SM73] Alberto J. Solari and Montrose J. Moses. The structure of the central region in the synaptonemal complexes of hamster and Cricket spermatocytes. Journal of Cell Biology, 56(1): 145–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/1/145.

Scheid:1974:MCD

[SM74a] C. R. Scheid and I. A. Macchi. Monolayer cultures derived from neonatal hamster pancreas. Journal of Cell Biology, 60 (3):685–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/3/685.

Skerrow:1974:CCI

[SM74b] Christine J. Skerrow and A. Gedeon Matoltsy. Chemical characterization of isolated epidermal desmosomes. Journal of Cell Biology, 63(2):524–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/524.

Skerrow:1974:IED

[SM74c] Christine J. Skerrow and A. Gedeon Matoltsy. Isolation of epidermal desmosomes. Journal of Cell Biology, 63(2):515– ??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/2/515.

Sachs:1973:TSC

[SMD73] Howard G. Sachs, Terence F. McDonald, and Robert L. De-Haan. Tetrodotoxin sensitivity of cultured embryonic heart cells depends on cell interactions. Journal of Cell Biology, 56(1):255–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/1/255.

Stahn:1970:NMP

[SMH70] R. Stahn, K.-P. Maier, and K. Hannig. A new method for the preparation of rat liver lysosomes. Journal of Cell Biology, 46(3):576–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/3/576.

Spicer:1972:JCA

[SMS72] S. S. Spicer, B. J. Martin, and J. V. Simson. The junctional complex associated body of human eccrine sweat gland. Journal of Cell Biology, 53(2):582–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/53/2/582.

Soslau:1971:EEB

[SN71] Gerald Soslau and Margit M. K. Nass. Effects of ethidium bromide on the cytochrome content and ultrastructure of L cell mitochondria. Journal of Cell Biology, 51(2):514– ??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/2/514.

Sanes:1972:IDS

[SO72a] Joshua R. Sanes and Lawrence M. Okun. Induction of DNA synthesis in cultured neurons by ultraviolet light or methyl methane sulfonate. Journal of Cell Biology, 53(2):587–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/2/587.

Stockdale:1972:RDS

[SO72b] Frank E. Stockdale and Michael C. O’neill. Repair DNA synthesis in differentiated embryonic muscle cells. Journal of Cell Biology, 52(3):589–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/3/589.

Simionescu:1971:DGP

[SP71a] Nicolae Simionescu and George E. Palade. Dextrans and glycogens as particulate tracers for studying capillary permeability. Journal of Cell Biology, 50(3):616–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/616.

Stevens:1971:RNL

[SP71b] A. R. Stevens and D. M. Prescott. Reformation of nucleoluslike bodies in the absence of postmitotic RNA synthesis. Journal of Cell Biology, 48(3):443–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/3/443.

Stevens:1973:RST

[SP73] A. R. Stevens and P. F. Pachler. RNA synthesis and turnover during density-inhibited growth and encystment of Acanthamoeba Castellanii. Journal of Cell Biology, 57(2):525–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/2/525.

Schabtach:1974:MSH

[SP74] Eric Schabtach and Terry A. Parkening. A method for sequential high resolution light and electron microscopy of selected areas of the same material. Journal of Cell Biology, 61 (1):261–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/1/261.

Spencer:1973:ETS

[SPAR73] Peter S. Spencer, Edith R. Peterson, Ricardo Madrid A., and Cedric S. Raine. Effects of thallium salts on neuronal mitochondria in organotypic cord-ganglia-muscle combination cultures. Journal of Cell Biology, 58(1):79–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/58/1/79.

Spitzer:1970:LRC

[Spi70] Nicholas C. Spitzer. Low resistance connections between cells in the developing anther of the Lily. Journal of Cell Biology, 45(3):565–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/3/565.

Sawicki:1974:IAC

[SRA74] Wojciech Sawicki, Jan Rowínski, and Jan Abramczuk. Image analysis of chromatin in cells of preimplantation mouse embryos. Journal of Cell Biology, 63(1):227–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/227.

Smith:1973:FES

[SRS73] Una Smith, J. W. Ryan, and D. S. Smith. Freeze-etch studies of the plasma membrane of pulmonary endothelial cells. Journal of Cell Biology, 56(2):492–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
2/492.

Sengel:1970:GOT

[SS70a] A. Sengel and P. Stoebner. Golgi origin of tubular inclusions in endothelial cells. Journal of Cell Biology, 44(1): 223–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/44/1/223.

Senger:1970:UMB

[SS70b] P. L. Senger and R. G. Saacke. Unusual mitochondria of the bovine oocyte. Journal of Cell Biology, 46(2):405–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
2/405.

Salpeter:1971:REM

[SS71] M. M. Salpeter and E. E. Salpeter. Resolution in electron microscope radioautography. Journal of Cell Biology, 50(2): 324–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/50/2/324.

Sacktor:1972:DCM

[SS72a] Bertram Sacktor and Yoshio Shimada. Degenerative changes in the mitochondria of flight muscle from aging blowflies. Journal of Cell Biology, 52(2):465–??, February 1972. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/52/2/
465.

Simson:1972:FPM

[SS72b] J. V. Simson and S. S. Spicer. Ferritin particles in macrophages and in associated mast cells. Journal of Cell Biology, 52(3):536–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/3/536.

Stambrook:1972:ICR

[SS72c] Peter J. Stambrook and Jesse E. Sisken. Induced changes in the rates of uridine-^{3}H uptake and incorporation during the G_{1} and S periods of synchronized Chinese hamster cells. Journal of Cell Biology, 52(3):514–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/3/514.

Stoner:1973:ANIa

[SS73a] Clinton D. Stoner and Howard D. Sirak. Adenine nucleotideinduced contraction of the inner mitochondrial membrane. Journal of Cell Biology, 56(1):51–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/1/51.

Stoner:1973:ANIb

[SS73b] Clinton D. Stoner and Howard D. Sirak. Adenine nucleotideinduced contraction of the inner mitochondrial membrane. Journal of Cell Biology, 56(1):65–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/56/1/65.

Sattler:1974:CMD

[SS74a] Carol A. Sattler and L. Andrew Staehelin. Ciliary membrane differentiations in Tetrahymena Pyriformis. Journal of Cell Biology, 62(2):473–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/2/473.

Sjolund:1974:FFS

[SS74b] Richard D. Sjolund and Donna D. Smith. Freeze-fracture studies of photosynthetically deficient “supergranal” chloroplasts in tissue cultures containing virus-like particles. Journal of Cell Biology, 60(1):285–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/1/285.

Seljelid:1973:EPL

[SSC73] Rolf Seljelid, Samuel C. Silverstein, and Zanvil A. Cohn. The effect of poly-L-lysine on the uptake of reovirus doublestranded RNA in macrophages in vitro. Journal of Cell Biology, 57(2):484–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/2/484.

Steinman:1974:PF

[SSC74] Ralph M. Steinman, Jonathan M. Silver, and Zanvil A. Cohn. Pinocytosis in fibroblasts. Journal of Cell Biology, 63(3): 949–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/949.

Somlyo:1974:EME

[SSD^{+}74] A. P. Somlyo, A. V. Somlyo, C. E. Devine, P. D. Peters, and T. A. Hall. Electron microscopy and electron probe analysis of mitochondrial cation accumulation in smooth muscle. Journal of Cell Biology, 61(3):723–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/3/723.

Saito:1974:MJI

[SSF74] Akitsugu Saito, Murray Smigel, and Sidney Fleischer. Membrane junctions in the intermembrane space of mitochondria from mammalian tissues. Journal of Cell Biology, 60(3): 653–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/3/653.

Seyberth:1973:CAM

[SSGJH73] H. W. Seyberth, H. Schmidt-Gayk, K. H. Jakobs, and E. Hackenthal. Cyclic adenosine monophosphate in phagocytizing granulocytes and alveolar macrophages. Journal of Cell Biology, 57(2):567–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/2/567.

Siegers:1974:DRD

[SSHS74] M. P. Siegers, J. C. Schaer, H. Hirsiger, and R. Schindler. Determination of rates of DNA synthesis in cultured mammalian cell populations. Journal of Cell Biology, 62(2): 305–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/305.

Smith-Sonneborn:1974:CDS

[SSK74] Joan Smith-Sonneborn and Michael Klass. Changes in the DNA synthesis pattern of paramecium with increased clonal age and interfission time. Journal of Cell Biology, 61(3): 591–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/3/591.

Simionescu:1972:PIC

[SSP72] Nicolae Simionescu, Maia Simionescu, and George E. Palade. Permeability of intestinal capillaries. Journal of Cell Biology, 53(2):365–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/365.

Simionescu:1973:PMC

[SSP73] Nicolae Simionescu, Maia Simionescu, and George E. Palade. Permeability of muscle capillaries to exogenous myoglobin. Journal of Cell Biology, 57(2):424–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/2/424.

Simionescu:1974:MDE

[SSP74] Maia Simionescu, Nicolae Simionescu, and George E. Palade. Morphometric data on the endothelium of blood capillaries. Journal of Cell Biology, 60(1):128–??, January 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/60/1/
128.

Satir:1973:MFM

[SSS73] Birgit Satir, Caroline Schooley, and Peter Satir. Membrane fusion in a model system. Journal of Cell Biology, 56(1): 153–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/1/153.

Stein:1974:CII

[SSS74] O. Stein, L. Sanger, and Y. Stein. Colchicine-induced inhibition of lipoprotein and protein secretion into the serum and lack of interference with secretion of biliary phospholipids and cholesterol by rat liver in vivo. Journal of Cell Biology, 62(1):90–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/1/90.

Solari:1970:TDR

[ST70] A. J. Solari and Laura L. Tres. The three-dimensional reconstruction of the XY chromosomal pair in human spermatocytes. Journal of Cell Biology, 45(1):43–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/45/1/43.

Stephens:1973:TAM

[Ste73] R. E. Stephens. A thermodynamic analysis of mitotic spindle equilibrium at active metaphase. Journal of Cell Biology, 57 (1):133–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/133.

Steck:1974:OPH

[Ste74] Theodore L. Steck. The organization of proteins in the human red blood cell membrane. Journal of Cell Biology, 62(1):1–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
62/1/1.

Stirling:1972:RLS

[Sti72] Charles E. Stirling. Radioautographic localization of sodium pump sites in rabbit intestine. Journal of Cell Biology, 53 (3):704–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/3/704.

Slavkin:1974:LHH

[STM^{+}74] Harold C. Slavkin, Gary N. Trump, Virginia Mansour, Pamela Matosian, and William G. Mino. Localization of h-2 histocompatibility alloantigens on mouse embryonic tooth epithelial and Mesenchymal cell surfaces. Journal of Cell Biology, 60(3):795–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/60/3/795.

Storrie:1973:ADA

[Sto73a] Brian Storrie. Antagonism by dibutyryl adenosine cyclic 3’,5’-monophosphate and testosterone of cell rounding reactions. Journal of Cell Biology, 59(2):471–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
2/471.

Stossel:1973:QSP

[Sto73b] Thomas P. Stossel. Quantitative studies of phagocytosis. Journal of Cell Biology, 58(2):346–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/2/346.

Storrie:1974:EDA

[Sto74] Brian Storrie. Effect of dibutyryl adenosine cyclic 3’,5’- monophosphate and testololactone on concanavalin a binding and cell Killing. Journal of Cell Biology, 62(1):247–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
1/247.

Straus:1971:CAC

[Str71a] Werner Straus. Comparative analysis of the concentration of injected horseradish peroxidase in cytoplasmic granules of the kidney cortex, in the blood, urine, and liver. Journal of Cell Biology, 48(3):620–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/48/3/620.

Strum:1971:UPP

[Str71b] Judy M. Strum. Unusual peroxidase-positive granules in the developing rat submaxillary gland. Journal of Cell Biology, 51(2):575–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/51/2/575.

Strauss:1972:ARH

[Str72] Phyllis R. Strauss. Acriflavin resistance in the hemoflagellate, Leishmania tarentolae. Journal of Cell Biology, 53(2):312–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/2/312.

Strauss:1974:ESM

[Str74] Phyllis R. Strauss. Effects of serum on membrane transport. Journal of Cell Biology, 60(3):571–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/3/571.

Stuart:1971:ERK

[Stu71] K. D. Stuart. Evidence for the retention of kinetoplast DNA in an acriflavine-induced dyskinetoplastic strain of trypanosoma brucei which replicates the altered central element of the kinetoplast. Journal of Cell Biology, 49(1):189–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/1/189.

Schenkein:1970:ISS

[SU70] Isaac Schenkein and Jonathan W. Uhr. Immunoglobulin synthesis and secretion. Journal of Cell Biology, 46(1):42–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
46/1/42.

Summers:1974:AIS

[Sum74] Keith Summers. ATP-induced sliding of microtubules in bull sperm flagella. Journal of Cell Biology, 60(1):321–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
60/1/321.

Spielholz:1973:LPC

[SV73] Neil I. Spielholz and William G. Van der Kloot. Localization and properties of the cholinesterase in crustacean muscle. Journal of Cell Biology, 59(2):407–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
2/407.

Speth:1970:MET

[SW70] Volker Speth and Frank Wunderlich. The macronuclear envelope of Tetrahymena Pyriformis Gl in different physiological states. Journal of Cell Biology, 47(3):772–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
3/772.

Saunders:1971:RMC

[SW71] G. C. Saunders and M. Wilder. Repetitive maturation cycles in a cultured mouse myeloma. Journal of Cell Biology, 51(1): 344–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/344.

Steinberg:1972:DMT

[SW72a] Malcolm S. Steinberg and Lawrence L. Wiseman. Do morphogenetic tissue rearrangements require active cell movements? Journal of Cell Biology, 55(3):606–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
3/606.

Stewart:1972:ELB

[SW72b] James R. Stewart and Robert A. Weisman. Exocytosis of latex beads during the encystment of Acanthamoeba. Journal of Cell Biology, 52(1):117–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/1/117.

Schuetz:1974:SIP

[SWD74] Allen W. Schuetz, Robin A. Wallace, and James N. Dumont. Steroid inhibition of protein incorporation by isolated amphibian oocytes. Journal of Cell Biology, 61(1):26–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
1/26.

Shortman:1971:SDC

[SWJ^{+}71] Ken Shortman, Neil Williams, Heather Jackson, Pamela Russell, Pauline Byrt, and E. Diener. The separation of different cell classes from lymphoid organs. Journal of Cell Biology, 48 (3):566–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/48/3/566.

Spooner:1970:DDV

[SWR70] Brian S. Spooner, Bernt T. Walther, and William J. Rutter. The development of the dorsal and ventral mammalian pancreas in vivo and in vitro. Journal of Cell Biology, 47(1): 235–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/1/235.

Spray:1974:PCA

[SWS74] T. L. Spray, R. A. Waugh, and J. R. Sommer. Peripheral couplings in adult vertebrate skeletal muscle. Journal of Cell Biology, 62(1):223–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/1/223.

Strum:1971:AFE

[SWSK71] Judy M. Strum, Janice Wicken, John R. Stanbury, and Morris J. Karnovsky. Appearance and function of endogenous peroxidase in fetal rat thyroid. Journal of Cell Biology, 51(1): 162–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/162.

Shahrabadi:1971:MSI

[SY71] M. S. Shahrabadi and T. Yamamoto. A method for staining intracellular antigens in thin sections with ferritin-labeled antibody. Journal of Cell Biology, 50(1):246–??, July 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/50/1/
246.

Sawai:1974:WSP

[SY74] Tsuyoshi Sawai and Mitsuki Yoneda. Wave of stiffness propagating along the surface of the newt egg during cleavage. Journal of Cell Biology, 60(1):1–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/1/1.

Spooner:1971:MCL

[SYW71] Brian S. Spooner, Kenneth M. Yamada, and Norman K. Wessells. Microfilaments and cell locomotion. Journal of Cell Biology, 49(3):595–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/595.

Szollosi:1970:CCF

[Szo70] Daniel Szollosi. Cortical cytoplasmic filaments of cleaving eggs: a structural element corresponding to the contractile ring. Journal of Cell Biology, 44(1):192–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
1/192.

Szubinska:1971:NMF

[Szu71] Barbara Szubinska. ”new membrane” formation in Amoeba Proteus upon injury of individual cells. Journal of Cell Biology, 49(3):747–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/747.

Tani:1970:SMB

[TA70] Eiichi Tani and Toshio Ametani. Substructure of microtubules in brain nerve cells as revealed by ruthenium red. Journal of Cell Biology, 46(1):159–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/1/159.

Touster:1970:IRL

[TADH70] Oscar Touster, N. N. Aronson, John T. Dulaney, and Herman Hendrickson. Isolation of rat liver plasma membranes. Journal of Cell Biology, 47(3):604–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/47/3/604.

Tamai:1974:MCS

[TAKS74] Yoichi Tamai, Shigeko Matsukawa Araki, Kimio Katsuta, and Mei Satake. Molecular composition of the submicrosomal membrane lipid of rat brain. Journal of Cell Biology, 63(3): 749–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/749.

Tamm:1972:CMP

[Tam72a] Sidney L. Tamm. Ciliary motion in paramecium. Journal of Cell Biology, 55(1):250–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/55/1/250.

Tamm:1972:FKA

[Tam72b] Sidney L. Tamm. Free kinetosomes in Australian flagellates. Journal of Cell Biology, 54(1):39–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/1/39.

Tanaka:1973:IMO

[Tan73] Kenji Tanaka. Intranuclear microtubule organizing center in early prophase nuclei of the plasmodium of the slime mold, Physarum Polycephalum. Journal of Cell Biology, 57 (1):220–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/220.

Tennyson:1970:AAD

[TB70] Virginia M. Tennyson and Miro Brzin. The appearance of acetylcholinesterase in the dorsal root neuroblast of the rabbit embryo. Journal of Cell Biology, 46(1):64–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/46/1/64.

Tilney:1973:MEP

[TBB^{+}73] Lewis G. Tilney, Joseph Bryan, Doris J. Bush, Keigi Fujiwara, Mark S. Mooseker, Douglas B. Murphy, and Daniel H. Snyder. Microtubules: Evidence for 13 protofilaments. Journal of Cell Biology, 59(2):267–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/2/267.

Trelease:1974:CLM

[TBB74] Richard N. Trelease, Wayne M. Becker, and John J. Burke. Cytochemical localization of malate synthase in glyoxysomes. Journal of Cell Biology, 60(2):483–??, February 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/60/2/

Tashjian:1970:PBP

[TBL70] Armen H. Tashjian, Frank C. Bancroft, and Lawrence Levine. Production of both prolactin and growth hormone by clonal strains of rat pituitary tumor cells. Journal of Cell Biology, 47(1):61–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/1/61.

Tennyson:1971:AAM

[TBS71] Virginia M. Tennyson, Miro Brzin, and Paul Slotwiner. The appearance of acetylcholinesterase in the myotome of the embryonic rabbit. Journal of Cell Biology, 51(3):703–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
51/3/703.

Tulkens:1974:AFH

[TBT74] Paul Tulkens, Henry Beaufay, and André Trouet. Analytical fractionation of homogenates from cultured rat embryo fibroblasts. Journal of Cell Biology, 63(2):383–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/383.

Thompson:1974:NDT

[TBW74] Guy A. Thompson, Lonnie C. Baugh, and Laurie F. Walker. Nonlethal deciliation of Tetrahymena by a local anesthetic and its utility as a tool for studying cilia regeneration. Journal of Cell Biology, 61(1):253–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/61/1/253.

Tilney:1970:FCR

[TC70] Lewis G. Tilney and Robert R. Cardell. Factors controlling the reassembly of the microvillous border of the small intestine of the Salamander. Journal of Cell Biology, 47(2): 408–??, November 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/2/408.

Trelstad:1971:MCS

[TC71] Robert L. Trelstad and Alfred J. Coulombre. Morphogenesis of the collagenous stroma in the chick cornea. Journal of Cell Biology, 50(3):840–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/50/3/840.

Tanguay:1972:SM

[TC72] R. Tanguay and K. D. Chaudhary. Studies on mitochondria. Journal of Cell Biology, 54(2):295–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/54/2/295.

Tobey:1972:MCE

[TCK72] Robert A. Tobey, Harry A. Crissman, and Paul M. Kraemer. A method for comparing effects of different synchronizing protocols on mammalian cell cycle traverse. Journal of Cell Biology, 54(3):638–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/3/638.

Taylor:1973:CBA

[TCMA73] D. L. Taylor, J. S. Condeelis, P. L. Moore, and R. D. Allen. The contractile basis of amoeboid movement. Journal of Cell Biology, 59(2):378–??, November 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/2/378.

Tranzer:1972:SHM

[TdPP72] J. P. Tranzer, M. da Prada, and A. Pletscher. Storage of 5-hydroxytryptamine in megakaryocytes. Journal of Cell Biology, 52(1):191–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/1/191.

Tennyson:1970:FSA

[Ten70] Virginia M. Tennyson. The fine structure of the axon and growth cone of the dorsal root neuroblast of the rabbit embryo. Journal of Cell Biology, 44(1):62–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/44/1/62.

Tilney:1970:NSA

[TG70] Lewis G. Tilney and Janette Goddard. Nucleating sites for the assembly of cytoplasmic microtubules in the ectodermal cells of blastulae of Arbacia Punctulata. Journal of Cell Biology, 46(3):564–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/3/564.

Trigg:1973:UBB

[TGW73] Monica J. Trigg and Salome Gluecksohn-Waelsch. Ultrastructural basis of biochemical effects in a series of lethal alleles in the mouse. Journal of Cell Biology, 58(3):549– ??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/3/549.

Taylor:1972:FAE

[TH72] John D. Taylor and Mac E. Hadley. The fate of amphibian egg melanosomes. Journal of Cell Biology, 52(2):493– ??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/493.

Tandler:1973:DGM

[TH73a] Bernard Tandler and Charles L. Hoppel. Division of giant mitochondria during recovery from cuprizone intoxication. Journal of Cell Biology, 56(1):266–??, January 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/56/1/
266.

Teichberg:1973:AAR

[TH73b] Saul Teichberg and Eric Holtzman. Axonal agranular reticulum and synaptic vesicles in cultured embryonic chick sympathetic neurons. Journal of Cell Biology, 57(1):88–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/88.

Tilney:1973:PAR

[THIM73] Lewis G. Tilney, Sadashi Hatano, Harunori Ishikawa, and Mark S. Mooseker. The polymerization of actin: Its role in the generation of the acrosomal process of certain echinoderm sperm. Journal of Cell Biology, 59(1):109–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/59/
1/109.

Trelstad:1974:ECG

[THT74] Robert L. Trelstad, Kimiko Hayashi, and Bryan P. Toole. Epithelial collagens and glycosaminoglycans in the embryonic cornea. Journal of Cell Biology, 62(3):815–??, September 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
3/815.

Tidwell:1970:MTR

[Tid70] T. Tidwell. The methylation of transfer ribonucleic acid during regeneration of the liver. Journal of Cell Biology, 46(2): 370–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/2/370.

Tilney:1971:HMP

[Til71] Lewis G. Tilney. How microtubule patterns are generated. Journal of Cell Biology, 51(3):837–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
3/837.

Tandler:1971:ICR

[TK71] C. J. Tandler and A. L. Kierszenbaum. Inorganic cations in rat kidney. Journal of Cell Biology, 50(3):830–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/830.

Tres:1972:ICC

[TKT72] Laura L. Tres, A. L. Kierszenbaum, and C. J. Tandler. Inorganic cations in the cell nucleus. Journal of Cell Biology, 53(2):483–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/2/483.

Tougard:1973:LEM

[TKTVJ73] C. Tougard, B. Kerdelhue, A. Tixier-Vidal, and M. Jutisz. Light and electron microscope localization of binding sites of antibodies against ovine luteinizing hormone and its two subunits in rat adenohypophysis using peroxidase-labeled antibody technique. Journal of Cell Biology, 58(3):503– ??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/58/3/503.

Tobey:1970:RID

[TL70a] R. A. Tobey and K. D. Ley. Regulation of initiation of DNA synthesis in Chinese hamster cells. Journal of Cell Biology, 46(1):151–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/1/151.

Togasaki:1970:CSF

[TL70b] R. K. Togasaki and R. P. Levine. Chloroplast structure and function in ac-20 , a mutant strain of Chlamydomonas reinhardi. Journal of Cell Biology, 44(3):531–??, March 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
3/531.

Tandler:1970:ILI

[TLS70] Carlos J. Tandler, César M. Libanati, and Carlos A. Sanchis. The intracellular localization of inorganic cations with potassium pyroantimonate. Journal of Cell Biology, 45(2):355–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
45/2/355.

Tillack:1970:DOS

[TM70] Thomas W. Tillack and Vincent T. Marchesi. Demonstration of the outer surface of freeze-etched red blood cell membranes. Journal of Cell Biology, 45(3):649–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
3/649.

Thomas:1973:GIM

[TML73] D. B. Thomas, Gay Medley, and C. A. Lingwood. Growth inhibition of murine tumor cells, in vitro, by puromycin, [^{6}n]O^{2′}-dibutyryl 3’,5’-adenosine monophosphate, or adenosine. Journal of Cell Biology, 57(2):397–??, May 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/57/2/

Tashiro:1970:SPS

[TO70] Yutaka Tashiro and Eiichi Otsuki. Studies on the posterior silk gland of the silkworm bombix mori. Journal of Cell Biology, 46(1):1–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/46/1/1.

Tokuyasu:1973:TUC

[Tok73] K. T. Tokuyasu. A technique for ultracryotomy of cell suspensions and tissues. Journal of Cell Biology, 57(2):551–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
57/2/551.

Tokuyasu:1974:SHS

[Tok74] K. T. Tokuyasu. Spoke heads in sperm tail of Drosophila Melanogaster. Journal of Cell Biology, 63(1):334–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/334.

Tripodi:1972:DST

[TPG72] G. Tripodi, P. Pizzolongo, and M. Giannattasio. A DNasesensitive twisted structure in the mitochondrial matrix of Polysiphonia (Rhodophyta). Journal of Cell Biology, 55(2): 530–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/530.

Telford:1973:MIC

[TR73] John N. Telford and Efraim Racker. A method for increasing contrast of mitochondrial inner membrane spheres in thin sections of epon-araldite embedded tissue. Journal of Cell Biology, 57(2):580–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/2/580.

Trelstad:1970:GAC

[Tre70] Robert L. Trelstad. The Golgi apparatus in chick corneal epithelium: Changes in intracellular position during development. Journal of Cell Biology, 45(1):34–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/45/1/34.

Trelstad:1971:VEC

[Tre71] Robert L. Trelstad. Vacuoles in the embryonic chick corneal epithelium, an epithelium which produces collagen. Journal of Cell Biology, 48(3):689–??, March 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/48/3/689.

Temmink:1974:RRS

[TS74] J. H. M. Temmink and H. Spiele. Ruthenium red staining of normal and transformed murine fibroblasts in vitro. Journal of Cell Biology, 61(2):548–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/61/2/548.

Tupper:1970:OEC

[TSE70] Joseph Tupper, John W. Saunders, and Charles Edwards. The onset of electrical communication between cells in the developing starfish embryo. Journal of Cell Biology, 46(1): 187–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/1/187.

Thompson:1971:HMD

[TT71] E. Brad Thompson and Gordon M. Tomkins. A histochemical method for the demonstration of tyrosine aminotransferase in tissue culture cells, and studies of this enzyme in hepatoma tissue culture cells. Journal of Cell Biology, 49(3):921–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/921.

Tucker:1974:MAC

[Tuc74] J. B. Tucker. Microtubule arms and cytoplasmic streaming and microtubule bending and stretching of intertubule links in the feeding tentacle of the suctorian ciliate tokophrya. Journal of Cell Biology, 62(2):424–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/62/2/424.

Turner:1970:ECS

[Tur70] F. Rudolf Turner. The effects of colchicine on spermatogenesis in nitella. Journal of Cell Biology, 46(2):220–??, August 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/46/
2/220.

Turner:1972:MMT

[Tur72] F. Rudolf Turner. Modified microtubules in the testis of the water strider, Gerris remigis (say). Journal of Cell Biology, 53(1):263–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/263.

Tixier-Vidal:1970:SRP

[TVG70] A. Tixier-Vidal and D. Gourdji. Synthesis and renewal of proteins in duck anterior hypophysis in organ culture. Journal of Cell Biology, 46(1):130–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/46/1/130.

Uehara:1970:DGJ

[UB70] Yasuo Uehara and Geoffrey Burnstock. Demonstration of “gap junctions” between smooth muscle cells. Journal of Cell Biology, 44(1):215–??, January 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/44/1/215.

Uehara:1972:PSS

[UB72] Yasuo Uehara and Geoffrey Burnstock. Postsynaptic specialization of smooth muscle at close neuromuscular junctions in the guinea pig sphincter pupillae. Journal of Cell Biology, 53 (3):849–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/3/849.

Upholt:1974:ARI

[UB74] William B. Upholt and Piet Borst. Accumulation of replicative intermediates of mitochondrial DNA in Tetrahymena Pyriformis grown in ethidium bromide. Journal of Cell Biology, 61(2):383–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/2/383.

Ukena:1974:ECC

[UBKB74] Thomas E. Ukena, Joan Z. Borysenko, Morris J. Karnovsky, and Richard D. Berlin. Effects of colchicine, cytochalasin B, and 2-deoxyglucose on the topographical organization of surface-bound concanavalin a in normal and transformed fibroblasts. Journal of Cell Biology, 61(1):70–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/61/1/70.

Uehara:1971:CFD

[UCB71] Y. Uehara, G. R. Campbell, and G. Burnstock. Cytoplasmic filaments in developing and adult vertebrate smooth muscle. Journal of Cell Biology, 50(2):484–??, August 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/50/2/484.

Ugel:1971:SIA

[Uge71] Arthur R. Ugel. Studies on isolated aggregating oligoribonucleoproteins of the epidermis with histochemical and morphological characteristics of keratohyalin. Journal of Cell Biology, 49(2):405–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/2/405.

Ugel:1972:FCB

[UI72] Arthur R. Ugel and William Idler. Further characterization of bovine keratohyalin. Journal of Cell Biology, 52(2): 453–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/52/2/453.

Umiel:1972:IPL

[UP72] N. Umiel and W. Plaut. Interaction of poly-L-lysine with chromatin. Journal of Cell Biology, 54(3):556–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
3/556.

Umiel:1973:IPL

[UP73] N. Umiel and W. Plaut. Interaction of poly-L-lysine with chromatin. Journal of Cell Biology, 56(1):139–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/56/
1/139.

Ulsamer:1971:PPM

[UWWK71] Andrew G. Ulsamer, Paddy L. Wright, Mary G. Wetzel, and Edward D. Korn. Plasma and phagosome membranes of Acanthamoeba Castellanii. Journal of Cell Biology, 51(1): 193–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/193.

Vracko:1970:CBL

[VB70] Rudolf Vracko and Earl P. Benditt. Capillary basal lamina thickening. Journal of Cell Biology, 47(1):281–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/47/
1/281.

Vracko:1972:BLS

[VB72] Rudolf Vracko and Earl P. Benditt. Basal lamina: the scaffold for orderly cell replacement. Journal of Cell Biology, 55(2): 406–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/2/406.

Vankin:1970:USR

[VBD70] G. Lawrence Vankin, E. Michael Brandt, and William De-Witt. Ultrastructural studies of red blood cells from thyroxintreated Rana Catesbeiana tadpoles. Journal of Cell Biology, 47(3):767–??, December 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/47/3/767.

Vassart:1971:ISA

[VDC71] G. M. Vassart, J. E. Dumont, and F. R. L. Cantraine. Interpretation of surface aspects of cell sections. Journal of Cell Biology, 49(1):210–??, April 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/1/210.

Vassar:1972:PEA

[VHB^{+}72] P. S. Vassar, J. M. Hards, D. E. Brooks, B. Hagenberger, and G. V. F. Seaman. Physicochemical effects of aldehydes on the human erythrocyte. Journal of Cell Biology, 53(3):809–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
53/3/809.

Vaughn:1974:QSS

[VHG74] James E. Vaughn, Charles K. Henrikson, and Judith A. Grieshaber. A quantitative study of synapses on motor neuron dendritic growth cones in developing mouse spinal cord. Journal of Cell Biology, 60(3):664–??, March 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/3/664.

Vigil:1970:CDC

[Vig70] Eugene L. Vigil. Cytochemical and developmental changes in microbodies (glyoxysomes) and related organelles of castor bean endosperm. Journal of Cell Biology, 46(3):435– ??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/435.

vanIterson:1971:DMC

[vIG71] Woutera van Iterson and J. B. Groen. DNA-membrane complexes of bacillus subtilis. Journal of Cell Biology, 49(2): 553–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/553.

Vladutiu:1974:IDP

[VR74] Georgirene D. Vladutiu and Noel R. Rose. Intracellular distribution of a primate-specific esterase in cultured cells and tissues. Journal of Cell Biology, 62(2):560–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/62/
2/560.

vanRijn:1974:RPC

[vRBvWW74] Han van Rijn, Marinus M. Bevers, Roeland van Wijk, and Wesley D. Wicks. Regulation of phosphoenolpyruvate carboxykinase and tyrosine transaminase in hepatoma cell cultures. Journal of Cell Biology, 60(1):181–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/60/
1/181.

Venkatachalam:1970:FSL

[VSF70] M. A. Venkatachalam, M. H. Soltani, and H. Dariush Fahimi. Fine structural localization of peroxidase activity in the epithelium of large intestine of rat. Journal of Cell Biology, 46 (1):168–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/1/168.

Vignais:1972:MCU

[VSHA72] Pierre V. Vignais, Barbara J. Stevens, Janine Huet, and Jean André. Mitoribosomes from Candida utilis. Journal of Cell Biology, 54(3):468–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/3/468.

Vaughn:1972:KSD

[VT72] Jack C. Vaughn and Lillian A. Thomson. A kinetic study of DNA and basic protein metabolism during spermatogenesis in the sand crab, emerita analoga. Journal of Cell Biology, 52(2):322–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/52/2/322.

Vonderhaar:1974:RCC

[VT74] Barbara K. Vonderhaar and Yale J. Topper. A role of the cell cycle in hormone-dependent differentiation. Journal of Cell Biology, 63(2):707–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/63/2/707.

Vergara:1974:ICL

[VZRE74] J. Vergara, F. Zambrano, J. D. Robertson, and H. Elrod. Isolation and characterization of luminal membranes from urinary bladder. Journal of Cell Biology, 61(1):83–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
1/83.

Wolk:1974:ALA

[WABG74] C. Peter Wolk, Sam M. Austin, John Bortins, and Aaron Galonsky. Autoradiographic localization of ^{13}N after fixation of ^{13}N-labeled nitrogen gas by a heterocyst-forming blue–green alga. Journal of Cell Biology, 61(2):440–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
2/440.

Wibo:1971:EME

[WACBB71] Maurice Wibo, Alain Amar-Costesec, Jacques Berthet, and Henri Beaufay. Electron microscope examination of subcellular fractions. Journal of Cell Biology, 51(1):52–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/52.

Wake:1974:DVR

[Wak74] Kenjiro Wake. Development of vitamin A-rich lipid droplets in multivesicular bodies of rat liver stellate cells. Journal of Cell Biology, 63(2):683–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/683.

Warner:1970:NOF

[War70] Fred D. Warner. New observations on flagellar fine structure. Journal of Cell Biology, 47(1):159–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/47/1/
159.

Warren:1974:MOE

[War74] Robert H. Warren. Microtubular organization in elongating myogenic cells. Journal of Cell Biology, 63(2):550–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
63/2/550.

Watson:1972:ORG

[Wat72] Kenneth Watson. The organization of ribosomal granules within mitochondrial structures of aerobic and anaerobic cells of Saccharomyces Cerevisiae. Journal of Cell Biology, 55(3): 721–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/55/3/721.

Woodcock:1970:EVQ

[WB70] Christopher L. F. Woodcock and Lawrence Bogorad. Evidence for variation in the quantity of DNA among plastids of acetabularia. Journal of Cell Biology, 44(2):361–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/44/
2/361.

Wunderlich:1974:RTA

[WBSW74] Frank Wunderlich, Werner Batz, Volker Speth, and Donald F. H. Wallach. Reversible, thermotropic alteration of nuclear membrane structure and nucleocytoplasmic RNA transport in Tetrahymena. Journal of Cell Biology, 61(3):633–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
3/633.

Witman:1972:CFa

[WCBR72] G. B. Witman, K. Carlson, J. Berliner, and Joel L. Rosenbaum. Chlamydomonas flagella. Journal of Cell Biology, 54(3):507–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/3/507.

Walker:1972:MA

[WCDI72] W. A. Walker, R. Cornell, L. M. Davenport, and K. J. Isselbacher. Macromolecular absorption. Journal of Cell Biology, 54(2):195–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/2/195.

Widmann:1972:MPK

[WCF72] Jean-Jacques Widmann, Ramzi S. Cotran, and H. Dariush Fahimi. Mononuclear phagocytes (kupffer cells) and endothelial cells. Journal of Cell Biology, 52(1):159–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
1/159.

Weintraub:1971:PEE

[WCH71] Harold Weintraub, Graham Le M. Campbell, and Howard Holtzer. Primitive erythropoiesis in early chick embryogenesis. Journal of Cell Biology, 50(3):652–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/50/
3/652.

Witman:1972:CFb

[WCR72] G. B. Witman, K. Carlson, and Joel L. Rosenbaum. Chlamydomonas flagella. Journal of Cell Biology, 54(3):540–??, September 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/54/3/540.

Wade:1971:EOF

[WD71] James B. Wade and Vincent A. Discala. The effect of osmotic flow on the distribution of horseradish peroxidase within the intercellular spaces of toad bladder epithelium. Journal of Cell Biology, 51(2):553–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/51/2/553.

Wanson:1972:RSR

[WD72] Jean-Claude Wanson and Pierre Drochmans. Role of the sarcoplasmic reticulum in glycogen metabolism. Journal of Cell Biology, 54(2):206–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/54/2/206.

Walther:1970:PID

[WE70] William G. Walther and Leland N. Edmunds. Periodic increase in deoxyribonuclease activity during the cell cycle in synchronized euglena. Journal of Cell Biology, 46(3):613– ??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/46/3/613.

Watrud:1973:CMD

[WE73] Lidia S. Watrud and Albert H. Ellingboe. Cobalt as a mitochondrial density marker in a study of cytoplasmic exchange during mating of schizophyllum commune. Journal of Cell Biology, 59(1):127–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/59/1/127.

Weber:1972:USB

[Web72] Nicholas E. Weber. Ultrastructural studies of beef heart mitochondria. Journal of Cell Biology, 55(2):457–??, November 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
2/457.

Weisenberg:1972:COT

[Wei72] Richard C. Weisenberg. Changes in the organization of tubulin during meiosis in the eggs of the surf clam, spisula solidissima. Journal of Cell Biology, 54(2):266–??, August 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
2/266.

Weller:1974:VCB

[Wel74] Nancy K. Weller. Visualization of concanavalin a-binding sites with scanning electron microscopy. Journal of Cell Biology, 63(2):699–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/63/2/699.

Wise:1970:CSG

[WF70] Gary E. Wise and Charles J. Flickinger. Cytochemical staining of the Golgi apparatus in Amoeba Proteus. Journal of Cell Biology, 46(3):620–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/3/620.

Williams:1973:RMTa

[WF73] Norman E. Williams and Joseph Frankel. Regulation of microtubules in Tetrahymena. Journal of Cell Biology, 56(2): 441–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/441.

Weinbaum:1970:MMN

[WFO70] George Weinbaum, Donald A. Fischman, and Shinichi Okuda. Membrane modifications in nutritionally induced filamentous Escherichia coli B. Journal of Cell Biology, 45 (3):493–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/3/493.

Wise:1973:EML

[WG73] Gary E. Wise and Lester Goldstein. Electron microscope localization of nuclear RNA’s that Shuttle between cytoplasm and nucleus and nuclear RNA’s that do not. Journal of Cell Biology, 56(1):129–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/1/129.

Weinstein:1970:DAS

[WH70] Roslyn B. Weinstein and Elizabeth D. Hay. Deoxyribonucleic acid synthesis and mitosis in differentiated cardiac muscle cells of chick embryos. Journal of Cell Biology, 47(1): 310–??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/1/310.

Wodzicki:1973:MPT

[WH73] Tomasz J. Wodzicki and Walter J. Humphreys. Maturing pine tracheids. Journal of Cell Biology, 56(1):263–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
56/1/263.

Weber:1974:DVL

[WHCFB74] G. H. Weber, U. Heine, M. Cottler-Fox, and G. Beaudreau. Direct visualization of labeled nucleotides in DNA using electron microscopy and autoradiography. Journal of Cell Biology, 61(1):257–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/61/1/257.

Wainberg:1973:ESV

[WHG73] M. A. Wainberg, C. Howe, and G. C. Godman. Enhancement of sendai virus-mediated cell fusion by cupric ions. Journal of Cell Biology, 57(2):388–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/57/2/388.

Whitmore:1972:IMR

[Whi72] P. G. Whitmore. Isolation of microfibrils from rat vibrissal follicles. Journal of Cell Biology, 52(1):174–??, January 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
1/174.

Watson:1970:BM

[WHL70] K. Watson, J. M. Haslam, and Anthony W. Linnane. Biogenesis of mitochondria. Journal of Cell Biology, 46(1):88–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
46/1/88.

Widnell:1972:CLN

[Wid72] Christopher C. Widnell. Cytochemical localization of 5’- nucleotidase in subcellular fractions isolated from rat liver. Journal of Cell Biology, 52(3):542–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/52/3/542.

Wade:1974:FFO

[WK74a] James B. Wade and Morris J. Karnovsky. Fracture faces of osmotically disrupted zonulae occludentes. Journal of Cell Biology, 62(2):344–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/62/2/344.

Wade:1974:SZO

[WK74b] James B. Wade and Morris J. Karnovsky. The structure of the zonula occludens. Journal of Cell Biology, 60(1): 168–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/1/168.

Wolstenholme:1973:SSC

[WKCF73] David R. Wolstenholme, Katsuro Koike, and Patricia Cochran-Fouts. Single Strand-containing replicating molecules of circular mitochondrial DNA. Journal of Cell Biology, 56 (1):230–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/56/1/230.

Wolstenholme:1972:HDS

[WKG72] David R. Wolstenholme, Robert G. Kirschner, and Nicholas J. Gross. Heat denaturation studies of rat liver mitochondrial DNA. Journal of Cell Biology, 53(2):393–??, May 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/53/
2/393.

Wood:1970:PAR

[WL70] Richard L. Wood and Peter G. Legg. Peroxidase activity in rat liver microbodies after amino-triazole inhibition. Journal of Cell Biology, 45(3):576–??, June 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/45/3/576.

Weinstock:1971:EMG

[WL71a] Alfred Weinstock and C. P. Leblond. Elaboration of the matrix glycoprotein of enamel by the secretory ameloblasts of the rat incisor as revealed by radioautography after galactose-3H injection. Journal of Cell Biology, 51(1):26–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/26.

Wood:1971:PIN

[WL71b] David D. Wood and David J. L. Luck. A paracrystalline inclusion in neurospora crassa. Journal of Cell Biology, 51(1): 249–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/1/249.

Weinstock:1973:RVD

[WL73a] Melvyn Weinstock and C. P. Leblond. Radioautographic visualization of the deposition of a phosphoprotein at the mineralization front in the dentin of the rat incisor. Journal of Cell Biology, 56(3):838–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/56/3/838.

Williams:1973:ZRD

[WL73b] Richard O. Williams and Lawrence A. Loeb. Zinc requirement for DNA replication in stimulated human lymphocytes. Journal of Cell Biology, 58(3):594–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
3/594.

Weinstock:1974:SMR

[WL74] Melvyn Weinstock and C. P. Leblond. Synthesis, migration, and release of precursor collagen by odontoblasts as visualized by radioautography after [^{3}H]proline administration. Journal of Cell Biology, 60(1):92–??, January 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/60/1/92.

Wen:1974:MGS

[WLH74] Wu-Nan Wen, Pedro E. León, and Donald R. Hague. Multiple gene sites for 5s and 18 + 28s RNA on chromosomes of glyptotendipes barbipes (staeger). Journal of Cell Biology, 62(1):132–??, July 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/1/132.

Wright:1972:MER

[WM72] Daniel G. Wright and Stephen E. Malawista. The mobilization and extracellular release of granular enzymes from human leukocytes during phagocytosis. Journal of Cell Biology, 53(3):788–??, June 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/3/788.

Williams:1973:DEA

[WM73a] J. G. Williams and I. A. MacPherson. The differential effect of actinomycin D in normal and virus-transformed cells. Journal of Cell Biology, 57(1):148–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/57/1/148.

Wilson:1973:MAC

[WM73b] Leslie Wilson and Isaura Meza. The mechanism of action of colchicine. Journal of Cell Biology, 58(3):709–??, September 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/58/
3/709.

Wood:1974:VCB

[WMB74] John G. Wood, Barbara J. McLaughlin, and Robert P. Barber. The visualization of concanavalin a binding sites in Purkinje cell somata and dendrites of rat cerebellum. Journal of Cell Biology, 63(2):541–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/2/541.

Whitlock:1974:IAH

[WMG74] James P. Whitlock, Haruko Miller, and Harry V. Gelboin. Induction of aryl hydrocarbon (benzo[a]pyrene) hydroxylase and tyrosine aminotransferase in hepatoma cells in culture. Journal of Cell Biology, 63(1):136–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
1/136.

Wolstenholme:1973:COC

[WMKJ73] David R. Wolstenholme, John D. McLaren, Katsuro Koike, and Elaine L. Jacobson. Catenated oligomeric circular DNA molecules from mitochondria of malignant and normal mouse and rat tissues. Journal of Cell Biology, 56(1):247–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
56/1/247.

Williams:1973:RMTb

[WN73] Norman E. Williams and E. Marlo Nelsen. Regulation of microtubules in Tetrahymena. Journal of Cell Biology, 56(2): 458–??, February 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/2/458.

Woolley:1971:SPF

[Woo71] D. M. Woolley. Striations in the peripheral fibers of rat and mouse spermatozoa. Journal of Cell Biology, 49(3):936–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/936.

Wood:1974:CPA

[Woo74] Richard L. Wood. A closely packed array of membrane intercalated particles at the free surface of hydra. Journal of Cell Biology, 62(2):556–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/62/2/556.

Wrathall:1973:SPM

[WOSE73] Jean R. Wrathall, Constance Oliver, Selma Silagi, and Edward Essner. Suppression of pigmentation in mouse melanoma cells by 5-bromodeoxyuridine. Journal of Cell Biology, 57(2):406–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/57/2/406.

Warner:1972:CBA

[WP72a] David A. Warner and James F. Perdue. Cytochalasin B and the adenosine triphosphate content of treated fibroblasts. Journal of Cell Biology, 55(1):242–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
1/242.

Wolff:1972:SHC

[WP72b] David A. Wolff and HÅkan Pertoft. Separation of HeLa cells by colloidal silica density gradient centrifugation. Journal of Cell Biology, 55(3):579–??, December 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/55/3/579.

Wibo:1974:PDC

[WP74a] Maurice Wibo and Brian Poole. Protein degradation in cultured cells. Journal of Cell Biology, 63(2):430–??, November 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
2/430.

Willingham:1974:CAM

[WP74b] Mark C. Willingham and Ira Pastan. Cyclic AMP mediates the concanavalin a agglutinability of mouse fibroblasts. Journal of Cell Biology, 63(1):288–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/1/288.

Waxman:1972:MCF

[WPB72] S. G. Waxman, G. D. Pappas, and M. V. L. Bennett. Morphological correlates of functional differentiation of nodes of ranvier along single fibers in the neurogenic electric organ of the knife fish sternarchus. Journal of Cell Biology, 53 (1):210–??, April 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/53/1/210.

Warshaw:1972:CGO

[WR72] Joseph B. Warshaw and Miriam D. Rosenthal. Changes in glucose oxidation during growth of embryonic heart cells in culture. Journal of Cell Biology, 52(2):283–??, February 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/52/
2/283.

Webb:1974:CCD

[WR74a] Glenda C. Webb and Stephen Roth. Cell contact dependence of surface galactosyltransferase activity as a function of the cell cycle. Journal of Cell Biology, 63(3):796–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/63/
3/796.

Wolf:1974:CMC

[WR74b] Barbara Anne Wolf and P. W. Robbins. Cell mitotic cycle synthesis of nil hamster glycolipids including the Forssman antigen. Journal of Cell Biology, 61(3):676–??, June 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
3/676.

Wolfe:1974:INR

[WR74c] Jason Wolfe and Lewis Rodriguez. Inhibition of nuclear reconstruction by dithiothreitol. Journal of Cell Biology, 60(2): 497–??, February 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/60/2/497.

Wagner:1971:ECL

[WRE71] Roger Wagner, Murray Rosenberg, and Richard Estensen. Endocytosis in Chang liver cells. Journal of Cell Biology, 50(3):804–??, September 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/50/3/804.

Wright:1973:EPC

[Wri73] J. A. Wright. Evidence for pleiotropic changes in lines of Chinese hamster ovary cells resistant to concanavalin a and phytohemagglutinin-p. Journal of Cell Biology, 56(3): 666–??, March 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/56/3/666.

Warner:1974:SBC

[WS74a] Fred D. Warner and Peter Satir. The structural basis of ciliary bend formation. Journal of Cell Biology, 63(1):35– ??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/1/35.

Waugh:1974:LJS

[WS74b] R. A. Waugh and J. R. Sommer. Lamellar junctional sarcoplasmic reticulum. Journal of Cell Biology, 63(1):337– ??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/1/337.

Winborn:1974:POD

[WS74c] William B. Winborn and Leonard L. Seelig. Pattern of osmium deposition in the parietal cells of the stomach. Journal of Cell Biology, 63(1):99–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/1/99.

Waugh:1973:FSR

[WSS73] R. A. Waugh, T. L. Spray, and J. R. Sommer. Fenestrations of sarcoplasmic reticulum. Journal of Cell Biology, 59(1): 254–??, October 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/59/1/254.

Warshaw:1970:CEM

[WT70] Joseph B. Warshaw and Mary L. Terry. Cellular energy metabolism during fetal development. Journal of Cell Biology, 44(2):354–??, February 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/44/2/354.

Wanson:1971:MBC

[WT71] Jean-Claude Wanson and Louis Tielemans. Morphological and biochemical characteristics of glycogen particles isolated from rabbit polymorphonuclear leukocytes. Journal of Cell Biology, 49(3):816–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/49/3/816.

Weidner:1973:ATE

[WT73a] E. Weidner and W. Trager. Adenosine triphosphate in the extracellular survival of an intracellular parasite (Nosema Michaelis, Microsporidia). Journal of Cell Biology, 57(2): 586–??, May 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/2/586.

Woodruff:1973:PIB

[WT73b] Richard I. Woodruff and William H. Telfer. Polarized intercellular bridges in ovarian follicles of the Cecropia moth. Journal of Cell Biology, 58(1):172–??, July 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/1/172.

Wu:1971:CSN

[WW71] Roy S. Wu and Jonathan R. Warner. Cytoplasmic synthesis of nuclear proteins. Journal of Cell Biology, 51(3): 643–??, December 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/51/3/643.

Williams:1972:CBP

[WW72] John A. Williams and J. Wolff. Colchicine-binding protein and the secretion of thyroid hormone. Journal of Cell Biology, 54(1):157–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/54/1/157.

Waaland:1971:GV

[WWB71] J. Robert Waaland, Susan Drury Waaland, and Daniel Branton. Gas vacuoles. Journal of Cell Biology, 48(1):212–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
48/1/212.

Wang:1974:GCC

[WWBG74] Wei-Yeh Wang, Wenan Lee Wang, John E. Boynton, and Nicholas W. Gillham. Genetic control of chlorophyll biosynthesis in chlamydomonas. Journal of Cell Biology, 63(3): 806–??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/806.

Waaland:1972:NPP

[WWC72] Susan Drury Waaland, J. Robert Waaland, and Robert Cleland. A new pattern of plant cell elongation: Bipolar band growth. Journal of Cell Biology, 54(1):184–??, July 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/54/
1/184.

Walig:1974:ULC

[WWvdN74] C. Walig, J. M. M. Walboomers, and J. van der Noordaa. Ultrastructural localization of concanavalin a receptors in somatic cell hybrids between normal mouse cells and simian virus 40-transformed rat cells. Journal of Cell Biology, 61 (2):553–??, May 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/61/2/553.

Westfall:1971:UEP

[WYE71] Jane A. Westfall, S. Yamataka, and Paul D. Enos. Ultrastructural evidence of polarized synapses in the nerve net of hydra. Journal of Cell Biology, 51(1):318–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/318.

Yip:1974:SFH

[YA74] Dominic K. Yip and Nelly Auersperg. Structural and functional heterogeneity of the surface of rat leukemia cells. Journal of Cell Biology, 63(1):109–??, October 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/63/1/109.

Yates:1970:ESN

[YCD70] R. D. Yates, I-Li Chen, and Donald Duncan. Effects of sinus nerve stimulation on carotid body glomus cells. Journal of Cell Biology, 46(3):544–??, September 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/3/544.

Yavin:1973:CDC

[YM73] Ephraim Yavin and John H. Menkes. The culture of dissociated cells from rat cerebral cortex. Journal of Cell Biology, 57(1):232–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/57/1/232.

Yoo:1970:UCC

[Yoo70] Bong Yul Yoo. Ultrastructural changes in cells of pea embryo radicles during germination. Journal of Cell Biology, 45 (1):158–??, April 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/45/1/158.

Young:1971:RRC

[You71a] Richard W. Young. The renewal of rod and cone outer segments in the Rhesus monkey. Journal of Cell Biology, 49 (2):303–??, May 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/49/2/303.

Youson:1971:PCM

[You71b] John H. Youson. Prismatic cristae and matrix granules in mitochondria of the kidneys of ammocoetes. Journal of Cell Biology, 48(1):189–??, January 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/48/1/189.

Young:1973:RGC

[You73] Richard W. Young. The role of the Golgi complex in sulfate metabolism. Journal of Cell Biology, 57(1):175–??, April 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/57/
1/175.

Yago:1972:GDM

[YSS^{+}72] Nagasumi Yago, Masatoshi Seki, Shigetaka Sekiyama, Shigeru Kobayashi, Hiromi Kurokawa, Yuko Iwai, Fumiaki Sato, and Akihiro Shiragai. Growth and differentiation of mitochondria in the regenerating rat adrenal cortex. Journal of Cell Biology, 52(3):503–??, March 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/52/3/503.

Yamada:1971:UFG

[YSW71] Kenneth M. Yamada, Brian S. Spooner, and Norman K. Wessells. Ultrastructure and function of growth cones and axons of cultured nerve cells. Journal of Cell Biology, 49(3):614–??, June 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
49/3/614.

Yavin:1974:ACD

[YY74] Ephraim Yavin and Ziva Yavin. Attachment and culture of dissociated cells from rat embryo cerebral hemispheres on polylysine-coated surface. Journal of Cell Biology, 62(2): 540–??, August 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/62/2/540.

Zirwer:1970:LUD

[ZBSW70] D. Zirwer, E. Buder, W. Schälike, and R. Wetzel. Lineardichroitische Untersuchungen zur DNS-Organisation in Spermienköpfen von Locusta Migratoria L. (German) [Linear dichroitics on DNA organization in sperm heads of Locusta Migratoria L]. Journal of Cell Biology, 45(2):431–??, May 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/45/
2/431.

Zweidler:1971:LSI

[ZC71] Alfred Zweidler and Leonard H. Cohen. Large-scale isolation and fractionation of organs of Drosophila Melanogaster larvae. Journal of Cell Biology, 51(1):240–??, October 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
1/240.

Zucker-Franklin:1970:SFH

[ZF70] Dorothea Zucker-Franklin. The submembranous fibrils of human blood platelets. Journal of Cell Biology, 47(1):293– ??, October 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/47/1/293.

Zucker-Franklin:1974:UAH

[ZFG74] Dorothea Zucker-Franklin and George Grusky. Ultrastructural analysis of hematopoietic colonies derived from human peripheral blood. Journal of Cell Biology, 63(3):855– ??, December 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/63/3/855.

Zurier:1973:MLE

[ZHW73] Robert B. Zurier, Sylvia Hoffstein, and Gerald Weissmann. Mechanisms of lysosomal enzyme release from human leukocytes. Journal of Cell Biology, 58(1):27–??, July 1973. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/58/1/27.

Zuckerman:1973:UMC

[ZKH73] Bert M. Zuckerman, Marian Kisiel, and Stanley Himmelhoch. Unusual mitochondrial cristae in the vinegar eelworm. Journal of Cell Biology, 58(2):476–??, August 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/58/2/476.

Zakai:1974:FEO

[ZLK74] Nehama Zakai, Abraham Loyter, and Richard G. Kulka. Fusion of erythrocytes and other cells with retention of erythrocyte cytoplasm. Journal of Cell Biology, 61(1):241–??, April 1974. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/61/
1/241.

Zobel:1973:ESC

[Zob73] C. Richard Zobel. Effect of solution composition and proteolysis on the conformation of axonemal components. Journal of Cell Biology, 59(3):573–??, December 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/59/3/573.

Zampighi:1973:FSS

[ZR73] Guido Zampighi and J. David Robertson. Fine structure of the synaptic discs separated from the goldfish medulla oblongata. Journal of Cell Biology, 56(1):92–??, January 1973. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/56/1/92.

Zalik:1972:CSC

[ZS72] Sara E. Zalik and Vi Scott. Cell surface changes during dedifferentiation in the metaplastic transformation of Iris into lens. Journal of Cell Biology, 55(1):134–??, October 1972. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/55/
1/134.

Zucker:1974:TPP

[ZTB74] Marjorie B. Zucker, Walter Troll, and Sidney Belman. The tumor-promoter phorbol ester (12-0-tetradecanoyl-phorbol-13-acetate), a potent aggregating agent for blood platelets. Journal of Cell Biology, 60(2):325–??, February 1974. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/60/2/
325.

Zagury:1970:ISS

[ZUJP70] Daniel Zagury, Jonathan W. Uhr, James D. Jamieson, and George E. Palade. Immunoglobulin synthesis and secretion. Journal of Cell Biology, 46(1):52–??, July 1970. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/46/1/52.

Zalta:1971:IN

[ZZS71] J. Zalta, J.-P. Zalta, and R. Simard. Isolation of nucleoli. Journal of Cell Biology, 51(2):563–??, November 1971. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/51/
2/563.